A Complete Bibliography of Publications in the Revue de l’Institut International de

Statistique / Review of the International Statistical Institute

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

30 October 2024 Version 1.00

Title word cross-reference

1, 2 [Tik67]. 2 × 2 [Har65]. n [Hab62]. r = 11 [Cla68b]. R^{2} [Mah69, Pra71]. s [Hab62]. ∑_{1}^{n} P_{h}^{i}σ_{h}^{j} [Ekm63]. T [vE61, Sid64]. W [Pur69].

-distribution [vE61].

1 [Iva69, Kri69, Lep68, Mor71b, Wil68]. 1919 [dW54]. 1931-’32 [Der53b]. 1933 [Ano33t]. 1934 [Dre36, G.34a]. 1936 [Sav36]. 1938 [Nix53]. 1940 [Der53b]. 1940-’41 [Der53b]. 1941 [Tin50]. 1946 [Hen54]. 1947 [Ric47b, Sto50]. 1949 [Ric49]. 1950 [Nix58a]. 1951 [Ric51, Dep57]. 1952 [Zwi53]. 1952-’53 [Zwi53]. 1953 [Ric53, Cal62b]. 1954 [Cal62a]. 1955 [Dar55]. 1956 [Der58]. 1957 [Ano57f, DG57, Nix61a]. 1958 [Dar58b, Cal62d, Gea63a, Cla61]. 1960 [Ano60k, Tin44, Nix62b]. 1961 [Ano60i, Nix64]. 1962 [Fou66]. 1963 [Ano62l, Ano62m, Mew65]. 1964 [Ric64, Haw67, Nix67]. 1965 [Min69, Nai67]. 1966 [Jar69]. 1968 [Ano69d].

2 [Bil65, Bil70, Gre67].

31.12.1953 [Für53]. ’32 [Der53b]. 32e [Ano60l]. 32ème [Bol60b]. 32nd [Ano60l, Bol60a, Bol60b, Bol60c]. 33ème [Bol62]. 33me [Ano60i]. 33nd [Ano60j]. 33rd [Ano60i, Bol62]. 34ème [Ano62l]. 34th [Ano62l, Ano62m]. 35ième [Ano64a]. 35th [Ano64a, Ano64b]. 36ième [Ano66a]. 36th [Ano66a, Ano66b]. 37ième [Ano68a]. 37th [Ano68a, Ano68b]. 38ième [Ano70a]. 38th [Ano70a, Ano70b].

’41 [Der53b].

50 [S.34]. ’53 [Zwi53]. 58 [Som62].

746/746a [Gou62]. 746a [Gou62]. 747/747a [Gou62]. 747a [Gou62]. 75 [Pét66]. 75th [Bol60a].

âges [Hub36b].

A. [Har67a]. Abgrenzung [Dre38]. Abramowitz [Hem65]. ábrázolásban [K.35]. Abrikosov [Wer67]. Absorption [EG61]. Abstract [RMK^{+}70a, Thi63c]. Abstracts [Jar69, Min69]. Acceptance [Rid64]. accident [vZ36]. Accidents [EG65, vZ36, vZ38]. According [Gin34, Liv36, Sau68, Sav33]. account [SL61]. Accounting [Der60, Ken66b, Lau64, SUD50, TI66]. Accounts [Bje69, Gil53, Les67, Der68]. Ackoff [Dem63]. Activities [Hun71, Per41]. Activity [Sau63a]. actuel [Coc50, Gin50b]. actuelles [Hos37]. Adam [Ham60]. Address [Bol60c]. Adelman [Zim69]. Adjusted [Mah69, Pra71]. Adjusting [Thi59]. Adjustment [Bor66, Tie69]. Adjustments [Gar71]. Administration [Ano67f, Ano67g, Ano67h, Ano68i, Ano68j, Ano68k, Ano69e, Ano69f, Ano69g, Ano70j, Ano70k, Ano70l, Ano71h, Ano71i, Ano71j, Ano33h, Ano34g, Ano34h, Ano34i, Ano35h, Ano36i, Ano37i, Ano38m, Ano38n, Ano38o, Ano39h, Ano39i, Ano40f, Ano41f, Ano41g, Ano43e, Ano43f, Ano44c, Ano47b, Ano48c, Ano49d, Ano49e, Ano50c, Ano50d, Ano51e, Ano51f, Ano51g, Ano52d, Ano53f, Ano53g, Ano54c, Ano55c, Ano56c, Ano57c, Ano58c, Ano59b, Ano60c, Ano60d, Ano61a, Ano61e, Ano61f, Ano61g, Ano62d, Ano62i, Ano62j, Ano63e, Ano63f, Ano63g, Ano64h, Ano64i, Ano65g, Ano65h, Ano65i, Ano66k, Ano66l, Ano66m, Bow46, Bun33, Bun37, Dun40, Dun42, Gin33, God37, God39, Hec34, Hec35, Hub38a, Ide36a, IM52, Mey42, Mor35a, Ric36, Ric45a]. administration [Stu35, WI37, dL38, dS36]. Admission [Bit71]. Adolf [Sae36]. Adolphe [Jul38b, Jul38c]. adopted [Ano36y, Ano37v, Bun38a]. adoptée [Bun38a]. adoptés [Ano36y, Ano37v]. adoption [Bun38a]. Advanced [Lan70b, Rao68, Wil59, Mor71d]. Advisory [Nai67]. affaires [Bar46, Bar43]. Affairs [Ano63q]. Africa [Sco67a]. African [Sha63]. after [Ano33i, Woo68]. Age [Ric59a, Liv36]. ages [Hub36b]. Aggregation [All55a, Gor67]. Aggregative [BJT63]. Agrarstatistik [Dal58]. agricole [Har36]. Agricultural [Gou63, Lou67, SEM^{+}70, Har36]. Agriculture [Cur70, JHH61, Dor36, Dal62, Pan63]. agronomie [Mar61]. aid [Wag38]. aide [Mor41]. Aids [AHW^{+}71]. Aitchison [Fin57, KG67]. ajustement [Thi59]. Akademie [Ver61]. al [Für61]. Albert [Jul38a, Weg57]. Alder [Rén61]. aléatoire [Fré48]. aléatoires [Dre69, Fré42, MK70, Gea38, Wol51]. Alessandro [Bar62]. Alexander [Lan62, Nix62a]. Alexandre [Thi63a]. Alfred [Wol70]. Algèbre [Dag64]. Algol [Sza64a]. Algorithms [ML69]. Allen [Der63]. Allgemeine [Gou62, Mad46, Sag58]. Allied [Hun67, Tra68]. Allocation [Kul63, Lou67]. almost [Fré58]. Aloys [Thi63b]. alte [Win38b]. Alternatives [Dow71]. Alterungsprozesses [Fla65]. America [Cro69]. American [G.48, Mos71c, Per41, Stu41, Ric59a]. Analyse [Ben66, BP50, Dar53, Dag67, Gri66]. Analyses [Tör68]. Analysis [All55b, AB66, Dal65, Dok66, Dur53, EG65, Fis62, Gna67, Gor67, HR68, Kap67, Key67, Kin64, Lau64, Lea68, Mok62, Muh65, Par67, Pra53, PA54, Pri36, Qua71b, Rao68, RMK^{+}70a, Sch71a, Sha56, Sto58, The53, Vos65, Wal67, Ben66, Dar53, KG67, Sto48, BP50, Nor67, And68, Fre71, Gou53d, Hen61, Mac70, Mit71, Rao58, Sch59, Tia68]. Analytic [Ple67]. analytique [Mal69]. Anderson [Whi69, F.36, Kel60]. Andree [Roy69]. Andrews [Ken61]. anfallende [WW38]. Annals [G.34a]. année [Pto34]. Années [S.34]. Anniversary [Bol60a]. Announcement [Ano53a, Ano56a, Ano57a, Ano58a, Ano66c]. Announcements [Ano48a, Ano49a, Ano51a, Ano52a, Ano53b, Ano53c, Ano54a, Ano55a, Les52, Ric47a]. Annuaire [Mew65]. anthropométrie [Aue34]. anthropometry [Aue34]. anticonceptionnelles [Gin42]. António [Gue62]. Anwendung [Gou62, Pét66]. août [Ano57f, Ano60i, Ano62l, DG57]. Apercu [Cla61, Ano35l]. aplicaciones [Zar62b]. appendice [For67]. Appennino [G.34c]. applicable [BMO^{+}70, Thi54]. applicables [BMO^{+}70]. Application [Bit71, Cra66, Day49, EG65, JHCS70, KC70, Par62, Vra64, Žar56, Fré46, Par62]. Applications [Bol64, Dal62, Gre67, Kap67, Kri69, RMK^{+}70a, Sco67b, Voi70, Dag67, Dre69, Oni36, Pre68, Thi68, Sav64, Bar61, Cac71, Dem54, Der71a, For68, KLM67, Kit62b, Wol55, vdP65, Dav70, Dri67, Lau64]. Applied [Bri64, Bri68, Bri71, Dem61, Dem63, Gil62, Sch71b, Som67b, Her42, dS33, Tia68]. appliquée [Bar43, Bar46, Her42]. appliquées [dS33]. Appraisal [Gou64, Lau64]. Approach [Lau64, Mad56, Ped69, Pfo66, Rob63, Rot63, Vin56]. Approaches [Vaj65]. Approximate [Bri64]. Approximation [Rév70]. approximations [vE61, Luk71]. Arab [Nic54]. Arbeitsgebiet [Für53]. arbitraires [G.38]. arbitrio [Vin36]. arbre [Dre69]. Areas [Man68a, Ric56, Som62, Zwi60]. aree [Für61]. Armengaud [vdB62]. Arrow [Gil62, Vad59]. Arthur [Gou45, Gou51c, All57]. Article [Gin39a, Wag38]. áruforgalma [dKT38]. Ascertain [Pri36]. Ash [Ios71, Upp68]. Ashton [vBW34]. Asimakopulos [Nix65]. ASN [Abr54]. Aspects [LNS^{+}70, Mah69]. Assay [Mar66]. Assays [Sen71]. Assessment [Ham67]. assessments [Gin45]. Association [Ano62r, Gil51, Les67, Mos71c, Nix62b, Nix67, Sha63, Nix64, Nix65]. Assured [SEM^{+}70]. Astronomy [Dar58a]. Asymptotic [Bri64, NvdP66]. Athènes [Stu36]. Athens [Stu36]. Atlante [Ano33a]. Atlas [Ano33a]. attached [Dre69, Gin39b, Gin39a]. attachées [Dre69]. Attempt [Kev67, Fré50]. attempts [Fré45]. attrazione [Für61]. Attributes [Abr54, Hal67, Ham50, Sha56]. Audience [HDPD71]. Aufgaben [Bur37, Ver58]. Aufhellung [Böh38]. Aufsatze [Wey38]. August [Ano57g, Ano57f, Ano60i, Ano62l, Ano62m, Ano69d, DG57, Ano60j]. Augusto [Lop56]. Ausführung [Dre36]. Austin [Mar69]. Auswerten [Ham54]. Auswertung [Ven68]. Auswirkungen [Fla65]. Authorship [Kha66]. automation [Ben64a]. automatique [Lin66b]. automatisme [Ben64a]. autoregressifs [Mal61]. autoregressive [Mal61]. Autres [Ano61a, SL61]. aux [Ano33b, Ano33c, Ano33d, Ano34a, Ano34b, Ano38a, Ano38b, Ano39a, Ano40a, Ano41a, Ano46a, Ano48a, Ano49a, Ano51a, Ano52a, Ano53b, Ano53a, Ano53c, Ano54a, Ano55a, Ano56a, Ano57a, Ano58a, Ano66c, Bar43, Bar46, Her44, Hub36b, Les52, Ric47a, Stu38, Thi54, Tin47b]. Available [Dix71a, Nix38]. Avant [Hub37, Met33a, Met33b, Met35, Met45]. Avant-propos [Hub37, Met33a, Met33b, Met35, Met45]. Average [Kev67, Ric56, Fré41, Her40, Ham53]. Avis [Ano33b, Ano33c, Ano33d, Ano34a, Ano34b, Ano36y, Ano38a, Ano38b, Ano39a, Ano40a, Ano41a, Ano46a, Ano48a, Ano49a, Ano51a, Ano52a, Ano53b, Ano53a, Ano53c, Ano54a, Ano55a, Ano56a, Ano57a, Ano58a, Ano66c, Les52, Ric47a, Ano37v]. Axiomatic [Pfo66]. ayant [Dug69]. Aylmer [Ano62k, Yat62]. Azorin [Zar62b].

B [Azo67, Ben67b, Der53a, Gei69, Gou56, Har65, Ios71, Joh71, Kit62a, Kit62b, Lea68, Lin61b, Mac70, Mor71d, Muh69, Nai67, Nel70, Pér71, Pey61, Pey62, Rao61, Rén61, Sun71, Tag57, Thi63c, Upp68, dW52]. Bachi [G.34b, Ped69]. Bacteriophage [Kri67]. Baggaley [Wet70]. Bailey [Sco67b]. Baisse [BL38, LS38, Win39]. balance [Ano34j, Der58, Tie69]. Balanced [CL68, Gué65, Sil64, Tak62, Cla68b, Mas67]. Bancroft [Fin54]. Band [Gou62, Ver61, vdB63]. Baranow [Gou51a]. Barberi [Tag57]. Barsy [Thi66b]. Bartlett [For68, Wol55]. Barton [Har71, Sve63, Tra68]. Bas [MT34, Stu38, Tin47b]. base [dDBL34]. Based [Dix71a, Dok66, Hal67, NT70, vE55, dDBL34, Har35]. basée [Har35]. bases [vZ38, vZ38]. Basic [Coh66, Gou56, Ney57, Sny34a, Zwi71, Ziz33, Der63, Ios71, Pfa68]. Basis [Rid64, Bra69, Lin56]. Bauer [Lep68]. Bayes [Kru71, Cor67a, NT70, Rob63].

Bayesian [Mar71, Rén67]. Bd [Lep68]. be [Dre69]. Beach [Nix62b]. bearbeiteten [Für53]. Beard [Hoe70]. Becker [Muh65]. Beckmann [Der71b]. Bedeutung [WW38]. before [Ano33i]. Begriff [Thi33]. Begriffe [Ziz38]. behalf [Ide38, Sch36]. Behavior [Ney57, Sch71a, Vaj65]. Behavioral [Som67b]. Behaviour [Bri64, Sch68b, The55]. being [Bun38a]. Beitrag [Wag38, Win42]. Beiträge [Mee35, Sae36]. Belgique [Jul34, Fed69, Pre68]. Belgium [Jul34]. Bemerkungen [Wey38, Win44]. Benes [Lin64]. Benini [Gin57b]. Benjamin [Mac70, Muh69]. Bennett [Har65]. Bericht [Böh38]. Berkeley [Dar58a, Gep58, Kha58, Le 69]. Bernard [Lin66b]. Berne [Ric49]. Bernstein [Gin57a]. Berücksichtigung [Gou62]. Beschlusses [Dre36]. Besondere [vdB63]. besonderer [Gou62]. Beta [Hem69]. Beta-Function [Hem69]. Betriebswirtschaft [Ham60]. Betteley [Hun67]. between [Gea63b, Ide38]. Beveridge [Cam64]. Bevölkerung [Ver61, Win42, Win44]. Bevölkerungsstatistik [Bur37, vdB63]. Bevölkerungswachstums [Win38b]. Bharucha [Bar61]. Bias [Dem61, Mos51, KG67]. Bibliographic [Ham67]. Bibliographie [Ano33e, Ano33f, Ano33g, Ano34c, Ano34d, Ano34e, Ano34f, Ano34k, Ano34l, Ano35a, Ano35b, Ano35c, Ano35d, Ano36a, Ano36b, Ano36c, Ano36d, Ano36j, Ano37a, Ano37b, Ano37c, Ano37d, Ano38c, Ano38d, Ano38e, Ano38f, Ano38g, Ano39b, Ano39c, Ano39d, Ano40b, Ano40c, Ano41b, Ano41c, Ano42a, Ano42b, Ano43a, Ano43b, Ano44a, Ano45a, Ano46b, Ano47a, Ano48b, Ano49b, Ano49c, Ano50a, Ano50b, Ano51b, Ano51c, Ano51d, Ano52b, Ano52c, Ano53d, Ano53e, Ano54b, Ano55b, Ano56b, Ano57b, Ano58b, Ano59a, Ano60a, Ano60b, Ano61b, Ano61c, Ano61d, Ano62a, Ano62b, Ano62c, Ano63a, Ano63b, Ano63p, Ano64c, Ano64d, Ano65a, Ano65b, Ano65c, Ano66d, Ano66e, Ano66f, Ano67e]. Bibliographies [Lan69, Lan70a, Lan71]. Bibliography [Ano47a, Ano48b, Ano49b, Ano49c, Ano50a, Ano50b, Ano51b, Ano51c, Ano51d, Ano52b, Ano52c, Ano53d, Ano53e, Ano54b, Ano55b, Ano56b, Ano57b, Ano58b, Ano59a, Ano60a, Ano60b, Ano61b, Ano61c, Ano61d, Ano62a, Ano62b, Ano62c, Ano63a, Ano63b, Ano63p, Ano64c, Ano64d, Ano65a, Ano65b, Ano65c, Ano66d, Ano66e, Ano66f, Ano67e, GV63, Lan69, Lan70a, Ano33e, Ano33f, Ano33g, Ano34c, Ano34d, Ano34e, Ano34f, Ano35a, Ano35b, Ano35c, Ano35d, Ano36a, Ano36b, Ano36c, Ano36d, Ano37a, Ano37b, Ano37c, Ano37d, Ano38c, Ano38d, Ano38e, Ano38f, Ano38g, Ano39b, Ano39c, Ano39d, Ano40b, Ano40c, Ano41b, Ano41c, Ano42a, Ano42b, Ano43a, Ano43b, Ano44a, Ano45a, Ano46b, Lan71, G.48, Gea63a, Zwi53]. bicentenaire [Hor69b]. bicentenary [Hor69b]. bien [Thi66a]. bilan [Hos38]. Bilder [K.35]. Billingsley [Kin69, Lin61a, Rév66]. Binomial [Blo65, Ham53, Luk71, Hos35]. binominales [Hos35]. Bio [Sen71]. Bio-Assays [Sen71]. Biologen [Ano71a, Wis71]. Biological [Gou63, Mar66, HT71, Fin66]. biologie [Mar61, Ves68b, Dri67, Gri66]. biologistes [Ves67]. Biologists [Man68b]. Biology [Gep58, Sch71b, Wil68, Fin54]. Biometrics [Gin50b, Nai71]. Biométrie [Coc50, Gin50b]. Biometrika [Gou55]. Biometry [Cur70]. Biostatistics [Bar69b]. Birkhoff [Elf66]. Birth [BL38, LS38, MH50, JL43a, JL43b, Win38a, Win39]. Birthday [Dun64, Rao63a]. births [Sav33]. Bivariate [Ben67a, HT71]. bivariate-compounded [HT71]. Blanc [Col64]. Blind [Bic68a, Sae36]. Bliss [Wil68]. Block [CL68, Gué65, RS68, Sil64, Tak62, Cla68b]. blocks [Mas67]. blocs [Gué65, Mas67]. Bloemena [Gan66]. Blumen [Ste57]. Bodio [Jul38b, Jul38c]. Boehm [Wag38]. Bojorquez [Loy69]. Bolch [Sun71]. Boldrini [Ugg69]. Book [All50, All55b, All55a, Amu68, And68, And44, Ano64g, Ano66j, Ano67d, Ano68c, Ano68h, Ano70h, Ano70i, Arm70, AB66, Azo67, Bar65a, Bar67, Bar69a, Bar68, Bar43, Bar46, Bar66a, Bar61, Bar69b, Bar66b, Bea69, Ben64b, Ben65, Ben67b, Bic68a, BF68, Bil64, Bil65, Bil70, Bir59, Bje62, Bje69, Blo65, Bor66, Box61, Box70, Bra67, Bra69, Bri68, Bri71, Bru70, Buc69, Buc71, Bun41, Cac71, Cal62a, Cal62c, Cal62d, Cal62b, Can55, Car69, Cla58, Cla61, Cla68a, Cla71, Coh66, Col64, Cra68, Cra69, Cur70, Dag64, Dag67, Dag70, Dag71, Dal58, Dal62, Dal63, Dal65, Dal68a, Dal69, Dar58a, Dav62, Dav67, Dav69, Dav70, Dem54, Dem61, Dem63, Dem66, Dem68a, Dem62, Dep57, Der53a, Der53b, Der54a, Der54b, Der57, Der58, Der60]. Book [Der61a, Der61b, Der63, Der65, Der68, Der71a, Der71b, Doo71, Dow71, Dri67, Dum70, Dur53, Dur58, Eic65, Elf66, Elf69, F.36, Fal69, Fed69, Fin54, Fin57, Fin66, Fla65, For67, For68, Fou66, Fre71, Für53, Für61, G.34a, G.34b, G.34c, G.38, G.48, Gan66, Gan67, Gan71, Gea38, Gea63a, Gea65, Gei69, Gep58, Gil51, Gil62, Gir66, Gna67, God68, Gol60, Gor67, Gou45, Gou51a, Gou51b, Gou51c, Gou53b, Gou53c, Gou53e, Gou53d, Gou53a, Gou54a, Gou54b, Gou54c, Gou55, Gou56, Gou57, Gou58, Gou61a, Gou62, Gou63, Gou65, Gou68, Gou70a, Gou70b, Gov66, Grd64, Gre67, Gri66, Gut68, Hal57, Ham54, Ham58, Ham60, Ham67, Har67a, Har70, Har68, Har67b, Har65, Har71, Haw67, Hay70, Hec69, Hem65, Hem69, Hen68]. Book [Hen61, Hen54, Hoe70, Hor69a, Hub42, Hub66, Hun67, Ill37, Ios71, Iva69, Jar69, Joh54, Joh71, Joh62, JSH65, Jos53, Jua71, K.35, Kád68, KLM67, Kap67, Kem71, Ken61, Ken65, Ken71, Ken67, Kha58, Kha66, Kin64, Kin69, Kit62a, Kit62b, Kit64, Kit68, Koc63, Kol57, Kol60, Kol65, Kre65, Kri67, Kri69, Kru71, Kün65, L.36, Lai67, Lam68, Lan62, Lan63, Lan70b, Lau64, Le 69, Lea68, Lep68, Les67, Lin56, Lin66a, Lin67, Lin69, Lin71a, Lin61b, Lin61a, Lin64, Lin66b, Loy70, Luk71, Lüt65, Mac70, Mad46, Mal68, Mal69, Man68b, Mar68, Mar71, Mar61, Mar66, Mar69, McC66, Mee40, Men70, Mew65, Mih69, Min69, Mit71, Mok62, Mor66, Mor69, Mor71b, Mor71c, Mor71a, Mor71d, Mos54, Mos71a, Muh65, Muh69, Mui65, Nai67, Nai71, Nel70]. Book [Neu67, Nix53, Nix54, Nix57, Nix58a, Nix61a, Nix62b, Nix64, Nix65, Nix67, Nor67, Nor66, Oak71, Ohl61, Oma67, Pai58, Pan63, Par66, Pea62, Ped69, Pér71, Per41, Pét66, Pey61, Pey62, Pey63, Pfa68, Ple67, Pre68, Qvi66, Rao58, Rao61, Rao68, Rao69, Rén61, Rén63, Rén67, Rév66, Rév70, Rid64, Rio65, Rop70, RR68, Roy69, Rut65, S.34, Sae36, Sag58, Sas71, Sau63a, Sav36, Sax66, Sch71a, Sch59, Sch67, Sch68a, Sch65, Sch66, Sch68b, Sco67b, Sha63, Shr65, Som62, Som67a, Som67b, Spr71, Sri70, Ste67, Ste57, Ste60, Sto50, Sto56, Sto58, Stu41, Sun71, Sve63, Syk68, Sza64a, Sza64b, T.36, Tag57, Teg62, dKT38, The69, Thi63c, Thi64, Tia68, Tie69, Tin47a, Tin50, Tör63a, Tör68, TI66, Tra68, Upp68, Vad59, Vaj65]. Book [Ven68, Ver58, Ver61, Ver63, Ves67, Ves68a, Ves68d, Ves68c, Ves68b, Ves68e, Ves68f, Ves68g, Ves70a, Ves70b, Vin56, Vos65, Vos69, Vos71, vBW34, Wal67, Wal66, Weg64, Wer67, Wet70, Whi66, Whi69, Wid67, Wij69, Wil59, Wil68, Win57, Wis71, Wol51, Wol55, Wol67, Wol68, Woo68, Zar62a, Zar62b, Zar66, Zar69, Zim69, Zwi53, Zwi60, Zwi71, dO67, dS71b, dS71c, dW52, dW54, vR70, vR68, vZ66, vdP65, vdP70, vdB69, vdB68, vdB61, vdB62, vdB63]. Boole [Dag64]. booléenne [Dag67]. Boot [Bra67]. Borel [Dar56]. born [Sau68]. bornes [Fré60]. Both [Aig70]. bourgeois [Cal62d]. bourgeoises [Cal62b]. Boustedt [Ver58]. Bowes [Zwi53]. Bowley [All57]. Bracken [Rut65]. Bradford [Mar69]. Brambilla [Der61a, Bar69a]. Branch [Ano63q]. Brazilian [Man68a]. Breakthroughs [Ney62]. Bresciani [Gia64]. Bretherton [Blo65]. Britain [Hen54]. British [Gou57, Thi63c, Joh62, Nix54]. broadcasting [Ide38, Ide36c]. Brookes [Hun67]. Brown [Fin57]. Brownian [Dug69]. Browniennes [Dug69]. Brownlee [Bar66b]. Brüschweiler [Kol56]. Brussels [Dar58b]. bruts [Her34]. Bruxelles [Dar58b]. Bruyn [Box70]. Buckland [Ham67]. Budapest [Fou66, K.35, dKT38, dKT38]. budgets [Hos37, Jul34, Nix36, All55b]. Bugg [Gou70a]. Building [Cra69, Nor70, Nys36, Nys38]. Bulmer [Cla68a, Lin66a]. Bundesamtes [Für53]. Bundesbehörden [Für53]. Bundesrepublik [Sch68a, Ver61, Win57]. Bundesstatistik [Für53]. Bunker [Arm70]. Bureau [Ano57g, Ano57f, Ano60m, Ano60k, Ano63l, Ano63h, Ano63q, Der65, Har67a, Hem65, Hor68a, Hub41, Nor70, Sto56, vdB61, Ano35l, MT34]. Bureaux [Ano35e, Ano35f, Ano35g, Ano36e, Ano36f, Ano36g, Ano36h, Ano37e, Ano37f, Ano37g, Ano37h, Ano38h, Ano38i, Ano38j, Ano38k, Ano38l, Ano39e, Ano39f, Ano39g, Ano40d, Ano40e, Ano41d, Ano41e, Ano42c, Ano42d, Ano43c, Ano43d, Ano44b, Ano45b, Ano46c]. Burgdoerfer [Hor68b]. Burgdörfer [Mee40]. Burkhardt [Pét66]. Business [And52, BJT63, Pri36, SS63, Sto56, Sun71, The52, Hoe70, Bil64]. Buttingha [Gou61b].

C [All50, Amu68, Bar69b, Box70, Bra67, Dal63, Der53b, Der71b, Dow71, Dri67, Dum70, Fed69, Fin57, Fin66, Gou53e, Gou53d, Gou54c, Ham58, Har68, Hun67, Joh62, KLM67, Kit68, Man68b, Nix62b, Nor67, Pér71, Qvi66, Rao63a, Ric47b, Sax66, Som67a, Som67b, Ste60, Vos69, Wil68, Wol68, vdP65]. cahier [And44]. calcolo [Mui65]. Calcul [And44, Ano68c, Dre69, Dug55, Fré41, Fré46, Hos35, Hub36b, Ide36b, Ves68d, Wol51, For67]. Calculating [Gin34, Pto34, Hub36b]. calculation [Dre69, Dug55, Fré41, Fré46, Hos35, Ide36b]. calculer [Pto34]. Calculs [Ste67]. Calcutta [Ric51]. called [Fré34, Fré50]. Calot [Ano68c, Ves68d].

Cambridge [Zwi53]. Campbell [Man68b]. can [Böh38, Dre69]. Canada [Mar51, Cla58]. Canadian [Nix64, Nix65, Nix67, Nix62b]. cancer [Böh38, Stu38, Stu38]. capable [Bun38a]. capacite [Hos38]. Capital [Der54c, Muh65]. caractéristique [Gin36]. Carl [Kol56]. Carlo [Ken65, Qvi66]. Carlström [Dem66]. Carolina [Haw67]. Carrés [Bar65b, BKB65]. Carried [Som62, dDBL34, Har36]. Carryover [Cra33]. carta [Für61]. Carvallo [Dag64, Dag67]. cas [Fré43, G.38, Gin39b, Gin39a]. Case [Wol58, Fré43, Gin39b, Gin39a]. Cases [Har67b]. Categories [Ric56]. Catholics [Nix61c]. Causal [Wol58]. Causalité [Mar33]. cause [Led55, SL61]. Causes [BL38, LS38, Hub38c, Led55, SL61, Win38a, Win39]. Celebration [Bol60a]. Census [Boy58, BW69, GF55, Hen54, Hor68a, Mar51, Voi70, Mor41, HU36]. Censuses [HW70, Dem68a]. Centenary [vBW34]. Center [Bli49, Lin62]. Central [MT34, Nor70, vE61, MT34]. Centrale [Sav36]. Centre [Gan66, Sch71b]. Centroids [Har71]. Certain [Cra66, Luk70, The51, vdP70, Dar45, Fré43, Fré58, Gal51, Her40]. certaine [Fré58]. certaines [Dar45, Fré43, Her40]. certains [Gal51, Thi66a]. ces [Div43]. ceux [Thi59]. Ceylon [Pai58]. Ch [Gol60]. chaîne [G.38, Oni36]. Chain [Bit71, Wol58, Oni36, Lin61b]. Chakravarti [Bri68]. Chance [Pfa68, Hos35, Sve63]. Chang [Oak71]. Change [She71, vdB61]. Changes [Cla58, Flu38, Der54b]. Chapman [Nix53]. character [Ziz33]. Characteristic [Ham67, JHCS70, Mor71d, Žar56, Gin36]. Characteristics [Nix61c]. Characterization [Luk70]. Charge [Hos38]. Charles [Her61, Nix58b, Rén60, Rue54]. Chartier [Gou58]. Chemical [Hal57, Hun67]. Chemistry [Hun67]. chercheurs [Dug65]. Chevry [Sau71, Ver63]. chez [Ves70a]. Chiang [Bar69b]. children [Sau68]. China [Wil37]. Chover [Mih69]. Christ [Der71b, Wol68]. circulation [vZ36, vZ38]. circulatory [Wag38]. Cities [Zwi71, Büc36]. città [G.34b]. City [Bow33]. Clark [Tin44]. Clarke [Joh71]. Class [MH50, Pur65, Sär70]. classic [Fré50]. Classical [Buc69]. Classification [Ano63p, DF68, GV63, GHM70, HR68, Div43]. Classifications [Bar71, dS38]. classifying [Div43]. classique [Fré50]. Classroom [Mos63]. Cloud [CAON70, MCC70]. Clumping [Sri70]. Cluster [Tör63b]. Clyde [Ben64b]. coût [Cal62a, Har35, Sau48, Thi54, vZ33]. Coale [Sch66, Wid67]. Cochran [Coc50, Dal68a, Gin50b, Kol65]. coefficient [Fré34, Fré36b, Fré36a, Fré50, Gin36, Pie37]. Cohen [Loy70]. Colin [Tin44]. Collected [Whi69]. Collection [Dag64]. collective [LS71]. College [Hun71]. Colonial [Nix61a]. Combination [Pur65, Spr71]. combinatoire [Sor71]. Combinatorial [Sch59, Tak64, Sor71, Sve63]. combinatorique [Dav70]. Combined [Ric56]. Comité [Hub39, Hub41]. comme [Gin36]. Comment [Hos46, Sav41, The53, Gin37b, Gin37a]. commentaires [Ste67]. commentées [Mor71a, Ves68g]. Comments [CL68, Der54c, HDPD71, Hot50, Win44]. Commerce [Hem65]. commerciale [Für61]. commercio [Für61]. Commission [Dor36, Ide38, Sch36, Bun38b, Fré36a, Böh38, Cra33, Dor36, Nys38, Sch36]. Committee [Hub39, Hub41, Mos71c, vdB61, Nix36]. Commodities [Cra33]. Commodity [Ric56]. Commonwealth [Hor68a]. Communications [Ano33h, Ano34g, Ano34h, Ano34i, Ano35h, Ano36i, Ano37i, Ano38m, Ano38n, Ano38o, Ano39h, Ano39i, Ano40f, Ano41f, Ano41g, Ano43e, Ano43f, Ano44c, Ano47b, Ano48c, Ano49d, Ano49e, Ano50c, Ano50d, Ano51e, Ano51f, Ano51g, Ano52d, Ano53f, Ano53g, Ano54c, Ano55c, Ano56c, Ano57c, Ano58c, Ano59b, Ano60c, Ano60d, Ano61e, Ano61f, Ano61g, Ano62d, Bow46, Bun33, Bun37, Dun40, Dun42, God37, God39, Hec34, Hec35, Hub38a, Ide36a, IM52, Mey42, Mor35a, Ric36, Ric45a, Stu35, WI37, dL38, dS36, Ano37v, Ano61a, Ano33h, Ano34g, Ano34h, Ano34i, Ano35h, Ano36i, Ano37i, Ano38m, Ano38n, Ano38o, Ano39h, Ano39i, Ano40f, Ano41f, Ano41g, Ano43e, Ano43f, Ano44c, Bow46, Bun33, Bun37, Dun40, Dun42, God37, God39, Hec34]. Communications [Hec35, Hub38a, Ide36a, Mey42, Mor35a, Ric36, Ric45a, Stu35, WI37, dL38, dS36]. como [dDBL34]. Comparability [Der54c, Nys36]. Comparaison [Fré40, vZ33, vZ38]. comparative [Hos37]. compared [Har36]. comparées [Har36, Hos37]. Comparing [Dur58, vE55, vZ38]. Comparison [Der53a, Gar71, Rao66, vZ33, Fré40]. Comparisons [Shr65]. Compendium [Per41]. Competition [Fra68]. compilation [Ide38, Sch36]. Complete [Gla62]. Components [Spj68, Zar65]. Composite [CS70, Her34]. composites [Her34]. compounded [HT71]. Comprendre [Lin66b]. compte [SL61]. Computation [Bli49]. Computations [Gou53c]. Computer [Dix71a, Ken66a, Nor67, Qua71b, Roy69]. Computerized [Dal71]. Computers [Qua71a, vH66]. con [Für61]. Concept [Abr54, Ney57, Thi33]. conceptions [Bol37]. Concepts [Gou56, Pfa68]. Conceptual [Ken66b]. concernant [For60]. concerning [Ekm69, Nix38, Per41]. concordance [Gin36, Pie37, Pie37]. Conditions [Bow33, Nic54, Hub42]. Conference [Nai67, Nix62b, Ohl61, Sau63a, Nix64, Nix67]. conferences [Ano35l, Ano35l, Nix65]. Confidence [Bri64, EG65, Gar71, Ham53, Nad68, Spj68]. Congested [Rot63]. Congestion [Haw67]. coniugata [Liv36]. Conjoncture [MT34]. conjuncture [MT34]. Connor [McC66]. connus [Thi66a]. Consequences [BL38, Fin56, LS38, LS38, Win39, Win39]. considerations [vZ33, vZ33]. consommation [Hub42]. constituée [Dor36, Sch36]. constituted [Dor36, Sch36]. Constitution [Nix35, Nix52]. constitutions [Gin37b, Gin37a]. Construction [GV63, Mas67, McC63]. Consumer [Kon58, Sch68b, The55, Tör68, dW54]. Consumption [Cra33, Mah69, Tay68, Sto56]. contagion [Div43]. Contagious [Buc69]. contamination [Div40, Div43]. contaminations [Div40]. Content [Kev67]. Continental [Mos63]. Contingency [Har65]. contingentements [Tin47b]. Continuing [Hun71]. Continuous [Buc71, Mit71]. contraceptive [Gin42]. contribuable [Hos38]. Contribution [Thi68, BPH52, Her52, Win42, Thi68]. Contributions [Dar58a, Gep58, God65, Kha58, Sav64, Mee35, Kit68, Rao61].

Control [Bic68b, Boy58, Cro69, DF68, Lai67, Min69, Nor66, TI66]. Controlled [SEM^{+}70]. convergence [Fré58, Kin69]. Conversational [Qua71b]. Convex [Wal66]. Cooley [Bea69]. Cooper [Nel70]. Coordination [Cro69, Gin43]. Corrado [Cas65]. Correcting [Bos66]. Correction [Ano68d, Ano70d, Ano70c, Ano71a]. Correlated [Ham58]. Correlation [Ben67a, DS71a, Vra64, Mar33, Fré34, Fré36b, Fré36a, Fré50, Gin36, Pie37, Fré34, Fré36b, Fré36a, Fré50, Fré57, Gin36, Fré57]. Correlational [Rao58]. correspondance [Jul38b, Jul38c]. correspondence [Jul38b, Jul38c]. Corso [L.36]. Cost [Dal69, Hen61, Lou67, Har35, Sau48, Thi54, vZ33]. Costantino [Gia64]. Costs [Dem66]. Coulmy [Teg62]. Council [Mos71c]. Countries [Ano69d, Gea65, Gil53, Ric56, Sto50, Zwi71, Ano33n, Ano33o, Ano33p, Ano33q, Ano33r, Ano33s, Ano34q, Ano34r, Ano34s, Ano34t, Ano34w, Ano34x, Ano34y, Ano34z, Ano35e, Ano35f, Ano35g, Ano35m, Ano35n, Ano35o, Ano35p, Ano35q, Ano36e, Ano36f, Ano36g, Ano36h, Ano36p, Ano36q, Ano36r, Ano36s, Ano37e, Ano37f, Ano37g, Ano37h, Ano37n, Ano37o, Ano37p, Ano37q, Ano38h, Ano38i, Ano38j, Ano38k, Ano38l, Ano38r, Ano38s, Ano38t, Ano38u, Ano38v, Ano39e, Ano39f, Ano39g, Ano39m, Ano39n, Ano39o, Ano40d, Ano40e, Ano40g, Ano40h, Ano41d, Ano41e, Ano41h, Ano41i, Ano42c, Ano42d, Ano42e, Ano42f, Ano43c, Ano43d, Ano43g, Ano43h, Ano44b, Ano44e, Ano45b, Ano45d, Ano46c, Ano46e, Cal66, HG43, HU36]. countries [Ide38, Sch36, vdB61]. Country [Cla58, Ano34j]. courbe [Mar64]. cours [Ano36y, Ano68c, Bar43, Bar46, Bun41]. Course [AHW^{+}71, Cur70, Lan63, Mos63, Qua71b, HDPD71, Kit62a]. Courses [BR66, Qua71a]. Courtin [Gru64]. covariation [Ano36k, Jul35]. Coverage [Gla62]. Covilagam [Cha64]. Cowden [Sun71]. Cox [AB66, Sch67, Kit64]. Cramer [Dag70, Ken71]. Creamer [Zwi53]. Credit [Sny34c, dDBL34]. credito [dDBL34]. Crime [HG43]. criminal [Sch36]. criminelle [HG43]. criminelles [Sch36]. Crisis [Ric66, Büc36]. critère [Div43]. criterion [Div43]. critical [BPH52, Her52, Gal51]. criticisms [Her44]. critique [BPH52, Gal51, Her52]. critiques [Her44]. Croft [Der60]. Croft-Murray [Der60]. Crop [Med65]. Croxton [Sun71]. Cumulants [CF38]. Cumulative [Nad68, Box70]. Currencies [Der53a, Dur58]. current [Gin50b, Hos37]. Curriculum [Mos71c]. Curso [Zar62b]. curva [Vin36]. curve [Mar64]. Curves [vR70]. Cushion [Bir59]. customs [Har35]. Cycle [Pri36, SS63]. cyclic [Tin36, Mas67]. cyclique [Mas67, Tin36]. Cygni [Mar64].

D [AB66, Bir59, Box61, Car69, Dav69, Der63, Gou70a, Har65, Har71, Hen54, Ken61, Kha66, Kit64, Lin56, Lüt65, Mar66, Pan63, Pey61, Pey62, Qvi66, Rén67, Ric47b, Sch67, Sch68b, Sun71, Sve63, Sza64a, Tör68, Tra68, Ves68b, Ves70a, Vos65, Zwi53, dO67, vdP65]. d’être [Bun38a]. d’achat [Sau48]. Dagnélie [Mor71b, Mor71c]. d’Agriculture [Dor36]. Dahlberg [Wah56]. daily [BMO^{+}70]. Dana [Pri60]. d’analyse [vdB68]. dans [Ano33q, Ano33r, Ano33s, Ano34w, Ano34x, Ano34y, Ano34z, Ano35n, Ano35o, Ano35p, Ano35q, Ano36p, Ano36q, Ano36r, Ano36s, Ano37n, Ano37o, Ano37p, Ano37q, Ano38r, Ano38s, Ano38t, Ano38u, Ano38v, Ano39m, Ano39n, Ano39o, Ano40g, Ano40h, Ano41h, Ano41i, Ano42e, Ano42f, Ano43g, Ano43h, Ano44e, Ano45d, Ano46e, Ano62r, Ben64a, Bol37, Cal66, Dug65, Dug69, Fré41, G.38, Gin39b, Gin39a, Hos35, Hos39, HU36, Ide36b, Ide38, Mal61, Sch36, Ves68f, Weg64, vZ36]. Dantzig [Hem60]. d’apprentissage [The69]. dargestellt [Nix57]. Darmois [Dug60]. Data [Ken66b, Mah69, Nor70, Pra71, RMK^{+}70b, Vin56, Voi70, Zwi71, vH66, Cal66, Mor41, Bor66, Har71, BF68]. d’autres [Gin39b, Gin39a]. David [Har71, Sve63, Tra68, Gan67, Gou53a, Hem60, Pea71, Shr65]. Davies [Hal57]. Davis [Zwi71]. Deane [Les67, Thi63c, Zwi53]. Death [Gin34, Hor69b, Rah33, SL61, Hub38c]. Deathrate [Gin34]. Deaths [Man68a, Led55]. debits [BMO^{+}70]. Debt [Fis34]. Debt-Deflation [Fis34]. December [Ric51, Nic54]. Decennale [Sav36, Hub38c]. décès [Hub38c, Led55, SL61]. Decision [Dav62, Kün65, Ney62, Dre36, Sve66]. Decisions [Kre65, vE55, Ves68e, Ves68f, Dem63]. Decline [BL38, LS38, JL43a, JL43b, Win38a, Win39]. Decomposition [NT70]. Definition [Pri36, Bun38a, Bun38b, Fré41, Bun38a, Bun38b, Fré41]. Definitions [Wil36]. Deflation [Fis34]. Dégager [Fré46]. degli [Mor35b]. degré [Gin36]. degree [Gin36]. del [Für61]. Delatour [Jul38a]. Delefortrie [Cal62c]. Delefortrie-Soubeyroux [Cal62c]. Delhi [Ric51]. delimitation [Dre38]. della [G.34b, Liv36, Vin36]. Demand [Dum70, KC70, Kon58, Tör68, Sto48, Dur53]. Demeny [Wid67]. Deming [Bor66, Syk68]. demografia [L.36]. demografischen [Fla65]. Demographic [Fou66, Nix61c, Nor70, Sco67a, Vin56, vH66, Cal66, Gin50a, Mor41, Sau50, Mac70, vdB61]. demographie [BPH52, Her34, Her52, BP51, Bun41, Dri67, vdB68]. démographique [Col64, Fed69, Gin50a, Sau50]. demographiques [Cal66, Mor41]. Demography [Pai58, BP51, BPH52, Her34, Her52, Stu41]. d’employés [Jul34]. Dempster [Mit71]. dénombrements [Thi59]. d’enquête [Hub42]. d’ensemble [Gué65]. Department [Dem66, Hem65]. dépendance [Oni36]. dependence [Oni36]. Dependent [Mal68]. dépenses [Hub42]. Depoid [Che68]. Depressions [Fis34]. d’équilibre [Tin36]. Derived [Sha56]. Derksen [Gin50c]. derniers [Sau68]. derniers-nés [Sau68]. Desabie [Bar67]. Desai [Der53b]. Describing [Tek65]. description [BMO^{+}70]. descriptive [Fré41]. Design [Bax71, Bil56, JHH61, KB71, Lin56, Lin67, Lin71a, Rao68, Zim69, Box61]. Designing [KB71]. Designs [CL68, Gué65, Mad56, RS68, Sil64, Tak62, Cla68b]. details [Her44]. d’états [G.38]. Detecting [Bos66]. Determinate [Zar65]. Determination [Ekm69, Ano36k, Jul35, Tin36, Tin36, Der65]. determining [Ano36k, Jul35]. dettaglio [Für61]. d’étude [Fré36a]. d’études [Hos37]. Deutsche [Joh54, Ano33i, Ano34j]. deutschen [Büc36, Bur37, Sae36]. Deutschland [Mee40, Nix57, Ver61, Sch68a, WW38, Win57]. deuxième [Ano60k, Gin37a, Gin39a]. Developed [vdB61]. Developing [Ano69d, Cal66]. Development [Bic68b, Cla58, Day49, Qua71a, Dep64, Zim69]. Developments [Arm66, Hoe68, Vos71, Gin50a, Sau50]. développé [Col64]. développement [Cal66]. développements [Gin50a, Sau50]. Deviates [Dok66, Ham58]. Devons [Gou57]. d’expériences [Dug65]. d’experts [Hub39, Hub41]. d’explication [Fré45]. d’Extrême [HU36]. d’Extrême-Orient [HU36]. Dezsö [Thi63d]. Diagnosis [Oak71]. diagrammes [Gin39b, Gin39a]. diagrams [Gin39b, Gin39a]. Dictionnaire [Ves68a]. dies [Sav41]. Diet [Kev67]. Difference [CL68, Que51, Sil64, Tak62, Žar56]. differences [Sor71, Sor71, Joh62]. Different [Gin43, Ziz38]. différentes [Ano35l]. Differentials [MH50]. difficultés [Hos37]. difficulties [Hos37]. Digest [Ano34u]. Digital [Oak71]. Dimensions [Lan63]. dimostrazione [Vin36]. d’indices [Ste67]. d’integrales [Dug69]. d’introduction [Fré41, Zah33]. Dios [Loy69]. Dips [Ric68]. Direct [Les66]. Direction [Sau63a]. Directives [Ide38, Sch36, Div43]. Directory [Ano34v]. dirigeants [Cal62c]. disappearance [SL61]. discipline [Hor69b]. Discours [Bol60b]. Discrete [Buc69, Kem71, KB71, Nad68, vZ66]. Discriminant [Cor67b]. Discussion [Ano69d, BKB65, RRK65, Gea63b]. diseases [Div43, Wag38]. Disney [Joh71]. disparition [SL61]. dispersion [Dar45, Fré40, Fré42]. Disputed [Kha66]. Distribution [Aig70, BT64, DS71a, Dok66, Fin57, KS53, Lan70b, Luk71, Mat62, MS70, Nix38, Rut65, Sid64, Sto68, Tek65, Wil59, Fré39, Fré45, Gum65, HT71, Led55, Sau68, Sor71, vE61, Blo65, Mar68, Pey61, Kem71]. Distribution-Free [BT64, Dok66, KS53]. Distributions [CF38, The51, Hab62, Buc69, Buc71, Kem71, Ken71, Buc71]. distribuzione [Der61a]. disuguaglianze [Mor34]. dit [Fré34, Fré36b, Fré50]. d’Italia [Für61]. divers [Ano33n, Ano33o, Ano33p, Ano33q, Ano33r, Ano33s, Ano34q, Ano34r, Ano34s, Ano34t, Ano34w, Ano34x, Ano34y, Ano34z, Ano35e, Ano35f, Ano35g, Ano35m, Ano35n, Ano35o, Ano35p, Ano35q, Ano36e, Ano36f, Ano36g, Ano36h, Ano36p, Ano36q, Ano36r, Ano36s, Ano37e, Ano37f, Ano37g, Ano37h, Ano37n, Ano37o, Ano37p, Ano37q, Ano38h, Ano38i, Ano38j, Ano38k, Ano38l, Ano38r, Ano38s, Ano38t, Ano38u, Ano38v, Ano39e, Ano39f, Ano39g, Ano39m, Ano39n, Ano39o, Ano40d, Ano40e, Ano40g, Ano40h, Ano41d, Ano41e, Ano41h, Ano41i, Ano42c, Ano42d, Ano42e, Ano42f, Ano43c, Ano43d, Ano43g, Ano43h, Ano44b, Ano44e, Ano45b, Ano45d, Ano46c, Ano46e, Ide38, Sch36]. diverses [Fré40]. Divisia [Roy64]. do [Böh38, Wür35]. Dobben [Box70]. Dobrovits [Thi63a]. document [Gea63b]. Dogio [Gue62]. Doig [Gea63a]. Dolányi [Thi63b]. donna [Liv36]. Donnant [Div43]. donnée [SL61]. données [Fré60, Mor41, Cal66]. dont [Fré60]. Doornbos [Bar68]. d’ordre [Fré48]. Dorothy [Rio65]. douane [Har35]. Double [Abr54, Put51]. Douglas [Tin42]. d’ouvriers [Jul34]. Dr [Wag38, Wey38]. Dr. [Dun64, Gin50c, Wag38]. Dr.Wilhelm [Wey38]. Draper [Tia68]. Drawn [Nad68]. driving [Hub33, Hub36a]. droits [Har35]. Duality [Vra64].

Dublin [Dun69a, Woo68]. Dubois [Rao58]. Dumas [dO67, Bar65a, Mor71a]. d’un [Fré42, Fré48, G.38, Gin37b, Gin37a]. d’une [BPH52, Bun38a, Div40, Har36, Her52, Jul34, Mor41, Put51, Sau68, Thi54]. Dupriez [Tin47a]. Durand [Pri60]. during [Ano36y, Som62]. duties [Har35]. Dwyer [Gou53c]. Dynamic [All50]. dynamics [Ben64a]. dynamique [Ben64a]. Dzyaloshinskii [Wer67].

Early [Elf69, Ken69]. Easterlin [Har70]. Eastern [HU36]. échantillonnage [Dre69, Put51]. échantillons [Fré43]. Econometric [Gou54c, NvdP66, TB62, Amu68, Bar66a, Wol68]. Econometricians [Mal68]. Econometrics [Som67a, Dag70, Dur53, Gou65, Gou53b]. Economic [All50, All55a, And52, Ano34u, Bow33, BL38, Cla58, Der60, Der65, Gor67, Gou57, Har67a, Har70, Joh62, Ken66b, Kre65, LS38, Les67, Men70, NWW68, Nix54, Rot63, Sny34b, Sto56, The52, The57, Tin41, Wol70, Zim69, vdP65, vdB69, vdB61, Büc36, Fré46, Mal61, WW38, Win38a, Win39, dS33, Sch71a, Tin50, vBW34]. Economically [Rut65]. Economics [BN69, Cra69, Fin57, Tin44, Ken67]. economiques [Mal61, Bar46, Fré46, Iva69, Tin47a, Win39, dS33, Ano34u, BL38, LS38, Dag64]. Economists [Gou54b, Pér71]. Economy [Ano36t]. Education [Ano49g, Gou64, Hot50, Hun71, Muh65, Der71a]. Edward [Pri60, Rom63]. Edwards [vZ66]. efectos [dDBL34]. effectif [Dep64]. effectiveness [Gin42]. Effects [NWW68, RS68, dDBL34, Win38a]. effectuée [Har36]. efficacité [Gin42]. Efficiency [Hus70, TvY56]. Ehrengabe [Mee40]. Eigenvalues [Dug69]. Eighth [Ric53]. Eighty [Woo68]. Eilon [Jar69, Min69]. Einaudi [For62]. einem [Wag38]. Einfluss [Büc36]. Einführung [Ham60, F.36, Kol57]. Einige [Win38b]. Eisen [Gan71]. ejecutados [dDBL34]. Ekambaram [Rid64]. Ekeblad [Bil64]. Eklund [Dem61]. elaboration [Ide38, Sch36]. Elderton [vR70]. Electronic [Ken66a]. Elementare [Gou62]. Elementary [Gou68, Mar71, vZ66, Fal69, Gei69, Ple67, Som67b]. Elementi [Mui65]. Elements [Sco67b, Wol51, Bar61, Fre71, God68, Lan62, Mit71, vdP65, Ves68c, Ves68b]. elettrici [Mor35b]. élevés [Hub36b]. Elteren [Pur69]. Emil [Gum67]. Émile [Dar56]. Emiliano [G.34c]. Emiliano-Tosco-Romagnolo [G.34c]. Emma [Eic65]. Empirical [Der71a, Muh65, Rob63, Dag70, Kru71]. emploi [Hos39, Tin47b]. employees [Jul34]. Employment [Nix58a, Tin47b]. end [Rah33]. endlich [dS71c]. endlichen [Kol60]. Energy [Kev67]. Engineering [BR66, Day49, Hun67, Whi66, Bar66b]. Engineers [Joh71]. England [Nix35, Nix52]. enquête [Fré46, Har35, Jul34]. enquêtes [Thi54, Ver63]. Enrick [Har67b, Lai67, Nor66]. Enrico [Gin56]. enseignement [Fré41, Ide36b]. entre [Ide38]. enumerations [Thi59]. Environment [Dal71]. épreuves [Mor71a]. Equal [DS71a]. Equation [Ano70c, KC70, Key67, NvdP66, Ram70, TB62]. Equations [Mal68]. équilibrés [Gué65, Mas67]. Equilibria [Tin41]. Equilibrium [Rot63, Tin36]. équipe [Dug65]. Equipment [Flu38]. Equivalent [CS70]. Erfahrungen [Joh54]. Ergodic [Rév66]. Erich [Bom71]. Erkentnisse [Joh54]. Erlander [vdP70]. Ernest [Her61]. Ernst [T.36, Lan56]. Errata [Ano34k, Ano34l, Ano36j]. Erratum [Ano36k]. Error [Bos66]. Errors [Aig70, Dur54, Fin56, Fis57, HW70, NvdP66, Pra53, Suk52, The53, The57, KG67]. Erstellung [Ven68]. especially [Hos39]. Essai [Div43, Fré50, LS71]. essais [Fré45]. essay [Wey38]. Essays [Kin64, Rao61, Som67a]. estadística [Can55]. estadisticas [dDBL34]. Estimate [Žar56]. Estimated [Mah69]. Estimates [Bri64, Gin49, Gin52, Ken66b, Sch66, The51, Dar45, Tin47b]. Estimating [Mat62, Cal66, Put51]. Estimation [Aig70, GF55, HR68, JHH61, KC70, Les66, Mal61, Med65, NT70, The57, Cal66, Gum65, Sau63b, Mal61]. estimations [Dar45, Tin47b]. Estimator [MS70]. Estimators [Ano70c, Fis57, Hus70, Ram70, KG67]. estimer [Put51]. età [Liv36]. état [Coc50, Gin50b]. ethics [LS71]. éthique [LS71]. Étienne [Thi64]. étoile [Mar64]. Ettan [Cha64]. étude [Dar53, Der58, Div40, Div43, Dug69, Fré46, Har36, Her42, Thi68, dS33, Har35, Del50, Mar64, Fed69]. études [Gin50a, Sau50]. étudiants [Ves68b]. Eulerian [Bar65b]. euleriens [Bar65b, BKB65]. europee [G.34b]. Evaluating [NWW68]. Evaluation [Gre68, ID40]. evaluations [Fré43, Gin45]. événements [G.38]. Events [AB66]. éventuellement [Div43]. Evidence [Pra71, Sär70]. èvolution [Dep64]. Ewens [Hay70]. Example [Ekm69]. Excerpts [Jul38b, Jul38c]. Executives [Rid64]. Exercices [Ves67, Ves68d]. Exercises [Gou54a, Rao69]. exhaustive [Mor71a, Mor71a]. expansion [Cla61]. Expectation [MH50]. Expectations [Sch71a]. Expected [vZ66]. Expenditure [Sch68b, Tay68, dW54, Bje62]. Experience [Voi70]. Experiment [Boy58]. Experimental [Box61, Dug65, Lin56, Lin71a]. Experimentalists [Nel70]. expérimentaux [Teg62]. Experiments [MSC^{+}70, NWW68, Lin67, Dal63]. Experts [Hub39, Hub41]. explain [Fré45]. Explicata [Sär70]. Exploitation [Teg62]. Exposé [And44]. Extension [HT58, Fré43]. extensive [Hab62]. Extraits [Jul38b, Jul38c].

F [Ano64j, Arm70, Bar69a, Bil64, Bri71, Cal62d, Dal65, Der61a, Der71b, Fla65, Gan67, Gou53a, Gou58, Gou63, Gut68, Har68, Har71, Kha66, Kit62a, Kit62b, L.36, Lin71a, Mad46, Mar68, Ric61, Stu41, Sun71, Sve63, Tra68, Wol68, Zar62b, dS71b, vdB69]. facteurs [JL43a, JL43b]. Factor [And68, Sha56, Dar53, Vos65]. Factorial [Mad56]. factorielle [Dar53]. factors [JL43a, JL43b]. facts [Hos35]. Fair [Dum70]. faire [Dre69]. faits [Hos35]. familiaux [Hos37]. Familienstande [Dep57]. Familienstatistik [Bur37]. Families [Nic54, Sau68]. famille [Sau68]. familles [Cal62b, Hub42, Sau68]. Family [All55b, Ben64b, Hen54, Nix35, Nix36, Hos37, Sau68, Bur37, Zar62a]. Far [HU36]. Fault [Oak71]. Favorable [Tho69]. fecondità [Liv36]. fécondité [Gin41, Hen53]. Federal [Cro69]. Federalist [Kha66]. Feeding [Oma67]. Felicitations [Rao63a]. Felix [Pét66, Gin57a]. Feller [Cra70, Gre67, Kri69]. femme [Gin41]. Ferguson [Gna67]. Ferschl [dS71b]. Fertigwarenpreise [Pla34]. Fertility [Fou66, Hen54, Sch66, Bur37, Gin41, Hen53, Liv36]. Festgabe [Sae36]. Festschrift [Pét66]. Fetal [Man68a]. fiabilité [Mor69]. Field [Wer67]. Fielding [Ric59b]. Fieller [Ham58]. Fifth [Le 69, Hub38c, Ric47b]. Files [Nor70]. Filtering [RMK^{+}70b]. Final [TB62]. Findlay [Wil55]. fini [G.38]. finished [Pla34]. Finite [God65, MS70, Tik67]. Finland [Har35, Har36]. Finlande [Har35, Har36]. Finney [Har65, Box61, Dav69, Lin56, Mar66, Pan63]. Firestone [Cla58]. Firms [Ano68d, Wol67]. First [AHW^{+}71, Cur70, Kit62a, Stu64, Wol70, Dar58b, Sor71, Fré46]. First-Stage [Stu64]. Fiscal [Cox60, Rao63b, Hos38]. fiscale [Hos38]. Fisher [Gou63, Ano62k, KG67, Lin67, Yat62]. Fisk [Mal68]. Fit [BT64, RMK^{+}70a]. Fitting [Dur60, HT58, TvY56]. Flamant [Ves70a]. Flament [Sax66]. Flaskämper [Sae36, vdB63]. Flow [Lau64]. Flow-of-Funds [Lau64]. flows [BMO^{+}70]. Fluctuations [Sto56, Tin50]. fonction [Thi54]. fonctions [G.38]. fondamentales [Her44]. Fondements [Hen53, Gin45]. fontes [Del50]. fonts [Del50]. Forbrig [Sch68a]. forces [Hub33, Hub36a]. Forecasting [Kev67, Kre65, Sch71a, Mal61, Bra69]. Forecasts [BJT63, Sch71a]. foreigner [Thi33]. Forest [Nai67, Dor33, Nai71]. forestière [Dor33, Dor36]. forestry [Dor36]. Foreword [Hub37, Met33a, Met33b, Met35, Met45]. Form [TB62]. Formation [Ano67m, Ano67n, Ano68o, Ano68p, Ano68q, Ano69k, Ano69l, Ano71n, BSH71, BH69, Der54c, HK70, MPDoA71, PBW70, Pur70]. Formeln [Kád68]. Formulae [Vos69]. Formulas [Hem65, Fré39]. Formulation [FG65]. formules [Fré39, Ste67]. Forrest [Arm70]. Forschung [Wag38]. Forschungsund [Ver61]. Foundation [Qua71a]. Foundations [Dem62, RRK65, Rén65, Gin45, Hen53]. Founder [Wil38]. Fourastié [Ste67]. Fourth [Ano63l, Ano63h, Lan71]. Fox [Men70]. Framework [Ken66b]. francaise [Dug65, Cal62c]. France [Der53a, Der58]. Francis [Ric64]. Franco [dM63]. Francois [Roy64]. Frankreich [Nix57]. Franz [Sae36]. Fraser [Car69]. Frauenarbeit [Eic65]. Fréchet [And44, G.38, Gea38, Gum65, Wol51]. Frederick [Nix58b]. Free [BT64, Dok66, KS53, Wey38, Win37]. Freeman [Som67b]. Freitas [Lop56]. Fremden [Thi33]. Fremdenverkehrsstatistik [Thi33]. French [Ano33b, Ano33c, Ano33d, Ano33e, Ano33f, Ano33g, Ano33h, Ano33n, Ano33o, Ano33p, Ano33t, Ano33q, Ano33r, Ano33s, Ano33u, Ano33v, Ano33w, Ano34a, Ano34b, Ano34c, Ano34d, Ano34e, Ano34f, Ano34g, Ano34h, Ano34i, Ano34q, Ano34r, Ano34s, Ano34t, Ano34v, Ano34u, Ano34w, Ano34x, Ano34y, Ano34z, Ano34-28, Ano34-29, Ano34-30, Ano35a, Ano35b, Ano35e, Ano35c, Ano35f, Ano35d, Ano35g, Ano35h, Ano35m, Ano35l, Ano35n, Ano35o, Ano35p, Ano35q, Ano35r, Ano35s, Ano35t, Ano35u, Ano36a, Ano36e, Ano36b, Ano36f, Ano36c, Ano36g, Ano36d, Ano36h, Ano36i, Ano36k, Ano36p, Ano36q, Ano36r, Ano36s, Ano36u, Ano36v, Ano36w, Ano36x, Ano36y, Ano37a, Ano37e, Ano37b, Ano37f, Ano37c, Ano37g, Ano37d, Ano37h].

French [Ano37i, Ano37n, Ano37o, Ano37p, Ano37q, Ano37r, Ano37s, Ano37t, Ano37u, Ano37v, Ano38a, Ano38b, Ano38c, Ano38h, Ano38d, Ano38i, Ano38e, Ano38j, Ano38f, Ano38k, Ano38g, Ano38l, Ano38m, Ano38n, Ano38o, Ano38r, Ano38s, Ano38t, Ano38u, Ano38v, Ano38w, Ano38x, Ano38y, Ano38z, Ano38-27, Ano39a, Ano39b, Ano39e, Ano39c, Ano39f, Ano39d, Ano39g, Ano39h, Ano39i, Ano39m, Ano39n, Ano39o, Ano39p, Ano39q, Ano39r, Ano40a, Ano40b, Ano40d, Ano40c, Ano40e, Ano40f, Ano40g, Ano40h, Ano40i, Ano40j, Ano41a, Ano41b, Ano41d, Ano41c, Ano41e, Ano41f, Ano41g, Ano41h, Ano41i, Ano41j, Ano41k, Ano42a, Ano42c, Ano42b, Ano42d, Ano42e, Ano42f, Ano42g, Ano43a, Ano43c, Ano43b, Ano43d, Ano43e, Ano43f]. French [Ano43g, Ano43h, Ano44a, Ano44b, Ano44c, Ano44e, Ano44f, Ano45a, Ano45b, Ano45d, Ano45e, Ano46a, Ano46b, Ano46c, Ano46e, Ano46f, Ano57f, Ano60i, Ano60k, Ano62l, Ano63h, Ano64a, Ano66a, Ano68a, Ano70a, Aue34, Bar65b, Ben64a, Ben66, BMO^{+}70, Bol37, Bol60b, Bol62, BP50, BP51, BPH52, Bow46, Bun33, Bun37, Bun38a, Bun38b, BL38, Cal66, Dar45, Dar53, Dar55, DG57, Dar58b, Del50, Dep64, Div40, Div43, Dor33, Dor36, Dre69, Dug55, Dug65, Dug69, Dun40, Dun42, For60, Fré34, Fré36b, Fré36a, Fré39, Fré40, Fré41, Fré42, Fré43, Fré45, Fré46, Fré48, Fré50, Fré57, Fré58, Fré60, Gal51, Gin33, Gin36, Gin37b, Gin37a, Gin39b, Gin39a, Gin41, Gin42, Gin45, Gin50b, Gin50a, God37, God39, Gou51d, Gué65, Har35, Har36]. French [Hec34, Hec35, Hen53, Her34, Her40, Her42, Her44, Her52, HG43, Hor69b, Hos35, Hos37, Hos39, Hos46, Hub33, Hub36b, HU36, Hub36a, Hub37, Hub38a, Hub38c, Hub38b, Hub39, Hub41, Ide36a, Ide36b, Ide36c, Ide38, Jul34, Jul35, Jul38b, Jul38c, JL43a, JL43b, LS38, Led55, LS71, Les35, Mal61, Mar64, Mas67, MK70, Met33a, Met33b, MT34, Met35, Met45, Mey42, Mic38, Mol36, Mor35a, Mor41, Oni36, Par62, Pie37, Pie40, Pie44, Pto34, Put51, Ric36, Ric45a, Sau48, Sau50, Sau68, Sav41, Sch36, SL61, Sor71, Stu34b, Stu34a, Stu35, Stu36, Stu38, Thi54, Thi59, Thi66a, Thi68, Tin36, Tin47b, Win39, WI37, Zah33, dL38, dS33, dS36, dS38, vZ33, vZ36, vZ38]. fréquences [Hos35]. frequencies [Hos35]. Frequency [Nad68, vR70]. Freudenberg [Dep57]. Freund [Gou68]. Friedman [Sas71]. Friedrich [Mee40, S.34, Hor68b]. Frisch [Wol70]. Front [Ano33j, Ano33k, Ano33l, Ano34m, Ano34n, Ano34o, Ano34p, Ano35i, Ano35j, Ano35k, Ano36l, Ano36m, Ano36n, Ano36o, Ano37j, Ano37k, Ano37l, Ano37m, Ano38p, Ano38q, Ano39j, Ano39k, Ano39l, Ano44d, Ano45c, Ano46d, Ano47c, Ano48d, Ano49f, Ano50e, Ano52e, Ano53h, Ano53i, Ano54d, Ano55d, Ano56d, Ano57d, Ano58d, Ano59c, Ano60e, Ano60f, Ano61h, Ano61i, Ano61j, Ano62e, Ano62f, Ano62g, Ano63c, Ano63d, Ano64e, Ano64f, Ano65d, Ano65e, Ano65f, Ano66g, Ano66h, Ano66i, Ano67a, Ano67b, Ano67c, Ano68e, Ano68f, Ano68g, Ano69a, Ano69b, Ano69c, Ano70e, Ano70f, Ano70g, Ano71b, Ano71c, Ano71d]. Fruchtbarkeitsmessung [Bur37]. Fuchs [Bru70, Ves68g, Wis71]. Function [Nad68, Neu67, Tin42, Tra68, Thi54, Hem69]. Functional [Fre71]. Functions [Cor67b, Hem65, Kon58, Luk71, Mor71d]. Fundamental [Pri36, Her44]. Funds [Lau64]. Furlan [Kul58]. Further [Ano61a, Pra71].

Furtherance [Ano49g, Hot50]. future [Dep64].

G [Ano68c, Bje62, Bra67, Bri68, Buc69, Cla68a, Dal68a, Dem61, Der60, Der63, Der68, Elf66, Fed69, Für61, Gea63a, Gei69, Gna67, God68, Gou53b, Gou54a, Grd64, Gut68, Hun67, Jua71, Kin64, Kit62a, Kit62b, Kol60, Kol65, Lan63, Lan70b, Lin66a, Mor69, Muh65, Nix58a, Nor67, Oak71, Pér71, Pey61, Pre68, Rao61, Rao68, Rén63, RR68, Sch68a, Teg62, Tra68, Ven68, Ver63, Ves68d, Ves68g, Vin56, Wil59, Wis71]. Gabriel [Sau71]. Gaetano [For61]. Gambling [Tho69]. Game [Vaj65]. Games [Tho69, Hos35]. Ganeshalingham [Har71]. Gani [Kri67]. Gaps [Ric68]. Garrigue [Iva69]. Geary [vdP65]. Geburtenrückgang [Win38a]. Geburtstages [Sae36]. Gegenwart [K.35]. Gel’fond [Ple67]. Gemüsestatistik [Sae35]. General [Bir59, Bri64, GV63, Gut68, HDPD71, Pur65, Fré34, Fré34, Dar53, Bir59, Elf66, Bri71, Lin64]. générale [Dar53, vdB62, Ves68b]. générales [Bar46]. Generalised [Bit71]. Généralités [Wol51, Gea38]. Generalization [Ben67a]. Generating [CL68, Sil64, Tak62]. Generation [Sto68, Sav33, Sav41, Sav41]. Generators [Azo67]. généraux [Tin47a]. generazione [Sav33]. Genetic [EG61]. Genetics [Hay70]. génétiques [Ves70b]. Geographical [Tek65]. Geological [RMK^{+}70a, RMK^{+}70b]. gèologie [MK70]. Geology [GHM70, MK70]. Geometrical [Rén63]. Geometry [Lan63]. George [Kin64, Ano52f, Wil55]. Georges [Dug60]. Gerhard [Gou54b]. Gerloff [Fla54]. German [Ano33i, Ano34j, Ano36t, Böh38, Büc36, Bur37, Dre36, Dre38, Mee35, Pla34, Rah33, Sae35, Thi33, WW38, Wag38, Wey38, Win37, Win38b, Win38a, Win42, Win44, Wür35, Zah34, ZH42, Ziz33, Ziz38]. Germany [Der53a, Tie69, WW38]. Gesamtheiten [Kol60]. Gesamtrechnung [Der57]. Geschichte [Eic65]. Ghurye [Rao61]. Gilbert [Der53a, Der54a]. Gilford [Rio65]. Gini [Cas65]. Giorgio [Kin68]. Girard [Bar55]. given [Fré60, SL61]. Giving [Div43]. Glass [Hen54]. Glauber [Kre65]. gleichnamigen [Wey38]. globaux [Her34]. Goals [TI66]. Godfrey [Ait55]. Goldberg [Les67]. Goldberger [Bar66a]. Goldsmith [Gol60, Jos53]. Goodness [BT64]. goods [Pla34]. Gor’kov [Wer67]. Göschen [Gou62]. Gothenburg [Dem66]. Goulden [Gou53d]. Government [Kün65, Cro69]. Govind [Suk69]. Gowen [Fin54]. Graf [Kád68]. grafica [Vin39]. grafikus [K.35]. Graph [Gan66]. graphes [Sax66]. Graphical [Ekm69, Ped69, Ped69]. Graphische [K.35]. Graphs [Fra68, Hem65]. Great [Bow33, Fis34, Hen54]. Greater [Has33]. Grebenik [Hen54]. Greece [Som62, Kev67]. Green [Gor67, Neu67]. Greenwood [Iss52]. Grenzen [Ven68]. Group [Nai67]. Grouping [Aig70, Ekm69, PA54]. Groups [vE55, Dre38]. Growing [Oma67]. Growth [Ano68d, Cla61, Koc63, Sny34a, Sny34b, Wol67, Win38b, Har70, Les67]. Grundbegriffe [Gou51a]. Grundcharakter [Ziz33]. Grundfragen [Gou51b]. Grundriss [Sch68a, vdB63]. Grundzüge [Ham60, Lep68]. Gruppen [Dre38]. Guide [Gna67, Sza64a, Ham67]. guidelines [Div43, Ide38, Sch36]. Guinea [BW69]. Gumbel [Gum67]. Gunnar [Wah56]. Gustave [Ill37]. Güterverkehr [dKT38]. Guttman [Whi66]. Gyula [Thi66b].

H [All55b, All55a, And68, Arm70, BF68, Blo65, Bra69, Bri71, Cam64, Dal58, Dur53, Gil62, Gor67, Gou53e, Gou53d, Gou55, Gou62, Har71, Hoe70, Hub66, Kád68, Ken61, Ken71, Kün65, Lan62, Lep68, Lin69, Lin61b, McC66, Mok62, Mos54, Myb63, Oak71, Rao58, Rao61, Rop70, Sch67, Sha63, Shr65, The69, Tia68, Tin47a, Tör68, Vad59, Ver58, Vos71, Wij69, dS71c]. Hague [Ano69d]. Haight [Mar68]. Half [McC69]. Halothane [Ano70d, Arm70]. Hammersley [Qvi66]. Handbook [Ano36t, Bri68, Gov66, Hem65, Mar68, Pey62]. Handbuch [Ano36t]. Handliche [Pea62]. Handscomb [Qvi66]. Hans [Jac64, Nix61b]. Happy [Dun64]. Harman [And68]. Harold [Rao61]. Harris [Lea68]. Harter [Har71]. Hartkemeier [BF68]. Hartley [Gou55]. hasard [Hos35]. Haskel [Gin56]. Hatvani [Thi64, Thi64]. Haupt [K.35, dKT38]. Haupt- [K.35, dKT38]. Haushalt [Der57]. Haven [Voi70]. having [Dug69]. Health [Lin62, Wil35, Gep58, Muh69]. Heavy [AMO^{+}70]. Hecht [Kai52]. Hemisphere [Per41]. Henderson [Gou70a]. Henning [Kád68]. Henri [Ide55]. Henripin [Nix65]. Hérault [Ves68c]. Herbert [Rio65, Atw65]. Herbold [Nor67]. Heron [Pea71]. Hersch [Win44, Nix55, Win44]. Heteroscedastic [Fis57, Pra53, The51, The53, KG67]. Heteroscedasticity [Fis62]. heutigen [Mee40]. hiérarchique [Ben67b]. Hilfsmittel [Wag38]. Hilsenrath [Nor67]. Hilton [Ait55]. Histoire [vdB62]. Historical [Thi63c, Nix61a]. History [vBW34, Ken69]. Hoeffding [Rao61]. Hofsten [Der54b]. Höhere [Gou62]. Holden [Gou70a]. hongroise [Thi64]. Honor [Rao61]. Honoraire [Ide53]. Honorary [Ide53, Ano62k, Ano64j]. Honour [Kin64]. Hood [Gou54c]. Hoover [Atw65]. Hope [Fal69]. Horváth [Thi64]. Hospital [Bit71]. Hotelling [Rao61]. Hours [Dum70]. Household [Som62]. Households [Nix52]. Housing [Nix35, Nix52, Nys36, Nys38]. Houthakker [Tör68, All55b]. Hsu [Har65]. Huber [Bar43, Bar46, Bun41]. Huitson [Wal67]. humaine [Ves68b]. humaines [Bol37, Weg64]. human [Bol37, Muh65]. Hunt [RR68]. Hurwicz [Vad59]. Hurwitz [Han69]. Hydrology [BAM^{+}70, JHCS70]. Hypergeometric [Luk71, Pey61]. Hypotheses [CS70, Rob63]. Hypothesis [Gut68, Pur69].

I.I.S. [Bol60b]. I.N.S.E.E. [Thi54]. I.S.I. [Bol60a, Bol60c]. IASI [Ric45b]. Identify [Fré46]. IFO [And52]. IFO-Institute [And52]. ihre [WW38]. ihrem [Mee40]. II [Gou62, Jos53, vdB63, Gov66, Sch68b]. III [Bar46, Der54a]. Ijiri [TI66]. IJzeren [Dur58]. Illustration [Žar56]. Illyefalvi [K.35, dKT38]. im [Gou62]. impianti [Mor35b]. Implementation [Dre36]. important [BPH52, Dug69, Gin37b, Gin37a, Her52]. importante [BPH52, Gin37b, Gin37a, Her52]. Imports [Ric56, Grd64]. Improved [Wil35]. Improvement [BJT63]. Improving [Nys36, SEM^{+}70]. Incertitudes [Ves68e]. Inch [Dix71b]. Income [Bje62, Der65, Gil51, Gin49, Gin50c, Gin52, ID40, Jos53, Les67, Ohl61, Sha63, Shi37, Sto50, SUD50, Sto58, Sto68, Zwi53, Ano33i, Fré39, Fré45, Der54a, Gil51, Jos53]. Incomplete [CL68, Gué65, Hem69, RS68, Sil64, Tak62, Cla68b, Mas67]. incomplets [Gué65, Mas67]. Increase [Wil37]. Increments [Wal62]. Independent [Pur65, Wal62]. indeterminate [Hos46]. indeterminée [Hos46, Led55]. Index [AFR^{+}63, Ano33m, Ano64g, Ano66j, Ano67d, Ano68h, Ano70h, Ano70i, Pfo66, Sha56, Ken69]. Index-Numbers [Ano33m]. Indexes [Der54b, Kon58, McC63]. Indexing [Bos66]. India [Der53b, Oma67, Shi37]. Indian [Dal63, Gea65, Mah69]. Indicators [SS63]. Indices [Ric56, Hos39, Hos46, vZ33]. Indifferent [Tin41]. individu [Gin37b, Gin37a]. Individual [Nor70, Div43, Gin37b, Gin37a, LS71]. individuel [Div43]. individuelle [LS71]. Inductive [Ney57]. Industrial [Cla61, GF55, Rid64, Ste57, Har36, Jar69, Koc63]. industrie [Cal62c]. industrielle [Cla61, Har36]. industriellen [Ham60]. industrielles [Ves68e]. Industries [Nix58a, Bru70]. Industriestatistik [Sch68a]. Industry [Hal57, Kün65, Nor66, Cra69]. infant [BP50, Pto34]. Infantile [Gin34, BP50, Pto34]. Inference [All50, KLM67, Ken66a, Lin61a, Pfa68, RS68, Mor71c, Car69, Dem62, Rén67, Rop70, Kha66]. infini [Fré48]. infinite [Fré48, Tik67]. influence [Tin47b, Büc36]. infondé [Fré50]. Information [Ano34-31, Ano35v, Ano35w, Ano36z, Ano37w, Ano38-28, Ano39s, Ano40k, Ano41l, Ano42h, Ano42i, Ano43i, Ano49j, Ano50i, Ano51n, Ano52k, Ano53n, Ano56g, Ano60p, Ano61r, Ano62u, Ano63u, Ano64q, Ano64r, Ano65q, Ano65r, Ano66v, Ano66w, Ano67r, Ano68t, Ano69p, Ano70s, Lam68, Mat62, NT70, Per41, RRK65, Rén65, Sär70, Rév66, Upp68]. Informations [Ano55e, Ano56e, Ano57e, Ano58e, Ano59d, Ano60g, Ano60h, Ano61k, Ano61l, Ano61m]. informationswissenschaftlichen [Ham60]. Ingenieure [Gou45, Gou51c, Gou61a]. Ingvar [Cow62]. Input [Ohl61, Osb66, Ven68, Ken67]. Input-Output [Ven68]. Inspection [Hal67, Ham50]. Institut [Dar58b, Dor36, S.34, Ano60i, Ano36y, Ano57f, Ano60k, Ano63h, Ano64a, Ano66a, Ano68a, Ano70a, Bol62, DG57, Dep64, Hub38b, Sch36, Stu34b, Stu34a, Stu36]. Institute [Ano36y, Ano49g, Ano57g, Ano60j, Ano60m, Ano62m, Ano63l, Ano64a, Ano64b, Ano66a, Ano66b, Ano68a, Ano68b, Ano70a, Ano70b, Dor36, Dre36, Nix54, Ric59a, Ric66, Stu36, And52, Ano57f, Ano60i, Ano60k, Ano62l, Ano62k, Ano63h, Ano64j, Bol62, DG57, Dar58b, Dep64, Gou64, Hub38b, Ide38, Sch36, Stu34b, Stu34a, Dal63, Gea65]. Institutes [Ano62h, Dre36]. Institutions [Ano34v, Ano35e, Ano35f, Ano35g, Ano36e, Ano36f, Ano36g, Ano36h, Ano37e, Ano37f, Ano37g, Ano37h, Ano38h, Ano38i, Ano38j, Ano38k, Ano38l, Ano39e, Ano39f, Ano39g, Ano40d, Ano40e, Ano41d, Ano41e, Ano42c, Ano42d, Ano43c, Ano43d, Ano44b, Ano45b, Ano46c]. Instituts [Ano62h]. Instructional [Qua71a]. Insurance [Med65]. Integral [Key67, Fré41, Par62]. intégrale [Fré41, Par62]. integrals [Dug69]. integration [Gou51d, Gou51d]. Inter [Ric59a]. Interdependent [Wol58]. Intermediate [Qua71b]. internacionales [dDBL34]. International [Ano34v, Ano34u, Ano36y, Ano47a, Ano48b, Ano49b, Ano49c, Ano49g, Ano50a, Ano50b, Ano51b, Ano51c, Ano51d, Ano52b, Ano52c, Ano53d, Ano53e, Ano54b, Ano55b, Ano56b, Ano57b, Ano57g, Ano57f, Ano58b, Ano59a, Ano60a, Ano60b, Ano60j, Ano60m, Ano60k, Ano61c, Ano61d, Ano62c, Ano62m, Ano62k, Ano62r, Ano63l, Ano63h, Ano63p, Ano64d, Ano64j, Ano64a, Ano64b, Ano65c, Ano66f, Ano66a, Ano66b, Ano67e, Ano68a, Ano68b, Ano70a, Ano70b, Bli49, Bol62, DG57, Dar58b, Der53a, Dor36, Dre36, Fou66, Gil51, Gou64, Hub38b, Hub41, Ide38, Les67, Mic38, Nai67, Nys36, Ric66, S.34, Sch36, Sha63, Stu34b, Stu34a, Stu36, Ano35l, Ano60i, Ano62l, dDBL34, Dep64, Gou51d, Mol36, Sae35, ZH42, vZ33, Ano33e, Ano33f, Ano33g, Ano34c]. International [Ano34d, Ano34e, Ano34f, Ano34v, Ano34u, Ano35a, Ano35b, Ano35c, Ano35d, Ano35l, Ano36a, Ano36b, Ano36c, Ano36d, Ano36y, Ano37a, Ano37b, Ano37c, Ano37d, Ano38c, Ano38d, Ano38e, Ano38f, Ano38g, Ano39b, Ano39c, Ano39d, Ano40b, Ano40c, Ano41b, Ano41c, Ano42a, Ano42b, Ano43a, Ano43b, Ano44a, Ano45a, Ano46b, Ano57f, Ano60i, Ano60k, Ano62l, Ano63h, Ano64a, Ano66a, Ano68a, Ano70a, Bol62, DG57, Dar58b, Dep64, Dor33, Dor36, Hub33, Hub36a, Hub38c, Hub38b, Ide38, Sch36, Stu34b, Stu34a, Stu36]. Internationale [Ide38, Sch36, Ano33e, Ano33f, Ano33g, Ano34c, Ano34d, Ano34e, Ano34f, Ano34k, Ano34l, Ano35a, Ano35b, Ano35c, Ano35d, Ano35l, Ano36a, Ano36b, Ano36c, Ano36d, Ano36j, Ano37a, Ano37b, Ano37c, Ano37d, Ano38c, Ano38d, Ano38e, Ano38f, Ano38g, Ano39b, Ano39c, Ano39d, Ano40b, Ano40c, Ano41b, Ano41c, Ano42a, Ano42b, Ano43a, Ano43b, Ano44a, Ano45a, Ano46b, Ano47a, Ano48b, Ano49b, Ano49c, Ano50a, Ano50b, Ano51b, Ano51c, Ano51d, Ano52b, Ano52c, Ano53d, Ano53e, Ano54b, Ano55b, Ano56b, Ano57b, Ano58b, Ano59a, Ano60a, Ano60b, Ano61c, Ano61d, Ano62c, Ano62r, Ano63p, Ano64d, Ano65c, Ano66f, Ano67e, Dor33, Dor36, Gou51d, Hub33, Hub36a, Hub38c, Mol36, Nix57, vZ33]. internationalement [Bun38a]. Internationalen [Dre36, Sae35, ZH42]. internationales [Ano35l]. internationally [Bun38a]. interno [G.34b]. Interpretation [Har67a, Ken66c]. Intervals [Gar71, Nad68, Spj68]. Interview [Mos51]. Introducción [Can55]. introducing [Fré41]. Introduction [Box61, Dal62, For68, Gou57, Gre67, KWT^{+}63, Kri69, Mar71, Pan63, Rid64, Sch59, Sza64b, Wol55, Zah33, Ano71e, Dav69, Fin66, Bar69b, Bea69, Dav70, Gan71, Gou65, Har68, Hec69, Rén61, Rén67, Rop70]. Introductory [Gor67, Mal68, Wet70, Whi66]. Inventaire [BMO^{+}70]. Inventive [Sau63a]. Inventories [Coh66]. Inventory [Rio65, Vad59, vdP70, BMO^{+}70, Min69]. Inversions [Gin47]. Investigating [Bow33]. investigation [Fré46, Ste60]. Investigations [vBW34]. Investment [Kre65]. Irene [Hem65]. Irish [Pra71]. Isaac [Nix54]. ISI [Ric45b]. Issued [Ano61q, Ano62s, Ano62t, Ano63q, Ano63r, Ano63s, Ano63t, Ano64o, Ano64p, Ano65p, Ano66t, Ano66u, Ano67o, Ano67p, Ano67q, Ano68r, Ano68s, Ano69m, Ano69n, Ano69o, Ano70p, Ano70q, Ano70r, Ano71f, Ano71o, Off61].

Istituto [Sav36]. István [Thi64]. Italia [G.34c]. Italian [Ano33a, Liv36, Pie40, Pie44, Sav33]. Italiano [Ano33a]. italienne [Pie40, Pie44]. Italy [Der53a]. Item [Mok62]. Iterative [Fis62, RMK^{+}70a, KG67]. Iterative-Fit [RMK^{+}70a]. IV [Bar46, G.34c, Le 69].

J [All55b, Arm70, Bar67, Box61, Bra67, Bri68, Cac71, Cal62a, Cla58, Cra68, Dag70, Dal65, Dar58a, Dav69, Dem62, Der61b, Dur58, Elf69, Fin54, Fin57, Fre71, Gep58, Gil62, Gor67, Gou62, Gou68, Gou70a, Gov66, Gum67, Har68, Har65, Har71, Hay70, Hen68, Hen61, Hun67, Jar69, Kád68, Kap67, Kha58, Kre65, Kri67, Kru71, Le 69, Lin56, Lin61b, Lin66b, Loy70, Mar66, McC66, Men70, Mih69, Neu67, Nix54, Nix58a, Nix62b, Nix65, Nor67, Pan63, Pey61, Qvi66, Rop70, Rut65, Sas71, Sch71a, Sch59, Sch66, Sco67b, Spr71, Ste67, Ste57, Sun71, Vad59, Ves67, Ves68e, Ves68f, Ves70a, Vin56, Wid67, Wij69, Wol67]. J. [Myb63]. Jackknife [Bri64, Mos71b]. Jacobs [Nix57]. Jacquard [Ves70b]. Jan [Gou61b, Wol70]. Jansson [Azo67]. Japan [Tie69]. jelene [K.35]. jeunesses [Har36]. jeux [Hos35]. Job [Har67a]. Johann [Hor69b]. John [Irw56]. Johnson [Buc71, Kem71, Vos71, vR70]. Johnston [Hen61]. Joint [Ide38, Dor36, Mos71c, Sch36, Cra68]. jointes [Gin39b, Gin39a]. Jordan [Rén60]. Jose [Ano52g]. Joseph [Bar55]. Jostock [Für65]. journaliers [BMO^{+}70]. Jr [Arm70, Hub66]. Jr. [Kit62a, Kit62b, Rut65, Vos71]. Ju [Ken65]. Juan [Loy69]. Jubilé [Ill37]. judge [Gin37b, Gin37a]. juger [Gin37b, Gin37a]. juillet [Dar55]. Juin [Ano33t, Ano60k, Dar55]. July [Dar55]. June [Ano60m, Ano60k, Dar55, Ano33t]. Juréen [Dur53].

kann [Böh38]. Karl [Für63]. Karlin [Gil62, Kap67, Vad59]. Kaufmann [Bil65, Bil70, Dav70, Fla65]. Keilson [Neu67]. Kellerer [Gou62, Mos54]. Kemperman [Lin61b]. Kempthorne [Fin54]. Kendall [Gea63a, Lan70b, Rao68, Rén63, Tra68, Wil59, Gou54a, Lan63]. Kendrick [Der61b]. Kérel [Cal62a]. kezdetei [Thi64]. Khintchine [Ken61]. kidnappings [dDBL34]. Kim [Har71]. King [Jar69]. Kingdom [Der53a, Grd64, Nix53, Sch68b, dW54]. Kingman [Har68]. Kiser [Ben64b]. Kish [Zar66]. Kizhakke [Cha64]. Klein [Tin50]. Klezl [Mad46]. Klezl-Norberg [Mad46]. Known [SS63, Fré36b, Thi66a]. Kogan [Ste57]. Koller [Weg57]. Kommission [Böh38]. Koopmans [All50, Der71b, Gou54c, Der71b]. Kotz [Buc71, Kem71, Lam68]. Kovero [Lin54]. Kravis [Der53a]. Krebskranken [Mey41]. Krebskranken- [Mey41]. Krebsproblems [Böh38]. Krebsstatistik [Böh38]. Krebssterbe [Mey41]. Krebssterbe- [Mey41]. Krebsvorsorgestatistik [Mey41]. Kreislaufkrankheiten [Wag38]. Krickeberg [vR68]. Kriege [Ano33i]. Kurth [Hec69]. Kurtosis [Zar65]. Kurzfristig [WW38]. Kutty [Cha64]. Kuznets [Jos53]. Kyburg [Hub66].

[Arm70, Bar69b, Buc71, Dav62, Dem63, Dem62, Dur53, Fin54, Gir66, Gou51a, Gou53e, Hal57, Har67b, Har71, Haw67, Hoe70, Joh71, K.35, Kem71, Kha66, Kol57, Lai67, Le 69, McC66, Mos71a, Nix53, Nor66, Pey63, Rén61, Sha63, Som67b, dKT38, Tör68, Vad59, Vos71, Wer67, Woo68, Zar66, vR70, vZ66]. Labor [Ano35l, Hub41, Ste57]. Laborforce [Har70]. Laha [Bri68]. Laky [Thi63d]. Lamperti [Hen68]. Lampkin [Min69]. L’analyse [Ben67b]. Ländern [Ano34j]. Landesplanung [Ver61]. Lange [Gou65, Rom66]. Langford [Bow52]. Langlebigkeit [ZH42]. Laplacian [Put51]. Laplacienne [Put51]. Large [Bos66, Gla62]. Larson [Rop70]. Larsson [Ben65]. l’Association [Ano62r]. last [Sau68]. last-born [Sau68]. Latent [NvdP66]. Latin [Bar65b]. latins [Bar65b, BKB65]. Latscha [Har65]. lattice [Sav64]. Law [Ano68d, Wol67]. Lawley [Vos65]. Lawrence [Tin50]. Laws [Pfa68]. Lazar [Ves68b, Ves67]. Lazard [Riv53]. League [Hub39]. Lebenshaltungskosten [Nix57]. Lebensjahre [Win44]. Lecam [Le 69]. l’économétrie [Dag71, Sch65]. Lecons [And44, Mal69]. legislation [Ano33h, Ano34g, Ano34h, Ano34i, Ano35h, Ano36i, Ano37i, Ano38m, Ano38n, Ano38o, Ano39h, Ano39i, Ano40f, Ano41f, Ano41g, Ano43e, Ano43f, Ano44c, Bow46, Bun33, Bun37, Dun40, Dun42, Gin33, God37, God39, Hec34, Hec35, Hub38a, Ide36a, Mey42, Mor35a, Ric36, Ric45a, Stu35, WI37, dL38, dS36, Ano33h, Ano34g, Ano34h, Ano34i, Ano35h, Ano36i, Ano37i, Ano38m, Ano38n, Ano38o, Ano39h, Ano39i, Ano40f, Ano41f, Ano41g, Ano43e, Ano43f, Ano44c, Bow46, Bun33, Bun37, Dun40, Dun42, God37, God39, Hec34, Hec35, Hub38a, Ide36a, Mey42, Mor35a, Ric36, Ric45a, Stu35, WI37, dL38, dS36]. lehrt [Wür35]. Lellouch [Ves70a]. Lemos [Gue62]. length [Hab62]. l’entreprise [Ves68f, Bar65a]. Léon [Tin47a]. Leontief [Ken67]. Leser [Amu68]. Leslie [dSR58]. L’essai [Ves70a]. L’Estimation [dO67]. leur [Jul38b, Jul38c]. Level [BR66]. Levels [Hal67, Kev67]. Lewis [Ham58, AB66]. l’homme [Ves70a]. liaisons [Dar53]. libero [Vin36]. Lieberman [Pey61]. Liebmann [Win44, Nix55, Win44]. Life [Dem66, Ham67, MH50, Wid67, Her40, Her42, Her44, Win44, Vin56]. Light [Ken66a, Böh38, Mar64]. Likelihood [Bri64, Fis57]. limit [Dug55]. Limited [Vin56]. limites [BP51, BPH52, Dar45, Dug55, Fré46, Her52]. Limits [Bri64, Ham53, BP51, BPH52, Dar45, Fré46, Fré60, Her52]. Linder [Gou45, Gou51c, Gou61a, Ham54, Pea62]. Lindley [Rén67]. linéaire [Dar53]. linéaires [Fré57]. Linear [Gut68, KC70, Sen71, ST71, Vad59, Zar65, vdP65, Dar53, Fré57, All55a, Gou53c, KLM67]. linéarité [Fré36b]. linearity [Fré36b]. Lines [HT58, TvY56]. l’information [For67, Lin66b]. links [Dar53]. Linnik [Mal69, Ple67]. l’Institut [Ano62l, Ide38]. List [Lan69, Lan70a, Lan71]. Literature [Gea63a]. Livi [Par70]. Living [Der53b, Nic54, Zar62a, Har35, Sau48, vZ33]. Livio [Par70]. Loadings [Sha56]. Logan [Bir59]. logical [Ziz33]. logische [Ziz33]. Logistics [JSH65]. Lognormal [Fin57]. Lohnes [Bea69]. London [Dre36, Stu34a]. Londres [Stu34a]. Long [Har70, Les67]. Long-Term [Les67]. Longevity [ZH42]. Longo [Mui65]. Lord [Cam64]. l’Organisation [Ano62d, Ano62i, Ano62j, Ano63e, Ano63f]. Lot [Ham50]. Lot-by-Lot [Ham50]. Louis [Sil54, Dun69a, Kul58]. Louis-Vladimir [Kul58]. Loveday [Nix62a]. Lovell [Bow52]. Loxton [Hun67]. loyers [Cal62d]. Lucas [Gou61b]. Lüdeke [Lüt65]. lui [Dre69]. Luigi [For62, Jul38b, Jul38c]. lumière [Mar64]. Lund [Gou70a]. Lundberg [Gil51]. l’Union [Ide38]. Lush [Fin54]. Lyon [All57].

M [Bar43, Bar46, Blo65, Bri68, Bun41, Cal62b, Cla68a, Dag64, Dag67, Dal62, Der53a, Der54a, Der71b, For68, G.38, Gan71, Gea38, Gea63a, Gou54a, Gou70a, Grd64, Gue62, Har65, Har71, Hub42, Hun67, Ill37, Ken61, Lan63, Lan70b, Lin66a, Mor69, Mor71a, Nai71, Qvi66, Rao68, Rén63, Rév70, Rio65, Sas71, Sch66, Ste57, Tie69, Tra68, Weg64, Wil59, Wol51, Wol55, Zar69, dW52, vdP65, vdB69, vdB62]. M. [BL38, Gin50a, Hus68, LS38, Mic38, Sau50, Win39]. Mack [Sto56]. Maddox [Fre71]. made [Dre69]. Madow [Rao61]. magyar [Thi64]. Mahalanobis [Dal63, Rao63a]. mai [Ano60k]. Main [Ano33n, Ano33o, Ano33p, Ano34q, Ano34r, Ano34s, Ano34t, Ano35m, Jul34]. Maizels [Koc63]. Major [Gin52, Iss52]. Making [MSC^{+}70, Ney62]. maladies [Div43]. Malinvaud [Dag71, Sch65]. Management [Gil62, Har67b, JSH65, Nor66, TI66]. Manchester [vBW34]. Mann [Rao61]. Manning [Oak71]. Manpower [Der71a]. Manual [Nor66]. Manuel [Col64]. Manuila [Spo65]. Marcello [Ugg69]. March [Ric64]. marges [Fré60]. margins [Fré60]. mariée [Gin41]. Mario [Bar58, Lop56]. Maritz [Kru71]. market [Sto48]. Markoff [dS71c]. Markov [Bar61, Lin61b, Lin61a, Bit71, RR68, Mih69]. Markovketten [dS71b]. Marnata [Cal62d]. married [Gin41, Liv36]. Martin [Vaj65]. Martingales [RR68]. Martti [Lin54]. Masculinité [Sau68]. Masculinity [Sau68]. Masstheorie [Lep68]. Masze [Win38b]. Matalon [Ben67b]. Material [GF55, JL43a, JL43b]. Materials [Dem50, Dix71a, Qua71a]. matériels [JL43a, JL43b]. Matérn [Nai67]. Mathematical [Dar58a, Gan66, Gan71, Gep58, Hem65, Hen68, Kha58, Kin64, Lan62, Le 69, Vad59, Cal66, Fré46, MK70, Cra69, Ken61]. Mathematicians [Dav69]. Mathematics [Fin54, Mos71c, Der63, Roy69, Gou54b, Hun67, Pér71, Wet70]. Mathematik [Ano71a, Wis71]. Mathématique [And44, Dag64, MK70]. mathématiques [Cal66, Fré46, Iva69]. mathematisch [Pea62]. mathematisch-statistischer [Pea62]. mathematische [F.36, Kol57, Lin69]. mathematischen [Kád68]. Mather [Gri66]. matières [Ano63p]. matrices [Dug69]. Matrix [Osb66]. Matter [Ano33j, Ano33k, Ano33l, Ano34m, Ano34n, Ano34o, Ano34p, Ano35i, Ano35j, Ano35k, Ano36l, Ano36m, Ano36n, Ano36o, Ano37j, Ano37k, Ano37l, Ano37m, Ano38p, Ano38q, Ano39j, Ano39k, Ano39l, Ano44d, Ano45c, Ano46d, Ano47c, Ano48d, Ano49f, Ano50e, Ano52e, Ano53h, Ano53i, Ano54d, Ano55d, Ano56d, Ano57d, Ano58d, Ano59c, Ano60e, Ano60f, Ano61h, Ano61i, Ano61j, Ano62e, Ano62f, Ano62g, Ano63c, Ano63d, Ano64e, Ano64f, Ano65d, Ano65e, Ano65f, Ano66g, Ano66h, Ano66i, Ano67a, Ano67b, Ano67c, Ano68e, Ano68f, Ano68g, Ano69a, Ano69b, Ano69c, Ano70e, Ano70f, Ano70g, Ano71b, Ano71c, Ano71d]. Matthai [Vos69]. Maurice [And44, Bje69]. Mauritius [Sil54]. Max [Riv53]. Maximizing [Ekm69]. Maximum [Bri64, Fis57]. Maxwell [Vos65]. May [Ano60m, Ano60k]. Maynard [Rio65]. McCarthy [Dal65, Ste57, vdP65]. McCracken [Sza64a]. McGarrah [JSH65]. mean [Put51]. meaning [Hos46]. Measure [Har68]. Measurement [Cro69, Fin56, Flu38, Gol60, Ohl61, Sch68b, Sto58, Suk52, The55, Bur37, Gin41, Par62, Win42]. Measurements [Har67a, Sny34b, Hen53]. measures [Fré40, Win38b, Kin69]. Measuring [Wil35, Gin42, Sau48, Mar71]. Meat [KC70]. médécine [Gin37b, Gin37a]. médecins [Ves67]. median [Fré42]. médiane [Fré42]. Medians [Har71]. Medical [Arm66, Gou63, Mar69, Sch71b, Wag38]. Medicine [Sch71b, Gin37b, Gin37a]. Medizin [Gou62]. Mediziner [Ano71a, Gou45, Gou51c, Gou61a, Wis71]. medizinischen [Wag38]. Meeting [Ohl61]. Member [Ric56]. Members [Ano48a, Ano49a, Ano51a, Ano52a, Ano53b, Ano53a, Ano53c, Ano54a, Ano55a, Ano56a, Ano57a, Ano58a, Ano66c, Les52, Ric47a, Ano33b, Ano33c, Ano33d, Ano34a, Ano34b, Ano38a, Ano38b, Ano39a, Ano40a, Ano41a, Ano46a]. membres [Ano33b, Ano33c, Ano33d, Ano34a, Ano34b, Ano38a, Ano38b, Ano39a, Ano40a, Ano41a, Ano46a, Ano48a, Ano49a, Ano51a, Ano52a, Ano53b, Ano53a, Ano53c, Ano54a, Ano55a, Ano56a, Ano57a, Ano58a, Ano66c, Les52, Ric47a]. Memorial [Wol70]. Menges [Kol60, Tör63a]. Menschen [Zah34]. Mensuel [Ano33t]. mercato [Für61]. Merrington [Har71]. Messen [Ham60]. Messina [Nor67]. mesure [Gin41, Gin42, Par62, Sau48]. mesures [Fré40, Hen53, Pre68]. Meteorology [FG65, NSW69]. Method [Bil64, Bri64, Dem63, Gin34, HT58, Mar66, Que51, Shr65, TvY56, Wil35, Gal51, Gin39b, Gin39a, Her42, Her44, Mee35, Put51, Gou54c, Ken65, Vos65, Ken65]. méthode [Gal51, Gin39b, Gin39a, Her42, Her44, Put51, Mee35, Bil70, G.38, Gou58]. Methoden [Dal58, Gou45, Gou51c, Gou61a, Gou62, Pét66, Ver58]. Methodenlehre [Gou62, Mad46]. Méthodes [Pto34, Ano36k, HU36, Jul35, Mor71b, Mor71c, dS33, Bil65, Dag71, Hub42, Mar61, Sch65, Weg64]. Methodik [Gou51a]. methodological [Pie40, Pie44]. Methodologie [Gin33, Ano33h, Ano34g, Ano34h, Ano34i, Ano35h, Ano36i, Ano37i, Ano38m, Ano38n, Ano38o, Ano39h, Ano39i, Ano40f, Ano41f, Ano41g, Ano43e, Ano43f, Ano44c, Bow46, Bun33, Bun37, Dun40, Dun42, God37, God39, Hec34, Hec35, Hub38a, Ide36a, Mey42, Mor35a, Ric36, Ric45a, Stu35, WI37, dL38, dS36]. methodologique [Pie40, Der58, Pie44]. methodologische [Nix57]. Methodology [Bar66b, MSC^{+}70, Ano33h, Ano34g, Ano34h, Ano34i, Ano35h, Ano36i, Ano37i, Ano38m, Ano38n, Ano38o, Ano39h, Ano39i, Ano40f, Ano41f, Ano41g, Ano43e, Ano43f, Ano44c, Bow46, Bun33, Bun37, Dun40, Dun42, Gin33, God37, God39, Hec34, Hec35, Hub38a, Ide36a, Mey42, Mor35a, Ric36, Ric45a, Stu35, WI37, dL38, dS36]. Methods [AHW^{+}71, Bow33, Bri68, BN69, Day49, Dem68a, Dur58, For68, Gou53e, Gou70a, Hal57, ID40, KS53, Ken61, NWW68, Neu67, NT70, Ple67, Ste60, SUD50, Tak64, Wer67, Wes34, Wol55, Ano36k, HU36, Jul35, Sto48, dS33, Pto34, Bje69, Dal68a, Gei69, Kru71, Par66, Qvi66, Wol68, Zar69, Gou53d]. Methorst [Ide55]. métodos [Can55]. Metropolitan [Ric56, Zwi60]. Metze [Oak71]. Mewes [dW52]. Meyer [Kre65]. Michaely [Tie69]. microbial [Div43]. microbiennes [Div43]. Microvariables [The52]. Migration [Nix54, Büc36]. militaire [Aue34, Gin37b, Gin37a]. military [Aue34, Gin37b, Gin37a]. Miller [Sch67]. Millham [Pér71]. Millikan [vdB69]. Millions [Oma67]. Milton [Hem65, Dow71]. Mincer [Sch71a]. Minimax [NT70]. Minor [Gin52]. Mises [Elf66, Elf66]. Mitchell [Thi63c]. Mitra [Vos69]. Mixed [Bil56, MS70]. mixte [Dor36, Ide38, Sch36]. mobilità [G.34b]. Mobility [Fou66, Ste57]. Mode [Qua71b, Cal62b, Fré41]. Model [And52, Cra69, Hus70, Nor70, Wid67, HT71]. modèles [Ben66, BMO^{+}70, Bil65, Bil70, Cal66, Mal61, The69]. Models [Der71a, Dur60, Gre65, Kri67, Rio65, She71, Spj68, Wol68, vdP70, Ben66, BMO^{+}70, Cal66, Mal61, All50, Der60, Men70, Cla71, Der65]. Modern [Wil37, And68, Gou68]. moderne [Ves68c, Dal58]. modernen [Gou62]. moderner [Gou51a]. modernes [Wol51]. Modification [CAON70]. Modified [Osb66]. Möglichkeit [Sae35]. Möglichkeiten [Ven68]. mois [Pto34]. Molenaar [Luk71]. Molinari [Bar62, Mic38]. Moments [Aig70, CF38, Dre69]. monde [Ano33n, Ano33o, Ano33p, Ano33q, Ano33r, Ano33s, Ano34q, Ano34r, Ano34s, Ano34t, Ano34w, Ano34x, Ano34y, Ano34z, Ano35e, Ano35f, Ano35g, Ano35m, Ano35n, Ano35o, Ano35p, Ano35q, Ano36e, Ano36f, Ano36g, Ano36h, Ano36p, Ano36q, Ano36r, Ano36s, Ano37e, Ano37f, Ano37g, Ano37h, Ano37n, Ano37o, Ano37p, Ano37q, Ano38h, Ano38i, Ano38j, Ano38k, Ano38l, Ano38r, Ano38s, Ano38t, Ano38u, Ano38v, Ano39e, Ano39f, Ano39g, Ano39m, Ano39n, Ano39o, Ano40d, Ano40e, Ano40g, Ano40h, Ano41d, Ano41e, Ano41h, Ano41i, Ano42c, Ano42d, Ano42e, Ano42f, Ano43c, Ano43d, Ano43g, Ano43h, Ano44b, Ano44e, Ano45b, Ano45d, Ano46c, Ano46e]. mondiale [vdB62]. Monetary [Sas71]. Monographie [Dag64]. montano [G.34c]. Monte [Ken65, Qvi66]. Month [Gin34, Pto34]. Monthly [Ano33t]. Moore [Jua71]. moral [JL43a, JL43b]. Moran [Rén63]. moraux [JL43a, JL43b]. Morbidity [Sco67a, Bir59]. Morice [Gou58, Ves68a]. Moritz [Kai52]. Morrell [McC66, Wij69]. mort [Hor69b]. mortalité [BP50, Hub36b, Pto34, SL61, Stu38]. mortality [BP50, Hub36b, Pto34, Stu38, Win42, Wür35, SL61, Ben65]. Mortara [Kin68]. Moser [Joh62, Ste60]. Moses [Pey63]. Most [Sag58]. Mosteller [Arm70, Kha66, Kit62a, Kit62b]. mot [Zah33]. Mothes [Ves68e, Ves68f]. motrices [Hub33, Hub36a]. mouvement [Her42, Mor41]. mouvements [Tin47a]. movement [Büc36, Her42, Mor41]. moyenne [Fré41, Her40, Put51]. Mr. [Gin50a, Mic38, Sau50]. muestreo [Zar62b]. Multi [Hus70]. Multi-Regression [Hus70]. Multinormal [DS71a]. multiphase [Tik67]. Multiple [HT58]. Multistage [Stu64, Rio65]. Multivariate [Mit71, Rao58]. multja [K.35]. Munich [And52, The52]. Murray [Der60]. Murthy [Zar69].

N [Buc71, Cal62c, Coh66, F.36, Gan67, Gou53a, Han69, Har67b, Har71, Kem71, Lai67, Nor66, Pai58, Rao69, Sco67b, Sve63, Tia68, Tra68, Vos65, Vos71, Zar69, vR70, Lan63]. nach [Ano33i, Ano34j, Dep57, Mee40]. name [Wey38]. Narrenspiegel [T.36]. nascite [Sav33]. Natalité [BL38, LS38, JL43a, JL43b, Pre68, Win39]. National [Ano70d, Arm70, Cla58, Der53a, Der65, Flu38, Gil53, Gin49, Gin50c, Gin52, Gol60, Har67a, Hem65, ID40, Les67, Lin62, Mos71c, Nix54, Qua71a, Shi37, SUD50, Sto56, vdB61, Ano33i, Gin45, Bje62, Bje69, Sto50, vdB69]. nationale [Gin45]. Nations [Ano61q, Ano62s, Ano62t, Ano63q, Ano63r, Ano63s, Ano63t, Ano64o, Ano64p, Ano65p, Ano66t, Ano66u, Ano67o, Ano67p, Ano67q, Ano68r, Ano68s, Ano69m, Ano69n, Ano69o, Ano70p, Ano70q, Ano70r, Ano71o, Off61, Per41, Sto58, Ano71f, Hub39, Per41, G.48]. Natural [Sco67b, Gin41, Hen53, Her42, Win38b]. naturel [Her42]. naturelle [Gin41, Hen53]. natürlichen [Win38b]. Naturwissenschaft [Gou62]. Naturwissenschafter [Gou45, Gou51c, Gou61a]. Need [SEM^{+}70]. Negative [Blo65]. Nemtchinov [Rja65]. Nerlove [Der71b]. nés [Sau68]. Netherlands [ID40, MT34, Stu38, Tin47b, Gan66, Tie69]. Neue [Bur37, Win38b]. neuvième [Dar55]. News [Ano47d, Ano48e, Ano49h, Ano50f, Ano51h, Ano51i, Ano51j, Ano52h, Ano53j, Ano53k, Ano54e, Ano55e, Ano56e, Ano57e, Ano58e, Ano59d, Ano60g, Ano60h, Ano61k, Ano61l, Ano61m]. Neyman [Cra68, Dar58a, Elf69, Gep58, Kha58, Le 69, Sve66, Elf69]. Nicolaas [Gou61b]. Nineteen [Ric59a]. Ninth [Dar55]. No [Ano33t]. Nobel [Wol70]. Noether [God68]. nom [Ide38, Sch36]. nombre [Fré42, Fré48, G.38, Sor71, vZ38]. Nomenclature [Hub38c]. Non [HW70, Vad59, Gin39b, Gin39a, Mor71a, Pur69, Thi66a, vE61]. non-central [vE61]. non-null [Pur69]. non-parametric [Thi66a]. Non-Sampling [HW70]. Nonlinear [vdP70]. Nonparametric [God68, Gov66, Hoe68, Gna67]. Norberg [Mad46]. Nordic [HG43]. nordiques [HG43]. Normal [Dok66, Dow71, EG65, Ham58, Rut65, Gin37b, Gin37a, Har71]. normale [Vin36]. normales [Gin37b, Gin37a]. Normung [Dre36]. North [Haw67]. Northeast [Man68a]. Northern [Osb66]. Norwegian [Nor70]. Note [Ano70c, Bar71, Bax71, KG67, NvdP66, Pra53, Ram70, Sil64, The53, Gin37a]. Notes [Der54c, Dix71a, Ham50]. Notice [Ano33b, Ano33c, Ano33d, Ano34a, Ano34b, Ano38a, Ano38b, Ano39a, Ano40a, Ano41a, Ano46a]. notion [Oni36]. Notions [Pri36, Her44]. nouveau [Fré41]. Nouveaux [Fré45]. Nouvelles [Ano47d, Ano48e, Ano49h, Ano50f, Ano51h, Ano51i, Ano51j, Ano52h, Ano53j, Ano53k, Ano54e]. November [Fou66, Nic54]. November-December [Nic54]. nul [Fré48]. null [Pur69]. Number [Azo67, Sha56, Fré42, Fré48, Sor71, vZ38]. Numbers [AFR^{+}63, Ano33m, Bos66, Pfo66, Ken69, Ziz33, Ple67]. Numerical [Nor67]. numériques [Ste67]. Numéro [Ill37]. numerous [Hab62]. nuova [Vin39]. nuptialité [Fed69]. Nutritional [Kev67].

O [Bra69, Cla58, F.36, Fin54, Gou55, Hal57, Ple67, Sag58, Ver58]. Oakford [Pey63]. obersten [Für53]. observation [WW38]. Observational [Suk52]. Observations [BL38, Gin50b, Gin50a, KB71, LS38, McC63, Mic38, Mol36, Nys36, PA54, Win39, dS33, vE55, dS38, BL38, Coc50, Gin50b, Gin50a, LS38, Mic38, Mol36, Sau50, Win39, dS33]. occasion [Bol60b]. Occasions [Kul63, Tik67]. Occupancy [Bit71]. Occupied [Nix38]. October [Ric64]. OEEC [Ric56]. oeuvre [Ano35l]. öffentliche [Der57]. Office [Ano71f, MT34, Ano35l, Ano61q, Ano62s, Ano62t, Ano63r, Ano63s, Ano63t, Ano64o, Ano64p, Ano65p, Ano66t, Ano66u, Ano67o, Ano67p, Ano67q, Ano68r, Ano68s, Ano69m, Ano69n, Ano69o, Ano70p, Ano70q, Ano70r, Ano71o, Hub41, Off61]. Offices [Ano35e, Ano35f, Ano35g, Ano36e, Ano36f, Ano36g, Ano36h, Ano37e, Ano37f, Ano37g, Ano37h, Ano38h, Ano38i, Ano38j, Ano38k, Ano38l, Ano39e, Ano39f, Ano39g, Ano40d, Ano40e, Ano41d, Ano41e, Ano42c, Ano42d, Ano43c, Ano43d, Ano44b, Ano45b, Ano46c]. Official [Dal68b, Ken66c]. Ogburn [Ric59b]. Ökonometrie [Gou51b, Lüt65, Tör63a]. old [Win38b]. older [Hub36b]. Olkin [Rao61]. OMNITAB [Nor67]. One [Spr71, Hab62, Ric68]. One-Sided [Spr71]. Ones [CS70]. Oosterhoff [Spr71]. Operation [Sch71b]. opérationnelle [Bil65, Bil70]. Opinions [Ano36y, Dal65, Ano37v]. Optimal [Ekm69, Tho69]. Optimization [Men70]. Optimizing [Dem63, vdP70]. Optimum [Bil56, Kul63, NT70, Rut65, Kün65]. Order [Dow71, KS53, Fré48, Sav64, Sav33]. Ordinal [Har71]. ordinate [Gin39b, Gin39a]. ordine [Sav33]. ordonnées [Gin39b, Gin39a]. Organisation [Ano63g, Ano64h, Ano64i, Ano65g, Ano65h, Ano65i, Ano66k, Ano66l, Ano66m, Ano67f, Ano67g, Ano67h, Ano68i, Ano68j, Ano68k, Ano69e, Ano69f, Ano69g, Ano70j, Ano70k, Ano70l, Ano71h, Ano71i, Ano71j, Ano33h, Ano34g, Ano34h, Ano34i, Ano35h, Ano36i, Ano37i, Ano38m, Ano38n, Ano38o, Ano39h, Ano39i, Ano40f, Ano41f, Ano41g, Ano43e, Ano43f, Ano44c, Ano47b, Ano48c, Ano49d, Ano49e, Ano50c, Ano50d, Ano51e, Ano51f, Ano51g, Ano52d, Ano53f, Ano53g, Ano54c, Ano55c, Ano56c, Ano57c, Ano58c, Ano59b, Ano60c, Ano60d, Ano61a, Ano61e, Ano61f, Ano61g, Bow46, Bun33, Bun37, Dun40, Dun42, Gin33, God37, God39, Hec34, Hec35, Hub38a, Ide36a, IM52, Mey42, Mor35a, Ric36, Ric45a, Stu35, WI37, dL38, dS36]. organiser [Lin66b]. Organization [Ano54c, Ano55c, Ano56c, Ano57c, Ano58c, Ano59b, Ano60c, Ano60d, Ano61a, Ano61e, Ano61f, Ano61g, Ano62d, Ano62i, Ano62j, Ano63e, Ano63f, Ano63g, Ano64h, Ano64i, Ano65g, Ano65h, Ano65i, Ano66k, Ano66l, Ano66m, Ano67f, Ano67g, Ano67h, Ano68i, Ano68j, Ano68k, Ano69e, Ano69f, Ano69g, Ano70j, Ano70k, Ano70l, Ano71h, Ano71i, Ano71j, Gil53, Lin62, Ano33h, Ano34g, Ano34h, Ano34i, Ano35h, Ano36i, Ano37i, Ano38m, Ano38n, Ano38o, Ano39h, Ano39i, Ano40f, Ano41f, Ano41g, Ano43e, Ano43f, Ano44c, Bow46, Bun33, Bun37, Dun40, Dun42, Gin33, God37, God39, Hec34, Hec35, Hub38a, Ide36a, Mey42, Mor35a, Ric36, Ric45a, Stu35, WI37, dL38, dS36]. Orient [HU36]. Origin [Ric56]. origines [Thi64]. Oskar [Gou65, Kel60, Rom66]. Ostle [Gou56]. Ostry [Nix67]. other [Gin39b, Gin39a, SL61]. Ottawa [Ano62l, Ano62m]. Otto [And57]. Outliers [Fer61]. outline [Sto48]. Output [Ken67, Ohl61, Osb66, Ven68]. ouvrières [Har36]. Overall [Her34]. overcome [Hos46]. Overview [Ano35l, Gué65]. Owen [Pey61, Pey62].

P [Arm70, AB66, Bea69, BF68, Bir59, Buc69, Cla71, Dal63, Dem54, Gou70a, Har65, Har71, Iva69, Jua71, Kin69, Les67, Lin61a, Mal68, Mit71, Mor71b, Mor71c, Nor67, Oma67, Rao63a, Rén63, Rév66, Sto56, Sto58, Sza64b, Ves68b, Weg64, Wer67, Wid67, vR70, vdB63]. pair [Hab62]. Paired [Shr65]. Pakistan [Der53b]. Panse [Suk69]. Papers [Cra68, Dav67, Der71b, Elf66, Elf69, Gan67, Ohl61, Lin71a, Whi69]. Papua [BW69]. Parallélisme [Mar33]. Parameters [Hus70, The51, Gum65]. parametric [Thi66a]. paramètriques [Thi66a]. Pareto [Wol67, Aig70, Ano68d]. Paris [Cal62a, Cal62d, Ano60i, Ano60j]. Partially [Gué65, Mas67]. particular [Dar53, Fré46]. particulier [Fré46]. particulière [Dar53]. partiellement [Gué65, Mas67]. Passage [Lin61b]. Past [Sto58]. paths [Bur37]. Patil [Buc69]. Pattern [Hen54]. Patterns [Ped69, Cla61]. Paul [Gou53c, Sae36, Für65]. payements [Der58]. Payments [Tie69, Ano34j]. pays [Ano33n, Ano33o, Ano33p, Ano33q, Ano33r, Ano33s, Ano34q, Ano34r, Ano34s, Ano34t, Ano34w, Ano34x, Ano34y, Ano34z, Ano35e, Ano35f, Ano35g, Ano35m, Ano35n, Ano35o, Ano35p, Ano35q, Ano36e, Ano36f, Ano36g, Ano36h, Ano36p, Ano36q, Ano36r, Ano36s, Ano37e, Ano37f, Ano37g, Ano37h, Ano37n, Ano37o, Ano37p, Ano37q, Ano38h, Ano38i, Ano38j, Ano38k, Ano38l, Ano38r, Ano38s, Ano38t, Ano38u, Ano38v, Ano39e, Ano39f, Ano39g, Ano39m, Ano39n, Ano39o, Ano40d, Ano40e, Ano40g, Ano40h, Ano41d, Ano41e, Ano41h, Ano41i, Ano42c, Ano42d, Ano42e, Ano42f, Ano43c, Ano43d, Ano43g, Ano43h, Ano44b, Ano44e, Ano45b, Ano45d, Ano46c, Ano46e, Cal66, Col64, HG43, HU36, Ide38, Sch36, MT34, Stu38, Tin47b]. Pays-Bas [MT34, Stu38, Tin47b]. Pearce [Fin66]. Pearson [Gou55, Bol64, Cra68, Dav67, Ham58, Hem69, Sve66, Cra68, Dav67]. Penal [Sch36]. Pénale [Sch36]. pénitentiaire [HG43, Sch36]. Pentikäinen [Hoe70]. people [Zah34]. Pèpe [Weg64]. percentage [vE61]. performance [Ide36b, Sch71a]. Permutations [Pey63]. Perrache [Mor69]. Perrier [Weg64]. Perrot [Cal62b]. Personal [Tay68]. Personnel [Per41]. Pesonen [Hoe70]. Peter [Hor69b]. petit [Fré42]. petits [Fré43]. Petropolis [Dar55]. peut [Dre69]. Pfanzagl [Gou62, Gou62]. Ph [Dal65, Rao58, Ste57, Ves67, Zwi53]. Phelps [Per41]. phenomena [Fré46]. phénomènes [Fré46]. Philosophy [Ney57]. Photograph [Ano62k, Ano64j].

Phyllis [Thi63c]. Physical [Ano62r, Dem50, HT71]. Physics [Dar58a, Wer67]. physiques [Ano62r]. Pierre [Che68]. Pietra [For61]. plan [Dug65]. Planen [Ham54]. Planning [Ben64b, Som67a, vdB69, Gea65]. Plans [Abr54, Bil56, Gué65, Rut65]. Platzer [Jac64]. Players [Dav71]. Poch [Zar62b]. point [Div43]. points [Gal51, vE61]. Poisson [Luk71, Mar68, HT71]. Policies [NWW68, Tie69]. Policy [Kre65]. Political [Nix62b, Nix64, Nix65, Nix67]. Pólya [Kin64]. popolazione [G.34b]. Population [Ano63q, BW69, Gin43, Has33, Hor69a, Key67, Nix38, Sch66, She71, Tek65, Wil37, vH66, Bun38a, Bun38b, Bur37, Her42, Les35, Mor41, Put51, Tik67, Win38b, Win42, Win44, vdB62, Les35, Har70, Hay70]. Populations [God65, Wid67, Ves70b]. Port [Sil54]. portée [Fré58]. position [Tin36]. Positive [DS71a, Sor71]. positives [Sor71]. possibilités [Fré46, Mor41]. Possibilities [vH66, Fré46, Mor41]. Possibility [Nys36, Sae35]. Possible [vE55]. possibles [G.38]. possibly [Div43]. Post [Hun71, Nix54]. Post-War [Nix54]. Potential [Mih69, BP51, BPH52, Her42, Her44, Her52]. potentials [Her40]. Potentiel [Win44, Win44]. Potentiel-vie [Win44, Win44]. potentielle [BP51, BPH52, Her52]. potentiels [Her40, Her42, Her44]. potentiels-vie [Her40, Her42, Her44]. Pour [Her34, Ano62r, Bun38b, Cal66, Der58, Div43, Gin37b, Gin37a, Ide38, Mol36, Pto34, Put51, Sch36]. pouvoir [Sau48]. Power [Der53a, Dur58, Sau48]. Prabhu [Coh66]. Practical [Gna67, Voi70, Fré58, Ano37v, Sun71]. Practice [Bir59, McC66, Que51, Box70]. practices [Gin42]. Prague [Hub38b]. Prais [All55b, KG67]. pratique [Ano37v, Bar67, Fré58, Ver63, vdB68]. pratiques [Gin42]. précisions [Her44]. Prediction [Mok62]. Preface [Ney70, Stu33]. préliminaire [And44]. Preliminary [Stu64, Nix36, Nys38]. Premier [And44]. Première [Fré46]. premières [Sor71]. préparatoires [Hub38c]. preparatory [Hub38c]. present [Sve66, Sto58]. Presentation [BJT63, Ped69]. Présenté [BL38, LS38, Mic38, Win39]. Presented [Bol60c, BL38, LS38, Mic38, Ohl61, Ano37v, Win39]. présentés [Ano37v]. President [Ano62k, Ano64j, Ide53, Ide53]. Presidential [Bol60b, Bol60c]. Presidentiel [Bol60b]. presque [Fré58]. Pressat [vdB68]. Prest [dW54]. prévision [Mal61]. Prévisions [Ves68f]. Price [AFR^{+}63, Kon58, McC63, Bow52, Der54b]. Prices [Sny34c, Pla34]. Primer [Gou53a]. Principaux [Ano33n, Ano33o, Ano33p, Ano34q, Ano34r, Ano34s, Ano34t, Ano35m, Der58, Jul34]. Principes [Dag67, Iva69]. Principles [Lin62, Cla68a, Jua71, Lin66a, Mar69]. Prior [NT70]. Prison [Sch36, HG43]. prix [Ste67]. Prize [Wol70]. probabilistes [Ben66]. probabilistic [Ben66]. probabilità [Mui65, Vin36]. probabilité [Ves68b]. probabilités [Dug55, For67, Fré41, Fré46, Gea38, Hos35, Wol51, And44, Ano68c, Ves68d, Ves68c]. Probabilities [Dow71, EG61, Ham53, vE55, Dug55, Fré46, Hos35]. Probability [Blo65, BR66, Cac71, Dar58a, Elf66, Gan71, Gep58, Gil62, Gre67, Ios71, Ken71, Kha58, Kin69, Kri69, Le 69, Mos63, Mos71a, Neu67, Pey61, Rao61, Rao69, Rén61, Rén67, Rop70, Ste57, Fré41, Ken69, Mos71c, Har68, Hub66, Rén63, Doo71, Hen68, Joh71, Kit62a, Kit62b, vR68]. Problem [CS70, Ekm69, Lin61b, MS70, Wey38, Win37, Böh38, Thi68]. problème [Thi68]. Problèmes [For60, Gir66, Dor33, Mal69]. Problems [BAM^{+}70, CAON70, Cra66, Gep58, Gin52, HR68, Ken66b, Kul63, LNS^{+}70, MCC70, Sen71, Dor33, For60, HT71, Amu68, Hor69a, Wij69]. Procedure [JHH61, Dal65]. Procedures [Sen71, Bea69]. Proceedings [Fou66, Dar58a, Gep58, Haw67, Kha58, Le 69, Nai67]. Process [AMO^{+}70, BMO^{+}70, Ste57]. Processes [Ano68d, Bar61, Bar69b, Bit71, Cra66, EG61, For68, Joh71, Lin64, Luk70, Mih69, Sco67b, Wal62, Wol55, Wol67, Zar65, Coh66, Lin61a, Mor66, Sch67]. Processing [BF68]. processus [BMO^{+}70, RR68, Gir66]. Prodan [Nai71]. Product [Cla58]. Production [GF55, Hal57, Sny34a, SEM^{+}70, Tin42, vH66, Ben64a, Vad59, Bru70, JSH65]. Productive [Flu38]. Productivity [Bru70, GF55, Ohl61, Der61b]. Products [Der53a]. Prof [BL38, LS38, Wey38, Win39]. Prof. [BL38, Coc50, Gin50b, Her44, LS38, Win39]. professeur [Thi64]. Professional [Hor68a]. Professor [Tör63a, Rao63a, Tin42]. professzor [Thi64]. profile [Gin39b, Gin39a]. profils [Gin39b, Gin39a]. Program [Nor67, Ano60i, Ano62l, Hos37]. Programme [Ano60i, Ano60j, Ano62l, Ano62m, Gou64, Hos37]. Programmed [Wet70]. programmes [Fré57, Ves67]. Programming [KWT^{+}63, Roy69, ST71, vdP65, Vad59, Bra67, Sza64a, Sza64b]. Programs [Gre68, Fré57]. Project [Qua71a]. Projections [Tör68]. Projects [Bic68b]. Properties [KS53, Del50]. proportion [Bol37]. Proportions [Gar71, Gin37b, Gin37a]. propos [Div40, Hub37, Met33a, Met33b, Met35, Met45]. Proposal [Bli49]. proposals [Mol36]. propositions [Mol36]. propres [Dug69]. propriétés [Del50]. Prospects [LNS^{+}70]. Protein [Kev67]. protestados [dDBL34]. Protestants [Nix61c]. protested [dDBL34]. Provisional [Ano60j, Ano60l, Ano62m, Ano60i, Ano62l]. provisoire [Ano60i, Ano60l, Ano62l]. Prozesse [dS71c]. Pseudo [McC69]. Pseudo-Replication [McC69]. psychologie [Ben66]. psychology [Ben66]. Public [Cro69, Kre65, Wil35]. Publications [Ano61q, Ano62s, Ano62t, Ano63q, Ano63r, Ano63s, Ano63t, Ano64o, Ano64p, Ano65p, Ano66t, Ano66u, Ano67o, Ano67p, Ano67q, Ano68r, Ano68s, Ano69m, Ano69n, Ano69o, Ano70p, Ano70q, Ano70r, Ano71f, Ano71o, Off61]. Published [McC66]. Purchasing [Der53a, Dur58, Sau48]. Pursuit [Gre65].

Quadratic [Bra67]. Quality [Dem66, Der54b, Hal67, Lai67]. Quantitative [Men70, Sen71]. Quantum [Wer67]. quatrième [Ano63h]. que [Dre69]. Quelques [Her44, Mor41, Tin47b, dS38, vZ33, Ano36k, Div43, Her40, Jul35, Pre68, Har36]. Quenouille [Ken61]. question [Gin37b, Gin37a, dS38]. Questions [Ken66c]. Quetelet [Jul38b, Jul38c]. Queue [Loy70]. Queueing [Ken61]. Queues [Tak64, Lin64, Coh66]. qui [Dre69]. Quick [GN68, Gum65]. quotas [Tin47b]. quote [Für61].

R [Arm70, AB66, Bar65a, Bar68, Bea69, Bje62, Bje69, Bri68, Bru70, Col64, Dal62, Dem63, Der53b, Der60, Der63, Dow71, Dri67, Dum70, Gan66, Gol60, Ham67, Har70, Har65, Hec69, Hoe70, Hub42, Ios71, Jar69, Joh71, JSH65, Jos53, KLM67, Kit62a, Kit62b, Kit64, Kit68, Kre65, Lin67, Mal68, Man68b, McC66, Myb63, Nor67, Ped69, Pey63, Roy69, Sch67, Sch68b, Som67a, Sto56, Thi63c, Thi64, Tia68, Tin50, Upp68, Ver63, Ves70a, Vos69, Wal66, Wet70, dM63, dW54, vdP65, vdB68, vdB62]. Råde [Mos71a]. radiodiffusion [Ide36c, Ide38, Ide38]. Ragnar [Wol70]. Rahman [Rao69]. Rain [LNS^{+}70, MSC^{+}70]. Rain-Making [MSC^{+}70]. Rainfall [AMO^{+}70]. Raja [Cha64]. Ramachandran [Mor71d]. Random [Ano68d, Ham58, Joh71, Mor66, Nad68, Pey63, Sri70, Stu64, Tör63b, Wal66, vZ66, Dre69, Fré42, Fré48, MK70, Azo67, Ken71, Wol67]. Randomness [ML69]. Rank [Ben67a, Sav64, Dow71]. Ranking [Dav71]. rankings [Hab62]. Ranz [Ver58]. Rao [Vos69, KLM67, Kit68, Som67a]. Rapport [Ano57f, Ano60k, Ano60l, Ano63h, Bol62, Bun38b, BL38, Dor36, Fré36a, Hub38c, LS38, Mic38, Coc50, Gin50b, Gin50a, Sau50, Win39, Ide38, Sch36]. rapports [Ano37v]. rappresentazione [Vin39]. Rate [BL38, LS38, Sau63a, JL43a, JL43b, SL61, Win39]. rates [Her34, Hub36b, Pto34, Win38a]. Rational [Ped69]. Ratios [Spj68]. Rauly [dO67]. Raum [Ver61]. Raumforschung [Ver61]. raw [Her34]. Raymond [Che58]. Rayner [Cur70]. razionale [Vin36]. Re [Gou64]. Re-Appraisal [Gou64]. real [vZ33]. recensement [HU36]. recensements [Mor41]. Recently [Ano61q, Ano62s, Ano62t, Ano63q, Ano63r, Ano63s, Ano63t, Ano64o, Ano64p, Ano65p, Ano66t, Ano66u, Ano67o, Ano67p, Ano67q, Ano68r, Ano68s, Ano69m, Ano69n, Ano69o, Ano70p, Ano70q, Ano70r, Ano71f, Ano71o, Off61]. récents [Gin50a, Sau50]. Recherche [Thi54, Ano67m, Ano67n, Ano68o, Ano68p, Ano68q, Ano69k, Ano69l, Ano71n, BSH71, Bil65, Bil70, BH69, Col64, HK70, MPDoA71, PBW70, Pur70]. Recherches [Ano62n, Ano62o, Ano63i, Ano63j, Ano63k, Ano64k, Ano64l, Ano65j, Ano65k, Ano65l, Ano66n, Ano66o, Ano66p, Ano67i, Har35, Ano62h, MT34, Wol51]. Recommendations [Ano71g]. reconstructing [Mor41]. reconstruction [Mor41]. rectangular [Sid64]. Recueil [Ano34u]. recul [JL43a, JL43b]. redditi [Der61a]. Reducing [Dal69]. Reeb [Ves68g]. réels [vZ33]. Reference [Cla58, Fin57, For68, Hal57, McC66, Muh65, Sen71, Wol55]. Referenten [Win39]. Reformvorschläge [Joh54]. Regard [Ken66c]. regarding [For60]. Regeln [Ham60]. Region [Man68a]. Regional [Wid67]. Regionale [Ver58]. regionaler [Ven68]. Regions [EG65, dS33, dS33, Zwi71]. Registers [vH66, Les35]. Registration [Man68a]. registres [Les35]. Regression [Cla71, EG65, Fis62, HT58, Hus70, MS70, Pra53, PA54, Sen71, The53, Tia68, KG67]. rehabilitation [Her34, Her34]. Reid [Bar61].

Reinhard [vdB62]. Rejection [Fer61]. Related [Cla71, Kin64, MS70, Roy69, Vaj65]. relating [MT34]. Relation [Ben67a, Sny34c]. Relations [Gin47, Gea63b, All55a]. Relationships [The57]. relative [Für61]. relatives [MT34]. relevés [Teg62]. Reliability [Cra66, Lai67]. Remarks [Gin50c, Gea63b, Wey38]. remédier [Hos46]. rendement [Ide36b]. René [Bun58, Gru64]. Renewal [Kit64]. Rényi [Doo71, For67]. répartition [Fré39, Fré45, Led55]. Répertoire [Ano34v]. Replacement [Rao66]. Replacing [CS70]. Replicated [Tör63b]. Replication [McC69]. replications [Cla68b]. Reply [Coc50]. Réponse [BPH52, Sau50, Her44]. Report [Ano49g, Ano57g, Ano60m, Ano60l, Ano63l, BL38, Cox60, Cra33, Gil53, Hen54, Hot50, LS38, Mic38, Nix36, Nys38, Rao63b, Gin50b, Gin50a, Sau50, Win39, Ano57f, Ano60k, Ano63h, Böh38, Bol62, Bun38b, Dor36, Fré36a, Hub38c, Ide38, Sch36, Sau63a]. Reporting [Cro69]. reports [Ano37v]. repräsentativen [Mee35]. representation [Dre69]. representative [Gal51, Mee35]. Representing [Vra64]. Research [And52, Ano62h, Ano62n, Ano62o, Ano63i, Ano63j, Ano63k, Ano64k, Ano64l, Ano65j, Ano65k, Ano65l, Ano66n, Ano66o, Ano66p, Ano67i, Ano67m, Ano67n, Ano68o, Ano68p, Ano68q, Ano69k, Ano69l, Ano71n, BSH71, Bic68b, BH69, Day49, Dem63, Gil51, Gin50c, Gna67, Gou56, Hal57, HW70, Har67a, HK70, Les67, MPDoA71, Ohl61, PBW70, Pur70, Sha63, Har35, Wag38, Der65, MT34, vdB61, Gou63, Nix54, Sto56, Ben64b, Gan67]. researchers [Dug65]. Residenzstadt [K.35, dKT38]. Response [BPH52, BN69, Sau50, Win39, Her44]. Restrictions [KC70]. Résultats [Jul34, Der58, Har36, Thi59]. Results [BJT63, Lam68, Har36, Jul34, Sto48, Thi59]. Revenu [Hub42]. Revenues [Ekm69]. revenus [Fré39, Fré45]. Review [All50, All55b, All55a, Amu68, And68, And44, Ano68c, Arm70, AB66, Azo67, Bar65a, Bar67, Bar69a, Bar68, Bar43, Bar46, Bar66a, Bar61, Bar69b, Bar66b, Bea69, Ben64b, Ben65, Ben67b, Bic68a, BF68, Bil64, Bil65, Bil70, Bir59, Bje62, Bje69, Blo65, Bor66, Box61, Box70, Bra67, Bra69, Bri68, Bri71, Bru70, Buc69, Buc71, Bun41, Cac71, Cal62a, Cal62c, Cal62d, Cal62b, Can55, Car69, Cla58, Cla61, Cla68a, Cla71, Coh66, Col64, Cra68, Cra69, Cur70, Dag64, Dag67, Dag70, Dag71, Dal58, Dal62, Dal63, Dal65, Dal68a, Dal69, Dar58a, Dav62, Dav67, Dav69, Dav70, Dem54, Dem61, Dem63, Dem66, Dem68a, Dem62, Dep57, Der53a, Der53b, Der54a, Der54b, Der57, Der58, Der60, Der61a, Der61b, Der63, Der65, Der68, Der71a, Der71b]. Review [Doo71, Dow71, Dri67, Dum70, Dur53, Dur58, Eic65, Elf66, Elf69, F.36, Fal69, Fed69, Fin54, Fin57, Fin66, Fla65, For67, For68, Fou66, Fre71, Für53, Für61, G.34a, G.34b, G.34c, G.38, G.48, Gan66, Gan67, Gan71, Gar71, Gea38, Gea63a, Gea65, Gei69, Gep58, Gil51, Gil62, Gir66, Gna67, God65, God68, Gol60, Gor67, Gou45, Gou51a, Gou51b, Gou51c, Gou53b, Gou53c, Gou53e, Gou53d, Gou53a, Gou54a, Gou54b, Gou54c, Gou55, Gou56, Gou57, Gou58, Gou61a, Gou63, Gou65, Gou68, Gou70a, Gou70b, Gov66, Grd64, Gre67, Gri66, Gut68, Hal57, Ham54, Ham58, Ham60, Ham67, Har67a, Har70, Har68, Har67b, Har65, Har71, Haw67, Hay70, Hec69, Hem65, Hem69, Hen68, Hen61, Hen54, Hor69a, Hub42, Hub66, Hun67, Ill37]. Review [Ios71, Iva69, Jar69, Joh54, Joh71, Joh62, JSH65, Jos53, Jua71, K.35, Kád68, KLM67, Kap67, Kem71, Ken61, Ken65, Ken71, Ken67, Kha58, Kha66, Kin64, Kin69, Kit62a, Kit62b, Kit64, Kit68, Koc63, Kol57, Kol60, Kol65, Kre65, Kri67, Kri69, Kru71, Kün65, L.36, Lai67, Lam68, Lan62, Lan63, Lan70b, Lau64, Le 69, Lea68, Lep68, Les67, Lin56, Lin66a, Lin67, Lin69, Lin71a, Lin61b, Lin61a, Lin64, Lin66b, Loy70, Luk71, Lüt65, Mac70, Mad46, Mal68, Mal69, Man68b, Mar68, Mar71, Mar61, Mar66, Mar69, McC66, Mee40, Men70, Mew65, Mih69, Min69, Mit71, Mok62, Mor66, Mor69, Mor71b, Mor71c, Mor71a, Mor71d, Mos54, Mos71a, Muh65, Muh69, Mui65, Nai67, Nai71, Nel70, Neu67, Nix53, Nix54, Nix57, Nix58a, Nix61a, Nix62b, Nix64, Nix65]. Review [Nix67, Nor67, Nor66, Oak71, Ohl61, Oma67, Pai58, Pan63, Par66, Pea62, Ped69, Pér71, Per41, Pét66, Pey61, Pey62, Pey63, Pfa68, Ple67, Pre68, Qvi66, Rao58, Rao61, Rao68, Rao69, Rén61, Rén63, Rén67, Rév66, Rév70, Rid64, Rio65, Rop70, RR68, Roy69, Rut65, S.34, Sae36, Sag58, Sas71, Sau63a, Sav36, Sax66, Sch71a, Sch59, Sch67, Sch65, Sch66, Sch68b, Sco67b, ST71, Sha63, She71, Shr65, Som62, Som67a, Som67b, Spr71, Sri70, Ste67, Ste57, Ste60, Sto50, Sto56, Sto58, Stu41, Sun71, Sve63, Syk68, Sza64a, Sza64b, T.36, Tag57, Teg62, dKT38, The69, Thi63c, Thi64, Tia68, Tie69, Tin47a, Tin50, Tör63a, Tör68, TI66, Tra68, Upp68, Vaj65, Ven68, Ver58, Ver61, Ver63, Ves67, Ves68a, Ves68d, Ves68c]. Review [Ves68b, Ves68e, Ves68f, Ves68g, Ves70a, Ves70b, Vin56, Vos65, Vos69, Vos71, vBW34, Wal67, Wal66, Weg64, Wer67, Wet70, Whi66, Whi69, Wid67, Wij69, Wil59, Wil68, Win57, Wis71, Wol51, Wol55, Wol67, Wol68, Woo68, Zar62a, Zar62b, Zar66, Zar69, Zim69, Zwi53, Zwi60, Zwi71, dO67, dS71b, dS71c, dW52, dW54, vR70, vR68, vZ66, vdP65, vdP70, vdB69, vdB68, vdB61, vdB62, vdB63, Gal51, Hub38c, Wet70]. Reviews [Ano64g, Ano66j, Ano67d, Ano68h, Ano70h, Ano70i, Gou62, Hoe70, Sch68a, Vad59]. Revision [Gal51, Hub38c]. Rhodesia [Osb66]. Rice [Dun64, Dun69b, Ide53]. Richard [Elf66]. richesse [Gin45]. rigueur [Sau63b]. Rilevazioni [Tag57]. Riordan [Sch59]. Rios [Can55]. Risk [Hoe70, Hoe70]. Risser [Dri67, Bun58]. Rist [Rue54]. river [BMO^{+}70]. Rivet [Che58]. rivières [BMO^{+}70]. Roach [Sri70]. Road [EG65, vZ38]. Robert [G.34b]. Robin [Dav71]. Robust [Sen71]. Rocher [Iva69]. Rodolfo [Gin57b]. Roessler [Rén61]. Role [Par67, Dug69]. Romagnolo [G.34c]. Rome [Ric53]. Ronald [Ano62k, Gou63, Yat62]. Roos [Nix58b]. Roots [NvdP66]. Rouanet [The69]. roughness [Par62]. Round [Ano69d, Ano71g, Dav71]. Round-Table [Ano69d]. Rounds [Boy58]. Rouquet [Iva69]. Rourke [Kit62a, Kit62b]. routière [vZ38]. Rowe [Sch68b]. Roy [Bri68]. Royal [G.34a]. rugosité [Par62]. Rules [Fer61, Kün65]. Run [Dum70]. rural [Bun38b, Bun38a]. rurale [Bun38a, Bun38b]. Rymes [Nix67].

S [All55b, Bar66a, Box70, Buc71, Can55, Car69, Cra68, Dag70, Dal69, Dav67, Dem68a, Fin66, For68, Gil62, Gou53c, Gou55, Ham58, Har67a, Har68, Har71, Hun67, Jar69, Jos53, Kap67, Kem71, Kru71, Lam68, Les67, Mar71, Min69, Muh65, Nix67, Par66, Rao61, Rid64, Sri70, Tör68, Vad59, Vos69, vBW34, Whi66, Whi69, Wol55, vdP70]. S. [Whi69]. sa [Ano36y, Fré42]. Sabin [Spo65]. Sagoroff [Win71]. Saibante [Bar58]. salaires [vZ33]. Salaries [Nix53]. Sales [Ekm69]. same [Wey38, Win42]. Sammlung [Gou62, Pea62]. Sampford [Dal62]. Sample [Ben67a, BW69, Dow71, JHH61, Man68a, MS70, Nad68, NT70, Pur65, Thi54]. Samples [McC69, Fré43, Sid64]. Sampling [Abr54, Aig70, Bil56, Dal62, Dal63, Dal71, Dem50, God65, Hal67, HW70, Kul63, McC63, Rao66, Rid64, Rut65, Sco67a, SUD50, Stu64, The51, Tör63b, Vos64, Dre69, Put51, Tik67, Syk68, Vos71, Zar66, Dal65, Dem54, Dem68a, Gan66, Kol65, Zar69]. Samuel [Coc64]. Samuels [Sha63]. sanitaire [Bun41]. sans [Sau63b]. Sarkar [Pai58]. Satya [Log62]. Saunders [Gol60]. sauver [Fré50]. Sauvy [Gin50a, Sau50]. Savage [Dem62]. save [Fré50]. Savorgnan [L.36, dM63]. Scarf [Rio65, Vad59, Gil62]. Schätzprobleme [Lüt65]. Scheduling [Jar69]. Scheme [Ano63p, GV63]. Schleifer [Rut65]. Schluss [Rah33]. Schlusswort [Win39]. Schmetterer [Kol57]. Schmitt [Mar71]. Schneider [Bom71]. Schott [Fla53]. Schwartz [Ves70a, Sas71, Ves68b]. Schwob [Mor69]. Science [Bar66b, Ken66c, Ney57, Nix62b, Nix64, Nix65, Nix67, Pan63, Qua71a, Gin37b, Gin37a, Wol70, Gil62, Som67b]. sciences [Ano62r, Fré46, Hos39, Weg64, Gou70a, Sco67b]. Scientific [Bra69, MCC70, Div43, Fré46, Dem63, Der71b]. scientifique [Div43, Fré46]. Scientist [Gou70b]. Scientists [Joh71]. Scores [Har71]. Scott [Grd64, Joh62]. scrutin [Thi68]. Seal [Hoe70]. Search [Thi54]. Seasonal [Les66]. Seber [Gut68]. Second [Lan69, Nai67, Ric66, Ano60k, Ano60m, Gin37a, Gin39a]. secondo [Liv36, Sav33]. secuestros [dDBL34]. Seeding [CAON70, MCC70]. seine [Win38a]. seines [Sae36]. Selected [Dav67, G.48, Lin71a, Elf66]. Selection [Bic68b, Dem61, Elf69, Fin56]. Selections [Vaj65]. selon [Pto34, Sau68]. Semi [dS71c]. Semi-Markoff-Prozesse [dS71c]. Sengupta [Men70]. September [Ano60i, Dar58b, Ric47b, Ric49, Ric53, Ano60j]. septembre [Ano60i, Dar58b]. Sequential [Mad56, vE55]. séquentielles [Mor71a]. Sergejevitch [Rja65]. Serial [Bos66]. Series [AB66, Der54a, Dur60, Gil51, Jos53, Par67, Gin39b, Gin39a, Gin39b, Gin39a, Lea68, Rao68]. Server [Loy70]. Service [Bru70, Nix58a]. Services [Ano33n, Ano33o, Ano33p, Ano34q, Ano34r, Ano34s, Ano34t, Ano35e, Ano35f, Ano35g, Ano35m, Ano36e, Ano36f, Ano36g, Ano36h, Ano37e, Ano37f, Ano37g, Ano37h, Ano38h, Ano38i, Ano38j, Ano38k, Ano38l, Ano39e, Ano39f, Ano39g, Ano40d, Ano40e, Ano41d, Ano41e, Ano42c, Ano42d, Ano43c, Ano43d, Ano44b, Ano45b, Ano46c, Per41]. ses [BL38, Dor33, Her44, LS38, Oni36, Win39, Hos37]. Session [Ano36y, Ano49g, Ano57g, Ano57f, Ano60i, Ano60j, Ano60m, Ano60k, Ano60l, Ano62m, Ano63l, Ano63h, Ano64a, Ano64b, Ano66a, Ano66b, Ano68a, Ano68b, Ano70a, Ano70b, Bol60a, Bol60b, Bol60c, Bol62, Dar55, DG57, Dar58b, Dre36, Hub38b, Ric47b, Ric49, Ric51, Ric53, Stu34b, Stu34a, Stu36, Ano62l, Dar55, Dar58b]. Set [KB71]. s’éteint [Sav41]. Sets [CL68, Sil64, Tak62]. Setting [Bri64]. Seventh [Ric51]. Seventieth [Rao63a]. sexes [Bol37, Sau68]. Shape [The52]. Shaul [Myb63]. shed [Böh38]. Shelley [Rio65]. Shewhart [Dem68b]. Shift [Dow71]. Shirras [Wil55]. Short [Dum70, Les67, WW38]. Short-Run [Dum70]. Short-Term [Les67, WW38]. Shreider [Ken65]. Shubik [Vaj65]. si [Gin37b, Gin37a]. Side [Dix71b]. Sided [Spr71]. Sigmund [Fla53]. Sign [DS71a]. Significance [Gar71, Har65, Fré58, WW38]. signification [Hos46]. Simple [BT64, CS70]. Simulation [NWW68]. Simultaneous [Ano70c, EG65, KC70, Ram70]. Simultaneous-Equation [KC70]. Singer [Cal62a]. Singer-Kérel [Cal62a]. Single [Ano70c, Hal67, Loy70, Ram70]. Sir [Ano62k, Yat62]. Situations [Žar56]. Sitzungsbericht [Ver61]. Sixth [Ric49]. Size [Nix35, Nix52]. Skorohod [Mor66]. Slawtscho [Win71]. Slippage [GN68, Bar68]. small [Fré42, Fré43]. Smith [Haw67, Vos71, Tia68]. Smokler [Hub66]. Snedecor [Dal68a]. Snow [Her61]. so-called [Fré34, Fré50]. Sobre [dDBL34]. Social [Ano63q, Bow33, BL38, Der68, Fou66, Gou70a, Gou70b, Gre68, Joh62, LS38, Lau64, Nix54, Sil54, Ste60, SUD50, Vaj65, vBW34, Div43, Fré46, Hos39, Win38a, Win39, Der60]. Sociales [BL38, LS38, Hos39, Win39, Sax66]. Socialist [Gea65]. sociaux [Fré46]. Société [Hub39]. Societes [Ano62p, Ano63m, Ano63n, Ano63o, Ano64m, Ano64n, Ano65n, Ano67j, Ano67k, Ano67l, Ano68l, Ano68m, Ano68n, Ano33q, Ano33r, Ano33s, Ano34w, Ano34x, Ano34y, Ano34z, Ano35n, Ano35o, Ano35p, Ano35q, Ano36p, Ano36q, Ano36r, Ano36s, Ano37n, Ano37o, Ano37p, Ano37q, Ano38r, Ano38s, Ano38t, Ano38u, Ano38v, Ano39m, Ano39n, Ano39o, Ano40g, Ano40h, Ano41h, Ano41i, Ano42e, Ano42f, Ano43g, Ano43h, Ano44e, Ano45d, Ano46e, Ano47e, Ano48f, Ano49i, Ano50g, Ano50h, Ano51k, Ano51l, Ano51m, Ano52i, Ano52j, Ano53l, Ano53m, Ano54f, Ano55f, Ano56f, Ano57h, Ano58f, Ano59e, Ano60n, Ano60o, Ano61n, Ano61o, Ano61p, Ano62q, Ano65m, Ano65o, Ano66q, Ano66r, Ano66s, Ano69h, Ano69i, Ano70m, Ano70n]. Sociétés [Ano70o, Ano71k, Ano71l, Ano71m]. Societies [Ano47e, Ano48f, Ano49i, Ano50g, Ano50h, Ano51k, Ano51l, Ano51m, Ano52i, Ano52j, Ano53l, Ano53m, Ano54f, Ano55f, Ano56f, Ano57h, Ano58f, Ano59e, Ano60n, Ano60o, Ano61n, Ano61o, Ano61p, Ano62p, Ano62q, Ano63m, Ano63n, Ano63o, Ano64m, Ano64n, Ano65m, Ano65n, Ano65o, Ano66q, Ano66r, Ano66s, Ano67j, Ano67k, Ano67l, Ano68l, Ano68m, Ano68n, Ano69j, Ano69h, Ano69i, Ano70m, Ano70n, Ano70o, Ano71k, Ano71l, Ano71m, Ano33q, Ano33r, Ano33s, Ano34w, Ano34x, Ano34y, Ano34z, Ano35n, Ano35o, Ano35p, Ano35q, Ano36p, Ano36q, Ano36r, Ano36s, Ano37n, Ano37o, Ano37p, Ano37q, Ano38r, Ano38s, Ano38t, Ano38u, Ano38v, Ano39m, Ano39n, Ano39o, Ano40g, Ano40h, Ano41h, Ano41i, Ano42e, Ano42f]. societies [Ano43g, Ano43h, Ano44e, Ano45d, Ano46e]. Society [G.34a, vBW34]. Socio [Nor70]. Socio-Demographic [Nor70]. Solomon [Mok62]. Solution [Ekm69]. Solutions [Fis62, Gir66]. Some [Bit71, Bol64, BAM^{+}70, Der54c, EG61, Gea63b, Ham50, Hoe68, Ken66c, Kul63, Lin62, Mah69, McC63, NWW68, Nix61c, Pri36, Žar56, vE61, Div43, Her40, Thi66a, Ano36k, Har36, Her44, Jul35, Mor41, Tin47b, Win38b, dS38, vZ33, Syk68]. son [Dep64]. sondage [Thi54]. sondages [Thi59, Bar67]. Sonnabend [Gin56]. sont [Dre69, Fré60, Gin37b, Gin37a]. Soubeyroux [Cal62c]. Sources [Bje69, G.48]. sous [Col64]. sous-développé [Col64]. soziale [Fla65]. sozialen [Win38a]. Sozialleben [Gou62]. Sozialwissenschaften [Gou62]. Sozialwissenschaftlichen [vdB63]. Space [NT70, Vos64]. Spanish [dDBL34]. Spatial [RMK^{+}70b]. Special [Cla58, Fin57, For68, Hal57, Ken66c, McC66, Muh65, Sen71, Wol55]. specialement [Hos39]. Spécialisation [Gou51d]. Specialization [Gou51d]. Specification [CF38, The57]. Spectral [Par67, Lea68]. Speech [Bol60b]. spopolamento [G.34c]. Sprent [Cla71]. squares [Bar65b]. SS [Mar64]. Stable [Wid67]. Städte [Joh54]. Städten [Büc36]. Städtestatistik [dW52]. Stadtregionen [Ver61]. Staehle [Nix61b]. Staff [Hor68a]. Stage [Hus70, Stu64]. Stand [Für53, Mee40]. Standard [DS71a, NvdP66, Der53b]. Standardisation [Ano35l]. Standardization [Ano35l, Dre36]. Standards [Nix35, Nix52, Hem65]. Standpunkt [Joh54]. Stange [Kád68]. Stanley [Coc64]. star [Mar64]. Start [Tör63b]. state [Gin50b, Sve66]. States [Cro69, Der53a, Jos53, Lin62, Nix61a, Tin50, Tör68, Wil35, Der61b, Sas71, Sch66]. stationäre [Win42]. Stationary [Lin61b, Wal62, Win42]. Statistica [Bar69a, Sav36, Sav33]. Statistical [Ano33a, Ano33n, Ano33o, Ano33p, Ano34q, Ano34r, Ano34s, Ano34t, Ano34v, Ano35e, Ano35f, Ano35g, Ano35m, Ano36e, Ano36f, Ano36g, Ano36h, Ano37e, Ano37f, Ano37g, Ano37h, Ano38h, Ano38i, Ano38j, Ano38k, Ano38l, Ano39e, Ano39f, Ano39g, Ano40d, Ano40e, Ano41d, Ano41e, Ano42c, Ano42d, Ano43c, Ano43d, Ano44b, Ano45b, Ano46c, Ano47a, Ano47b, Ano47d, Ano47e, Ano48b, Ano48c, Ano48e, Ano48f, Ano49b, Ano49c, Ano49d, Ano49e, Ano49g, Ano49h, Ano49i, Ano50a, Ano50b, Ano50c, Ano50d, Ano50f, Ano50g, Ano50h, Ano51b, Ano51c, Ano51d, Ano51e, Ano51f, Ano51g, Ano51h, Ano51i, Ano51j, Ano51k, Ano51l, Ano51m, Ano52b, Ano52c, Ano52d, Ano52h, Ano52i, Ano52j, Ano53d, Ano53e, Ano53f, Ano53g, Ano53j]. Statistical [Ano53k, Ano53l, Ano53m, Ano54b, Ano54c, Ano54e, Ano54f, Ano55b, Ano55c, Ano55e, Ano55f, Ano56b, Ano56c, Ano56e, Ano56f, Ano57b, Ano57c, Ano57e, Ano57g, Ano57f, Ano57h, Ano58b, Ano58c, Ano58e, Ano58f, Ano59a, Ano59b, Ano59d, Ano59e, Ano60a, Ano60b, Ano60c, Ano60d, Ano60g, Ano60h, Ano60i, Ano60j, Ano60m, Ano60k, Ano60n, Ano60o, Ano61a, Ano61c, Ano61d, Ano61e, Ano61f, Ano61g, Ano61k, Ano61l, Ano61m, Ano61n, Ano61o, Ano61p, Ano61q, Ano62c, Ano62d, Ano62h, Ano62i, Ano62j, Ano62l, Ano62m, Ano62k, Ano62n, Ano62o, Ano62p, Ano62q, Ano62s, Ano62t, Ano63e, Ano63f, Ano63g, Ano63l, Ano63h, Ano63i, Ano63j, Ano63k, Ano63m, Ano63n, Ano63o, Ano63p, Ano63r, Ano63s, Ano63t, Ano64d, Ano64h, Ano64i]. Statistical [Ano64j, Ano64k, Ano64l, Ano64b, Ano64m, Ano64n, Ano64o, Ano64p, Ano65c, Ano65g, Ano65h, Ano65i, Ano65j, Ano65k, Ano65l, Ano65m, Ano65n, Ano65o, Ano65p, Ano66f, Ano66k, Ano66l, Ano66m, Ano66n, Ano66o, Ano66p, Ano66b, Ano66q, Ano66r, Ano66s, Ano66t, Ano66u, Ano67e, Ano67f, Ano67g, Ano67h, Ano67i, Ano67j, Ano67k, Ano67l, Ano67m, Ano67n, Ano67o, Ano67p, Ano67q, Ano68b, Ano68i, Ano68j, Ano68k, Ano68l, Ano68m, Ano68n, Ano68o, Ano68p, Ano68q, Ano68r, Ano68s, Ano69e, Ano69f, Ano69g, Ano69j, Ano69h, Ano69i, Ano69k, Ano69l, Ano69m, Ano69n, Ano69o, Ano70b, Ano70j, Ano70k, Ano70l, Ano70m, Ano70n, Ano70o, Ano70p, Ano70q, Ano70r, Ano71h, Ano71i, Ano71j, Ano71k, Ano71l, Ano71m]. Statistical [Ano71n, Ano71o, AB66, AHW^{+}71, Bea69, BSH71, Bil64, Bol62, BAM^{+}70, BH69, Cra68, Dal63, DG57, Dar58b, Day49, Dem62, Dep64, Dix71a, Elf69, Flu38, FG65, G.34a, G.48, Gea63a, Gea65, Gei69, Gin47, Gla62, Gou53d, Gou53a, GV63, Gou64, Ham71, HK70, Hot50, Hub38b, Hub39, Hub41, Ide38, IM52, Joh62, Jua71, KLM67, KWT^{+}63, Ken65, Ken66a, Kit62b, Lan69, Lan70a, Lin56, Lin62, LNS^{+}70, MPDoA71, MT34, Mos71c, Ney62, Nor67, Pan63, PBW70, Per41, Pey62, Pfa68, Pur70, Qua71b, Ric59a, Ric66, Rid64, Rob63, Rop70, Sch36, Sen71, Sny34c, Spr71, Stu34b, Stu34a, Voi70, Vos65, Vos69, vBW34, Wer67, Wes34, Ano33e, Ano33f, Ano33g, Ano33h, Ano34c, Ano34d, Ano34e, Ano34f, Ano34g, Ano34h, Ano34i, Ano35a, Ano35b, Ano35c, Ano35d]. statistical [Ano35h, Ano36a, Ano36b, Ano36c, Ano36d, Ano36i, Ano37a, Ano37b, Ano37c, Ano37d, Ano37i, Ano38c, Ano38d, Ano38e, Ano38f, Ano38g, Ano38m, Ano38n, Ano38o, Ano39b, Ano39c, Ano39d, Ano39h, Ano39i, Ano40b, Ano40c, Ano40f, Ano41b, Ano41c, Ano41f, Ano41g, Ano42a, Ano42b, Ano43a, Ano43b, Ano43e, Ano43f, Ano44a, Ano44c, Ano45a, Ano46b, Bow46, Bun33, Bun37, Div40, Div43, Dre36, Dre38, Dun40, Dun42, Fré43, God37, God39, Har36, Hec34, Hec35, Hor69b, Hos46, Ide36a, Lan71, Mey42, Mor35a, Ric36, Ric45a, Stu35, Thi68, Tin36, WI37, Ziz33, dL38, dS36, dS38, Ano33q, Ano33r, Ano33s, Ano33u, Ano33v, Ano33w, Ano34w, Ano34x, Ano34y, Ano34z, Ano34-28, Ano34-29, Ano34-30, Ano35n, Ano35o, Ano35p, Ano35q]. Statistical [Ano35r, Ano35s, Ano35t, Ano35u, Ano36t, Ano36p, Ano36q, Ano36r, Ano36s, Ano36u, Ano36v, Ano36w, Ano36x, Ano37n, Ano37o, Ano37p, Ano37q, Ano37r, Ano37s, Ano37t, Ano37u, Ano38r, Ano38s, Ano38t, Ano38u, Ano38v, Ano38w, Ano38x, Ano38y, Ano38z, Ano38-27, Ano39m, Ano39n, Ano39o, Ano39p, Ano39q, Ano39r, Ano40g, Ano40h, Ano40i, Ano40j, Ano41h, Ano41i, Ano41j, Ano41k, Ano42e, Ano42f, Ano42g, Ano43g, Ano43h, Ano44e, Ano44f, Ano45d, Ano45e, Ano46e, Ano46f, Ano71f, Ben64a, Ben66, For60, Gin33, Gou64, Har35, Hos35, Hub38a, Ide36b, Mar64, Oni36, All50, Bar66b, Bor66, Dal68a, Dav62, Gou63, Gou70a, Hal57, Ham67, Hen61, Lin61a, Mar66, Par66, Per41]. statistiche [Mor34, Mor35b, Tag57]. Statisticians [Gou55, Nai67, Ano35l]. statisticiens [Ano35l, Hub39, Hub41]. Statistico [Ano33a]. Statistics [Ano34u, Ano35l, Ano62r, Ano69d, Arm66, Bic68b, Bir59, Bje69, BR66, Buc71, Cra33, Dal68b, Dar58a, Dem66, Der54c, Dok66, Fin66, Gea65, Gep58, Gil53, Gou54b, Gov66, HDPD71, Hoe68, Hor68a, Hun71, Hun67, JHCS70, Kem71, Ken66c, Kha58, Lan70b, Le 69, Mos63, Mos71c, NSW69, Nix38, Nix61a, Nix62b, Nix64, Nix65, Nix67, Nor70, Nys36, Pér71, Qua71a, Rao61, Rao68, Rén67, Ric68, Sas71, Sch71b, Som67b, Sto50, SUD50, Wet70, Wil59, Wil36, Wil38, Yos71, Böh38, dDBL34, Bur37, Dor33, Dor36, Elf66, Gin50a, Gou51d, HG43, Hub33, Hub36a, Ide36c, Ide38, Ken69, Mol36, Pie40, Pie44, Pla34, Sae35, Sau50, Sav64, Sch36, Thi33, WW38, Wey38, Win37, Wür35, ZH42, Ziz38, vZ36, Ano36y, Ano64a, Ano66a, Ano68a, Ano70a, Böh38, Dor36]. Statistics [Lin62, Mic38, Nys38, Sav33, Stu36, Wag38, Zah34, Bri68, Bri71, Cla68a, Dem61, Fal69, Gan67, God68, Gou53e, Gou54a, Gou57, Gou68, Har67b, Hec69, Kap67, Kha58, Kit68, Lan62, Lin66a, Mar71, Mar69, McC66, Mos71a, Muh69, Ped69, Rao69, Rén61, Sun71, Thi63c, Whi66, vZ66, Dav69, Fin54, Gou56, Gou70b, Man68b, McC66, Nel70, Wil68]. Statistik [Böh38, Gou62, Mee40, Pla34, Sae36, Wag38, Wey38, Ziz38, vdB63, Wag38, Win37, Zah34, ZH42, F.36, Kád68, Kol57, Lin69, Mad46, Sag58, T.36, Gou62]. Statistiken [Für53]. Statistique [And44, Ano33q, Ano33r, Ano33s, Ano34w, Ano34x, Ano34y, Ano34z, Ano35n, Ano35o, Ano35p, Ano35q, Ano36p, Ano36q, Ano36r, Ano36s, Ano37n, Ano37o, Ano37p, Ano37q, Ano38r, Ano38s, Ano38t, Ano38u, Ano38v, Ano39m, Ano39n, Ano39o, Ano40g, Ano40h, Ano41h, Ano41i, Ano42e, Ano42f, Ano43g, Ano43h, Ano44e, Ano45d, Ano46e, Ano47e, Ano48f, Ano49i, Ano50g, Ano50h, Ano51k, Ano51l, Ano51m, Ano52i, Ano52j, Ano53l, Ano53m, Ano54f, Ano55f, Ano56f, Ano57h, Ano58f, Ano59e, Ano60n, Ano60o, Ano61n, Ano61o, Ano61p, Ano62q, Ano65m, Ano65o, Ano66q, Ano66r, Ano66s, Ano69h, Ano69i, Ano70m, Ano70n, Ano71k, Ano71l, Ano71m, Bar46, Dar58b, Ide36c, MT34, Ano33e, Ano33f, Ano33g, Ano34c, Ano34d, Ano34e]. statistique [Ano34f, Ano34k, Ano34l, Ano35a, Ano35b, Ano35c, Ano35d, Ano36a, Ano36b, Ano36c, Ano36d, Ano36j, Ano37a, Ano37b, Ano37c, Ano37d, Ano38c, Ano38d, Ano38e, Ano38f, Ano38g, Ano39b, Ano39c, Ano39d, Ano40b, Ano40c, Ano41b, Ano41c, Ano42a, Ano42b, Ano43a, Ano43b, Ano44a, Ano45a, Ano46b, Ano47a, Ano48b, Ano49b, Ano49c, Ano50a, Ano50b, Ano51b, Ano51c, Ano51d, Ano52b, Ano52c, Ano53d, Ano53e, Ano54b, Ano55b, Ano56b, Ano57b, Ano58b, Ano59a, Ano60a, Ano60b, Ano60i, Ano61c, Ano61d, Ano62c, Ano62l, Ano62r, Ano62p, Ano63m, Ano63n, Ano63o, Ano63p, Ano64d, Ano64m, Ano64n, Ano65c, Ano65n, Ano66f, Ano67e, Ano67j, Ano67k, Ano67l, Ano68l, Ano68m, Ano68n, Ano69j, Bar43, Ben64a, Ben66, Bun41]. statistique [Del50, Div40, Div43, Dor33, Dor36, Dri67, Gin50a, Gou51d, Gri66, Har36, HG43, Hor69b, Hub33, Hub36a, Ide36b, Mal69, Mar64, Mew65, Mol36, Oni36, Pie40, Pie44, Sau50, Thi64, Tin36, Ves67, Ves68b, vZ36, Ano36y, Ano57f, Ano60k, Ano63h, Ano64a, Ano66a, Ano68a, Ano70a, Bol62, DG57, Hub38b, Ide38, Sch36, Stu34b, Stu34a, Stu36, Dep64, Gou58, S.34, Ves68a, Bar65a, Mor71c, dO67]. Statistiques [Ano33n, Ano33o, Ano33p, Ano34q, Ano34r, Ano34s, Ano34t, Ano34v, Ano34u, Ano35e, Ano35f, Ano35g, Ano35m, Ano36e, Ano36f, Ano36g, Ano36h, Ano37e, Ano37f, Ano37g, Ano37h, Ano38h, Ano38i, Ano38j, Ano38k, Ano38l, Ano39e, Ano39f, Ano39g, Ano40d, Ano40e, Ano41d, Ano41e, Ano42c, Ano42d, Ano43c, Ano43d, Ano44b, Ano45b, Ano46c, Ano57e, Ano58e, Ano59d, Ano60g, Ano60h, Ano62o, Ano63j, Ano63k, Ano64k, Ano64l, Ano70o, Bar46, Ano33h, Ano33u, Ano33v, Ano33w, Ano34g, Ano34h, Ano34i, Ano34-27, Ano34-28, Ano34-29, Ano34-30, Ano35h, Ano35l, Ano35r, Ano35s, Ano35t, Ano35u, Ano36i, Ano36u, Ano36v, Ano36w, Ano36x, Ano37i, Ano37r, Ano37s, Ano37t, Ano37u, Ano38m, Ano38n, Ano38o, Ano38w, Ano38x, Ano38y]. statistiques [Ano38z, Ano38-27, Ano39h, Ano39i, Ano39p, Ano39q, Ano39r, Ano40f, Ano40i, Ano40j, Ano41f, Ano41g, Ano41j, Ano41k, Ano42g, Ano43e, Ano43f, Ano44c, Ano44f, Ano45e, Ano46f, Ano47b, Ano47d, Ano48c, Ano48e, Ano49d, Ano49e, Ano49h, Ano50c, Ano50d, Ano50f, Ano51e, Ano51f, Ano51g, Ano51h, Ano51i, Ano51j, Ano52d, Ano52h, Ano53f, Ano53g, Ano53j, Ano53k, Ano54c, Ano54e, Ano55c, Ano55e, Ano56c, Ano56e, Ano57c, Ano58c, Ano59b, Ano60c, Ano60d, Ano61a, Ano61e, Ano61f, Ano61g, Ano61k, Ano61l, Ano61m, Ano62d, Ano62h, Ano62i, Ano62j, Ano62n, Ano63e, Ano63f, Ano63g, Ano63i, Ano64h, Ano64i, Ano65g, Ano65h, Ano65i, Ano65j, Ano65k, Ano65l, Ano66k, Ano66l, Ano66m, Ano66n, Ano66o, Ano66p, Ano67f]. statistiques [Ano67g, Ano67h, Ano67i, Ano67m, Ano67n, Ano68i, Ano68j, Ano68k, Ano68o, Ano68p, Ano68q, Ano69e, Ano69f, Ano69g, Ano69k, Ano69l, Ano70j, Ano70k, Ano70l, Ano71h, Ano71i, Ano71j, Ano71n, BSH71, Bow46, BH69, Bun33, Bun37, Dag71, Dun40, Dun42, For60, Fré43, Gin33, God37, God39, Har35, Hec34, Hec35, HK70, Hos35, Hos46, Hub38a, Ide36a, Ide38, IM52, MPDoA71, Mar61, Mey42, Mor35a, Mor71b, Mor71c, PBW70, Pur70, Ric36, Ric45a, Sch36, Sch65, Stu35, Thi68, Ves68f, Weg64, WI37, dL38, dS36, dS38, Mic38, Ver63, Ves68g]. Statistische [Gou45, Gou51c, Gou61a]. Statistischen [Dre36, Für53, Ziz33]. statistischer [Dre36, Dre38, Gou51a, Pea62, Pét66]. Statistisches [Ano36t]. statisztikai [Thi64]. Statutes [Ano62r]. Statuts [Ano62r]. Stegun [Hem65]. Steiger [Eic65]. Steindl [Wol67]. Stellar [Hec69]. Stephan [Dal65]. Sterbetafeln [Rah33]. Sterblichkeit [Dep57]. Sterblichkeitsmessung [Win42]. Sterblichkeitsstatistik [Wür35]. Stevens [dSR58]. Stichproben [Kol60]. Stichprobenverfahrens [Mos54]. Stieltjes [Fré41]. Stigler [Nix58a]. Stimulation [LNS^{+}70]. Stochastic [AMO^{+}70, Bar69b, Coh66, Cra66, For68, Fra68, Gre65, Lin64, Luk70, Sch67, Sco67b, ST71, Wal62, Wol55, BMO^{+}70, Dar53, Gea63b, Par62, Hoe70, Kri67, Rév70]. Stochastically [Mal68]. stochastique [Par62]. stochastiques [BMO^{+}70, Dar53, Gir66, The69]. Stockholm [Ano57g, Nai67, Ano57f, DG57]. Stocks [Wij69]. Stolnitz [Vin56]. Stone [Bje62, Der60, Sch68b, Bje62]. Storage [Wij69]. Störmer [dS71c]. Straight [HT58, TvY56]. Strassert [Ven68]. Stratification [Ekm69, Gar71, Stu64]. Stratified [Bax71, KB71, Put51]. stratifiée [Put51]. Strecker [Dal58]. Streets [Rot63]. Structure [Cac71, Car69, Jos53]. Structures [MK70, Sax66, Ves70b]. Struktur [Ver58]. Struktur- [Ver58].

Stuart [Dun69b, Ide53, Lan70b, Rao68, Wil59]. Studden [Kap67]. Studenski [Sto58]. Student [Rut65, Sid64]. Students [BR66, Hun67, Som67b, Cur70]. Studies [Ken69, Sha63, Gin50a, Hos37, Sau50, Vad59, Dem61, Fou66, Gil62, Gou54c, Hub66, Kin64, Les67, Mok62, Mor66, Ohl61, Stu41, Vad59, Zar62a]. Study [Ano68d, Coh66, Cra33, Der53a, Dur53, Gla62, Grd64, Joh62, Voi70, Wol58, Wol67, Dar53, Del50, Div40, Div43, Dug69, Fré36a, Fré46, Har36, Her42, Mar64, Thi68, dS33, Ano70d, Har35, Arm70]. studystatistical [Del50]. Stuvel [Der68]. subaree [Für61]. subclass [Cla68b]. Subject [Aig70, Fin56]. Subjective [Bar71, Hub66]. Subjects [Hun67]. successive [Tik67]. Südenhorst [And57]. Sudman [Dal69]. sue [Für61]. sujets [vdB68]. Sukhatme [Dem54, Oma67]. Sulla [Liv36]. Sulle [Mor34, Mor35b]. Sum [Box70, Ekm63]. Supplement [Ano33t, Ano33t]. Supplementary [Mat62]. Supply [SEM^{+}70]. Support [Dix71a]. Supported [Qua71a]. Surface [BN69, CMK^{+}70, RMK^{+}70a, Par62]. Survey [Ano33m, Dal65, Dal71, Gin50c, Hen68, JHH61, KB71, Man68a, Nic54, Sil54, Som62, Vos71, Har35, Jul34, Thi59, Gor67, Ste60, Zar66]. Surveys [Dem54, HW70, Lou67, Sco67a, Suk52, Thi54, Dal69]. susceptible [Bun38a]. Süssmilch [Hor69b]. Sverdrup [Pfa68]. Swaroop [Log62]. Sweden [Med65, Tay68, Ben65, Dem66]. Swings [Har70]. Switzerland [Nix61c]. Symbolic [Sza64b]. Symmetric [Tra68]. Symposium [Dar58a, Fou66, Gep58, Haw67, Kha58, Le 69]. Systematic [Tör63b]. systematique [Ano63p]. Systems [Ano70c, Bit71, Bos66, Kap67, NvdP66, Ram70, TB62, Tho69, Tin41, Wol58, Oak71, vdP70, Der68, vR70]. Szabady [Fou66, Hor69a]. Szegö [Kin64]. székesföváros [dKT38, K.35]. Sztrem [Rom63]. Szturm [Rom63]. Szulc [Par66].

T [All50, Bar61, Ben65, Der71a, Fin54, Gou54c, Ham58, Hoe70, Nix67, Rév70, Sco67b, vBW34, Whi69]. Tabellen [Dre36, Kád68, Ven68]. Table [Ano63p, Ano69d, Ano71g, CL68, Sil64, Tak62, Har65]. Tableau [Ano34-27]. Tableaux [Ano33u, Ano33v, Ano33w, Ano34-28, Ano34-29, Ano34-30, Ano35r, Ano35s, Ano35t, Ano35u, Ano36u, Ano36v, Ano36w, Ano36x, Ano37r, Ano37s, Ano37t, Ano37u, Ano38w, Ano38x, Ano38y, Ano38z, Ano38-27, Ano39p, Ano39q, Ano39r, Ano40i, Ano40j, Ano41j, Ano41k, Ano42g, Ano44f, Ano45e, Ano46f, Fré57, Fré60]. Tables [Gou55, Gou63, Hem65, Vin56, Vos69, Wid67, Ano33u, Ano33v, Ano33w, Ano34-28, Ano34-29, Ano34-30, Ano35r, Ano35s, Ano35t, Ano35u, Ano36u, Ano36v, Ano36w, Ano36x, Ano37r, Ano37s, Ano37t, Ano37u, Ano38w, Ano38x, Ano38y, Ano38z, Ano38-27, Ano39p, Ano39q, Ano39r, Ano40i, Ano40j, Ano41j, Ano41k, Ano42g, Ano44f, Ano45e, Ano46f, Dre36, Fré57, Fré60, Rah33, Pey62, Tra68, Blo65, Har65, Hem69, Mor71a, Pey61, Pey63, Rut65]. Tafeln [Pea62]. Tagliacarne [Für61]. Takács [Gir66]. Takeuchi [CL68, Sil64]. Taking [Mar51, SL61]. tasks [Bur37]. Taux [Her34, SL61, Hub36b, Pto34]. Taylor [Har68, Tör68]. Tchebycheff [Kap67]. teach [Wür35]. Teachers [Mos71c]. Teaching [Ano69d, AHW^{+}71, Dix71a, Ham71, Mos71a, Yos71, Fré41, Ide36b]. team [Dug65]. Technik [Mos54, Gou62, Kol60]. Technique [CMK^{+}70]. Techniques [Amu68, Gou56, Ham71, Men70, Qua71a, Jua71, Kol65, Rio65]. Technology [Hun67]. Teil [vdB63]. Teixeira [Lop56]. Television [Mos63]. Ten [Hub38c]. Ten-Year [Hub38c]. tenant [SL61]. teniendo [dDBL34]. Term [Les67, WW38]. terms [Ziz38]. Territory [Sil54, BW69]. Test [And52, Ben67a, BJT63, GN68, The52, vE55, Pur69, Div43]. Testing [Dem66, Har65, Ken65, Rob63]. Tests [BT64, Box70, Gar71, Pur65, Spj68, Thi66a, Bar68, Spr71]. Text [Mal68]. Textile [Nor66]. Theil [All55a, Kün65, MS70]. Their [Bar61, Bit71, Coh66, Dal68b, Gin47, Joh62, Nor70, The51, Jul38b, Jul38c, WW38]. Theorem [Cor67a]. théorèmes [Dug55]. Theorems [Luk70, Dug55]. Theoretical [And52, Der71a, Gou54a, Muh65, Wer67, Hen53]. Theorie [Kol60, G.38, Div43, For67, Ves68c, Mos54, Bar67, Mor71b, Mor71c, Sax66]. Theories [Gin43]. théoriques [Hen53, Wol51, Ste67]. Theory [Bar61, Bar66b, Ben67a, Box61, Box70, Boy58, Cac71, Dal62, Dem54, Fis34, God65, Gre65, Gre67, Hal67, Haw67, Hoe70, Ken61, Ken66a, Kha58, Kri69, Lan70b, Lin64, McC66, Mor66, Mor71d, Ney62, Ple67, Que51, RRK65, Rao68, Rén65, Rév66, Rop70, Sch67, Sri70, Sto58, Syk68, Tak64, Tik67, Tör63b, Vad59, Vaj65, Wil59, Zar69, Zim69, Div43, Sav64, Sve66, Bar66a, Dav62, Doo71, Gan71, Gut68, Hen68, Hoe70, Ios71, Kit64, Lam68, Lan70b, Mih69, Neu67, Upp68, Wil59, vR68]. Therapeutic [LS71]. therapeutique [LS71, Ves70a]. These [Per41, Div43]. Third [Dar58a, Gep58, Kha58, Lan70a]. Thirring [Ill37]. Thirtieth [DG57, Ano57f, Ano57g]. Thirty [Ano63l, Ano60k, Ano63h, Dar58b, Ano60m]. thirty-first [Dar58b]. Thirty-Fourth [Ano63l, Ano63h]. thirty-second [Ano60k, Ano60m]. Thomas [Kit62a, Kit62b]. Thomasian [Cac71]. Thomson [Ait55]. Thonstad [Der71a]. Thorbecke [Zim69]. those [Thi59]. Three [vE55, Dur58]. ties [Hab62]. Time [Dur60, Lea68, Par67, Rao68, The52, Vos64, Win42]. Time-Series [Dur60]. Times [Nix61a]. Tinbergen [Wol70]. Tintner [Pér71, Gou53b, Gou54b]. Tippett [Gou53e]. Tisserand [Weg64]. Tisserand-Perrier [Weg64]. titles [dDBL34]. titulos [dDBL34]. Tjalling [Der71b]. Together [Per41]. Tokio [Ano60k, Ano60l]. Tokyo [Ano60m, Ano60k, Ano60l, Bol60a, Has33]. Tome [Bil65, Bil70, Iva69]. Topics [Kin64, Cla71]. Tosco [G.34c]. Tourism [Mic38, Mol36, Thi33]. Tourisme [Mic38, Mol36]. Tournament [Dav71]. Tovar [Gue62]. Towns [Joh62]. Trade [Sny34a, Sny34b, Sny34c, Koc63]. traditional [Hos46]. traditionnels [Hos46]. Traffic [Rot63, For60, vZ36, vZ38]. trafics [For60]. Training [Ano67m, Ano67n, Ano68o, Ano68p, Ano68q, Ano69k, Ano69l, Ano71n, BSH71, BH69, HK70, Hor68a, MPDoA71, PBW70, Pur70]. Traité [Mor69]. traitement [Lin66b]. Transformations [Bol64, Wal66]. Transportation [vdP70]. Trattato [Bar69a]. travail [Ano35l, Ano35l, Hub41]. travaux [Dor36, Hub38c]. Traynard [Dri67]. Treasurer [Cox60, Rao63b]. Treatment [RS68]. tree [Dre69]. Treillis [Dag64]. Trend [CMK^{+}70, Gna67, Hen54, RMK^{+}70a, Ric68]. Trend-Surface [RMK^{+}70a]. Trends [Der61b, Jos53, Nix58a]. Trente [vdB68, Ano60k, Ano63h, Dar58b]. trente-deuxième [Ano60k]. trente-quatrième [Ano63h]. trentième [Ano57f, DG57]. trial [LS71]. Tripolitania [Nic54]. tuberculeuses [Div40]. tuberculosis [Div40]. tudomány [Thi64]. Tukey [Bri64]. tun [Böh38]. Turkey [Gil53]. Turroni [Gia64]. TV [HDPD71]. TV-Course [HDPD71]. Twenty [Dar55, Ric47b, Ric49, Ric51, Ric53]. Twenty-Eighth [Ric53]. Twenty-Fifth [Ric47b]. Twenty-Ninth [Dar55]. Twenty-Seventh [Ric51]. Twenty-Sixth [Ric49]. Two [Ben67a, BT64, Dow71, Hal67, Hus70, Kul63, Ney62, Pur65, vE55]. Two-Sample [Dow71]. Two-Stage [Hus70]. Typical [Fré48]. typiques [Fré48].

U [Coh66, Dem66, Har67a, Kád68]. U.S [Hem65]. Überalterung [Fla65]. Udny [Ano52f]. Ueber [Rah33]. Umrisse [Bic68a]. Unadjusted [Mah69, Pra71]. Unbalanced [Spj68]. Uncertainty [Mar71]. Underdeveloped [Gil53, Sil54]. Underlying [Lin62]. undetermined [Led55]. unfounded [Fré50]. unième [Dar58b]. Unified [God65]. uniformité [vZ36]. Uniformity [vZ36]. Union [Ide38]. Unit [Bax71, KB71]. Unit-Stratified [Bax71, KB71]. United [Ano71f, Der53a, Der61b, Grd64, Jos53, Nix53, Nix61a, Sas71, Sch66, Sch68b, Tin50, Tör68, dW54, Ano61q, Ano62s, Ano62t, Ano63q, Ano63r, Ano63s, Ano63t, Ano64o, Ano64p, Ano65p, Ano66t, Ano66u, Ano67o, Ano67p, Ano67q, Ano68r, Ano68s, Ano69m, Ano69n, Ano69o, Ano70p, Ano70q, Ano70r, Ano71o, Cro69, Lin62, Off61, Wil35]. Units [Gla62, Stu64]. Univ [Wey38]. Univ. [Wey38]. Univariate [Buc71]. universe [Sid64]. Universities [vdB61]. University [Ano69d, BR66, Dem66, Haw67]. Unknown [Rut65, SS63, vE55]. unordered [Gin39b, Gin39a]. Unstable [Tin41]. Unternehmensforschung [Ham60]. Untersuchung [Nix57]. Upward [Ric68]. Urban [Kev67, Som62]. Urbanization [Zwi71]. Ursachen [Fla65, Win38a]. USA [JHH61]. usage [Fré36a, Ves67, Ves68b]. Use [AHW^{+}71, Fin57, GF55, Mat62, Qua71a, SUD50, Sto68, Voi70, Ano37v, Fré36a, Hos39, Aue34, Cal66]. Users [CAON70]. Uses [Dal68b]. Using [NWW68, Qua71b, Mor41]. Utilisation [Ano37v, Aue34, Cal66]. Utilization [Nor70]. Uzawa [Vad59].

V [Amu68, Ben64b, Bru70, Dem54, Hen54, Iva69, Lin64, Mal69, Mor66, Oma67, Pey63, Ple67, Rén67, Roy69]. Vacancies [Har67a]. valeur [Fré41]. Valeurs [Dug69, Fré48]. Valleron [Ves67]. Value [Kev67, Ric56, Fré41, Zah34]. Values [vZ66, Fré48]. Vandam [Arm70]. Vandellós [Ano52g]. variabilitá [Bar69a]. Variable [Nad68, Mar64].

Variables [DS71a, Dur54, Gea38, Ken71, Mat62, vZ66, Dre69, Fré42, Gea63b, Wal66]. Variance [Dok66, HR68, Spj68, The51, Wal67]. Variances [Rut65]. Variate [Fin56, Que51]. Variate-Difference [Que51]. Variation [Les66]. Variations [Pfa68]. variétés [Her40]. varieties [Her40]. Various [Sto50, Ano33n, Ano33o, Ano33p, Ano33q, Ano33r, Ano33s, Ano34q, Ano34r, Ano34s, Ano34t, Ano34w, Ano34x, Ano34y, Ano34z, Ano35e, Ano35f, Ano35g, Ano35m, Ano35l, Ano35n, Ano35o, Ano35p, Ano35q, Ano36e, Ano36f, Ano36g, Ano36h, Ano36p, Ano36q, Ano36r, Ano36s, Ano37e, Ano37f, Ano37g, Ano37h, Ano37n, Ano37o, Ano37p, Ano37q, Ano38h, Ano38i, Ano38j, Ano38k, Ano38l, Ano38r, Ano38s, Ano38t, Ano38u, Ano38v, Ano39e, Ano39f, Ano39g, Ano39m, Ano39n, Ano39o, Ano40d, Ano40e, Ano40g, Ano40h, Ano41d, Ano41e, Ano41h, Ano41i, Ano42c, Ano42d, Ano42e, Ano42f, Ano43c, Ano43d, Ano43g, Ano43h, Ano44b, Ano44e, Ano45b, Ano45d, Ano46c, Ano46e, Fré40, Ide38, Sch36]. Vasily [Rja65]. vegetable [Sae35]. Vème [Hub38c]. Vergangenheit [K.35]. Vergleich [Nix57]. Verlauf [Fla65]. verschiedenen [Ziz38]. Versuchen [Ham54]. versus [Wol58]. Verwaltung [dW52]. Vessereau [Mar61]. vie [Cal62a, Cal62b, Har35, Her40, Her42, Her44, Hub42, Sau48, Win44, vZ33, Win44]. vielen [dS71c]. Vienna [Wey38]. view [Div43]. Viewpoint [Rén67]. Views [Hor69a]. Vinayak [Suk69]. vingt [Dar55]. vingt-neuvième [Dar55]. Vital [Muh69]. Vladimir [Kul58]. Voeux [Ano36y, Ano37v]. voie [Cal66]. Vol [Dar58a, Elf66, Gep58, Gre67, Kha58, Kri69, Le 69, Mor71b, Mor71c, Ohl61, Rao68, Sch68b, Wil68, Zwi71]. Volkseinkommen [Ano33i]. volkswirtschaftlichen [Der57]. Volume [Ano34-31, Ano35v, Ano35w, Ano36z, Ano37w, Ano38-28, Ano39s, Ano40k, Ano41l, Ano42h, Ano42i, Ano43i, Ano49j, Ano50i, Ano51n, Ano52k, Ano53n, Ano56g, Ano60p, Ano61r, Ano62u, Ano63u, Ano64q, Ano64r, Ano65q, Ano65r, Ano66v, Ano66w, Ano67r, Ano68t, Ano69p, Ano70s, Lan70b, Ric56]. vom [Joh54]. vor [Ano33i]. voting [Thi68]. vs [Sny34a]. Vue [Gué65, Dav70, Div43].

W [Arm70, AB66, Bea69, Bir59, Bor66, Dal68a, Der61b, Fin54, Gou51b, Gou54c, Gre67, Ham67, Haw67, Hay70, Joh62, Kap67, Ken67, Kol65, Kri69, Lan62, Loy70, Luk71, Min69, Rao61, Rio65, Sun71, Syk68, Wal66, Whi69, vR70]. Wagemann [Lan56, T.36]. Wages [Ano33m, vZ33, Nix53]. Wagner [Für63]. Wahrscheinlichkeitstheorie [Lep68]. Wald [HT58, TvY56]. Wales [Nix35, Nix52]. Wallace [Kha66]. Walley [Har71]. Walsh [Nor67, Gov66]. Walter [Stu41, Ano64j, Dem68b, Lin71b, Ric61, Ric64]. Wanderungsbewegung [Büc36]. War [Nix54, Ano33i]. Was [Böh38, Wür35]. Wasan [Rév70]. Washington [Ric47b]. Water [SEM^{+}70]. Wattelar [Fed69]. Wealth [Der54a, Gil51, Jos53, Ken66b, Ohl61, Sha63, Zwi53, Gin45, Les67, Cla58, Gil51, Gol60, Sha63, Ohl61]. Weather [CAON70]. Wedervang [Cow62]. Wege [Bur37]. Wegmüller [Lin71b].

Wegner [Sza64b]. Weight [Sär70]. Weiss [Dav62]. well [Thi66a]. well-known [Thi66a]. Welldon [Nix62b]. Weltwirtschaft [Ano36t]. Werner [Sch68a]. Wert [Zah34]. Westdeutschland [Dep57]. Western [Per41]. Wetherill [Gei69]. whether [Gin37b, Gin37a]. Which [Pri36]. whose [Fré60]. Wichers [Gou61b]. Wiederkehr [Sae36]. Wien [Wey38]. Wilcoxon [Ben67a]. Wilfred [dSR58]. Wilhelm [Her44, Wey38, Fla54]. Wilk [Whi69]. Wilkinson [Haw67]. Wilks [Coc64, Whi66, Whi69]. will [Wey38, Win37]. Willcox [Ric64, Ano64j, Ric61, Stu41]. Willem [Ide55]. Willensfreiheit [Wey38, Win37]. William [Cam64, Cra70, Han69, Ric59b]. Williamson [Blo65]. Winkler [Wey38, BL38, Gou51b, Her44, LS38, Win39]. Winners [Wol70]. Wirkungen [Win38a]. wirtschaftliche [Fla65]. wirtschaftlichen [Win38a]. Wirtschafts [Gou62]. Wirtschafts- [Gou62]. Wirtschaftsbeobachtung [WW38]. Wirtschaftsforschung [Ver58]. Wirtschaftskrise [Büc36]. Wirtschaftskunde [Win57]. Wirtschaftsstatistik [Bic68a]. Wirtschaftsstatistiken [WW38]. Wishart [Irw56]. wishes [Ano37v, Ano36y]. Wissenschaft [dW52]. within [Hor68a]. without [Rao66]. Witting [Lin69]. Wold [Dur53, Bra69]. Wolfgang [Jac64]. women [Gin41, Liv36]. Woodbury [Hub42]. word [Zah33]. work [Ano35l, Dor36, Hub38c, Vos69]. Workers [Dum70, Jul34, Gna67, Gou56]. workforce [Dep64]. working [Har36]. World [Ano36t, Koc63, Nix38, Ric68, Sny34a, Zwi60, Ano33n, Ano33o, Ano33p, Ano33q, Ano33r, Ano33s, Ano34q, Ano34r, Ano34s, Ano34t, Ano34w, Ano34x, Ano34y, Ano34z, Ano35e, Ano35f, Ano35g, Ano35m, Ano35n, Ano35o, Ano35p, Ano35q, Ano36e, Ano36f, Ano36g, Ano36h, Ano36p, Ano36q, Ano36r, Ano36s, Ano37e, Ano37f, Ano37g, Ano37h, Ano37n, Ano37o, Ano37p, Ano37q, Ano38h, Ano38i, Ano38j, Ano38k, Ano38l, Ano38r, Ano38s, Ano38t, Ano38u, Ano38v, Ano39e, Ano39f, Ano39g, Ano39m, Ano39n, Ano39o, Ano40d, Ano40e, Ano40g, Ano40h, Ano41d, Ano41e, Ano41h, Ano41i, Ano42c, Ano42d, Ano42e, Ano42f, Ano43c, Ano43d, Ano43g, Ano43h, Ano44b, Ano44e, Ano45b, Ano45d, Ano46c, Ano46e, Hor69a, Zwi71]. Wunsch [Fed69, Pre68].

X [Fla65]. XIV [Ver61]. XXI [Ken69]. XXIe [Stu34b]. XXII [Dre36, Dre36]. XXIIe [Stu34a]. XXIII [Stu36]. XXIIIe [Ano36y]. XXIIIème [Stu36]. XXIIIrd [Ano36y]. XXIInd [Stu34a]. XXIst [Stu34b]. XXIVe [Hub38b]. XXIVth [Hub38b]. XXVIth [Ano49g].

Yates [Gou63, Lin71a]. Ye [Wer67]. Year [Hub38c, Pto34]. Years [Cox60, Rao63b, Win44, Woo68]. Yeomans [Gou70b]. Yield [Med65]. youth [Har36]. Yu [Mal69, Ple67]. Yule [Ano52f].

Zackrisson [Dem66]. Zahlen [Ziz33]. Zahlungsbilanz [Ano34j]. Zahn [S.34, Mee40]. Zarkovich [Dem68a]. Zealand [KC70]. Zelnik [Sch66]. Zensus [Joh54]. zero [Fré48]. Ziauddin [Hus68]. Ziegler [Nor67]. Zizek [Sae36].

Zugleich [Win42]. zum [Pét66]. zur [Böh38, Mee35, Sae36, Win42, Pla34, Pét66]. Zürich [Eic65]. Zuständen [dS71c]. Zuwaylif [Bri71]. Zwet [Wal66]. Zwiedineck [And57]. Zwiedineck-Südenhorst [And57].

References

Arnold:1966:BRS

[AB66] B. C. Arnold and T. A. Bancroft. Book review: The Statistical Analysis of Series of Events by D. R. Cox, P. A. W. Lewis. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(3):448, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401265.

Abruzzi:1954:OAC

[Abr54] Adam Abruzzi. The \overline{asn}conceptinattributesdoublesamplingplans.Revuedel′Institutinternationaldestatistique = ReviewoftheInternationalStatisticalInstitute, 22 (1/3) : 48 − −56, 1954. CODENIST RDP. ISSN0373 − 1138(print), 2212−1846(electronic).URLhttpswwwjstororgstable1401919.

Allen:1963:PIN

[AFR^{+}63] R. G. D. Allen, G. M. W. Fürst, R. Roy, P. J. Loftus, V. Castellano, B. Barberi, and S. H. Khamis. Price index numbers. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(3):281–306, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401443.

Austwick:1971:FCS

[AHW^{+}71] K. Austwick, Joan Hine, G. B. Wetherill, F. Mosteller, B. Wetherill, T. Dalenius, J. S. Hunter, and W. Dixon. The first course in statistical methods and the use of teaching aids. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(3):287–306, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1402263.

Aigner:1970:EMP

[Aig70] Dennis J. Aigner. The estimation of moments for a Pareto distribution subject to both sampling and grouping errors. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):210–219, 1970. CODEN ISTRDP.

ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402144.

Aitken:1955:GHT

[Ait55] A. C. Aitken. Godfrey Hilton Thomson, 1881–1955. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 23(1/3):143, 1955. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1402023.

Allen:1950:BRS

[All50] R. G. D. Allen. Book review: Statistical Inference in Dynamic Economic Models by T. C. Koopmans. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18(3/4):260, 1950. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401040.

Allen:1955:BRL

[All55a] R. G. D. Allen. Book review: Linear Aggregation of Economic Relations by H. Theil. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 23(1/3):126– 127, 1955. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1402015.

Allen:1955:BRA

[All55b] R. G. D. Allen. Book review: The Analysis of Family Budgets by S. J. Prais, H. S. Houthakker. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 23(1/3):126, 1955. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402014.

Allen:1957:ALB

[All57] R. G. D. Allen. Arthur Lyon Bowley, 1869–1957. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):181–182, 1957. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401690.

Alexander:1970:HRS

[AMO^{+}70] G. N. Alexander, G. Morlat, P. E. O’Connell, Terence O’Donnell, R. Sneyers, P. J. Delaporte, R. D. Elston, and L. E. Borgman.

Heavy rainfall as a stochastic process. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):62–81 + 97–104, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1402325.

Amundsen:1968:BRE

[Amu68] A. Amundsen. Book review: Econometric Techniques and Problems by C. E. V. Leser. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):358– 359, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401487.

Anderson:1944:BRL

[And44] O. Anderson. Book review: Leçons de Statistique Mathématique. Premier cahier: Exposé préliminaire de Calcul des probabilités by Maurice Fréchet. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 12(1/4):76–77, 1944. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401073.

Anderson:1952:BTI

[And52] Oskar Anderson, Jr. The business test of the IFO-Institute for Economic Research, Munich, and its theoretical model. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 20(1):1–17, 1952. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401826.

Anderson:1957:OZS

[And57] O. Anderson. Otto von Zwiedineck-Südenhorst, 1871–1957. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):182–184, 1957. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401691.

Andersen:1968:BRM

[And68] E. B. Andersen. Book review: Modern Factor Analysis by H. H. Harman. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):355–357, 1968. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401485.

Anonymous:1933:ASI

[Ano33a] Anonymous. Atlante statistico italiano. (Italian) [Italian statistical atlas]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(2):79, July 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401176.

Anonymous:1933:AMFa

[Ano33b] Anonymous. Avis aux membres. (French) [Notice to members]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(1):54–55, 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401084.

Anonymous:1933:AMFb

[Ano33c] Anonymous. Avis aux membres. (French) [Notice to members]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(2):107, July 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401180.

Anonymous:1933:AMFc

[Ano33d] Anonymous. Avis aux membres. (French) [Notice to members]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(3):119–120, October 1933. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400783.

Anonymous:1933:BSIa

[Ano33e] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(1):68–77 + 79–80, 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401086.

Anonymous:1933:BSIb

[Ano33f] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(2):64–78 + 80–81, July 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401175.

Anonymous:1933:BSIc

[Ano33g] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(3):72–92, October 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1400779.

Anonymous:1933:CML

[Ano33h] Anonymous. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(3):66–71, October 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400778.

Anonymous:1933:DVV

[Ano33i] Anonymous. Das deutsche Volkseinkommen vor und nach dem Kriege. (German) [German national income before and after the war]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(1):78, 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401087.

Anonymous:1933:FMa

[Ano33j] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(1): 1–3, 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401078.

Anonymous:1933:FMb

[Ano33k] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1 (2):1–3, July 1933. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401167.

Anonymous:1933:FMc

[Ano33l] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(3): 1–3, October 1933. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400771.

Anonymous:1933:INW

[Ano33m] Anonymous. Index-numbers of wages: A survey. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(1):39–53, 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401083.

Anonymous:1933:PSSa

[Ano33n] Anonymous. Principaux services statistiques des divers pays du monde. (French) [Main statistical services of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(1):81–95, 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401088.

Anonymous:1933:PSSb

[Ano33o] Anonymous. Principaux services statistiques des divers pays du monde. (French) [Main statistical services of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(2):82–83, July 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401177.

Anonymous:1933:PSSc

[Ano33p] Anonymous. Principaux services statistiques des divers pays du monde. (French) [Main statistical services of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(3):93, October 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400780.

Anonymous:1933:SSDa

[Ano33q] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(1):96–97, 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401089.

Anonymous:1933:SSDb

[Ano33r] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(2):84–88, July 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401178.

Anonymous:1933:SSDc

[Ano33s] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(3):94–100, October 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1400781.

Anonymous:1933:SMN

[Ano33t] Anonymous. Supplément mensuel no. 1, juin 1933. (French) [Monthly supplement no. 1, June 1933]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(1):1–8, 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401091.

Anonymous:1933:TSFa

[Ano33u] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(1):98–116, 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401090.

Anonymous:1933:TSFb

[Ano33v] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(2):89–106, July 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401179.

Anonymous:1933:TSFc

[Ano33w] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(3):101–118, October 1933. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400782.

Anonymous:1934:AMFa

[Ano34a] Anonymous. Avis aux membres. (French) [Notice to members]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(4):139, January 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401109.

Anonymous:1934:AMFb

[Ano34b] Anonymous. Avis aux membres. (French) [Notice to members]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(1):159, April 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401023.

Anonymous:1934:BSIa

[Ano34c] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(4):72–95 + 98–99, January 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401102. See errata [Ano34k].

Anonymous:1934:BSIb

[Ano34d] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(1):106–128 + 131–132, April 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401018.

Anonymous:1934:BSIc

[Ano34e] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(2):208–231 + 236–237, July 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400999. See errata [Ano34l].

Anonymous:1934:BSId

[Ano34f] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(3):308–333, October 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1400989.

Anonymous:1934:CMLa

[Ano34g] Anonymous. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(4):66–71, January 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401101.

Anonymous:1934:CMLb

[Ano34h] Anonymous. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(2):199–207, July 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1400998.

Anonymous:1934:CMLc

[Ano34i] Anonymous. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(3):301–307, October 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400988.

Anonymous:1934:DZN

[Ano34j] Anonymous. Die deutsche Zahlungsbilanz nach Ländern. (German) [The German balance of payments by country]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(1):129–130, April 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401019.

Anonymous:1934:EBSa

[Ano34k] Anonymous. Errata: Bibliographie statistique internationale. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(4):95, January 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401103. See [Ano34c].

Anonymous:1934:EBSb

[Ano34l] Anonymous. Errata: Bibliographie statistique internationale. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(2):231, July 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401000. See [Ano34c].

Anonymous:1934:FMa

[Ano34m] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(4): 1–3, January 1934. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401094.

Anonymous:1934:FMb

[Ano34n] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2 (1):1–3, April 1934. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401010.

Anonymous:1934:FMc

[Ano34o] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2 (2):??, July 1934. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400994.

Anonymous:1934:FMd

[Ano34p] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(3): ??, October 1934. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400984.

Anonymous:1934:PSSa

[Ano34q] Anonymous. Principaux services statistiques des divers pays du monde. (French) [Main statistical services of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(4):100, January 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401106.

Anonymous:1934:PSSb

[Ano34r] Anonymous. Principaux services statistiques des divers pays du monde. (French) [Main statistical services of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(1):133, April 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401020.

Anonymous:1934:PSSc

[Ano34s] Anonymous. Principaux services statistiques des divers pays du monde. (French) [Main statistical services of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(2):238, July 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401006.

Anonymous:1934:PSSd

[Ano34t] Anonymous. Principaux services statistiques des divers pays du monde. (French) [Main statistical services of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(3):334–335, October 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400990.

Anonymous:1934:RIS

[Ano34u] Anonymous. Recueil international de statistiques Économiques, 1919–1930. (French) [International digest of economic statistics, 1919–1930]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(2):235, July 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401004.

Anonymous:1934:RII

[Ano34v] Anonymous. Répertoire international des institutions statistiques. (French) [International directory of statistical institutions]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(2):235, July 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401005.

Anonymous:1934:SSDa

[Ano34w] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(4):100–121, January 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401107.

Anonymous:1934:SSDb

[Ano34x] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(1):134–140, April 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401021.

Anonymous:1934:SSDc

[Ano34y] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(2):239–245, July 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401007.

Anonymous:1934:SSDd

[Ano34z] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(3):336–341, October 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400991.

Anonymous:1934:TS

[Ano34-27] Anonymous. Tableau statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1 (4):122–138, January 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401108.

Anonymous:1934:TSFa

[Ano34-28] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(1):141–158, April 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401022.

Anonymous:1934:TSFb

[Ano34-29] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(2):246–262, July 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401008.

Anonymous:1934:TSFc

[Ano34-30] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(3):342–358, October 1934. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400992.

Anonymous:1934:VI

[Ano34-31] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1 (4):??, January 1934. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401093.

Anonymous:1935:BSIa

[Ano35a] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(4):409–436 + 439–448, January 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400821.

Anonymous:1935:BSIb

[Ano35b] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(1):72–100, April 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401050.

Anonymous:1935:BSId

[Ano35c] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(2):194–222, July 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400856.

Anonymous:1935:BSIf

[Ano35d] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(3):311–335, October 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1400797.

Anonymous:1935:BSIc

[Ano35e] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(1):101, April 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401051.

Anonymous:1935:BSIe

[Ano35f] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(2):223, July 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400857.

Anonymous:1935:BSIg

[Ano35g] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(3):336, October 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400798.

Anonymous:1935:CML

[Ano35h] Anonymous. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(1):60–71, April 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401049.

Anonymous:1935:FMa

[Ano35i] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(4): i–iv, January 1935. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400816.

Anonymous:1935:FMb

[Ano35j] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3 (1):1–4, April 1935. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401045.

Anonymous:1935:FMc

[Ano35k] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(3): ??, October 1935. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400793.

Anonymous:1935:SIS

[Ano35l] Anonymous. La standardisation internationale des statistiques du travail. Aperçu de l’oeuvre du Bureau international du Travail et des différentes conférences internationales de statisticiens. (French)

[International standardization of labor statistics. Overview of the work of the International Labor Office and the various international conferences of statisticians]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(4): 437, January 1935. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400822.

Anonymous:1935:PSS

[Ano35m] Anonymous. Principaux services statistiques des divers pays du monde. (French) [Main statistical services of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(4):449, January 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400824.

Anonymous:1935:SSDa

[Ano35n] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(4):450–461, January 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400825.

Anonymous:1935:SSDb

[Ano35o] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(1):102–109, April 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401052.

Anonymous:1935:SSDc

[Ano35p] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(2):224–231, July 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1400858.

Anonymous:1935:SSDd the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(3):337–344, October 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400799.

Anonymous:1935:TSFa

[Ano35r] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(4):462–478, January 1935. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400826.

Anonymous:1935:TSFb

[Ano35s] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(1):110–126, April 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401053.

Anonymous:1935:TSFc

[Ano35t] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(2):232–248, July 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400859.

Anonymous:1935:TSFd

[Ano35u] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(3):345–362, October 1935. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400800.

Anonymous:1935:VIa

[Ano35v] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2 (4):??, January 1935. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400815.

Anonymous:1935:VIb

[Ano35w] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3 (1):??, April 1935. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401044.

Anonymous:1936:BSIa

[Ano36a] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(4):448–472 + 474–475, January 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401222.

Anonymous:1936:BSIc

[Ano36b] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(1):101–135, April 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/ stable/1401136. See errata [Ano36j].

Anonymous:1936:BSIe

[Ano36c] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(2):287–314, July 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400732.

Anonymous:1936:BSIg

[Ano36d] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(3):404–430, October 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1400808.

Anonymous:1936:BSIb

[Ano36e] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(4):476–477, January 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401224.

Anonymous:1936:BSId

[Ano36f] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(1):138, April 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401140.

Anonymous:1936:BSIf

[Ano36g] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(2):317, July 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400734.

Anonymous:1936:BSIh

[Ano36h] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(3):436, October 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400811.

Anonymous:1936:CML

[Ano36i] Anonymous. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(3):396–403, October 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400807.

Anonymous:1936:EBS

[Ano36j] Anonymous. Errata: Bibliographie statistique internationale. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(1):133, April 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401137. See [Ano36b].

Anonymous:1936:EQM

[Ano36k] Anonymous. Erratum: De quelques méthodes de determination de la covariation. (French) [Some methods for determining covariation]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(4):399, January 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401218. See [Jul35].

Anonymous:1936:FMa

[Ano36l] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(4): ??, January 1936. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401214.

Anonymous:1936:FMb

[Ano36m] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4 (1):??, April 1936. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401128.

Anonymous:1936:FMc

[Ano36n] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4 (2):??, July 1936. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400720.

Anonymous:1936:FMd

[Ano36o] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(3): ??, October 1936. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400802.

Anonymous:1936:SSDa

[Ano36p] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(4):478–489, January 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401225.

Anonymous:1936:SSDb

[Ano36q] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(1):139–148, April 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401141.

Anonymous:1936:SSDc

[Ano36r] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(2):318–326, July 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1400735.

Anonymous:1936:SSDd

[Ano36s] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(3):437–446, October 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400812.

Anonymous:1936:SHW

[Ano36t] Anonymous. Statistisches Handbuch der Weltwirtschaft. (German) [Statistical Handbook of the World Economy]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(1):136–137, April 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401139.

Anonymous:1936:TSFa

[Ano36u] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(4):490–506, January 1936. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401226.

Anonymous:1936:TSFb

[Ano36v] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(1):149–166, April 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401142.

Anonymous:1936:TSFc

[Ano36w] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(2):327–344, July 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400736.

Anonymous:1936:TSFd

[Ano36x] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(3):447–464, October 1936. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400813.

Anonymous:1936:VAA

[Ano36y] Anonymous. Voeux et avis adoptés par l’Institut International de Statistique au cours de sa XXIIIe session. (French) [Wishes and opinions adopted by the International Institute of Statistics during its XXIIIrd session]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4 (3):378–395, October 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400806.

Anonymous:1936:VI

[Ano36z] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4 (1):??, April 1936. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401127.

Anonymous:1937:BSIa

[Ano37a] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(4):546–575, January 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401061.

Anonymous:1937:BSIc

[Ano37b] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(1):48–75, April 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400924.

Anonymous:1937:BSIe

[Ano37c] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(2):147–175, July 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400832.

Anonymous:1937:BSIg

[Ano37d] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(3):240–271, October 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1400955.

Anonymous:1937:BSIb

[Ano37e] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(4):576, January 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401062.

Anonymous:1937:BSId

[Ano37f] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(1):76, April 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400925.

Anonymous:1937:BSIf

[Ano37g] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(2):176, July 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400833.

Anonymous:1937:BSIh

[Ano37h] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(3):274, October 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400957.

Anonymous:1937:CML

[Ano37i] Anonymous. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(1):37–47, April 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1400923.

Anonymous:1937:FMa

[Ano37j] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(4): ??, January 1937. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401055.

Anonymous:1937:FMb

[Ano37k] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5 (1):??, April 1937. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400919.

Anonymous:1937:FMc

[Ano37l] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5 (2):??, July 1937. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400828.

Anonymous:1937:FMd

[Ano37m] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(3): ??, October 1937. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400951.

Anonymous:1937:SSDa

[Ano37n] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(4):577–584, January 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401063.

Anonymous:1937:SSDb

[Ano37o] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(1):77–82, April 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400926.

Anonymous:1937:SSDc

[Ano37p] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(2):177–182, July 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1400834.

Anonymous:1937:SSDd the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(3):275–282, October 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400958.

Anonymous:1937:TSFa

[Ano37r] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(4):585–602, January 1937. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401064.

Anonymous:1937:TSFb

[Ano37s] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(1):83–100, April 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400927.

Anonymous:1937:TSFc

[Ano37t] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(2):183–200, July 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400835.

Anonymous:1937:TSFd

[Ano37u] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(3):283–300, October 1937. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400959.

Anonymous:1937:UPR

[Ano37v] Anonymous. Utilisation pratique des rapports et communications présentés et des voeux et avis adoptés. (French) [Practical use of the reports and communications presented and the wishes and opinions adopted]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(4):516–517, January 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401059.

Anonymous:1937:VI

[Ano37w] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5 (1):??, April 1937. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400918.

Anonymous:1938:AMFa

[Ano38a] Anonymous. Avis aux membres. (French) [Notice to members]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(4):433–434, January 1938. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400916.

Anonymous:1938:AMFb

[Ano38b] Anonymous. Avis aux membres. (French) [Notice to members]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(1):189–190, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400850.

Anonymous:1938:BSIa

[Ano38c] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(4):380–405, January 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1400911.

Anonymous:1938:BSIc

[Ano38d] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(1):133–160, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400845.

Anonymous:1938:BSIe

[Ano38e] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(2):321–349, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400749.

Anonymous:1938:BSIg

[Ano38f] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(3):463–487, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401197.

Anonymous:1938:BSIi

[Ano38g] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(4):572–601, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400888.

Anonymous:1938:BSIb

[Ano38h] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(4):407, January 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400913.

Anonymous:1938:BSId

[Ano38i] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(1):163, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1400847.

Anonymous:1938:BSIf

[Ano38j] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(2):350, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1400750.

Anonymous:1938:BSIh

[Ano38k] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(3):490, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401199.

Anonymous:1938:BSIj

[Ano38l] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(4):602, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1400889.

Anonymous:1938:CMLa

[Ano38m] Anonymous. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(1):129–132, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1400844.

Anonymous:1938:CMLb

[Ano38n] Anonymous. Communications sur la méthodologie la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(2):316–320, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1400748.

Anonymous:1938:CMLc

[Ano38o] Anonymous. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(4):566–571, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1400887.

Anonymous:1938:FMa

[Ano38p] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(4): ??, January 1938. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400903.

Anonymous:1938:FMb

[Ano38q] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(1): ??, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400838.

Anonymous:1938:SSDa

[Ano38r] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(4):408–414, January 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400914.

Anonymous:1938:SSDb

[Ano38s] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(1):164–170, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400848.

Anonymous:1938:SSDc

[Ano38t] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(2):351–357, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400751.

Anonymous:1938:SSDd

[Ano38u] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(3):491–499, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401200.

Anonymous:1938:SSDe

[Ano38v] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(4):603–608, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400890.

Anonymous:1938:TSFa

[Ano38w] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(4):415–432, January 1938. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400915.

Anonymous:1938:TSFb

[Ano38x] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(1):171–188, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400849.

Anonymous:1938:TSFc

[Ano38y] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(2):358–374, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400752.

Anonymous:1938:TSFd

[Ano38z] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(3):500–516, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401201.

Anonymous:1938:TSFe

[Ano38-27] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(4):609–626, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400891.

Anonymous:1938:VI

[Ano38-28] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6 (1):??, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400837.

Anonymous:1939:AMF

[Ano39a] Anonymous. Avis aux membres. (French) [Notice to members]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(1):111, 1939. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400938.

Anonymous:1939:BSIa

[Ano39b] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(1):47–80, 1939. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400934.

Anonymous:1939:BSIc

[Ano39c] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(2/3):171–195, 1939. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401029.

Anonymous:1939:BSIe

[Ano39d] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(4):274–301, 1939. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400966.

Anonymous:1939:BSIb

[Ano39e] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(1):81, 1939. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1400935.

Anonymous:1939:BSId

[Ano39f] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(2/3):196, 1939. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401030.

Anonymous:1939:BSIf

[Ano39g] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(4):302, 1939. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1400967.

Anonymous:1939:CMLa

[Ano39h] Anonymous. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(1):39–46, 1939. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400933.

Anonymous:1939:CMLb

[Ano39i] Anonymous. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(4):269–273, 1939. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1400965.

Anonymous:1939:FMa

[Ano39j] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(1): ??, 1939. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400930.

Anonymous:1939:FMb

[Ano39k] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(2/3):??, 1939. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401025.

Anonymous:1939:FMc

[Ano39l] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(4): ??, 1939. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400961.

Anonymous:1939:SSDa

[Ano39m] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(1):82–92, 1939. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400936.

Anonymous:1939:SSDb

[Ano39n] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(2/3):197–205, 1939. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401031.

Anonymous:1939:SSDc

[Ano39o] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(4):302–307, 1939.

CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400968.

Anonymous:1939:TSFa

[Ano39p] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(1):93–110, 1939. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400937.

Anonymous:1939:TSFb

[Ano39q] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(2/3):206–222, 1939. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401032.

Anonymous:1939:TSFc

[Ano39r] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(4):308–324, 1939. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400969.

Anonymous:1939:VI

[Ano39s] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7 (1):??, 1939. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400929.

Anonymous:1940:AMF

[Ano40a] Anonymous. Avis aux membres. (French) [Notice to members]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 8(3/4):244–245, 1940. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401165.

Anonymous:1940:BSIa

[Ano40b] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 8(1/2):50–97, 1940. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400715.

Anonymous:1940:BSIc

[Ano40c] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 8(3/4):199–234, 1940. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401160.

Anonymous:1940:BSIb

[Ano40d] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 8(1/2):98, 1940. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1400716.

Anonymous:1940:BSId

[Ano40e] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 8(3/4):238, 1940. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401162.

Anonymous:1940:CML

[Ano40f] Anonymous. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 8(3/4):195–198, 1940. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401159.

Anonymous:1940:SSDa

[Ano40g] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 8(1/2):99–103, 1940.

CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400717.

Anonymous:1940:SSDb

[Ano40h] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 8(3/4):239–240, 1940. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401163.

Anonymous:1940:TSFa

[Ano40i] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 8(1/2):104–109, 1940. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400718.

Anonymous:1940:TSFb

[Ano40j] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 8(3/4):241–243, 1940. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401164.

Anonymous:1940:VI

[Ano40k] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 8(1/2):??, 1940. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400711.

Anonymous:1941:AMF

[Ano41a] Anonymous. Avis aux membres. (French) [Notice to members]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 9(1/2):87, 1941. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400949.

Anonymous:1941:BSIa

[Ano41b] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 9(1/2):55–78, 1941. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400945.

Anonymous:1941:BSIc

[Ano41c] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 9(3/4):139–164, 1941. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400876.

Anonymous:1941:BSIb

[Ano41d] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 9(1/2):79, 1941. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1400946.

Anonymous:1941:BSId

[Ano41e] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 9(3/4):170, 1941. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400880.

Anonymous:1941:CMLa

[Ano41f] Anonymous. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 9(1/2):51–54, 1941. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1400944.

Anonymous:1941:CMLb

[Ano41g] Anonymous. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 9(3/4):133–138, 1941. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1400875.

Anonymous:1941:SSDa

[Ano41h] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 9(1/2):79–82, 1941. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400947.

Anonymous:1941:SSDb

[Ano41i] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 9(3/4):171–175, 1941. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400881.

Anonymous:1941:TSFa

[Ano41j] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 9(1/2):83–86, 1941. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400948.

Anonymous:1941:TSFb

[Ano41k] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 9(3/4):176–178, 1941. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400882.

Anonymous:1941:VI

[Ano41l] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 9(1/2):??, 1941. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400940.

Anonymous:1942:BSIa

[Ano42a] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 10(1/2):84–107, 1942. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401187.

Anonymous:1942:BSIc

[Ano42b] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 10(3/4):201–224, 1942. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400898.

Anonymous:1942:BSIb

[Ano42c] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 10(1/2):108, 1942. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401188.

Anonymous:1942:BSId

[Ano42d] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 10(3/4):226, 1942. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400900.

Anonymous:1942:SSDa

[Ano42e] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 10(1/2):108–113, 1942. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401189.

Anonymous:1942:SSDb

[Ano42f] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 10(3/4):226–230, 1942. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400901.

Anonymous:1942:TSF

[Ano42g] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 10(1/2):114–116, 1942. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401190.

Anonymous:1942:VIa

[Ano42h] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 10(1/2):??, 1942. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401182.

Anonymous:1942:VIb

[Ano42i] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 10(3/4):1–6, 1942. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400893.

Anonymous:1943:BSIa

[Ano43a] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 11(1/2):75–97, 1943. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400789.

Anonymous:1943:BSIc

[Ano43b] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 11(3/4):213–242, 1943. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401116.

Anonymous:1943:BSIb

[Ano43c] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 11(1/2):98, 1943. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400790.

Anonymous:1943:BSId

[Ano43d] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 11(3/4):245, 1943. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401118.

Anonymous:1943:CMLa

[Ano43e] Anonymous. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 11(1/2):68–74, 1943. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1400788.

Anonymous:1943:CMLb

[Ano43f] Anonymous. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 11(3/4):206–212, 1943. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401115.

Anonymous:1943:SSDa

[Ano43g] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 11(1/2):98–109, 1943. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400791.

Anonymous:1943:SSDb

[Ano43h] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 11(3/4):245–255, 1943. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401119.

Anonymous:1943:VI

[Ano43i] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 11(1/2):??, 1943. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400785.

Anonymous:1944:BSIa

[Ano44a] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 12(1/4):50–75, 1944. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401072.

Anonymous:1944:BSIb

[Ano44b] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 12(1/4):78, 1944. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401074.

Anonymous:1944:CML

[Ano44c] Anonymous. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 12(1/4):46–49, 1944. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401071.

Anonymous:1944:FM

[Ano44d] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 12(1/4):??, 1944. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401066.

Anonymous:1944:SSD

[Ano44e] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 12(1/4):78–81, 1944. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401075.

Anonymous:1944:TSF

[Ano44f] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 12(1/4):82–84, 1944. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401076.

Anonymous:1945:BSIa

[Ano45a] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 13(1/4):94–157, 1945. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400978.

Anonymous:1945:BSIb

[Ano45b] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 13(1/4):160–163, 1945. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400980.

Anonymous:1945:FM

[Ano45c] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 13 (1/4):1–3, 1945. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400971.

Anonymous:1945:SSD

[Ano45d] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 13(1/4):164–171, 1945. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400981.

Anonymous:1945:TSF

[Ano45e] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 13(1/4):172–177, 1945. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400982.

Anonymous:1946:AMF

[Ano46a] Anonymous. Avis aux membres. (French) [Notice to members]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 14(1/4):134–135, 1946. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401154.

Anonymous:1946:BSIa

[Ano46b] Anonymous. Bibliographie statistique internationale. (French) [International statistical bibliography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 14(1/4):73–118, 1946. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401148.

Anonymous:1946:BSIb

[Ano46c] Anonymous. Bureaux, services et institutions statistiques des divers pays du monde. (French) [Offices, services and statistical institutions of various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 14(1/4):123–124, 1946. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401151.

Anonymous:1946:FM

[Ano46d] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 14(1/4):??, 1946. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401144.

Anonymous:1946:SSD

[Ano46e] Anonymous. Sociétés de statistique dans les divers pays du monde. (French) [Statistical societies in various countries around the world]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 14(1/4):125–127, 1946. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401152.

Anonymous:1946:TSF

[Ano46f] Anonymous. Tableaux statistiques. (French) [Statistical tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 14(1/4):128–133, 1946. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401153.

Anonymous:1947:BSI

[Ano47a] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 15 (1/4):63–126, 1947. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401208.

Anonymous:1947:COA

[Ano47b] Anonymous. Communications sur l’organisation et l’administration statistiques / communications on statistical organisation and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 15(1/4):43–62, 1947. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401207.

Anonymous:1947:FM

[Ano47c] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 15(1/4):??, 1947. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401203.

Anonymous:1947:NSS

[Ano47d] Anonymous. Nouvelles statistiques / statistical news. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 15(1/4):135–136, 1947. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401211.

Anonymous:1947:SSS

[Ano47e] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 15(1/4):130–134, 1947. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401210.

Anonymous:1948:AMA

[Ano48a] Anonymous. Avis aux membres / announcements to members. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 16(1/4):140–143, 1948. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401237.

Anonymous:1948:BSI

[Ano48b] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 16 (1/4):65–129, 1948. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401233.

Anonymous:1948:COA

[Ano48c] Anonymous. Communications sur l’organisation et l’administration statistiques / communications on statistical organisation and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 16(1/4):45–64, 1948. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401232.

Anonymous:1948:FM 16(1/4):??, 1948. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401228.

Anonymous:1948:NSS

[Ano48e] Anonymous. Nouvelles statistiques / statistical news. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 16(1/4):138–139, 1948. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401236.

Anonymous:1948:SSS

[Ano48f] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 16(1/4):131–137, 1948. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401235.

Anonymous:1949:AMA

[Ano49a] Anonymous. Avis aux membres / announcements to members. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 17(3/4):243 + 251–270, 1949. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400869.

Anonymous:1949:BSIa

[Ano49b] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 17 (1/2):53–110, 1949. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401125.

Anonymous:1949:BSIb

[Ano49c] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 17 (3/4):175–231, 1949. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400866.

Anonymous:1949:COAa

[Ano49d] Anonymous. Communications sur l’organisation et l’administration statistiques / communications on statistical organisation and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 17(1/2):42–52, 1949. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401124.

Anonymous:1949:COAb

[Ano49e] Anonymous. Communications sur l’organisation et l’administration statistiques / communications on statistical organisation and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 17(3/4):156–174, 1949. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1400865.

Anonymous:1949:FM

[Ano49f] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 17 (3/4):244–250, 1949. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400861.

Anonymous:1949:FSE

[Ano49g] Anonymous. Furtherance of statistical education: Report to the XXVIth session of the international statistical institute. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 17(1/2):1–34, 1949. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401122.

Anonymous:1949:NSS

[Ano49h] Anonymous. Nouvelles statistiques / statistical news. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 17(3/4):241–242, 1949. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400868.

Anonymous:1949:SSS

[Ano49i] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 17(3/4):232–240, 1949. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400867.

Anonymous:1949:VI

[Ano49j] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 17(1/2):??, 1949. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401121.

Anonymous:1950:BSIa

[Ano50a] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18 (1/2):93–129, 1950. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400766.

Anonymous:1950:BSIb

[Ano50b] Anonymous. Bibliographie statistique internationale. / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18 (3/4):212–259, 1950. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401039.

Anonymous:1950:COAa

[Ano50c] Anonymous. Communications sur l’organisation et l’administration statistiques / communications on statistical organisation and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18(1/2):78–92, 1950. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400765.

Anonymous:1950:COAb

[Ano50d] Anonymous. Communications sur l’organisation et l’administration statistiques / communications on statistical organisation and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18(3/4):201–211, 1950. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401038.

Anonymous:1950:FM

[Ano50e] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18(1/2):??, 1950. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400755.

Anonymous:1950:NSS

[Ano50f] Anonymous. Nouvelles statistiques / statistical news. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18(1/2):143–149, 1950. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400769.

Anonymous:1950:SSSa

[Ano50g] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18(1/2):132–142, 1950. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400768.

Anonymous:1950:SSSb

[Ano50h] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18(3/4):263–265, 1950. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401042.

Anonymous:1950:VI

[Ano50i] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18(1/2):??, 1950. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400754.

Anonymous:1951:AMA

[Ano51a] Anonymous. Avis aux membres / announcements to members. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(2):206, 1951. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401247.

Anonymous:1951:BSIa

[Ano51b] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(1):56–108, 1951. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401502.

Anonymous:1951:BSIb

[Ano51c] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(2):155–198, 1951. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401243.

Anonymous:1951:BSIc

[Ano51d] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(3):252–282, 1951. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401998.

Anonymous:1951:COAa

[Ano51e] Anonymous. Communications sur l’organisation et l’administration statistiques / communications on statistical organisation and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(1):41–55, 1951. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401501.

Anonymous:1951:COAb

[Ano51f] Anonymous. Communications sur l’organisation et l’administration statistiques / communications on statistical organisation and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(2):148–154, 1951. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401242.

Anonymous:1951:COAc

[Ano51g] Anonymous. Communications sur l’organisation et l’administration statistiques / communications on statistical organisation und administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(3):239–251, 1951. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401997.

Anonymous:1951:NSSa

[Ano51h] Anonymous. Nouvelles statistiques / statistical news. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(1):119, 1951. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401506.

Anonymous:1951:NSSb

[Ano51i] Anonymous. Nouvelles statistiques / statistical news. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(2):206, 1951. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401246.

Anonymous:1951:NSSc

[Ano51j] Anonymous. Nouvelles statistiques / statistical news. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(3):288–290, 1951. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1402002.

Anonymous:1951:SSSa

[Ano51k] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(1):111–118, 1951. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401505.

Anonymous:1951:SSSb

[Ano51l] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(2):201–205, 1951. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401245.

Anonymous:1951:SSSc

[Ano51m] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(3):284–287, 1951. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402001.

Anonymous:1951:VI

[Ano51n] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19 (1):??, 1951. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401497.

Anonymous:1952:AMA

[Ano52a] Anonymous. Avis aux membres / announcements to members. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 20(1):89–90, 1952. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401835.

Anonymous:1952:BSIa

[Ano52b] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 20(1):51–81, 1952. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401831.

Anonymous:1952:BSIb

[Ano52c] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 20(2):162–194, 1952. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401820.

Anonymous:1952:COA

[Ano52d] Anonymous. Communications sur l’organisation et l’administration statistiques / communications on statistical organisation and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 20(1):44–50, 1952. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401830.

Anonymous:1952:FM

[Ano52e] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 20 (1):??, 1952. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401825.

Anonymous:1952:GUY

[Ano52f] Anonymous. George Udny Yule, 1871–1951. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 20(1):92–93, 1952. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401839.

Anonymous:1952:JV

[Ano52g] Anonymous. Jose A. Vandellós, 1899–1950. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 20(1):91, 1952. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401838.

Anonymous:1952:NSS

[Ano52h] Anonymous. Nouvelles statistiques / statistical news. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 20(1):87–88, 1952. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1401834.

Anonymous:1952:SSSa

[Ano52i] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 20(1):84–86, 1952. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401833.

Anonymous:1952:SSSb

[Ano52j] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 20(2):195–200, 1952. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401821.

Anonymous:1952:VI

[Ano52k] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 20 (1):??, 1952. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401824.

Anonymous:1953:AMAb

[Ano53a] Anonymous. Avis aux membres / announcement to members. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(3):207–209, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401437.

Anonymous:1953:AMAa

[Ano53b] Anonymous. Avis aux membres / announcements to members. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(1/2):91–117, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401525.

Anonymous:1953:AMAc

[Ano53c] Anonymous. Avis aux membres / announcements to members. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(3):223–240, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401440.

Anonymous:1953:BSIa

[Ano53d] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(1/2):36–80, 1953. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401517.

Anonymous:1953:BSIb

[Ano53e] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(3):160–195, 1953. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401427.

Anonymous:1953:COAa

[Ano53f] Anonymous. Communications sur l’organisation et l’administration statistiques / communications on statistical organisation and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(1/2):30–35, 1953.

CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401516.

Anonymous:1953:COAb

[Ano53g] Anonymous. Communications sur l’organisation et l’administration statistiques / communications on statistical organisation and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(3):152–159, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401426.

Anonymous:1953:FMa

[Ano53h] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(1/2):??, 1953. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401509.

Anonymous:1953:FMb

[Ano53i] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21 (3):212–222, 1953. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401422.

Anonymous:1953:NSSa

[Ano53j] Anonymous. Nouvelles statistiques / statistical news. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(1/2):89–90, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1401524.

Anonymous:1953:NSSb

[Ano53k] Anonymous. Nouvelles statistiques / statistical news. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(3):206, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401436.

Anonymous:1953:SSSa

[Ano53l] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(1/2):85–88, 1953. CODEN ISTRDP.

ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401523.

Anonymous:1953:SSSb

[Ano53m] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(3):204–205, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401435.

Anonymous:1953:VI

[Ano53n] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(1/2):??, 1953. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401508.

Anonymous:1954:AMA

[Ano54a] Anonymous. Avis aux membres / announcements to members. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22(1/3):192–193 + 196, 1954. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401940.

Anonymous:1954:BSI

[Ano54b] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22 (1/3):114–169, 1954. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401924.

Anonymous:1954:COA

[Ano54c] Anonymous. Communications sur l’organisation et l’administration statistiques / communications on statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22(1/3):95–113, 1954. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401923.

Anonymous:1954:FM 22(1/3):??, 1954. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401915.

Anonymous:1954:NSS

[Ano54e] Anonymous. Nouvelles statistiques / statistical news. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22(1/3):191, 1954. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1401939.

Anonymous:1954:SSS

[Ano54f] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22(1/3):182–190, 1954. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401938.

Anonymous:1955:AMA

[Ano55a] Anonymous. Avis aux membres / announcements to members. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 23(1/3):138–139 + 159–194, 1955. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402020.

Anonymous:1955:BSI

[Ano55b] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 23 (1/3):70–124, 1955. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402012.

Anonymous:1955:COA

[Ano55c] Anonymous. Communications sur l’organisation et l’administration statistiques / communications on statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 23(1/3):58–69, 1955. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402011.

Anonymous:1955:FM

[Ano55d] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 23 (1/3):146–158, 1955. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402004.

Anonymous:1955:ISS

[Ano55e] Anonymous. Informations statistiques / statistical news. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 23(1/3):137, 1955. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402019.

Anonymous:1955:SSS

[Ano55f] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 23(1/3):129–136, 1955. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402018.

Anonymous:1956:AMA

[Ano56a] Anonymous. Avis aux membres / announcement to members. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 24(1/3):152, 1956. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401290.

Anonymous:1956:BSI

[Ano56b] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 24 (1/3):90–140, 1956. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401283.

Anonymous:1956:COA

[Ano56c] Anonymous. Communications sur l’organisation et l’administration statistiques / communications on statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 24(1/3):77–89, 1956. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401282.

Anonymous:1956:FM 24(1/3):??, 1956. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401274.

Anonymous:1956:ISS

[Ano56e] Anonymous. Informations statistiques / statistical news. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 24(1/3):149–151, 1956. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401289.

Anonymous:1956:SSS

[Ano56f] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 24(1/3):144–148, 1956. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401288.

Anonymous:1956:VI

[Ano56g] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 24(1/3):1–19, 1956. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401273.

Anonymous:1957:AMA

[Ano57a] Anonymous. Avis aux membres / announcement to members. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):180, 1957. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401689.

Anonymous:1957:BSI

[Ano57b] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25 (1/3):74–150, 1957. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401676.

Anonymous:1957:COA

[Ano57c] Anonymous. Communications sur l’organisation et l’administration statistiques / communications on statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):56–73, 1957. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401675.

Anonymous:1957:FM

[Ano57d] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25 (1/3):192–205, 1957. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401669.

Anonymous:1957:ISS

[Ano57e] Anonymous. Informations statistiques / statistical news. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):178–179, 1957. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401688.

Anonymous:1957:RDB

[Ano57f] Anonymous. Rapport du bureau à la trentième session de l’Institut International de Statistique (Stockholm, 8–15 août 1957). (French) [Report of the bureau to the thirtieth session of the International Statistical Institute (Stockholm, August 8–15, 1957)]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):206–221, 1957. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401695.

Anonymous:1957:RBT

[Ano57g] Anonymous. Report of the bureau to the Thirtieth Session of the International Statistical Institute (Stockholm, 8–15 August, 1957). Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):222–236, 1957. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401696.

Anonymous:1957:SSS

[Ano57h] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):159–177, 1957. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401687.

Anonymous:1958:AMA

[Ano58a] Anonymous. Avis aux membres / announcement to members. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):167–168, 1958. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401592.

Anonymous:1958:BSI

[Ano58b] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26 (1/3):68–132, 1958. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401574.

Anonymous:1958:COA

[Ano58c] Anonymous. Communications sur l’organisation et l’administration statistiques / communications on statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):56–67, 1958. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401573.

Anonymous:1958:FM

[Ano58d] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26 (1/3):1–9, 1958. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401566.

Anonymous:1958:ISS

[Ano58e] Anonymous. Informations statistiques / statistical news. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):163–166, 1958. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401591.

Anonymous:1958:SSS

[Ano58f] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):143–162, 1958. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401590.

Anonymous:1959:BSI

[Ano59a] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 27(1/3):38–70, 1959. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402057.

Anonymous:1959:COA

[Ano59b] Anonymous. Communications sur l’organisation et l’administration statistiques / communications on statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 27(1/3):26–37, 1959. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402056.

Anonymous:1959:FM

[Ano59c] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 27(1/3):??, 1959. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402053.

Anonymous:1959:ISS

[Ano59d] Anonymous. Informations statistiques / statistical news. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 27(1/3):93–95, 1959. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402063.

Anonymous:1959:SSS

[Ano59e] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 27(1/3):74–92, 1959. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402062.

Anonymous:1960:BSIa

[Ano60a] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28 (1/2):56–104, 1960. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401848.

Anonymous:1960:BSIb

[Ano60b] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(3):259–295, 1960. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401324.

Anonymous:1960:COAa

[Ano60c] Anonymous. Communications sur l’organisation et l’administration statistiques / communications on statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(1/2):33–55, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401847.

Anonymous:1960:COAb

[Ano60d] Anonymous. Communications sur l’organisation et l’administration statistiques / communications on statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(3):245–258, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401323.

Anonymous:1960:FMa

[Ano60e] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28 (1/2):126–139, 1960. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401843.

Anonymous:1960:FMb

[Ano60f] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28 (3):??, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401318.

Anonymous:1960:ISSa

[Ano60g] Anonymous. Informations statistiques / statistical news. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(1/2):111–116, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401852.

Anonymous:1960:ISSb

[Ano60h] Anonymous. Informations statistiques / statistical news. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(3):303–308, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401332.

Anonymous:1960:PPSa

[Ano60i] Anonymous. Programme provisoire de la 33me session de l’institut international de statistique, Paris 29 août–7 septembre 1961. (French) [Provisional program of the 33rd session of the International Statistical Institute, Paris August 29–September 7, 1961]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(3):300, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401329.

Anonymous:1960:PPSb

[Ano60j] Anonymous. Provisional programme of the 33nd session of the international statistical institute, Paris, 29 August to 7 September 1961. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(3):301, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401330.

Anonymous:1960:RDB

[Ano60k] Anonymous. Rapport du bureau à la trente-deuxième session de l’Institut International de Statistique (Tokio, 30 mai–9 juin 1960). (French) [Report of the bureau to the thirty-second session of the International Statistical Institute (Tokyo, May 30–June 9, 1960)]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(1/2):140–160, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401858.

Anonymous:1960:RPS

[Ano60l] Anonymous. Rapport provisoire de la 32e session, Tokio 1960 / provisional report of the 32nd session, Tokyo 1960. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(3):208–227, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401320.

Anonymous:1960:RBT

[Ano60m] Anonymous. Report of the bureau to the Thirty-Second Session of the International Statistical Institute (Tokyo, 30 May to 9 June 1960). Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(1/2):161–180, 1960. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401859.

Anonymous:1960:SSSa

[Ano60n] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(1/2):107–110, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401851.

Anonymous:1960:SSSb

[Ano60o] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(3):302, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401331.

Anonymous:1960:VI

[Ano60p] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(1/2):1–21, 1960. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401842.

Anonymous:1961:ACO

[Ano61a] Anonymous. Autres communications sur l’organisation et l’administration statistiques / further communications on statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(1):81–110, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401554.

Anonymous:1961:BB

[Ano61b] Anonymous. Bibliographie / bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(3):69–87, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401951.

Anonymous:1961:BSIa

[Ano61c] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(1):111–146, 1961. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401555.

Anonymous:1961:BSIb

[Ano61d] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(2):57–81, 1961. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401658.

Anonymous:1961:COAa

[Ano61e] Anonymous. Communications sur l’organisation et l’administration statistiques / communications on statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(1):32–80, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401553.

Anonymous:1961:COAb

[Ano61f] Anonymous. Communications sur l’organisation et l’administration statistiques / communications on statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(2):42–56, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401657.

Anonymous:1961:COAc

[Ano61g] Anonymous. Communications sur l’organisation et l’administration statistiques / communications on statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(3):57–68, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401950.

Anonymous:1961:FMa

[Ano61h] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29 (1):??, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401550.

Anonymous:1961:FMb

[Ano61i] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29 (2):??, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401654.

Anonymous:1961:FMc

[Ano61j] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29 (3):100–102, 1961. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401945.

Anonymous:1961:ISSa

[Ano61k] Anonymous. Informations statistiques / statistical news. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(1):170–172, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401564.

Anonymous:1961:ISSb

[Ano61l] Anonymous. Informations statistiques / statistical news. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(2):91–93, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401665.

Anonymous:1961:ISSc

[Ano61m] Anonymous. Informations statistiques / statistical news. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(3):103–108, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401963.

Anonymous:1961:SSSa

[Ano61n] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(1):152–169, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401563.

Anonymous:1961:SSSb

[Ano61o] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(2):89–90, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401664.

Anonymous:1961:SSSc

[Ano61p] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(3):98–99, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401962.

Anonymous:1961:UNS

[Ano61q] Anonymous. United Nations Statistical Office publications recently issued. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(3):96–97, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401961.

Anonymous:1961:VI

[Ano61r] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29 (1):??, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401549.

Anonymous:1962:BBa

[Ano62a] Anonymous. Bibliographie / bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(2):248–270, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401900.

Anonymous:1962:BBb

[Ano62b] Anonymous. Bibliographie / bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(3):380–400, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401748.

Anonymous:1962:BSI

[Ano62c] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(1):60–83, 1962. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402073.

Anonymous:1962:CLA

[Ano62d] Anonymous. Communications sur l’Organisation et l’administration statistiques / communications on statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(1):42–59, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402072.

Anonymous:1962:FMa

[Ano62e] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30 (1):104–106, 1962. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402067.

Anonymous:1962:FMb

[Ano62f] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30 (2):245–247, 1962. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401892.

Anonymous:1962:FMc

[Ano62g] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30 (3):377–379, 1962. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401741.

Anonymous:1962:IRS

[Ano62h] Anonymous. Instituts de recherches statistiques / statistical research institutes. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(1):107–110, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1402083.

Anonymous:1962:LASa

[Ano62i] Anonymous. l’Organisation et l’administration statistiques / statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(2):160–236, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401897.

Anonymous:1962:LASb

[Ano62j] Anonymous. l’Organisation et l’administration statistiques / statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(3):367–376, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401746.

Anonymous:1962:PSR

[Ano62k] Anonymous. [Photograph]: Sir Ronald Aylmer Fisher, 1890–1962. Honorary President of the International Statistical Institute. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(2):??, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401893.

Anonymous:1962:PPSa

[Ano62l] Anonymous. Programme provisoire de la 34ème session de l’Institut international de statistique, Ottawa 21–29 août 1963. (French) [Provisional program of the 34th session of the International Statistical Institute, Ottawa, August 21–29, 1963]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(1):102, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402081.

Anonymous:1962:PPSb

[Ano62m] Anonymous. Provisional programme of the 34th session of the international statistical institute, Ottawa 21–29 August 1963. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(1):103, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402082.

Anonymous:1962:RSSa

[Ano62n] Anonymous. Recherches statistiques / statistical research. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(2):237–243, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401898.

Anonymous:1962:RSSb

[Ano62o] Anonymous. Recherches statistiques / statistical research. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(3):376, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401747.

Anonymous:1962:SSSa

[Ano62p] Anonymous. Societes de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(1):91–94, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402079.

Anonymous:1962:SSSb

[Ano62q] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(2):244, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401899.

Anonymous:1962:SLI

[Ano62r] Anonymous. Statuts de l’Association internationale pour la statistique dans les sciences physiques / statutes of the international association for statistics in physical sciences. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(1):95–101, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402080.

Anonymous:1962:UNSa

[Ano62s] Anonymous. United Nations Statistical Office publications recently issued. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(1):89–90, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402078.

Anonymous:1962:UNSb

[Ano62t] Anonymous. United Nations Statistical Office publications recently issued. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(3):407, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401757.

Anonymous:1962:VI

[Ano62u] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30 (1):??, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1402066.

Anonymous:1963:BBa

[Ano63a] Anonymous. Bibliographie / bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(2):235–259, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401378.

Anonymous:1963:BBb

[Ano63b] Anonymous. Bibliographie / bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(3):422–447, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401450.

Anonymous:1963:FMa

[Ano63c] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31 (2):225–228, 1963. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401369.

Anonymous:1963:FMb

[Ano63d] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31 (3):397–421, 1963. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401442.

Anonymous:1963:LASa

[Ano63e] Anonymous. l’Organisation et l’administration statistiques / statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(1):81–95, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401730.

Anonymous:1963:LASb

[Ano63f] Anonymous. l’Organisation et l’administration statistiques / statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(2):209–219, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401374.

Anonymous:1963:OAS

[Ano63g] Anonymous. Organisation et administration statistiques / statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(3):398–405, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401447.

Anonymous:1963:RDB

[Ano63h] Anonymous. Rapport du bureau à la trente-quatrième session de l’Institut International de Statistique. (French) [Report of the bureau to the thirty-fourth session of the International Statistical Institute]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(3):458–467, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401460.

Anonymous:1963:RSSa

[Ano63i] Anonymous. Recherches statistiques / statistical research. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(1):96–98, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401731.

Anonymous:1963:RSSb

[Ano63j] Anonymous. Recherches statistiques / statistical research. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(2):220–221, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401375.

Anonymous:1963:RSSc

[Ano63k] Anonymous. Recherches statistiques / statistical research. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(3):406–408, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401448.

Anonymous:1963:RBT

[Ano63l] Anonymous. Report of the bureau to the thirty-fourth session of the international statistical institute. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(3):468–477, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401461.

Anonymous:1963:SSSa

[Ano63m] Anonymous. Societes de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(1):99–113, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401732.

Anonymous:1963:SSSb

[Ano63n] Anonymous. Societes de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(2):222–224, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401376.

Anonymous:1963:SSSc

[Ano63o] Anonymous. Societes de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(3):409–417, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401449.

Anonymous:1963:TSM

[Ano63p] Anonymous. Table systematique des matières de la bibliographie statistique internationale / classification scheme of the international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(1):114–139, 1963. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401733.

Anonymous:1963:UNP

[Ano63q] Anonymous. United Nations Population Branch, Bureau of Social Affairs, publications recently issued. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(2):266, 1963. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401387.

Anonymous:1963:UNSa

[Ano63r] Anonymous. United Nations Statistical Office publications recently issued. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(1):144–145, 1963. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401739.

Anonymous:1963:UNSb

[Ano63s] Anonymous. United Nations Statistical Office publications recently issued. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(2):265–266, 1963. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401386.

Anonymous:1963:UNSc

[Ano63t] Anonymous. United Nations Statistical Office publications recently issued. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(3):452–453, 1963. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401456.

Anonymous:1963:VI

[Ano63u] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31 (1):??, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401723.

Anonymous:1964:SII

[Ano64a] Anonymous. 35ième session de l’Institut International de Statistique. (French) [35th session of the International Institute of Statistics]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(1/2):130, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401980.

Anonymous:1964:SIS

[Ano64b] Anonymous. 35th session of the international statistical institute. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(1/2):131, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401981.

Anonymous:1964:BB

[Ano64c] Anonymous. Bibliographie / bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):281–323, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401874.

Anonymous:1964:BSI

[Ano64d] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32 (1/2):132–180, 1964. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401982.

Anonymous:1964:FMa

[Ano64e] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32 (1/2):127–129, 1964. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401968.

Anonymous:1964:FMb

[Ano64f] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32 (3):278–280, 1964. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401862.

Anonymous:1964:IBR

[Ano64g] Anonymous. Index of book reviews. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):333–339, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401884.

Anonymous:1964:OASa

[Ano64h] Anonymous. Organisation et administration statistiques / statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(1/2):104–119, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401977.

Anonymous:1964:OASb

[Ano64i] Anonymous. Organisation et administration statistiques / statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):252–267, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401871.

Anonymous:1964:PWF

[Ano64j] Anonymous. [Photograph]: Walter F. Willcox, 1861–1964. Honorary President of the International Statistical Institute. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):??, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401863.

Anonymous:1964:RSSa

[Ano64k] Anonymous. Recherches statistiques / statistical research. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(1/2):120–124, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401978.

Anonymous:1964:RSSb

[Ano64l] Anonymous. Recherches statistiques / statistical research. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):268–274, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401872.

Anonymous:1964:SSSa

[Ano64m] Anonymous. Societes de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(1/2):125–126, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401979.

Anonymous:1964:SSSb

[Ano64n] Anonymous. Societes de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):275–277, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401873.

Anonymous:1964:UNSa

[Ano64o] Anonymous. United Nations Statistical Office publications recently issued. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(1/2):186–187, 1964. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401990.

Anonymous:1964:UNSb

[Ano64p] Anonymous. United Nations Statistical Office publications recently issued. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):330–332, 1964. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401883.

Anonymous:1964:VIa

[Ano64q] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(1/2):??, 1964. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401967.

Anonymous:1964:VIb

[Ano64r] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):1–16, 1964. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401861.

Anonymous:1965:BBa

[Ano65a] Anonymous. Bibliographie / bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(2):290–322, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402035.

Anonymous:1965:BBb

[Ano65b] Anonymous. Bibliographie / bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(3):521–554, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401707.

Anonymous:1965:BSI

[Ano65c] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(1):107–172, 1965. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401311.

Anonymous:1965:FMa

[Ano65d] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33 (1):103–106, 1965. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401300.

Anonymous:1965:FMb

[Ano65e] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33 (2):286–289, 1965. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402027.

Anonymous:1965:FMc

[Ano65f] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33 (3):518–520, 1965. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401699.

Anonymous:1965:OASa

[Ano65g] Anonymous. Organisation et administration statistiques / statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(1):87–95, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401308.

Anonymous:1965:OASb

[Ano65h] Anonymous. Organisation et administration statistiques / statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(2):270–279, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402032.

Anonymous:1965:OASc

[Ano65i] Anonymous. Organisation et administration statistiques / statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(3):443–512, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401704.

Anonymous:1965:RSSa

[Ano65j] Anonymous. Recherches statistiques / statistical research. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(1):96–100, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401309.

Anonymous:1965:RSSb

[Ano65k] Anonymous. Recherches statistiques / statistical research. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(2):280–283, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402033.

Anonymous:1965:RSSc

[Ano65l] Anonymous. Recherches statistiques / statistical research. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(3):513–515, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401705.

Anonymous:1965:SSSa

[Ano65m] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(1):101–102, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401310.

Anonymous:1965:SSSb

[Ano65n] Anonymous. Societes de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(2):284–285, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402034.

Anonymous:1965:SSSc

[Ano65o] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(3):516–517, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401706.

Anonymous:1965:UNS

[Ano65p] Anonymous. United Nations Statistical Office publications recently issued. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(3):565–567, 1965. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401720.

Anonymous:1965:VIa

[Ano65q] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33 (1):??, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401299.

Anonymous:1965:VIb

[Ano65r] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(3):1–20, 1965. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401698.

Anonymous:1966:SII

[Ano66a] Anonymous. 36ième session de l’Institut International de Statistique. (French) [36th session of the International Institute of Statistics]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):236, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401772.

Anonymous:1966:SIS

[Ano66b] Anonymous. 36th session of the international statistical institute. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):237, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401773.

Anonymous:1966:AMA

[Ano66c] Anonymous. Avis aux membres / announcement to members. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):219–232, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401771.

Anonymous:1966:BBa

[Ano66d] Anonymous. Bibliographie / bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):238–270, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401774.

Anonymous:1966:BBb

[Ano66e] Anonymous. Bibliographie / bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(3):411–445, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401261.

Anonymous:1966:BSI

[Ano66f] Anonymous. Bibliographie statistique internationale / international statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(1):87–110, 1966. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401537.

Anonymous:1966:FMa

[Ano66g] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34 (1):64–86, 1966. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401528.

Anonymous:1966:FMb

[Ano66h] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34 (2):233–235, 1966. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401760.

Anonymous:1966:FMc

[Ano66i] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34 (3):408–410, 1966. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401250.

Anonymous:1966:IBR

[Ano66j] Anonymous. Index of book reviews. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):286–293, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401789.

Anonymous:1966:OASa

[Ano66k] Anonymous. Organisation et administration statistiques / statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(1):65–74, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401534.

Anonymous:1966:OASb

[Ano66l] Anonymous. Organisation et administration statistiques / statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):195–212, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401768.

Anonymous:1966:OASc

[Ano66m] Anonymous. Organisation et administration statistiques / statistical organization and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(3):389–403, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401258.

Anonymous:1966:RSSa

[Ano66n] Anonymous. Recherches statistiques / statistical research. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(1):75–80, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401535.

Anonymous:1966:RSSb

[Ano66o] Anonymous. Recherches statistiques / statistical research. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):213–217, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401769.

Anonymous:1966:RSSc

[Ano66p] Anonymous. Recherches statistiques / statistical research. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(3):404–406, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401259.

Anonymous:1966:SSSa

[Ano66q] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(1):81–83, 1966. CODEN ISTRDP.

ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401536.

Anonymous:1966:SSSb

[Ano66r] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):218–219, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401770.

Anonymous:1966:SSSc

[Ano66s] Anonymous. Sociétés de statistique / statistical societies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(3):407, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401260.

Anonymous:1966:UNSa

[Ano66t] Anonymous. United Nations Statistical Office publications recently issued. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(1):116–117, 1966. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401547.

Anonymous:1966:UNSb

[Ano66u] Anonymous. United Nations Statistical Office publications recently issued. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):284–285, 1966. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401788.

Anonymous:1966:VIa

[Ano66v] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34 (1):??, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401527.

Anonymous:1966:VIb

[Ano66w] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(3):1–16, 1966. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401249.

Anonymous:1967:FMa

[Ano67a] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35 (1):87–90, 1967. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401631.

Anonymous:1967:FMb

[Ano67b] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35 (2):190–192, 1967. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401393.

Anonymous:1967:FMc

[Ano67c] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35 (3):319–320, 1967. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401793.

Anonymous:1967:IBR

[Ano67d] Anonymous. Index of book reviews. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(3):334–342, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401814.

Anonymous:1967:ISB

[Ano67e] Anonymous. International statistical bibliography / bibliographie statistique internationale. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35 (1):91, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401639.

Anonymous:1967:SOAa

[Ano67f] Anonymous. Statistical organization and administration / organisation et administration statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(1):59–71, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401636.

Anonymous:1967:SOAb

[Ano67g] Anonymous. Statistical organization and administration / organisation et administration statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):172–181, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401398.

Anonymous:1967:SOAc

[Ano67h] Anonymous. Statistical organization and administration / organisation et administration statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(3):296–308, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401798.

Anonymous:1967:SRR

[Ano67i] Anonymous. Statistical research / recherches statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(1):72–76, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401637.

Anonymous:1967:SSSa

[Ano67j] Anonymous. Statistical societies / societes de statistique. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(1):77–86, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401638.

Anonymous:1967:SSSb

[Ano67k] Anonymous. Statistical societies / societes de statistique. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):188–189, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401400.

Anonymous:1967:SSSc

[Ano67l] Anonymous. Statistical societies / societes de statistique. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(3):317–318, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401800.

Anonymous:1967:STRa

[Ano67m] Anonymous. Statistical training and research / formation et recherche statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):182– 187, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401399.

Anonymous:1967:STRb

[Ano67n] Anonymous. Statistical training and research / formation et recherche statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(3):309– 316, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401799.

Anonymous:1967:UNSa

[Ano67o] Anonymous. United Nations Statistical Office publications recently issued. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(1):102–103, 1967. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401651.

Anonymous:1967:UNSb

[Ano67p] Anonymous. United Nations Statistical Office publications recently issued. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):210, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401420.

Anonymous:1967:UNSc

[Ano67q] Anonymous. United Nations Statistical Office publications recently issued. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(3):331–333, 1967. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401813.

Anonymous:1967:VI

[Ano67r] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35 (1):??, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401630.

Anonymous:1968:SII

[Ano68a] Anonymous. 37ième session de l’Institut International de Statistique. (French) [37th session of the International Institute of Statistics]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):226, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401611.

Anonymous:1968:SIS

[Ano68b] Anonymous. 37th session of the international statistical institute. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):227, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401612.

Anonymous:1968:BRC

[Ano68c] Anonymous. Book review: Cours de calcul des probabilités by G. Calot. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):99, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401347.

Anonymous:1968:CRP

[Ano68d] Anonymous. [Correction]: Random processes and the growth of firms. A study of the Pareto law. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):113, 1968. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/

1401362. See [Wol67].

Anonymous:1968:FMa

[Ano68e] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36 (1):68–94, 1968. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401334.

Anonymous:1968:FMb

[Ano68f] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36 (2):228–230, 1968. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401600.

Anonymous:1968:FMc

[Ano68g] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36 (3):344–349, 1968. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401465.

Anonymous:1968:IBR

[Ano68h] Anonymous. Index of book reviews. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):364–371, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401494.

Anonymous:1968:SOAa

[Ano68i] Anonymous. Statistical organization and administration / organisation et administration statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):69–82, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401341.

Anonymous:1968:SOAb

[Ano68j] Anonymous. Statistical organization and administration / organisation et administration statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):207–217, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401608.

Anonymous:1968:SOAc

[Ano68k] Anonymous. Statistical organization and administration / organisation et administration statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):327–338, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401474.

Anonymous:1968:SSSa

[Ano68l] Anonymous. Statistical societies / societes de statistique. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):88–91, 1968. CODEN ISTRDP.

ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401343.

Anonymous:1968:SSSb

[Ano68m] Anonymous. Statistical societies / societes de statistique. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):224–225, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401610.

Anonymous:1968:SSSc

[Ano68n] Anonymous. Statistical societies / societes de statistique. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):345–346, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401476.

Anonymous:1968:STRa

[Ano68o] Anonymous. Statistical training and research / formation et recherche statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):83– 87, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401342.

Anonymous:1968:STRb

[Ano68p] Anonymous. Statistical training and research / formation et recherche statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):218– 223, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401609.

Anonymous:1968:STRc

[Ano68q] Anonymous. Statistical training and research / formation et recherche statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):339– 343, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401475.

Anonymous:1968:UNSa the International Statistical Institute, 36(1):116–117, 1968. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401366.

Anonymous:1968:UNSb

[Ano68s] Anonymous. United Nations Statistical Office publications recently issued. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):240, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401625.

Anonymous:1968:VI

[Ano68t] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36 (3):??, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401464.

Anonymous:1969:FMa

[Ano69a] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37 (1):103–105, 1969. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402089.

Anonymous:1969:FMb

[Ano69b] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37 (2):217–219, 1969. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402286.

Anonymous:1969:FMc

[Ano69c] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37 (3):318–320, 1969. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402115.

Anonymous:1969:RTD

[Ano69d] Anonymous. Round-table discussion on the university teaching of statistics in developing countries (The Hague, 26–29 August 1968). Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(1):86–98, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402098.

Anonymous:1969:SOAa

[Ano69e] Anonymous. Statistical organization and administration / organisation et administration statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(1):68–78, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402096.

Anonymous:1969:SOAb

[Ano69f] Anonymous. Statistical organization and administration / organisation et administration statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):195–207, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402292.

Anonymous:1969:SOAc

[Ano69g] Anonymous. Statistical organization and administration / organisation et administration statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(3):303–309, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402120.

Anonymous:1969:SSSb

[Ano69h] Anonymous. Statistical societies / sociétés de statistique. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):214–216, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402294.

Anonymous:1969:SSSc

[Ano69i] Anonymous. Statistical societies / sociétés de statistique. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(3):316–317, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402122.

Anonymous:1969:SSSa

[Ano69j] Anonymous. Statistical societies / societies de statistique. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(1):99–102, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402099.

Anonymous:1969:STRa

[Ano69k] Anonymous. Statistical training and research / formation et recherche statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(1):79– 85, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1402097.

Anonymous:1969:STRb

[Ano69l] Anonymous. Statistical training and research / formation et recherche statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):208– 213, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1402293.

Anonymous:1969:UNSa

[Ano69m] Anonymous. United Nations Statistical Office publications recently issued. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(1):115–117, 1969. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402112.

Anonymous:1969:UNSb

[Ano69n] Anonymous. United Nations Statistical Office publications recently issued. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):235–237, 1969. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402312.

Anonymous:1969:UNSc

[Ano69o] Anonymous. United Nations Statistical Office publications recently issued. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(3):330–331, 1969. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402134.

Anonymous:1969:VI

[Ano69p] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37 (3):??, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1402114.

Anonymous:1970:SII

[Ano70a] Anonymous. 38ième session de l’Institut International de Statistique. (French) [38th session of the International Institute of Statistics]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(3):412, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402207.

Anonymous:1970:SIS

[Ano70b] Anonymous. 38th session of the international statistical institute. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(3):413, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402208.

Anonymous:1970:CNS

[Ano70c] Anonymous. [Correction]: A note on single equation estimators in simultaneous equation systems. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(3):390, 1970. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/

1402203. See [Ram70].

Anonymous:1970:CNH

[Ano70d] Anonymous. [Correction]: The National Halothane Study. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):297, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402156. See [Arm70].

Anonymous:1970:FMa

[Ano70e] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38 (1):182–184, 1970. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402315.

Anonymous:1970:FMb

[Ano70f] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38 (2):268–293, 1970. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402141.

Anonymous:1970:FMc

[Ano70g] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38 (3):414–415, 1970. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402197.

Anonymous:1970:IBRa

[Ano70h] Anonymous. Index of book reviews. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):308–315, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402166.

Anonymous:1970:IBRb

[Ano70i] Anonymous. Index of book reviews. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(3):427–434, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402219.

Anonymous:1970:SOAa

[Ano70j] Anonymous. Statistical organization and administration / organisation et administration statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):171–176, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402332.

Anonymous:1970:SOAb

[Ano70k] Anonymous. Statistical organization and administration / organisation et administration statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):269–281, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402150.

Anonymous:1970:SOAc

[Ano70l] Anonymous. Statistical organization and administration / organisation et administration statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(3):391–401, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402204.

Anonymous:1970:SSSa

[Ano70m] Anonymous. Statistical societies / sociétés de statistique. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):180–181, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402334.

Anonymous:1970:SSSb

[Ano70n] Anonymous. Statistical societies / sociétés de statistique. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):288–290, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402152.

Anonymous:1970:SSSc

[Ano70o] Anonymous. Statistical societies / sociétés de statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(3):409–411, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402206.

Anonymous:1970:UNSa

[Ano70p] Anonymous. United Nations Statistical Office publications recently issued. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):189, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402341.

Anonymous:1970:UNSb

[Ano70q] Anonymous. United Nations Statistical Office publications recently issued. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):307, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402165.

Anonymous:1970:UNSc the International Statistical Institute, 38(3):424–426, 1970. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402218.

Anonymous:1970:VI

[Ano70s] Anonymous. Volume information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38 (3):??, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1402196.

Anonymous:1971:CMM

[Ano71a] Anonymous. [Correction]: Mathematik für Mediziner und Biologen. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(3):406, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402280. See [Wis71].

Anonymous:1971:FMa

[Ano71b] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39 (1):102–104, 1971. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402222.

Anonymous:1971:FMb

[Ano71c] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39 (2):235–237, 1971. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402168.

Anonymous:1971:FMc

[Ano71d] Anonymous. Front matter. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39 (3):402–404, 1971. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402260.

Anonymous:1971:I

[Ano71e] Anonymous. [Introduction]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39 (3):253–256, 1971. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402261.

Anonymous:1971:PRI

[Ano71f] Anonymous. Publications recently issued for the Statistical Office of the United Nations. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):249, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1402193.

Anonymous:1971:RTR

[Ano71g] Anonymous. Round table recommendations. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(3):369–372, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402274.

Anonymous:1971:SOAa

[Ano71h] Anonymous. Statistical organization and administration / organisation et administration statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):74–85, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402232.

Anonymous:1971:SOAb

[Ano71i] Anonymous. Statistical organization and administration / organisation et administration statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):224–229, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402175.

Anonymous:1971:SOAc

[Ano71j] Anonymous. Statistical organization and administration / organisation et administration statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(3):398, 1971. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402276.

Anonymous:1971:SSSa

[Ano71k] Anonymous. Statistical societies / sociétés de statistique. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):95–101, 1971. CODEN ISTRDP.

ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402234.

Anonymous:1971:SSSb

[Ano71l] Anonymous. Statistical societies / sociétés de statistique. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):232–234, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402177.

Anonymous:1971:SSSc

[Ano71m] Anonymous. Statistical societies / sociétés de statistique. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(3):401, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1402278.

Anonymous:1971:STR

[Ano71n] Anonymous. Statistical training and research / formation et recherche statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):230– 231, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1402176.

Anonymous:1971:UNS

[Ano71o] Anonymous. United Nations Statistical Office publications recently issued. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):119–120, 1971. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402254.

Armitage:1966:RDM

[Arm66] P. Armitage. Recent developments in medical statistics. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(1):27–42, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401531.

Armitage:1970:BRN

[Arm70] P. Armitage. Book review: The National Halothane Study by J. P. Bunker, W. H. Forrest, Jr., F. Mosteller, L. R. Vandam. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):185, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor. org/stable/1402335. See correction [Ano70d].

Atwood:1965:HH

[Atw65] A. W. Atwood. Herbert Hoover, 1874–1964. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(2):335–337, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402048.

Auerhan:1934:UAM

[Aue34] J. Auerhan. Utilisation de l’anthropométrie militaire. (French) [Use of military anthropometry]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2 (1):78–95, April 1934. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401016.

Azorin:1967:BRR

[Azo67] F. Azorín. Book review: Random Number Generators by B. Jansson. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(1):95–96, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401645.

Borgman:1970:SSP

[BAM^{+}70] L. E. Borgman, J. Amorocho, G. Morlat, P. E. O’Connell, Terence O’Donnell, R. Sneyers, P. J. Delaporte, and R. D. Elston. Some statistical problems in hydrology. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):82–104, 1970. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402326.

Barriol:1943:BRC

[Bar43] A. Barriol. Book review: Cours de statistique appliquée aux affaires by M. Huber. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 11(3/4):243– 244, 1943. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401117.

Barriol:1946:BRC

[Bar46] A. Barriol. Book review: Cours de Statistique appliquée aux affaires. III et IV: Statistiques économiques générales by M. Huber. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 14(1/4):119–121, 1946. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401149.

Barriol:1955:JG

[Bar55] A. Barriol. Joseph Girard, 1878–1955. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 23(1/3):142, 1955. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402022.

Barberi:1958:MS

[Bar58] B. Barberi. Mario Saibante, 1902–1958. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):171–172, 1958. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401596.

Bartlett:1961:BRE

[Bar61] M. S. Bartlett. Book review: Elements of the Theory of Markov Processes and Their Applications by A. T. Bharucha Reid. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(1):149, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401559.

Barberi:1962:AM

[Bar62] B. Barberi. Alessandro Molinari, 1898–1962. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(3):408, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401758.

Barberi:1965:BRL

[Bar65a] B. Barberi. Book review: L’entreprise et la statistique by R. Dumas. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(1):174, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401314.

Barra:1965:CLE

[Bar65b] J. R. Barra. Carrés latins et euleriens. (French) [Latin and Eulerian squares]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(1):16–23, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401304.

Barten:1966:BRE

[Bar66a] A. P. Barten. Book review: Econometric Theory by A. S. Goldberger. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(3):449–450, 1966. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401267.

Barton:1966:BRS

[Bar66b] D. E. Barton. Book review: Statistical Theory and Methodology in Science and Engineering by K. A. Brownlee. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(1):111, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401538.

Barberi:1967:BRT

[Bar67] B. Barberi. Book review: Théorie et pratique des sondages by J. Desabie. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):195–196, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401405.

Barlow:1968:BRS

[Bar68] R. E. Barlow. Book review: Slippage Tests by R. Doornbos. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):352–353, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401480.

Barberi:1969:BRT

[Bar69a] B. Barberi. Book review: Trattato di Statistica; I: La variabilitá by F. Brambilla. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):225–226, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402300.

Bartlett:1969:BRI

[Bar69b] M. S. Bartlett. Book review: Introduction to Stochastic Processes in Biostatistics by C. L. Chiang. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):226–227, 1969. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402301.

Bartoszynski:1971:NSC

[Bar71] Robert Bartoszyński. A note on subjective classifications. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):39–45, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402227.

Baxi:1971:NUS

[Bax71] H. R. Shiledar Baxi. A note on unit-stratified design. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):57–59, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402229.

Beale:1969:BRI

[Bea69] E. M. L. Beale. Book review: Introduction to Statistical Procedures by P. R. Lohnes, W. W. Cooley. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(3):324, 1969. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402126.

Benes:1964:DSA

[Ben64a] Jiří Beneš. La dynamique statistique dans l’automatisme de la production. (French) [Statistical dynamics in production automation]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):220–225, 1964. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401867.

Benjamin:1964:BRR

[Ben64b] B. Benjamin. Book review: Research in Family Planning by Clyde V. Kiser. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(1/2):183–185, 1964.

CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401987.

Benjamin:1965:BRM

[Ben65] B. Benjamin. Book review: Mortality in Sweden by T. Larsson. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(3):561, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401715.

Benzecri:1966:ASM

[Ben66] J. P. Benzécri. Analyse statistique et modèles probabilistes en psychologie. (French) [Statistical analysis and probabilistic models in psychology]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):139–155, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401763.

Bennett:1967:BGW

[Ben67a] B. M. Bennett. On the bivariate generalization of the Wilcoxon two sample test and its relation to rank correlation theory. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(1):30–33, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401633.

Benzecri:1967:BRL

[Ben67b] J. P. Benzécri. Book review: L’analyse hiérarchique by B. Matalon. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):205–206, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401416.

Billeter-Frey:1968:BRD

[BF68] E. P. Billeter-Frey. Book review: Data Processing by H. P. Hartkemeier. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):107, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401354.

Buckland:1969:STR

[BH69] W. R. Buckland and U. Hjorth. Statistical training and research / formation et recherche statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(3):310–315, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402121.

Bickel:1968:BRU

[Bic68a] W. Bickel. Book review: Umrisse einer Wirtschaftsstatistik by A. Blind. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):231, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401613.

Bicking:1968:SSC

[Bic68b] Ch. A. Bicking. Statistics in the selection and control of research and development projects. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36 (2):158–175, 1968. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401604.

Billeter:1956:ODM

[Bil56] E. P. Billeter. Optimum design in mixed sampling plans. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 24(1/3):73–76, 1956. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401281.

Billeter:1964:BRS

[Bil64] E. P. Billeter. Book review: The Statistical Method in Business by F. A. Ekeblad. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(1/2):182– 183, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401985.

Billeter:1965:BRM

[Bil65] E. Billeter. Book review: Méthodes et modèles de la recherche opérationnelle. Tome 2 by A. Kaufmann. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(2):324–325, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402038.

Billeter:1970:BRM

[Bil70] E. Billeter. Book review: Méthode et modèles de la recherche opérationnelle; tome 2 by A. Kaufmann. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):186, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402337.

Biraud:1959:BRM

[Bir59] Y. Biraud. Book review: Morbidity Statistics from General Practice. I. General by W. P. D. Logan, A. A. Cushion. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 27(1/3):72–73, 1959. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1402061.

Bithell:1971:SGM

[Bit71] J. F. Bithell. Some generalised Markov chain occupancy processes and their application to hospital admission systems. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):170–184, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402172.

Bjerke:1962:BRN

[Bje62] K. Bjerke. Book review: National Income and Expenditure by R. Stone, G. Stone. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(1):87–88, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402077.

Bjerke:1969:BRN

[Bje69] K. Bjerke. Book review: National Accounts Statistics: Sources and Methods by R. Maurice. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37 (3):324–325, 1969. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402127.

Bramness:1963:PAB

[BJT63] G. Bramness, D. B. Jochems, and H. Theil. On the presentation of aggregative business test results and the improvement of business test forecasts. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(1):58–66, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401728.

Bush:1965:CLE

[BKB65] K. A. Bush, M. G. Kendall, and J. R. Barra. [Carrès latins et euleriens]: Discussion. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(1):23, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401305.

Burgdorfer:1938:ORP

[BL38] Friedrich Burgdörfer and Livio Livi. Observations sur le rapport présenté par M. le Prof. Winkler: “La baisse de la natalité, ses causes, et ses consequences Économiques et sociales”. (French) [Observations on the report presented by Prof. Winkler: “The decline in the birth rate, its causes, and its economic and social consequences”]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(3):419–427, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401195.

Bliss:1949:PIC

[Bli49] C. I. Bliss. A proposal for an international computation center. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 17(1/2):35–41, 1949. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401123.

Blom:1965:BRT

[Blo65] G. Blom. Book review: Tables of the Negative Binomial Probability Distribution by E. Williamson, M. H. Bretherton. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(3):564–565, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401719.

Bernier:1970:IMP

[BMO^{+}70] J. Bernier, G. Morlat, P. E. O’Connell, Terence O’Donnell, R. Sneyers, P. J. Delaporte, R. D. Elston, and L. E. Borgman. Inventaire des modèles de processus stochastiques applicables à la description des debits journaliers des rivières. (French) [Inventory of stochastic process models applicable to the description of daily river flows]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):49–61 + 97– 104, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1402324.

Burdick:1969:RSM

[BN69] D. S. Burdick and T. H. Naylor. Response surface methods in economics. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(1):18–35, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402092.

Bohmert:1938:WKS

[Böh38] Wilhelm Böhmert. Was kann die Statistik zur Aufhellung des Krebsproblems tun? Bericht der Kommission für Krebsstatistik. (German) [What can statistics do to shed light on the problem of cancer? Report of the Commission on Cancer Statistics]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(4):335–346, January 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400908.

Boldrini:1937:PSC

[Bol37] Marcello Boldrini. La proportion des sexes dans les conceptions humaines. (French) [The proportion of the sexes in human conceptions]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(4):484–499, January 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401057.

Boldrini:1960:CAS

[Bol60a] M. Boldrini. The celebration of the 75th anniversary of the I.S.I. and its 32nd session, Tokyo: 1960. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(3):194–207, 1960. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401319.

Boldrini:1960:DPO

[Bol60b] M. Boldrini. Discours presidentiel à l’occasion de la 32ème session de l’I.I.S.. (French) [Presidential speech on the occasion of the 32nd session of the I.I.S.]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(1/2): 5–9, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401845.

Boldrini:1960:PAP

[Bol60c] M. Boldrini. The Presidential Address presented to the 32nd session of the I.S.I. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(1/2):1–4, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401844.

Boldrini:1962:RSI

[Bol62] M. Boldrini. Rapport sur la 33ème session de l’Institut International de Statistique. (French) [Report on the 33rd session of the International Statistical Institute]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(1):1–10, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402068.

Bolshev:1964:SAP

[Bol64] L. N. Bol’shev. Some applications of Pearson transformations. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(1/2):14–16, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401971.

Bombach:1971:ES

[Bom71] G. E. Bombach. Erich Schneider, 1900–1970. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):250–251, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402194.

Borenius:1966:BRS

[Bor66] G. Borenius. Book review: Statistical Adjustment of Data by W. E. Deming. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(1):112–113, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401541.

Bose:1966:EDE

[Bos66] R. C. Bose. Error detecting and error correcting indexing systems for large serial numbers. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(3):334– 340, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401253.

Bowley:1933:MIE

[Bow33] Arthur L. Bowley. Methods of investigating the economic and social conditions of a great city. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1 (2):7–18, July 1933. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401169.

Bowley:1946:CML

[Bow46] A. L. Bowley. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 14(1/4):52–72, 1946. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401147.

Bowley:1952:LLP

[Bow52] A. L. Bowley. Langford Lovell Price, 1862–1950. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 20(1):91, 1952. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401837.

Box:1961:BRI

[Box61] G. E. P. Box. Book review: An Introduction to the Theory of Experimental Design by D. J. Finney. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29 (3):89, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401953.

Box:1970:BRC

[Box70] G. E. P. Box. Book review: Cumulative Sum Tests: Theory and Practice by C. S. Van Dobben de Bruyn. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):305–306, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402164.

Boyarsky:1958:ETC

[Boy58] A. Y. Boyarsky. An experiment in the theory of a census with control rounds. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):48–55, 1958. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401572.

Bourgeois-Pichat:1950:AMI

[BP50] J. Bourgeois-Pichat. Analyse de la mortalité infantile. (French) [Analysis of infant mortality]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18 (1/2):45–68, 1950. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400758.

Bourgeois-Pichat:1951:LDP

[BP51] J. Bourgeois-Pichat. Les limites de la démographie potentielle. (French) [The limits of potential demography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(1):13–27, 1951. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401499.

Bourgeois-Pichat:1952:LIC

[BPH52] J. Bourgeois-Pichat and L. Hersch. [Les limites d’une importante contribution critique à la demographie potentielle]: Réponse. (French) [[The limits of an important critical contribution to potential demography]: Response]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 20(1):39–43, 1952. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401829.

Blom:1966:PSC

[BR66] G. Blom and L. Råde. Probability and statistics courses for students of engineering at the university level. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):165–173, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401765.

Brambilla:1967:BRQ

[Bra67] F. Brambilla. Book review: Quadratic Programming by J. C. G. Boot. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):193, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401401.

Brauers:1969:BRF

[Bra69] W. K. Brauers. Book review: Forecasting on a Scientific Basis by H. O. Wold. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(3):329–330, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402133.

Brillinger:1964:ABT

[Bri64] D. R. Brillinger. The asymptotic behaviour of Tukey’s general method of setting approximate confidence limits (the jackknife) when applied to maximum likelihood estimates. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):202–206, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401865.

Brillinger:1968:BRH

[Bri68] D. R. Brillinger. Book review: Handbook of Methods of Applied Statistics by I. M. Chakravarti, R. G. Laha, J. Roy. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):232–233, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1401615.

Brillinger:1971:BRG

[Bri71] David R. Brillinger. Book review: General Applied Statistics by F. H. Zuwaylif. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):249, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402192.

Bruckmann:1970:BRP

[Bru70] G. Bruckmann. Book review: Production and Productivity in the Service Industries by V. R. Fuchs. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):301–303, 1970. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402160.

Bhat:1971:STR

[BSH71] B. R. Bhat and Saw Swee-Hock. Statistical training and research / formation et recherche statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):86–94, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402233.

Birnbaum:1964:TSD

[BT64] Z. W. Birnbaum and Victor Kuang-Tao Tang. Two simple distribution-free tests of goodness of fit. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(1/2):2–13, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401970.

Buchner:1936:EWW

[Büc36] Oskar Büchner. Der Einfluss der Wirtschaftskrise auf die Wanderungsbewegung in den deutschen Städten. (German) [Influence of the economic crisis on the migration movement in German cities]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(1):1–26, April 1936. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401129.

Buckland:1969:BRC

[Buc69] W. R. Buckland. Book review: Classical and Contagious Discrete Distributions by G. P. Patil. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37 (1):110–111, 1969. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402106.

Buckland:1971:BRD

[Buc71] W. R. Buckland. Book review: Distributions in Statistics: Continuous Univariate Distributions by N. L. Johnson, S. Kotz. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(3):406–407, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402281.

Bunle:1933:CML

[Bun33] Henri Bunle. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(2):58–63, July 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401174.

Bunle:1937:CML

[Bun37] Henri Bunle. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(2):132–146, July 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1400831.

Bunle:1938:PRA

[Bun38a] Henri Bunle. La population rurale. Sur l’adoption d’une définition susceptible d’être internationalement adoptée. (French) [Rural population. On the adoption of a definition capable of being adopted internationally]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5 (4):347–357, January 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400909.

Bunle:1938:RCP

[Bun38b] M. Henri Bunle. Rapport de la commission pour la définition de la “population rurale”. (French) [Report of the commission for the definition of the “rural population”]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(2):229–234, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400740.

Bunle:1941:BRC

[Bun41] Henri Bunle. Book review: Cours de démographie et de statistique sanitaire by M. Huber. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 9 (3/4):166–167, 1941. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400878.

Bunle:1958:RR

[Bun58] H. Bunle. René Risser, 1869–1958. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26 (1/3):168–169, 1958. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401593.

Burgdorfer:1937:FFN

[Bur37] Friedrich Burgdörfer. Familienstatistik und Fruchtbarkeitsmessung. Neue Aufgaben und neue Wege der deutschen Bevölkerungsstatistik. (German) [Family statistics and fertility measurement. New tasks and new paths for German population statistics]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(3):212–226, October 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400953.

Brewer:1969:SPC

[BW69] K. R. W. Brewer and A. J. Whittington. The 1966 sample population census of the Territory of Papua and New Guinea. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):149–165, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402288.

Cacoullos:1971:BRS

[Cac71] T. Cacoullos. Book review: The Structure of Probability Theory with Applications by A. J. Thomasian. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):116–117, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402251.

Calitsunakis:1962:BRC

[Cal62a] D. E. Calitsunakis. Book review: Le coût de la vie à Paris de 1840 à 1954 by J. Singer-Kérel. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(2):273–274, 1962. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401904.

Calitsunakis:1962:BRM

[Cal62b] D. E. Calitsunakis. Book review: Le mode de vie des familles bourgeoises, 1873–1953 by M. Perrot. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(2):274–275, 1962. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401905.

Calitsunakis:1962:BRD

[Cal62c] D. E. Calitsunakis. Book review: Les dirigeants de l’industrie française by N. Delefortrie-Soubeyroux. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(2):275–276, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401907.

Calitsunakis:1962:BRL

[Cal62d] D. E. Calitsunakis. Book review: Les loyers des bourgeois de Paris, 1860–1958 by F. Marnata. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(2):275, 1962. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401906.

Callies:1966:UMM

[Cal66] J. M. Callies. Utilisation de modèles mathématiques pour l’estimation des donnèes demographiques dans les pays en voie de développement. (French) [Use of mathematical models for estimating demographic data in developing countries]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(3):341–359, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401254.

Campion:1964:LWH

[Cam64] Harry Campion. Lord William H. Beveridge, 1879–1963. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):346–347, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401886.

Cansado:1955:BRI

[Can55] E. Cansado. Book review: Introducción a los métodos de la estadística by S. Rios. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 23(1/3): 128, 1955. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1402017.

Cocheme:1970:PUW

[CAON70] Jacques Cocheme, G. N. Alexander, B. C. V. Oddie, and J. Neyman. Problems of the users of weather modification by cloud seeding. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):140–148, 1970. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402329.

Carner:1969:BRS

[Car69] J. B. Carner. Book review: The Structure of Inference by D. A. S. Fraser. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(3):322–323, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402125.

Castellano:1965:CG

[Cas65] V. Castellano. Corrado Gini, 1884–1965. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(2):337–344, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402049.

Cornish:1938:MCS

[CF38] E. A. Cornish and R. A. Fisher. Moments and cumulants in the specification of distributions. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5 (4):307–320, January 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400905.

Chandrasekaran:1964:KCK

[Cha64] C. Chandrasekaran. Kizhakke Covilagam Kutty Ettan Raja, 1893– 1963. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):351–352, 1964. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401889.

Chevry:1958:RR

[Che58] G. R. Chevry. Raymond Rivet, 1906–1958. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):172–173, 1958. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401597.

Chevry:1968:PD

[Che68] G. R. Chevry. Pierre Depoid, 1909–1968. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):241–242, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401626.

Clatworthy:1968:CTT

[CL68] W. H. Clatworthy and R. J. Lewyckyj. Comments on Takeuchi’s table of difference sets generating balanced incomplete block designs. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):12–18, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401337.

Clark:1958:BRC

[Cla58] C. Clark. Book review: Canada’s Economic Development 1867– 1953; With Special Reference to Changes in the Country’s National Product and National Wealth by O. J. Firestone. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):141–142, 1958. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401588.

Clark:1961:BRP

[Cla61] C. Clark. Book review: Patterns of Industrial Growth, 1938–58 / Aperçu de l’expansion industrielle, 1938–1958. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(1):149–150, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401560.

Clatworthy:1968:BRP

[Cla68a] W. H. Clatworthy. Book review: Principles of Statistics by M. G. Bulmer. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):350, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401478.

Clatworthy:1968:SBI

[Cla68b] W. H. Clatworthy. The subclass of balanced incomplete block designs with r = 11 replications. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):7–11, 1968. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401336.

Clatworthy:1971:BRM

[Cla71] Willard H. Clatworthy. Book review: Models in Regression and Related Topics by P. Sprent. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39 (3):409–410, 1971. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402284.

Cole:1970:TST

[CMK^{+}70] A. J. Cole, Matthias Matschinski, D. G. Kendall, F. Chayes, D. M. Hawkins, W. A. Read, E. L. Scott, E. H. Timothy Whitten, and Graham Lea. A trend surface technique. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):12–18 + 32–34, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1402319.

Cochran:1950:OAR

[Coc50] W. G. Cochran. [Observations au rapport du Prof. Cochran sur l’état actuel de la biométrie]: Reply. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18(1/2):72, 1950. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/

1400761. See [Gin50b].

Cochran:1964:SSW

[Coc64] W. G. Cochran. Samuel Stanley Wilks, 1906–1964. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(1/2):189–191, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401992.

Cohen:1966:BRQ

[Coh66] J. W. Cohen. Book review: Queues and Inventories, a Study of Their Basic Stochastic Processes by N. U. Prabhu. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):279, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401782.

Colombo:1964:BRM

[Col64] B. Colombo. Book review: Manuel de recherche démographique en pays sous-développé by R. Blanc. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(1/2):181, 1964. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401983.

Cornfield:1967:BT

[Cor67a] J. Cornfield. Bayes theorem. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(1):34–49, 1967. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401634.

Cornfield:1967:DF

[Cor67b] J. Cornfield. Discriminant functions. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):142–153, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401396.

Coward:1962:IW

[Cow62] D. Coward. Ingvar Wedervang, 1891–1961. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(2):283–286, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401913.

Cox:1960:RTF

[Cox60] Gertrude M. Cox. Report of the Treasurer for the fiscal years 1957, 1958 and 1959. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(1/2):181– 192, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401860.

Craig:1933:RCS

[Cra33] J. I. Craig. Report of the Commission for the Study of Statistics of Consumption and of Carryover of Commodities. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(3):47–58, October 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400776.

Cramer:1966:ASP

[Cra66] H. Cramér. The application of stochastic processes to certain reliability problems. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):119–124, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401761.

Cramer:1968:BRJ

[Cra68] H. Cramér. Book review: Joint Statistical Papers of J. Neyman and E. S. Pearson by J. Neyman, E. S. Pearson. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):235–236, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401619.

Cramer:1969:BRM

[Cra69] J. S. Cramer. Book review: Mathematical Model Building in Economics and Industry. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(1):109, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1402104.

Cramer:1970:WF

[Cra70] Harald Cramér. William Feller, 1906–1970. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(3):435–436, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402220.

Crowder:1969:PRF

[Cro69] E. T. Crowder. Public reporting to the Federal Government in the United States of America: Control, coordination, and measurement. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):176–185, 1969. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402290.

Csorgo:1970:PRC

[CS70] M. Csörgö and V. Seshadri. On the problem of replacing composite hypotheses by equivalent simple ones. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(3):351–368, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402201.

Curnov:1970:BRF

[Cur70] R. N. Curnov. Book review: A First Course in Biometry for Agriculture Students by A. A. Rayner. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):305, 1970. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402163.

Dagum:1964:BRM

[Dag64] C. Dagum. Book review: Monographie des Treillis et Algèbre de Boole. Collection de mathématique économiques by M. Carvallo. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):324–325, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401876.

Dagum:1967:BRP

[Dag67] C. Dagum. Book review: Principes et applications de l’analyse booléenne by M. Carvallo. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):194, 1967. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401403.

Dagum:1970:BRE

[Dag70] C. Dagum. Book review: Empirical Econometrics by J. S. Cramer. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(3):418, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402212.

Dagum:1971:BRM

[Dag71] Camilo Dagum. Book review: Méthodes statistiques de l’économétrie by E. Malinvaud. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):244–245, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1402187.

Dalenius:1958:BRM

[Dal58] T. E. Dalenius. Book review: Moderne Methoden in der Agrarstatistik by H. Strecker. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):133, 1958. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401575.

Dalenius:1962:BRI

[Dal62] T. E. Dalenius. Book review: An Introduction to Sampling Theory with Applications to Agriculture by M. R. Sampford. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(3):406–407, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401755.

Dalenius:1963:BRE

[Dal63] Tore Dalenius. Book review: Experiments in Statistical Sampling in the Indian Statistical Institute by P. C. Mahalanobis. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(2):264, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401385.

Dalenius:1965:BRS

[Dal65] T. E. Dalenius. Book review: Sampling Opinions. An Analysis of Survey Procedure by F. F. Stephan, Ph. J. McCarthy. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(1):175–176, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401316.

Dalenius:1968:BRS

[Dal68a] T. Dalenius. Book review: Statistical Methods by G. W. Snedecor, W. G. Cochran. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):361–362, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401492.

Dalenius:1968:OST

[Dal68b] T. E. Dalenius. Official statistics and their uses. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):121–140, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401601.

Dalenius:1969:BRR

[Dal69] T. Dalenius. Book review: Reducing the Cost of Surveys by S. Sudman. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(1):113, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402109.

Dalenius:1971:SSC

[Dal71] Tore Dalenius. Survey sampling in a computerized environment. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(3):373–397, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402275.

Darmois:1945:LDC

[Dar45] G. Darmois. Sur les limites de la dispersion de certaines estimations. (French) [On the limits of the dispersion of certain estimates]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 13(1/4):9–15, 1945. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400974.

Darmois:1953:AGL

[Dar53] G. Darmois. Analyse générale des liaisons stochastiques: étude particulière de l’analyse factorielle linéaire. (French) [General analysis of stochastic links: particular study of linear factor analysis]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(1/2):2–8, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401511.

Darmois:1955:VNS

[Dar55] G. Darmois. La vingt-neuvième session: Petropolis, 24 juin–2 juillet 1955. (French) [The Twenty-Ninth Session: Petropolis, June 24–July 2, 1955]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 23(1/3):1–5, 1955. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402005.

Darmois:1956:EB

[Dar56] G. Darmois. Émile Borel, 1871–1956. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 24(1/3):154–156, 1956. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401292.

Darmois:1958:BRP

[Dar58a] G. Darmois. Book review: Proceedings of the Third Berkeley Symposium on Mathematical Statistics and Probability. Vol. 3. Contributions to Astronomy and Physics by J. Neyman. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):134–135, 1958. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401577.

Darmois:1958:TUS

[Dar58b] G. Darmois. La trente et unième Session de l’Institut International de Statistique, Bruxelles, 3–8 septembre 1958. (French) [The thirty-first session of the International Statistical Institute, Brussels, September 3–8, 1958]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):1–4, 1958. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401567.

David:1962:BRS

[Dav62] F. N. David. Book review: Statistical Decision Theory by L. Weiss. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(3):407, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401756.

David:1967:BRS

[Dav67] H. A. David. Book review: The Selected Papers of E. S. Pearson by E. S. Pearson. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(3):330, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401812.

David:1969:BRS

[Dav69] H. A. David. Book review: Statistics for Mathematicians — an Introduction by D. J. Finney. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37 (1):106–107, 1969. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402101.

David:1970:BRI

[Dav70] F. N. David. Book review: Introduction à la combinatorique en vue des applications by A. Kaufmann. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):187, 1970. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402338.

David:1971:RPR

[Dav71] H. A. David. Ranking the players in a round Robin tournament. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):137–147, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402170.

Day:1949:ASM

[Day49] Besse B. Day. Application of statistical methods to research and development in engineering. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 17 (3/4):129–155, 1949. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400864.

Bojorquez:1934:EIE

[dDBL34] J. de D. Bojorquez and Gilberto Loyo. Sobre estadisticas internacionales de los efectos protestados y de los secuestros ejecutados teniendo como base titulos de credito. (Spanish) [On international statistics of the protested effects and kidnappings carried out based on credit titles]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(4):38–47, January 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401099.

Delaporte:1950:ESP

[Del50] Pierre Delaporte. Étude statistique sur les propriétés des fontes. (French) [studystatistical study on the properties of fonts]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18(3/4):161–178, 1950. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401035.

Deming:1950:SPM

[Dem50] W. Edward’s Deming. On the sampling of physical materials. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18(1/2):1–20, 1950. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400756.

Deming:1954:BRS

[Dem54] W. E. Deming. Book review: Sampling Theory of Surveys with Applications by P. V. Sukhatme. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22 (1/3):172–173, 1954. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401927.

Deming:1961:BRS

[Dem61] W. E. Deming. Book review: Studies of Selection Bias in Applied Statistics by G. Eklund. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(3):92– 94, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401958.

Dempster:1962:BRF

[Dem62] A. P. Dempster. Book review: The Foundations of Statistical Inference by L. J. Savage. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(3):403–404, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401752.

Deming:1963:BRS

[Dem63] W. E. Deming. Book review: Scientific Method: Optimizing Applied Research Decisions by R. L. Ackoff. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(3):448–449, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401451.

Deming:1966:BRQ

[Dem66] W. E. Deming. Book review: On Quality and Costs in Life Testing. Department of Statistics, University of Gothenburg, Sweden by E. Carlström, U. Zackrisson. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(1):112, 1966. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401540.

Deming:1968:BRS

[Dem68a] W. E. Deming. Book review: Sampling Methods and Censuses by S. S. Zarkovich. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):114–115, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401365.

Deming:1968:WS

[Dem68b] W. Edwards Deming. Walter A. Shewhart, 1891–1967. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):372–375, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401495.

Depoid:1957:BRS

[Dep57] P. Depoid. Book review: Die Sterblichkeit nach dem Familienstande in Westdeutschland 1949–1951 by K. Freudenberg. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):153, 1957. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401680.

Depoid:1964:LLI

[Dep64] P. Depoid. L’effectif de l’Institut international de statistique et son èvolution future. (French) [The workforce of the International Statistical Institute and its future development]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(1/2):72–87, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401975.

Derksen:1953:BRI

[Der53a] J. B. D. Derksen. Book review: An International Comparison of National Products and the Purchasing Power of Currencies; A Study of the United States, the United Kingdom, France, Germany and Italy by M. Gilbert, I. B. Kravis. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(3):199–201, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401430.

Derksen:1953:BRS

[Der53b] J. B. D. Derksen. Book review: Standard of Living in India and Pakistan 1931-’32 to 1940-’41 by R. C. Desai. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(3):197–199, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401429.

Derksen:1954:BRI

[Der54a] J. B. D. Derksen. Book review: Income and Wealth, Series III by M. Gilbert. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22(1/3):173–174, 1954. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401929.

Derksen:1954:BRP

[Der54b] J. B. D. Derksen. Book review: Price Indexes and Quality Changes by E. von Hofsten. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22(1/3):179– 180, 1954. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401934.

Derksen:1954:CSC

[Der54c] J. B. D. Derksen. Comparability of statistics of capital formation: Some notes and comments. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22 (1/3):57–67, 1954. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401920.

Derksen:1957:BRO

[Der57] J. B. D. Derksen. Book review: Der öffentliche Haushalt in der volkswirtschaftlichen Gesamtrechnung. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):156–157, 1957. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401684.

Derksen:1958:BRB

[Der58] J. B. D. Derksen. Book review: La balance des payements; étude méthodologique. Principaux résultats pour la France de 1910 à 1956. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):140–141, 1958. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401586.

Derksen:1960:BRS

[Der60] J. B. D. Derksen. Book review: Social Accounting and Economic Models by R. Stone, G. Croft-Murray. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(1/2):105, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401849.

Derksen:1961:BRD

[Der61a] J. B. D. Derksen. Book review: La distribuzione dei redditi by F. Brambilla. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(1):148, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401558.

Derksen:1961:BRP

[Der61b] J. B. D. Derksen. Book review: Productivity Trends in the United States by J. W. Kendrick. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29 (3):91–92, 1961. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401957.

Derksen:1963:BRB

[Der63] J. B. D. Derksen. Book review: Basic Mathematics by R. G. D. Allen. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(2):262, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401382.

Derksen:1965:BRM

[Der65] J. B. D. Derksen. Book review: Models of Income Determination by National Bureau of Economic Research. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(3):563–564, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401718.

Derksen:1968:BRS

[Der68] J. B. D. Derksen. Book review: Systems of Social Accounts by G. Stuvel. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):113–114, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401363.

Derksen:1971:BRE

[Der71a] J. B. D. Derksen. Book review: Education and Manpower: Theoretical Models and Empirical Applications by T. Thonstad. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):117–118, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402252.

Derksen:1971:BRS

[Der71b] J. B. D. Derksen. Book review: Scientific Papers of Tjalling C. Koopmans by M. Beckmann, C. F. Christ, M. Nerlove, Tjalling C. Koopmans. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):105, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402235.

Dalenius:1968:CC

[DF68] T. E. Dalenius and O. Frank. Control of classification. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):279–295, 1968. CODEN ISTRDP.

ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401469.

Darmois:1957:TSI

[DG57] G. Darmois and R. C. Geary. La trentième session de l’Institut International de Statistique, Stockholm, 8–15 août 1957. (French) [The thirtieth session of the International Statistical Institute, Stockholm, August 8–15, 1957]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):1–6, 1957. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401670.

Divisia:1940:PES

[Div40] F. Divisia. À propos d’une étude statistique des contaminations tuberculeuses. (French) [About a statistical study of tuberculosis contamination]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 8(3/4):113–127, 1940. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401156.

Divisia:1943:ETS

[Div43] F. Divisia. Essai de théorie statistique de la contagion et de la contamination. Donnant un critère de classification des maladies microbiennes au point de vue social et au point de vue individuel et, éventuellement, quelques directives pour l’étude scientifique de ces maladies. (French) [Test of statistical theory of contagion and contamination. Giving a criterion for classifying microbial diseases from a social and individual point of view and, possibly, some guidelines for the scientific study of these diseases]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 11(3/4):150–169, 1943. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401112.

Dixon:1971:NAM

[Dix71a] W. J. Dixon. Notes on available materials to support computer based statistical teaching. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39 (3):257–286, 1971. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402262.

Dixon:1971:SI

[Dix71b] W. J. Dixon. The side inch. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39 (3):315–339, 1971. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402266.

Thege:1938:BRB

[dKT38] J. de Konkoly Thege. Book review: Budapest székesföváros áruforgalma. Güterverkehr der Haupt- und Residenzstadt Budapest by I. L. Illyefalvi. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(4):406, January 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400912.

deLaky:1938:CML

[dL38] Désiré de Laky. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5 (4):358–379, January 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400910.

deMeo:1963:FRS

[dM63] Giuseppe de Meo. Franco R. Savorgnan, 1879–1963. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(3):456–457, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1401459.

deOliveira:1967:BRL

[dO67] J. Tiago de Oliveira. Book review: L’Estimation statistique by D. Dumas de Rauly. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(1):94, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401643.

Doksum:1966:DFS

[Dok66] K. Doksum. Distribution-free statistics based on normal deviates in analysis of variance. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(3):376– 388, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401257.

Doob:1971:BRP

[Doo71] J. L. Doob. Book review: Probability Theory by A. Rényi. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(3):409, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1402283.

Dore:1933:SFI

[Dor33] Valentino Dore. La statistique forestière internationale et ses problèmes. (French) [International forest statistics and its problems]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(3):59–65, October 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400777.

Dore:1936:SFI

[Dor36] Valentino Dore. La statistique forestière internationale. Rapport sur les travaux de la Commission mixte constituée par l’Institut International de Statistique et l’Institut International d’Agriculture. (French) [International forestry statistics. Report on the work of the Joint Commission constituted by the International Institute of Statistics and the International Institute of Agriculture]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(2):206–211, July 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400724.

Downton:1971:BRR

[Dow71] F. Downton. Book review: Rank Order Probabilities: Two-Sample Normal Shift Alternatives by R. C. Milton. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):111–112, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402245.

Drexel:1936:NST

[Dre36] Karl Drexel. Normung statistischer Tabellen. Ausführung des Beschlusses der XXII. Session des Internationalen Statistischen Institutes, London 1934. (German) [Standardization of statistical tables. Implementation of the decision of the XXII Session of the International Statistical Institute, London 1934]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(2):232–237, July 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1400726.

Drexel:1938:ASG

[Dre38] Karl Drexel. Die Abgrenzung statistischer Gruppen. (German) [The delimitation of statistical groups]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(2):288–293, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400745.

Drevon:1969:REP

[Dre69] Marc Drevon. Sur la representation de l’échantillonnage par un arbre et les applications que l’on peut en faire au calcul de moments de variables aléatoires qui lui sont attachées. (French) [On the representation of sampling by a tree and the applications that can be made to the calculation of moments of random variables attached to it]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(1):45–56, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402094.

Drion:1967:BRA

[Dri67] E. F. Drion. Book review: Applications de la statistique à la démographie et à la biologie by R. Risser, C. E. Traynard. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):206, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401417.

deSztrem:1933:OMA

[dS33] D’Édouard Szturm de Sztrem. Observations sur les méthodes appliquées à l’étude des régions économiques. (French) [Observations on the methods applied to the study of economic regions]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(2):46–57, July 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401173.

deSztrem:1936:CML

[dS36] E. Szturm de Sztrem. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(1): 86–100, April 1936. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401135.

deSztrem:1938:QOQ

[dS38] E. Szturm de Sztrem. Quelques observations sur la question des classifications statistiques. (French) [Some observations on the question of statistical classifications]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(1):50–55, 1938. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400841.

David:1971:SDS

[DS71a] H. A. David and F. B. Six. Sign distribution of standard multinormal variables with equal positive correlation. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):1–3, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402223.

deSmit:1971:BRM

[dS71b] Jos H. A. de Smit. Book review: Markovketten by F. Ferschl. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):242, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402184.

deSmit:1971:BRS

[dS71c] Jos H. A. de Smit. Book review: Semi-Markoff-Prozesse mit endlich vielen Zuständen by H. Störmer. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):248, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402190.

Rodrigues:1958:WLS

[dSR58] M. C. da Silva Rodrigues. Wilfred Leslie Stevens, 1911–1958. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):173–174, 1958. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401598.

Dugue:1955:TLD

[Dug55] D. Dugué. Sur les théorèmes limites du calcul des probabilités. (French) [On the limit theorems of the calculation of probabilities]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 23(1/3):29–35, 1955. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402008.

Dugue:1960:GD

[Dug60] D. Dugué. Georges Darmois (1888–1960). Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(1/2):122–123, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401856.

Dugue:1965:LFC

[Dug65] D. Dugué. L’équipe française de chercheurs dans le plan d’expériences. (French) [The French team of researchers in the experimental plan]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(1):15, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401303.

Dugue:1969:VPM

[Dug69] D. Dugué. Valeurs propres de matrices ayant un role important dans l’étude d’integrales Browniennes. (French) [Eigenvalues of matrices having an important role in the study of Brownian integrals]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(1):13–17, 1969. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402091.

Dumard:1970:BRS

[Dum70] J. Dumard. Book review: The Short-Run Demand for Workers and Hours by R. C. Fair. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):300–301, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1402159.

Dunn:1940:CML

[Dun40] Halbert L. Dunn. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 8 (1/2):32–49, 1940. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400714.

Dunn:1942:CML

[Dun42] Halbert L. Dunn. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 10 (1/2):75–83, 1942. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401186.

Dunn:1964:HBD

[Dun64] Halbert L. Dunn. Happy birthday, Dr. Rice! Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(1/2):1, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401969.

Dunn:1969:LD

[Dun69a] H. L. Dunn. Louis I. Dublin, 1882–1969. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(3):344–347, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402138.

Dunn:1969:SR

[Dun69b] H. L. Dunn. Stuart A. Rice, 1889–1969. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(3):332–334, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402135.

Durbin:1953:BRD

[Dur53] J. Durbin. Book review: Demand Analysis: A Study in Econometrics by H. Wold, L. Juréen. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(1/2):81–82, 1953. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401518.

Durbin:1954:EV

[Dur54] J. Durbin. Errors in variables. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22(1/3):23–32, 1954. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401917.

Durbin:1958:BRT

[Dur58] J. Durbin. Book review: Three Methods of Comparing the Purchasing Power of Currencies by J. van IJzeren. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):137–138, 1958. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401582.

Durbin:1960:FTS

[Dur60] J. Durbin. The fitting of time-series models. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(3):233–244, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401322.

deWolff:1952:BRS

[dW52] P. de Wolff. Book review: Städtestatistik in Verwaltung und Wissenschaft by M. B. Mewes. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 20 (1):82–83, 1952. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401832.

deWolff:1954:BRC

[dW54] P. de Wolff. Book review: Consumers’ Expenditure in the United Kingdom, 1900–1919 by A. R. Prest. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22(1/3):176–177, 1954. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401932.

Ewens:1961:APS

[EG61] W. J. Ewens and J. Gani. Absorption probabilities in some genetic processes. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(2):33–41, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401656.

Erlander:1965:SCR

[EG65] S. Erlander and J. Gustavsson. Simultaneous confidence regions in normal regression analysis with an application to road accidents. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(3):364–377, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401701.

Eicher:1965:BRG

[Eic65] F. Eicher. Book review: Geschichte der Frauenarbeit in Zürich by Emma Steiger. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(2):333, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402046.

Ekman:1963:S

[Ekm63] G. Ekman. On the sum ∑_{1}^{n} P_{h}^{i}σ_{h}^{j}. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(1):67–80, 1963. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401729.

Ekman:1969:GSD

[Ekm69] G. Ekman. A graphical solution for the determination of an optimal stratification or grouping, with an example concerning a problem of maximizing sales revenues. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):186–194, 1969. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402291.

Elfving:1966:BRS

[Elf66] G. Elfving. Book review: Selected Papers of Richard von Mises, Vol. 2. Probability and statistics, General by G. Birkhoff, Richard von Mises. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):280–281, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401783.

Elfving:1969:BRS

[Elf69] G. Elfving. Book review: A Selection of Early Statistical Papers of J. Neyman by J. Neyman. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37 (1):112–113, 1969. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402108.

F:1936:BRE

[F.36] M. F. Book review: Einführung in die mathematische Statistik by O. N. Anderson. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(2):315–316, July 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400733.

Faleschini:1969:BRE

[Fal69] L. Faleschini. Book review: Elementary Statistics by K. Hope. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):230, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402305.

Federici:1969:BRE

[Fed69] N. Federici. Book review: Étude démographique de la nuptialité en Belgique by C. Wattelar, G. Wunsch. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(1):113–114, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402110.

Ferguson:1961:RRO

[Fer61] Thomas S. Ferguson. Rules for rejection of outliers. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(3):29–43, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1401948.

Freiberger:1965:FSM

[FG65] W. Freiberger and U. Grenander. On the formulation of statistical meteorology. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(1):59–86, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401307.

Finney:1954:BRS

[Fin54] D. J. Finney. Book review: Statistics and Mathematics in Biology by O. Kempthorne, T. A. Bancroft, J. W. Gowen, J. L. Lush. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22(1/3):170–171, 1954. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401925.

Finney:1956:CSV

[Fin56] D. J. Finney. The consequences of selection for a variate subject to errors of measurement. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 24(1/3): 1–10, 1956. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401275.

Finney:1957:BRL

[Fin57] D. J. Finney. Book review: The Lognormal Distribution (With Special Reference to Its Use in Economics) by J. Aitchison, J. A. C. Brown. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):154, 1957. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401681.

Finney:1966:BRB

[Fin66] D. J. Finney. Book review: Biological Statistics: An Introduction by S. C. Pearce. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(1):114–115, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401544.

Fisher:1934:DDT

[Fis34] Irving Fisher. The debt-deflation theory of great depressions. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(4):48–65, January 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401100.

Fisher:1957:MLE

[Fis57] G. R. Fisher. Maximum likelihood estimators with heteroscedastic errors. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):52–55, 1957. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401674.

Fisher:1962:ISH

[Fis62] G. R. Fisher. Iterative solutions and heteroscedasticity in regression analysis. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(2):153–159, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401896.

Flaskamper:1953:SS

[Fla53] P. Flaskämper. Sigmund Schott, 1868–1953. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(3):211, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401439.

Flaskamper:1954:WG

[Fla54] P. Flaskämper. Wilhelm Gerloff, 1880–1954. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22(1/3):194–195, 1954. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401942.

Flaskamper:1965:BRU

[Fla65] P. Flaskämper. Book review: Die Überalterung (Ursachen, Verlauf, wirtschaftliche und soziale Auswirkungen des demografischen Alterungsprozesses) by F. X. Kaufmann. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(2):325–326, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402039.

Flux:1938:SMC

[Flu38] Alfred W. Flux. The statistical measurement of changes in national productive equipment. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(2):235– 240, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400741.

Fortet:1960:PSC

[For60] R. Fortet. Problèmes statistiques concernant des trafics. (French) [Statistical problems regarding traffic]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(3):228–232, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401321.

Fortunati:1961:GP

[For61] P. Fortunati. Gaetano Pietra, 1879–1961. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(3):109–110, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401964.

Forte:1962:LE

[For62] F. Forte. Luigi Einaudi, 1874–1961. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(1):111–112, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402084.

Fortet:1967:BRC

[For67] R. Fortet. Book review: Calcul des probabilités, avec un appendice sur la théorie de l’information by A. Rényi. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(3):329, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401810.

Fortet:1968:BRI

[For68] R. Fortet. Book review: An Introduction to Stochastic Processes with Special Reference to Methods and Applications by M. S. Bartlett. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):350, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401477.

Fougstedt:1966:BRS

[Fou66] G. Fougstedt. Book review: Studies on Fertility and Social Mobility. Proceedings of the International Demographic Symposium, Budapest, November 1962 by E. Szabady. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(1):115–116, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401545.

Frank:1968:SCG

[Fra68] O. Frank. Stochastic competition graphs. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):319–326, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401473.

Frechet:1934:CDC

[Fré34] Maurice Fréchet. Sur le coefficient, dit de corrélation et sur la corrélation en général. (French) [On the so-called correlation coefficient and on correlation in general]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(4):16–23, January 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401096.

Frechet:1936:RCE

[Fré36a] M. Fréchet. Rapport de la commission d’étude de l’usage du coefficient de corrélation. (French) [Report of the study commission on the use of the correlation coefficient]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(2):238–242, July 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1400727.

Frechet:1936:CLD

[Fré36b] Maurice Fréchet. Sur le coefficient de linéarité, dit de corrélation. (French) [On the coefficient of linearity, known as correlation]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(4):365–379, January 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401215.

Frechet:1939:FRR

[Fré39] Maurice Fréchet. Sur les formules de répartition des revenus. (French) [On income distribution formulas]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(1):32–38, 1939. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400932.

Frechet:1940:CDM

[Fré40] Maurice Fréchet. Comparaison des diverses mesures de la dispersion. (French) [Comparison of various measures of dispersion]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 8(1/2):1–12, 1940. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400712.

Frechet:1941:DDV

[Fré41] Maurice Fréchet. Sur une définition descriptive de la valeur moyenne et sur un mode nouveau d’introduction de l’intégrale de Stieltjes dans l’enseignement du calcul des probabilités. (French) [On a descriptive definition of the average value and on a new mode of introducing the Stieltjes integral in the teaching of probability calculation]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 9(3/4):113–121, 1941. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1400873.

Frechet:1942:MPN

[Fré42] Maurice Fréchet. La médiane d’un petit nombre de variables aléatoires et sa dispersion. (French) [The median of a small number of random variables and its dispersion]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 10(3/4):139–151, 1942. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400895.

Frechet:1943:ECE

[Fré43] Maurice Fréchet. Sur l’extension de certaines evaluations statistiques au cas de petits échantillons. (French) [On the extension of certain statistical evaluations to the case of small samples]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 11(3/4):182–205, 1943. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401114.

Frechet:1945:NEE

[Fré45] Maurice Fréchet. Nouveaux essais d’explication de la répartition des revenus. (French) [New attempts to explain income distribution]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 13(1/4):16–32, 1945. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400975.

Frechet:1946:DPL

[Fré46] M. Maurice Fréchet. Dégager les possibilités et les limites de l’application des sciences mathématiques (et en particulier du calcul des probabilités) à l’étude des phénomènes économiques et sociaux. (Première enquête scientifique). (French) [Identify the possibilities and limits of the application of mathematical sciences (and in particular the calculation of probabilities) to the study of economic and social phenomena. (First scientific investigation)]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 14(1/4):16–51, 1946. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401146.

Frechet:1948:VTO

[Fré48] Maurice Fréchet. Les valeurs typiques d’ordre nul ou infini d’un nombre aléatoire. (French) [Typical zero or infinite order values of a random number]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 16(1/4): 1–22, 1948. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401229.

Frechet:1950:EIS

[Fré50] Maurice Fréchet. Sur un essai infondé de sauver le coefficient classique dit de corrélation. (French) [On an unfounded attempt to save the classic so-called correlation coefficient]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18(3/4):157–160, 1950. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401034.

Frechet:1957:TCP

[Fré57] M. Fréchet. Les tableaux de corrélation et les programmes linéaires. (French) [Correlation tables and linear programs]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):23–40, 1957. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1401672.

Frechet:1958:PPC

[Fré58] M. Fréchet. Sur la portée pratique de la convergence presque certaine. (French) [On the practical significance of almost certain convergence]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):26–28, 1958. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401569.

Frechet:1960:TDM

[Fré60] M. Fréchet. Sur les tableaux dont les marges et des bornes sont données. (French) [On tables whose margins and limits are given]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(1/2):10–32, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401846.

Freudenthal:1971:BRE

[Fre71] H. Freudenthal. Book review: Elements of Functional Analysis by I. J. Maddox. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):110–111, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402243.

Furst:1953:BRB

[Für53] G. Fürst. Book review: Die Bundesstatistik; das Arbeitsgebiet des Statistischen Bundesamtes und die von den obersten Bundesbehörden bearbeiteten Statistiken, Stand 31.12.1953. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(3):202–203, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1401432.

Furst:1961:BRC

[Für61] G. M. W. Fürst. Book review: La carta commerciale d’Italia, con le sue 414 aree e subaree di attrazione del commercio al dettaglio e le relative quote di mercato by G. Tagliacarne. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(1):147, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401556.

Furst:1963:KW

[Für63] Gerhard Fürst. Karl Wagner, 1893–1963. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(2):274, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401391.

Furst:1965:PJ

[Für65] G. Fürst. Paul Jostock, 1895–1965. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(2):345–346, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402051.

G:1934:BRA

[G.34a] E. H. G. Book review: Annals of the Royal Statistical Society, 1834–1934. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(2):233–234, July 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401002.

G:1934:BRM

[G.34b] U. G. Book review: La mobilità della popolazione all’ interno delle città europee by Robert Bachi. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1 (4):96, January 1934. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401104.

G:1934:BRS

[G.34c] U. G. Book review: Lo spopolamento montano in Italia, IV l’Appennino Emiliano-Tosco-Romagnolo. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(4):97, January 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401105.

G:1938:BRM

[G.38] R. C. G. Book review: Méthode des fonctions arbitraires. Théorie des événements en chaîne dans le cas d’un nombre fini d’états possibles by M. Fréchet. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(3):488–489, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401198.

G:1948:BRB

[G.48] G. G. Book review: Bibliography of Selected Statistical Sources of the American Nations. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 16(1/4): 130, 1948. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401234.

Galvani:1951:RCC

[Gal51] Luigi Galvani. Revision critique de certains points de la méthode representative. (French) [Critical review of certain points of the representative method]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(1):1– 12, 1951. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401498.

Gani:1966:BRS

[Gan66] J. Gani. Book review: Sampling from a Graph. Mathematical Centre, Netherlands by A. R. Bloemena. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):271, 1966. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401775.

Gani:1967:BRR

[Gan67] J. Gani. Book review: Research Papers in Statistics by F. N. David. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(3):322–323, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401803.

Gani:1971:BRI

[Gan71] J. Gani. Book review: Introduction to Mathematical Probability Theory by M. Eisen. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):108–109, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1402240.

Gart:1971:CPR

[Gar71] John J. Gart. The comparison of proportions: A review of significance tests, confidence intervals and adjustments for stratification. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):148–169, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402171.

Geary:1938:BRG

[Gea38] R. C. Geary. Book review: Généralités sur les probabilités. Variables aléatoires by M. Fréchet. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(1):161–162, 1938. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400846.

Geary:1963:BRB

[Gea63a] R. C. Geary. Book review: Bibliography of Statistical Literature, 1950–1958 by M. G. Kendall, A. G. Doig. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(1):140–141, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401734.

Geary:1963:SRA

[Gea63b] R. C. Geary. Some remarks about relations between stochastic variables: a discussion document. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(2):163–181, 1963. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401371.

Geary:1965:BRP

[Gea65] R. C. Geary. Book review: Planning and Statistics in Socialist Countries. The Indian Statistical Institute. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(2):331–333, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402045.

Geisser:1969:BRE

[Gei69] S. Geisser. Book review: Elementary Statistical Methods by G. B. Wetherill. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(1):114–115, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402111.

Geppert:1958:BRP

[Gep58] M. P. Geppert. Book review: Proceedings of the Third Berkeley Symposium on Mathematical Statistics and Probability. Vol. 4. Contributions to Biology and Problems of Health by J. Neyman. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):135–136, 1958. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401578.

Geary:1955:UCI

[GF55] R. C. Geary and K. G. Forecast. The use of census of industrial production material for the estimation of productivity. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 23(1/3):6–19, 1955. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402006.

Gower:1970:CG

[GHM70] J. C. Gower, R. J. Howarth, and Matthias Matschinski. Classification and geology. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):35–41, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402322.

Giannone:1964:CBT

[Gia64] Antonino Giannone. Costantino Bresciani Turroni, 1882–19963. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):347–348, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401887.

Gilbert:1951:BRI

[Gil51] Milton Gilbert. Book review: Income and Wealth, Series I; International Association for Research in Income and Wealth by E.

Lundberg. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(2):199–200, 1951. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401244.

Gilbert:1953:ONA

[Gil53] Milton Gilbert. The organization of national accounts statistics in underdeveloped countries: A report on Turkey. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(1/2):9–16, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401512.

Gillis:1962:BRS

[Gil62] P. P. Gillis. Book review: Studies in Applied Probability and Management Science by K. J. Arrow, S. Karlin, H. Scarf. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(2):272–273, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401903.

Gini:1933:MLA

[Gin33] Corrado Gini. Methodologie, legislation, administration et organisation statistiques. (French) [Statistical methodology, legislation, administration and organization]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(1):56–67, 1933. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401085.

Gini:1934:MCI

[Gin34] Corrado Gini. On a method for calculating the infantile deathrate according to the month of death. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(3):292–300, October 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400987.

Gini:1936:CCC

[Gin36] Corrado Gini. Sur le coefficient de corrélation comme caractéristique du degré de la concordance. (French) [On the correlation coefficient as a characteristic of the degree of concordance].

Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(3):355–366, October 1936. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400804.

Gini:1937:QIPb

[Gin37a] Corrado Gini. Une question importante pour la science des constitutions et pour la médécine militaire: comment juger si les proportions d’un individu sont normales? Deuxième note. (French) [An important question for the science of constitutions and for military medicine: how to judge whether the proportions of an individual are normal? Second note]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(3):203–211, October 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400952.

Gini:1937:QIPa

[Gin37b] Corrado Gini. Une question importante pour la science des constitutions et pour la médécine militaire: comment juger si les proportions d’un individu sont normales?. (French) [An important question for the science of constitutions and for military medicine: how to judge whether the proportions of an individual are normal?]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(2):107–114, July 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400829.

Gini:1939:MPAb

[Gin39a] Corrado Gini. Sur la “méthode des profils” et sur d’autres diagrammes à ordonnées jointes, dans le cas de séries non ordonnées. Deuxième article. (French) [On the “profile method” and on other attached ordinate diagrams, in the case of unordered series. Second article]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(4):225–246, 1939. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400962.

Gini:1939:MPAa

[Gin39b] Corrado Gini. Sur la “méthode des profils” et sur d’autres diagrammes à ordonnées jointes, dans le cas de séries non ordonnées. (French) [On the “profile method” and on other attached ordinate diagrams, in the case of unordered series]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(1):1–31, 1939. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400931.

Gini:1941:MFN

[Gin41] Corrado Gini. Sur la mesure de la fécondité naturelle de la femme mariée. (French) [On the measurement of the natural fertility of married women]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 9(1/2):1–20, 1941. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400941.

Gini:1942:MEP

[Gin42] Corrado Gini. Sur la mesure de l’efficacité des pratiques anticonceptionnelles. (French) [On measuring the effectiveness of contraceptive practices]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 10(1/2):1–36, 1942. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401183.

Gini:1943:CDP

[Gin43] Corrado Gini. A coordination of the different population theories. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 11(1/2):35–67, 1943. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400787.

Gini:1945:FER

[Gin45] Corrado Gini. Sur les fondements des évaluations de la richesse nationale. (French) [On the foundations of assessments of national wealth]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 13(1/4):33–61, 1945. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400976.

Gini:1947:SRT

[Gin47] Corrado Gini. Statistical relations and their inversions. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 15(1/4):24–42, 1947. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401206.

Gini:1949:NIE

[Gin49] Corrado Gini. National income estimates. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 17(3/4):119–128, 1949. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1400863.

Gini:1950:OARb

[Gin50a] Corrado Gini. Observations au rapport de M. Sauvy sur les développements récents des études de statistique démographique. (French) [Observations on the report of Mr. Sauvy on recent developments in demographic statistics studies]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18(1/2):74–77, 1950. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400763.

Gini:1950:OARa

[Gin50b] Corrado Gini. Observations au rapport du Prof. Cochran sur l’état actuel de la biométrie. (French) [Observations to the report of Prof. Cochran on the current state of biometrics]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18(1/2):70–72, 1950. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/ stable/1400760. See reply [Coc50].

Gini:1950:RDD

[Gin50c] Corrado Gini. Remarks on Dr. Derksen’s survey of national income research. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18(1/2):72–74, 1950. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400762.

Gini:1952:MMP

[Gin52] Corrado Gini. Major and minor problems in the estimates of national income. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 20(1):18–27, 1952. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401827.

Gini:1956:EHS

[Gin56] Corrado Gini. Enrico Haskel Sonnabend, 1901–1956. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 24(1/3):162–164, 1956. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401297.

Gini:1957:FB

[Gin57a] C. Gini. Felix Bernstein, 1878–1956. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):185–186, 1957. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401692.

Gini:1957:RB

[Gin57b] C. Gini. Rodolfo Benini, 1862–1956. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):186–190, 1957. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401693.

Girault:1966:BRP

[Gir66] M. Girault. Book review: Processus stochastiques. Problèmes et solutions by L. Takács. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):283, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401786.

Glasser:1962:CCL

[Gla62] G. J. Glasser. On the complete coverage of large units in a statistical study. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(1):28–32, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402070.

Granger:1968:QTS

[GN68] C. W. J. Granger and H. R. Neave. A quick test for slippage. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):309–312, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401471.

Gnanadesikan:1967:BRN

[Gna67] R. Gnanadesikan. Book review: Nonparametric Trend Analysis. A Practical Guide for Research Workers by G. A. Ferguson. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):198–199, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401408.

Godfrey:1937:CML

[God37] Ernest H. Godfrey. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4 (4):518–545, January 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401060.

Godfrey:1939:CML

[God39] Ernest H. Godfrey. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7 (2/3):145–170, 1939. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401028.

Godambe:1965:RCT

[God65] V. P. Godambe. A review of the contributions towards a unified theory of sampling from finite populations. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(2):242–258, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402030.

Godambe:1968:BRE

[God68] V. P. Godambe. Book review: Elements of Nonparametric Statistics by G. E. Noether. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):360, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401490.

Goldberg:1960:BRM

[Gol60] S. A. Goldberg. Book review: The Measurement of National Wealth by R. Goldsmith, Ch. Saunders. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(3):296–297, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401326.

Gorman:1967:BRA

[Gor67] W. M. Gorman. Book review: Aggregation in Economic Analysis: An Introductory Survey by H. A. J. Green. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):202, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401413.

Goudswaard:1945:BRS

[Gou45] G. Goudswaard. Book review: Statistische Methoden für Naturwissenschafter, Mediziner und Ingenieure by Arthur Linder. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 13(1/4):158–159, 1945. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400979.

Goudswaard:1951:BRGa

[Gou51a] G. Goudswaard. Book review: Grundbegriffe moderner statistischer Methodik by L. von Baranow. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(1):109–110, 1951. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401504.

Goudswaard:1951:BRGb

[Gou51b] G. Goudswaard. Book review: Grundfragen der Ökonometrie by W. Winkler. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(3):283, 1951. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402000.

Goudswaard:1951:BRS

[Gou51c] G. Goudswaard. Book review: Statistische Methoden für Naturwissenschafter. Mediziner und Ingenieure by Arthur Linder. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(3):283, 1951. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401999.

Goudswaard:1951:SIS

[Gou51d] G. Goudswaard. Spécialisation, intégration et statistique internationale. (French) [Specialization, integration and international statistics]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(2):130–140, 1951. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401240.

Goudswaard:1953:BRS

[Gou53a] G. Goudswaard. Book review: A Statistical Primer by F. N. David. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(3):203, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401434.

Goudswaard:1953:BRE

[Gou53b] G. Goudswaard. Book review: Econometrics by G. Tintner. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(1/2):83, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401520.

Goudswaard:1953:BRL

[Gou53c] G. Goudswaard. Book review: Linear Computations by Paul S. Dwyer. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(1/2):84, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401522.

Goudswaard:1953:BRMb

[Gou53d] G. Goudswaard. Book review: Methods of Statistical Analysis by C. H. Goulden. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(3):203, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401433.

Goudswaard:1953:BRMa

[Gou53e] G. Goudswaard. Book review: The Methods of Statistics by L. H. C. Tippett. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(1/2):83–84, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401521.

Goudswaard:1954:BRE

[Gou54a] G. Goudswaard. Book review: Exercises in Theoretical Statistics by M. G. Kendall. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22(1/3):180, 1954. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401935.

Goudswaard:1954:BRM

[Gou54b] G. Goudswaard. Book review: Mathematics and Statistics for Economists by Gerhard Tintner. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22(1/3):180, 1954. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401936.

Goudswaard:1954:BRS

[Gou54c] G. Goudswaard. Book review: Studies in Econometric Method by W. C. Hood, T. C. Koopmans. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22(1/3):181, 1954. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401937.

Goudswaard:1955:BRB

[Gou55] G. Goudswaard. Book review: Biometrika Tables for Statisticians, I by E. S. Pearson, H. O. Hartley. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 23 (1/3):127–128, 1955. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402016.

Goudswaard:1956:BRS

[Gou56] G. Goudswaard. Book review: Statistics in Research, Basic Concepts and Techniques for Research Workers by B. Ostle. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 24(1/3):143–144, 1956. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401287.

Goudswaard:1957:BRI

[Gou57] G. Goudswaard. Book review: An Introduction to British Economic Statistics by E. Devons. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):158, 1957. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401686.

Goudswaard:1958:BRM

[Gou58] G. Goudswaard. Book review: Méthode Statistique by E. Morice, F. Chartier. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):137, 1958. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401581.

Goudswaard:1961:BRS

[Gou61a] G. Goudswaard. Book review: Statistische Methoden für Naturwissenschafter, Mediziner und Ingenieure by A. Linder. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(3):90, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401954.

Goudswaard:1961:NLJ

[Gou61b] G. Goudswaard. Nicolaas Lucas Jan van Buttingha Wichers, 1894– 1961. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(3):110–111, 1961. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401965.

Goudswaard:1962:BRS

[Gou62] G. Goudswaard. Book reviews: Statistik im modernen Wirtschaftsund Sozialleben by H. Kellerer; Allgemeine Methodenlehre der Statistik: I. Elementare Methoden unter besonderer Berücksichtigung der Anwendung in den Wirtschafts- und Sozialwissenschaften. Sammlung Göschen, Band 746/746a by J. Pfanzagl; Allgemeine Methodenlehre der Statistik: II. Höhere Methoden unter besonderer Berücksichtigung der Anwendung in Naturwissenschaft, Medizin und Technik. Sammlung Göschen, Band 747/747a by J. Pfanzagl. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(2):271–272, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401902.

Goudswaard:1963:BRS

[Gou63] G. Goudswaard. Book review: Statistical Tables for Biological, Agricultural and Medical Research by Ronald A. Fisher, F. Yates.

Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(3):449, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401452.

Goudswaard:1964:RAS

[Gou64] G. Goudswaard. A re-appraisal of the Statistical Education Programme of the International Statistical Institute. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(1/2):88–103, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1401976.

Goudswaard:1965:BRI

[Gou65] G. Goudswaard. Book review: Introduction to Econometrics by Oskar Lange. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(1):173–174, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401313.

Goudswaard:1968:BRM

[Gou68] G. Goudswaard. Book review: Modern Elementary Statistics by J. E. Freund. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):355, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401483.

Goudswaard:1970:BRSa

[Gou70a] G. Goudswaard. Book review: Statistical Methods in the Social Sciences by D. D. Bugg, M. A. Henderson, K. Holden, P. J. Lund. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(3):416, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402209.

Goudswaard:1970:BRSb

[Gou70b] G. Goudswaard. Book review: Statistics for the Social Scientist by K. A. Yeomans. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(3):424, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402217.

Govindarajulu:1966:BRH

[Gov66] Z. Govindarajulu. Book review: Handbook of Nonparametric Statistics, II by J. E. Walsh. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(3):452–453, 1966. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401271.

Grdjic:1964:BRS

[Grd64] G. Grdjić. Book review: A Study of United Kingdom Imports by M. G. Scott. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(1/2):185– 186, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401989.

Grenander:1965:SMT

[Gre65] U. Grenander. Stochastic models in the theory of pursuit. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(2):177–228, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402028.

Grenander:1967:BRI

[Gre67] U. Grenander. Book review: An Introduction to Probability Theory and Its Applications, Vol. 2 by W. Feller. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(3):325–326, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401805.

Greenberg:1968:ESP

[Gre68] B. G. Greenberg. Evaluation of social programs. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):260–278, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401468.

Grimm:1966:BRA

[Gri66] H. Grimm. Book review: Analyse statistique en biologie by K. Mather. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(1):114, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401543.

Gruson:1964:RC

[Gru64] C. Gruson. René Courtin, 1900–19964. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):352–353, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401890.

Guerreiro:1962:AMD

[Gue62] A. D. Guerreiro. António M. Dogio Tovar de Lemos, 1891–1961. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(1):112–113, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402085.

Guerin:1965:VEP

[Gué65] R. Guérin. Vue d’ensemble sur les plans en blocs incomplets équilibrés et partiellement équilibrés. (French) [Overview of balanced and partially balanced incomplete block designs]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(1):24–58, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401306.

Gumbel:1965:QEP

[Gum65] E. J. Gumbel. A quick estimation of the parameters in Fréchet’s distribution. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(3):349–363, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401700.

Gumbel:1967:EJG

[Gum67] H. Gumbel. Emil J. Gumbel, 1891–1966. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(1):104–105, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401652.

Guttman:1968:BRL

[Gut68] I. Guttman. Book review: The Linear Hypothesis: A General Theory by G. A. F. Seber. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):238, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401622.

Goudswaard:1963:CCS

[GV63] G. Goudswaard and J. Ch. W. Verstege. The construction of a classification scheme for a general statistical bibliography. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(2):229–234, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401377.

Haberman:1962:DNE

[Hab62] S. Haberman. The distributions of “s” for the numerous and extensive ties in one of a pair of rankings of length n. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(3):336–351, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401743.

Hald:1957:BRS

[Hal57] A. Hald. Book review: Statistical Methods in Research and Production; With Special Reference to the Chemical Industry by O. L. Davies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):151, 1957. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401677.

Hald:1967:TSS

[Hal67] A. Hald. On the theory of single sampling inspection by attributes based on two quality levels. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(1):1–29, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401632.

Hamaker:1950:SNL

[Ham50] H. C. Hamaker. Some notes on lot-by-lot inspection by attributes. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18(3/4):179–196, 1950. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401036.

Hamaker:1953:ACL

[Ham53] H. C. Hamaker. “Average confidence” limits for binomial probabilities. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(1/2):17–27, 1953. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401513.

Hamaker:1954:BRP

[Ham54] H. C. Hamaker. Book review: Planen und Auswerten von Versuchen by A. Linder. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22(1/3):171– 172, 1954. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401926.

Hamaker:1958:BRC

[Ham58] H. C. Hamaker. Book review: Correlated Random Normal Deviates by E. C. Fieller, T. Lewis, E. S. Pearson. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):136, 1958. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401579.

Hamaker:1960:BRM

[Ham60] H. C. Hamaker. Book review: Messen und Regeln in der Betriebswirtschaft; Einführung in die informationswissenschaftlichen Grundzüge der industriellen Unternehmensforschung by A. Adam. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(3):298–299, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401328.

Hamaker:1967:BRS

[Ham67] H. C. Hamaker. Book review: Statistical Assessment of the Life Characteristic. A Bibliographic Guide by W. R. Buckland. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(1):92–93, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401641.

Hamaker:1971:NTS

[Ham71] H. C. Hamaker. New techniques of statistical teaching. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(3):351–360, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402271.

Hansen:1969:WH

[Han69] M. H. Hansen. William N. Hurwitz, 1908–1969. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(3):343–344, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1402137.

Harmaja:1935:RSD

[Har35] Leo Harmaja. Recherches statistiques sur les droits de douane en Finlande. Étude basée sur l’enquête sur le coût de la vie. (French) [Statistical research on customs duties in Finland. Study based on the cost of living survey]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(2): 163–179, July 1935. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400854.

Harmaja:1936:JOA

[Har36] Leo Harmaja. Les jeunesses ouvrières agricole et industrielle comparées. Quelques résultats d’une étude statistique effectuée en Finlande. (French) [Agricultural and industrial working youth compared. Some results of a statistical study carried out in Finland]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(1):40–44, April 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401131.

Hartley:1965:BRT

[Har65] H. O. Hartley. Book review: Tables for Testing Significance in a 2× 2 Contingency Table by D. J. Finney, R. Latscha, B. M. Bennett, P. Hsu. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(3):560–561, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401713.

Harewood:1967:BRM

[Har67a] R. J. Harewood. Book review: The Measurements and Interpretation of Job Vacancies by National Bureau of Economic Research U. S. A. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):206–208, 1967. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401418.

Harrison:1967:BRC

[Har67b] T. K. Harrison. Book review: Cases in Management Statistics by N. L. Enrick. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):196–197, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401406.

Harris:1968:BRI

[Har68] B. Harris. Book review: Introduction to Measure & Probability by J. F. C. Kingman, S. J. Taylor. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):109–110, 1968. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401357.

Harewood:1970:BRP

[Har70] J. Harewood. Book review: Population, Laborforce, and Long Swings in Economic Growth by R. A. Easterlin. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):297–298, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1402157.

Hartley:1971:BRN

[Har71] H. O. Hartley. Book review: Normal Centroids and Medians and Scores for Ordinal Data by F. N. David, D. E. Barton, S. Ganeshalingham, H. L. Harter, P. J. Kim, M. Merrington, D. Walley. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):107, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402238.

Hasegawa:1933:PGT

[Has33] T. Hasegawa. The population of Greater Tokyo. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(2):30–39, July 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401171.

Hawkes:1967:BRP

[Haw67] A. G. Hawkes. Book review: Proceedings of the Symposium on Congestion Theory, University of North Carolina, 1964 by W. L.

Smith, W. E. Wilkinson. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(1):100– 101, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401649.

Hayashi:1970:BRP

[Hay70] C. Hayashi. Book review: Population Genetics by W. J. Ewens. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):186, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402336.

Hemelrijk:1971:CAG

[HDPD71] J. Hemelrijk, T. Dalenius, G. Palm, and W. Dixon. Comments about a general audience TV-Course on statistics. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(3):312–314, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1402265.

Hecht:1934:CML

[Hec34] M. Hecht. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(1):96–105, April 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401017.

Hecht:1935:CML

[Hec35] M. Hecht. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(2):180–193, July 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1400855.

Heckmann:1969:BRI

[Hec69] O. Heckmann. Book review: Introduction to Stellar Statistics by R. Kurth. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(1):107, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402102.

Hemelrijk:1960:DVD

[Hem60] J. Hemelrijk. David van Dantzig (1900–1959). Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(1/2):124–125, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401857.

Hemelrijk:1965:BRH

[Hem65] J. Hemelrijk. Book review: Handbook of Mathematical Functions with Formulas, Graphs and Mathematical Tables. U.S. Department of Commerce, National Bureau of Standards by Milton Abramowitz, Irene A. Stegun. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(2):324, 1965. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402037.

Hemelrijk:1969:BRT

[Hem69] J. Hemelrijk. Book review: Tables of the Incomplete Beta-Function by K. Pearson. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(1):111, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402107.

Henry:1953:FTM

[Hen53] L. Henry. Fondements théoriques des mesures de la fécondité naturelle. (French) [Theoretical foundations of natural fertility measurements]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(3):135–151, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401425.

Henry:1954:BRT

[Hen54] Louis Henry. Book review: The Trend and Pattern of Fertility in Great Britain; A Report on the Family Census of 1946 by D. V. Glass, E. Grebenik. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22(1/3):175– 176, 1954. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401931.

Henon:1961:BRS

[Hen61] R. A. A. J. Henon. Book review: Statistical Cost Analysis by J. Johnston. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(2):82–83, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401659.

Hennequin:1968:BRP

[Hen68] P. L. Hennequin. Book review: Probability; A Survey of the Mathematical Theory by J. Lamperti. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):237, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401620.

Hersch:1934:TGT

[Her34] L. Hersch. Taux globaux et taux composites en demographie (pour la réhabilitation des taux “bruts”). (French) [Overall rates and composite rates in demography (for the rehabilitation of “raw” rates)]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(1):5–25, April 1934. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401011.

Hersch:1940:QPV

[Her40] Liebmann Hersch. De quelques potentiels-vie et de certaines variétés de vie moyenne. (French) []. (French) [Of some life potentials and certain varieties of average life]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 8(3/4):128–162, 1940. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401157.

Hersch:1942:MPV

[Her42] Liebmann Hersch. La méthode des potentiels-vie appliquée à l’étude du mouvement naturel de la population. (French) [The life potential method applied to the study of the natural movement of the population]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 10(3/4):152–183, 1942. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1400896.

Hersch:1944:QPM

[Her44] Liebmann Hersch. Quelques précisions sur la méthode des potentiels-vie et ses notions fondamentales. (En réponse aux critiques du Prof. Wilhelm Winkler). (French) [Some details on the life potential method and its fundamental notions. (In response to criticisms from Prof. Wilhelm Winkler)]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 12(1/4):23–35, 1944. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401069.

Hersch:1952:LIC

[Her52] L. Hersch. Les limites d’une importante contribution critique à la demographie potentielle. (French) [The limits of an important critical contribution to potential demography]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 20(1):28–39, 1952. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401828.

Heron:1961:ECS

[Her61] D. Heron. Ernest Charles Snow (1886–1959). Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(2):94, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401666.

Hoijer:1943:SCP

[HG43] E. J. Höijer and Otto Grönlund. La statistique criminelle et pénitentiaire des pays nordiques. (French) [Crime and prison statistics in the Nordic countries]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 11 (3/4):170–181, 1943. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401113.

Hodson:1970:STR

[HK70] F. R. Hodson and P. R. Krishnaiah. Statistical training and research / formation et recherche statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):282–287, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402151.

Hoeffding:1968:SRD

[Hoe68] W. Hoeffding. Some recent developments in nonparametric statistics. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):176–183, 1968. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401605.

Hoem:1970:BRR

[Hoe70] J. M. Hoem. Book reviews: Risk Theory by R. E. Beard, T. Pentikäinen, E. Pesonen; Stochastic Theory of a Risk Business by H. L. Seal. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):294–296, 1970. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402154.

Horner:1968:TPS

[Hor68a] F. B. Horner. Training of professional staff within the Commonwealth bureau of census and statistics. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):201–206, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401607.

Horstmann:1968:FB

[Hor68b] K. Horstmann. Friedrich Burgdoerfer, 1890–1967. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):243–244, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1401628.

Horstmann:1969:BRW

[Hor69a] K. Horstmann. Book review: World Views on Population Problems by E. Szabady. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(3):328–329, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1402132.

Horvath:1969:BMJ

[Hor69b] R. A. Horváth. Le bicentenaire de la mort de Johann Peter Süssmilch (1717–67) et la discipline statistique. (French) [The bicentenary of the death of Johann Peter Süssmilch (1717–67) and the statistical discipline]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(1):36–44, 1969. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402093.

Hostelet:1935:FSC

[Hos35] Georges Hostelet. Les faits statistiques, le calcul des fréquences binominales et les probabilités dans les jeux de hasard. (French) [Statistical facts, the calculation of binomial frequencies and probabilities in games of chance]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3 (3):278–300, October 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400795.

Hostelet:1937:PEC

[Hos37] Georges Hostelet. Programme d’études comparées des budgets familiaux. Ses difficultés actuelles. (French) [Program of comparative studies of family budgets. Its current difficulties]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(1):15–36, April 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400922.

Hostelet:1938:CFB

[Hos38] Georges Hostelet. Charge fiscale, bilan fiscal, capacite fiscale du contribuable. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(3):394–409, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401193.

Hostelet:1939:LIS

[Hos39] Georges Hostelet. L’emploi des indices dans les sciences et specialement dans les sciences sociales. (French) [The use of indices in the sciences and especially in the social sciences]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(4):247–258, 1939. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400963.

Hostelet:1946:CRS

[Hos46] Georges Hostelet. Comment remédier à la signification indeterminée des indices statistiques traditionnels. (French) [How to overcome the indeterminate meaning of traditional statistical indices].

Revue de l’Institut international de statistique = Review of the International Statistical Institute, 14(1/4):1–15, 1946. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401145.

Hotelling:1950:CRF

[Hot50] Harold Hotelling. Comments on the report “Furtherance of statistical education”. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18(1/2):69–70, 1950. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1400759.

Hartley:1968:CEA

[HR68] H. O. Hartley and J. N. K. Rao. Classification and estimation in analysis of variance problems. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):141–147, 1968. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401602.

Hooper:1958:EWM

[HT58] J. W. Hooper and H. Theil. The extension of Wald’s method of fitting straight lines to multiple regression. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):37–47, 1958. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401571.

Hamdan:1971:MPB

[HT71] M. A. Hamdan and C. P. Tsokos. A model for physical and biological problems: The bivariate-compounded Poisson distribution. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):60–63, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402230.

Huber:1936:MRP

[HU36] Michel Huber and Henri Ulmer. Les méthodes de recensement dans les pays d’Extrême-Orient. (French) [Census methods in Far Eastern countries]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(4):412–426, January 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401220.

Huber:1933:SIF

[Hub33] Michel Huber. La statistique internationale des forces motrices. (French) [International statistics of driving forces]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(2):40–45, July 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401172.

Huber:1936:SIF

[Hub36a] Michel Huber. La statistique internationale des forces motrices. (French) [International statistics of driving forces]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(1):62–70, April 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401133.

Huber:1936:CTM

[Hub36b] Michel Huber. Le calcul des taux de mortalité aux âges élevés. (French) [Calculating mortality rates at older ages]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(4):400–411, January 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401219.

Huber:1937:APF

[Hub37] M. Huber. Avant-propos. (French) [Foreword]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(1):1, April 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400920.

Huber:1938:CML

[Hub38a] M. Huber. Communications sur la méthodologie. La législation, l’organisation et l’administration statistiques. (French) [Communications on methodology. Statistical legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(3):428–462, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401196.

Huber:1938:XSI

[Hub38b] Michel Huber. La XXIVe session de l’Institut International de Statistique, Prague, 1938. (French) [The XXIVth session of the International Statistical Institute, Prague, 1938]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(3):410–418, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401194.

Huber:1938:RTP

[Hub38c] Michel Huber. Rapport sur les travaux préparatoires à vème révision décennale de la nomenclature internationale des causes de décès. (French) [Report on the preparatory work for the Fifth Ten-Year Review of the International Nomenclature of Causes of Death]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(1):103–128, 1938. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400843.

Huber:1939:CES

[Hub39] Michel Huber. Le Comité d’experts statisticiens de la Société des Nations (1931–1939). (French) [The Committee of Statistical Experts of the League of Nations (1931–1939)]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(2/3):117–137, 1939. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401026.

Huber:1941:CES

[Hub41] Michel Huber. Le Comité d’experts statisticiens du Bureau International du Travail (1933–1939). (French) [The Committee of Statistical Experts of the International Labor Office (1933–1939)]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 9(3/4):104–112, 1941. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400872.

Huber:1942:BRM

[Hub42] M. Huber. Book review: Méthodes d’enquête sur les conditions de vie des familles. Revenu—dépenses—consommation by R. M. Woodbury. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 10(3/4):225, 1942. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400899.

Huber:1966:BRS

[Hub66] P. J. Huber. Book review: Studies in Subjective Probability by H. E. Kyburg, Jr., H. E. Smokler. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(3):452, 1966. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401270.

Huntsberger:1967:BRM

[Hun67] D. V. Huntsberger. Book review: Mathematics and Statistics for Students of Chemistry, Chemical Engineering, Chemical Technology, and Allied Subjects by C. J. Brookes, I. G. Betteley, S. M. Loxton. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):193–194, 1967. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401402.

Hunter:1971:PCC

[Hun71] J. Stuart Hunter. Post College continuing education activities in statistics. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(3):307–311, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402264.

Husain:1968:MZ

[Hus68] Q. M. Husain. M. Ziauddin, 1906–1967. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):118–119, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401367.

Hussain:1970:ETS

[Hus70] Ashiq Hussain. Efficiency of two-stage estimators for the parameters of a multi-regression model. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):220–222, 1970. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402145.

Hansen:1970:RNS

[HW70] Morris H. Hansen and Joseph Waksberg. Research on nonsampling errors in censuses and surveys. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(3):317–332, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402198.

Idenburg:1940:MEN

[ID40] Ph. J. Idenburg and J. B. D. Derksen. Methods of evaluation of the national income of the Netherlands, 1921–1938. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 8(1/2):13–31, 1940. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1400713.

Idenburg:1936:CML

[Ide36a] Philip J. Idenburg. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(2): 279–286, July 1936. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400731.

Idenburg:1936:CSD

[Ide36b] Philip J. Idenburg. Le calcul statistique du rendement dans l’enseignement. (French) [Statistical calculation of performance in teaching]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(1):27–39, April 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401130.

Idenburg:1936:SRF

[Ide36c] Philip J. Idenburg. Statistique de la radiodiffusion. (French) [Broadcasting statistics]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(2):266– 278, July 1936. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400730.

Idenburg:1938:DPE

[Ide38] Ph. J. Idenburg. Directives pour l’elaboration des statistiques de la radiodiffusion dans les divers pays. Rapport au nom de la commission mixte entre l’Institut International de Statistique et l’Union Internationale de Radiodiffusion. (French) [Guidelines for the compilation of broadcasting statistics in various countries. Report on behalf of the joint commission between the International Statistical Institute and the International Broadcasting Union]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(2):241–250, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400742.

Idenburg:1953:SRP

[Ide53] Ph. J. Idenburg. Stuart a. Rice, président honoraire / Stuart a. Rice, honorary President. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(1/2):1, 1953. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401510.

Idenburg:1955:HWM

[Ide55] Ph. J. Idenburg. Henri Willem Methorst, 1868–1955. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 23(1/3):140–142, 1955. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402021.

Illyefalvi:1937:BRN

[Ill37] Louis I. Illyefalvi. Book review: Numéro du Jubilé de M. Gustave Thirring. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(3):272–273, October 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400956.

Idenburg:1952:COA

[IM52] Ph. J. Idenburg and J. T. Marshall. Communications sur l’organisation et l’administration statistiques / communications on statistical organisation and administration. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 20(2):148–161, 1952. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401819.

Iosifescu:1971:BRB

[Ios71] Marius Iosifescu. Book review: Basic Probability Theory by R. B. Ash. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):238, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402178.

Irwin:1956:JW

[Irw56] J. O. Irwin. John Wishart, 1898–1956. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 24(1/3):161–162, 1956. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401296.

Isserlis:1952:MG

[Iss52] L. Isserlis. Major Greenwood, 1880–1949. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 20(1):90–91, 1952. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401836.

Ivanovic:1969:BRP

[Iva69]  B. Ivanović.  Book review: Principes de mathématiques

économiques; tome 1 by V. Rouquet La Garrigue, P. Rocher. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(3):327–328, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402130.

Jacobs:1964:WHP

[Jac64] A. Jacobs. Wolfgang Hans Platzer, 1879–1961. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(1/2):188–189, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401991.

Jaros:1969:BRI

[Jar69] F. Jaroš. Book review: Industrial Scheduling Abstracts, 1950– 1966 by S. Eilon, J. R. King. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):228–229, 1969. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402303.

Jensen:1970:CAS

[JHCS70] Arne Jensen, H. C. Hamaker, H. Cramér, and Eeva Stene. A characteristic application of statistics in hydrology. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):42–48, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402323.

Jabine:1961:SDE

[JHH61] T. B. Jabine, R. Hurley, and W. N. Hurwitz. Sample design and estimation procedure for the 1960 sample survey of agriculture in the USA. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(3):1–12, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401946.

Julin:1943:FMMa

[JL43a] Arm. Julin and Edm. Lesoir. Les facteurs moraux et matériels du recul de la natalité. (French) [The moral and material factors of the decline in the birth rate]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 11(1/2):1–34, 1943. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400786.

Julin:1943:FMMb

[JL43b] Arm. Julin and Edm. Lesoir. Les facteurs moraux et matériels du recul de la natalité. (French) [The moral and material factors of the decline in the birth rate]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 11 (3/4):113–149, 1943. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401111.

Johansen:1954:BRD

[Joh54] K. Johansen. Book review: Der Deutsche Zensus 1950 vom Standpunkt der Städte; Erfahrungen, Erkentnisse und Reformvorschläge. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22(1/3):177–179, 1954. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401933.

Johanssen:1962:BRB

[Joh62] K. Johanssen. Book review: British Towns. A Statistical Study of Their Social and Economic Differences by C. A. Moser, W. Scott. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(3):401, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401749.

Johansen:1971:BRP

[Joh71] Søren Johansen. Book review: Probability and Random Processes for Engineers and Scientists by A. B. Clarke, R. L. Disney. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):238–240, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402180.

Jostock:1953:BRI

[Jos53] Paul Jostock. Book review: Income and Wealth, Series II: Income and Wealth of the United States; Trends and Structure by S. Kuznets, R. Goldsmith. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(3):201– 202, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401431.

John:1965:BRP

[JSH65] B. John, K. Stange, and H. C. Hamaker. Book review: Production and Logistics Management by R. E. McGarrah. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(1):175, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401315.

Juan:1971:BRP

[Jua71] M. Jerez Juan. Book review: Principles of Statistical Techniques by P. G. Moore. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):113, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402247.

Julin:1934:RPE

[Jul34] Armand Julin. Résultats principaux d’une enquête sur les budgets d’ouvriers et d’employés en Belgique (1928–1929). (French) [Main results of a survey on the budgets of workers and employees in Belgium (1928–1929)]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(1): 56–71, April 1934. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401014.

Julin:1935:QMD

[Jul35] A. Julin. De quelques méthodes de determination de la covariation. (French) [Some methods for determining covariation]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(3):251–277, October 1935. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400794. See erratum

[Ano36k].

Julin:1938:AD

[Jul38a] Armand Julin. Albert Delatour. (1858–1938). Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(4):I–II, January 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1400904.

Julin:1938:LBAa

[Jul38b] Armand Julin. Luigi Bodio et Adolphe Quetelet. Extraits de leur correspondance (1868–1874). (French) [Luigi Bodio and Adolphe Quetelet. Excerpts from their correspondence (1868–1874)]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(1):1–24, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400839.

Julin:1938:LBAb

[Jul38c] Armand Julin. Luigi Bodio et Adolphe Quetelet. Extraits de leur correspondance (1868–1874). (French) [Luigi Bodio and Adolphe Quetelet. Excerpts from their correspondence (1868–1874)]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(2):195–218, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400738.

K:1935:BRS

[K.35] A. K. Book review: Székesföváros multja és jelene grafikus ábrázolásban (Graphische Bilder aus der Vergangenheit und Gegenwart der Haupt- und Residenzstadt Budapest. [1873–1933]) by L. I. A. Illyefalvi. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(4):437–438, January 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400823.

Kadas:1968:BRF

[Kád68] K. C. Kádas. Book review: Formeln und Tabellen der mathematischen Statistik by U. Graf, H. J. Henning, K. Stange. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):106–107, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401353.

Kaiser:1952:MH

[Kai52] W. Kaiser. Moritz Hecht, 1868–1952. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 20(1):93, 1952. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401840.

Kappos:1967:BRT

[Kap67] D. A. Kappos. Book review: Tchebycheff Systems: With Applications in Analysis and Statistics by S. J. Karlin, W. J. Studden. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(3):327, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401807.

Kossack:1971:DUS

[KB71] Carl F. Kossack and H. R. Shiledar Baxi. On designing of a unitstratified survey design for a discrete set of observations. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):46–56, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402228.

Kakwani:1970:SEE

[KC70] N. C. Kakwani and R. Court. Simultaneous-equation estimation under linear restrictions with application to new Zealand meat demand. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):244–257, 1970. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402148.

Kellerer:1960:OA

[Kel60] H. Kellerer. Oskar Anderson, 1887–1960. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(1/2):120–122, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401855.

Kemp:1971:BRD

[Kem71] C. D. Kemp. Book review: Distribution in Statistics: Discrete Distributions by N. L. Johnson, S. Kotz. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):243–244, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402186.

Kendall:1961:BRM

[Ken61] D. G. Kendall. Book review: Mathematical Methods in the Theory of Queueing by A. Y. Khintchine, D. M. Andrews, M. H. Quenouille. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(2):85–86, 1961. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401662.

Kendall:1965:BRM

[Ken65] D. G. Kendall. Book review: Method of Statistical Testing. Monte Carlo Method by Ju. A. Shreider. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(2):327–328, 1965. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402042.

Kendall:1966:SIL

[Ken66a] M. G. Kendall. Statistical inference in the light of the theory of the electronic computer. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(1):1– 12, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401529.

Kendrick:1966:CDP

[Ken66b] J. W. Kendrick. Conceptual and data problems of wealth estimates in an economic accounting framework. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(3):360–368, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401255.

Kenessey:1966:SQI

[Ken66c] Z. Kenessey. Some questions of the interpretation of statistics as a science with special regard to official statistics. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):156–164, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401764.

Kenessey:1967:BRI

[Ken67] Z. Kenessey. Book review: Input–Output Economics by W. Leontief. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(3):327–328, 1967. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401808.

Kendall:1969:SHP

[Ken69] M. G. Kendall. Studies in the history of probability and statistics. XXI. The early history of index numbers. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(1):1–12, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402090.

Kendall:1971:BRR

[Ken71] David Kendall. Book review: Random Variables and Probability Distributions by H. Cramér. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):240, 1971. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402181.

Kevork:1967:AFN

[Kev67] C. E. Kevork. An attempt at forecasting the nutritional levels: Energy value and protein content of the average urban diet in Greece. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(1):50–58, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401635.

Keyfitz:1967:IEP

[Key67] N. Keyfitz. The integral equation of population analysis. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(3):213–246, 1967. CODEN ISTRDP.

ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401794.

Kakwani:1967:NBP

[KG67] N. C. Kakwani and D. B. Gupta. Note on the bias of the Prais and Aitchison’s and Fisher’s iterative estimators in regression analysis with heteroscedastic errors. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35 (3):291–295, 1967. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401797.

Khamis:1958:BRP

[Kha58] S. H. Khamis. Book review: Proceedings of the Third Berkeley Symposium on Mathematical Statistics and Probability. Vol.

1. Contributions to the Theory of Statistics by J. Neyman. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):133–134, 1958. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401576.

Khamis:1966:BRI

[Kha66] S. K. Khamis. Book review: Inference and Disputed Authorship: The Federalist by F. Mosteller, D. L. Wallace. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):277–279, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401781.

Kingman:1964:BRS

[Kin64] J. F. C. Kingman. Book review: Studies in Mathematical Analysis and Related Topics. Essays in Honour of George Pólya by G. Szegö. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):325, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401877.

Kingston:1968:GM

[Kin68] J. Kingston. Giorgio Mortara, 1885–1967. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):242–243, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401627.

Kingman:1969:BRC

[Kin69] J. F. C. Kingman. Book review: Convergence of Probability Measures by P. Billingsley. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(3):322, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1402124.

Kitagawa:1962:BRPa

[Kit62a] T. Kitagawa. Book review: Probability: A First Course by F. Mosteller, R. E. K. Rourke, G. B. Thomas, Jr. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(2):277, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401909.

Kitagawa:1962:BRPb

[Kit62b] T. Kitagawa. Book review: Probability with Statistical Applications by F. Mosteller, R. E. K. Rourke, G. B. Thomas, Jr. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(2):277–278, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401910.

Kitagawa:1964:BRR

[Kit64] T. Kitagawa. Book review: Renewal Theory by D. R. Cox. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(1/2):181–182, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401984.

Kitagawa:1968:BRC

[Kit68] T. Kitagawa. Book review: Contributions to Statistics by C. R. Rao. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):233, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401616.

Kagan:1967:BRL

[KLM67] A. M. Kagan, Yu. V. Linnik, and N. M. Mitrofanova. Book review: Linear Statistical Inference and Its Applications by C. R. Rao. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(1):99–100, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401648.

Kock:1963:BRI

[Koc63] Karin Kock. Book review: Industrial Growth and World Trade by A. Maizels. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(3):450–451, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401454.

Koller:1956:CB

[Kol56] A. Koller. Carl Brüschweiler, 1878–1956. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 24(1/3):159–160, 1956. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401294.

Koller:1957:BRE

[Kol57] S. Koller. Book review: Einführung in die mathematische Statistik by L. Schmetterer. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):154– 155, 1957. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401682.

Koller:1960:BRS

[Kol60] S. Koller. Book review: Stichproben aus endlichen Gesamtheiten. Theorie und Technik by G. Menges. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(1/2):105–106, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401850.

Koller:1965:BRS

[Kol65] S. Koller. Book review: Sampling Techniques by W. G. Cochran. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(3):558–559, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401711.

Konus:1958:CPI

[Kon58] A. A. Konüs. Consumer price indexes and demand functions. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):29–36, 1958. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401570.

Krelle:1965:BRI

[Kre65] W. Krelle. Book review: Investment Decisions, Economic Forecasting, and Public Policy by J. R. Meyer, R. R. Glauber. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(2):328–329, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402043.

Krickeberg:1967:BRS

[Kri67] K. Krickeberg. Book review: Stochastic Models for Bacteriophage by J. Gani. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(3):326, 1967. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401806.

Krickeberg:1969:BRI

[Kri69] K. Krickeberg. Book review: An Introduction to Probability Theory and Its Applications; Vol. 1 by W. Feller. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(1):106, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402100.

Krutchkoff:1971:BRE

[Kru71] R. G. Krutchkoff. Book review: Empirical Bayes Methods by J. S. Maritz. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):111, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402244.

Kendall:1953:DFM

[KS53] M. G. Kendall and R. M. Sundrum. Distribution-free methods and order properties. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(3):124–134, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401424.

Kull:1958:LVF

[Kul58] W. Kull. Louis-Vladimir Furlan, 1886–1955. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):169–170, 1958. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401594.

Kulldorff:1963:SPO

[Kul63] G. Kulldorff. Some problems of optimum allocation for sampling on two occasions. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(1):24–57, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401727.

Kunzi:1965:BRO

[Kün65] H. P. Künzi. Book review: Optimum Decision Rules for Government and Industry by H. Theil. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(2):333–334, 1965. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402047.

Kendall:1963:ISP

[KWT^{+}63] M. G. Kendall, P. Wegner, M. E. Terry, I. Padunchewit, E. H. Jebe, and R. R. Perez. An introduction to statistical programming. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(3):307–360, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401444.

L:1936:BRC

[L.36] L. L. Book review: Corso di demografia by F. Savorgnan. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(1):136, April 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401138.

Laininen:1967:BRQ

[Lai67] P. Laininen. Book review: Quality Control and Reliability by N. L. Enrick. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):197–198, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401407.

Lampard:1968:BRR

[Lam68] D. G. Lampard. Book review: Recent Results in Information Theory by S. Kotz. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):110, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401358.

Langelutke:1956:EW

[Lan56] H. Langelütke. Ernst Wagemann, 1884–1956. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 24(1/3):157–159, 1956. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401293.

Lancaster:1962:BRE

[Lan62] H. O. Lancaster. Book review: Elements of Mathematical Statistics by H. W. Alexander. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(2):278–279, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401911.

Lancaster:1963:BRC

[Lan63] H. O. Lancaster. Book review: A Course in the Geometry of n Dimensions by M. G. Kendall. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(2):262, 1963. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401381.

Lancaster:1969:BSB

[Lan69] H. O. Lancaster. A bibliography of statistical bibliographies: A second list. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(1):57–67, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402095.

Lancaster:1970:BSB

[Lan70a] H. O. Lancaster. A bibliography of statistical bibliographies: A third list. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):258–267, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402149.

Lancaster:1970:BRA

[Lan70b] H. O. Lancaster. Book review: The Advanced Theory of Statistics; Volume 1: Distribution Theory by M. G. Kendall, A. Stuart. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):303–304, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402161.

Lancaster:1971:BSB

[Lan71] H. O. Lancaster. A bibliography of statistical bibliographies: a fourth list. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):64–73, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402231.

Laurila:1964:BRF

[Lau64] E. H. Laurila. Book review: The Flow-of-Funds Approach to Social Accounting. Appraisal, Analysis, and Applications. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):328–329, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1401881.

LeRoy:1969:BRP

[Le 69] H. L. Le Roy. Book review: Proceedings of the Fifth Berkeley Symposium on Mathematical Statistics and Probability; Vol. IV by L. Lecam, J. Neyman. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):230–231, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1402306.

Leadbetter:1968:BRS

[Lea68] M. R. Leadbetter. Book review: Spectral Analysis of Time Series by B. Harris. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):357, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401486.

Ledermann:1955:RDC

[Led55] S. Ledermann. La répartition des décès de cause “indéterminée”. (French) [The distribution of deaths from “undetermined” causes]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 23(1/3):47–57, 1955. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402010.

Leppink:1968:BRW

[Lep68] G. J. Leppink. Book review: Wahrscheinlichkeitstheorie und Grundzüge der Masstheorie; Bd. 1 by H. Bauer. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):95–96, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401345.

Lesoir:1935:RP

[Les35] Edm. Lesoir. Les registres de population. (French) [Population registers]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(1):14–39, April 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401047.

Lesoir:1952:AMA

[Les52] E. Lesoir. Avis aux membres / announcements to members. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 20(2):201–202, 1952. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401822.

Leser:1966:DES

[Les66] C. E. V. Leser. Direct estimation of seasonal variation. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(3):369–375, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401256.

Leser:1967:BRS

[Les67] C. E. V. Leser. Book review: Studies in Short-Term National Accounts and Long-Term Economic Growth by S. Goldberg, P. Deane, International Association for Research in Income and Wealth. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):201–202, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401412.

Lindberg:1954:MK

[Lin54] Valter Lindberg. Martti Kovero, 1882–1955. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22(1/3):195, 1954. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401943.

Linder:1956:BRE

[Lin56] A. Linder. Book review: Experimental Design and Its Statistical Basis by D. J. Finney. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 24(1/3): 141, 1956. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401284.

Lindley:1961:BRS

[Lin61a] D. V. Lindley. Book review: Statistical Inference for Markov Processes by P. Billingsley. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(3):91, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401956.

Lindley:1961:BRP

[Lin61b] D. V. Lindley. Book review: The Passage Problem for a Stationary Markov Chain by J. H. B. Kemperman. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(3):90–91, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401955.

Linder:1962:SPS

[Lin62] F. E. Linder. Some principles of statistical organization underlying the United States National Center for Health Statistics. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(1):33–41, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402071.

Lindley:1964:BRG

[Lin64] D. V. Lindley. Book review: General Stochastic Processes in the Theory of Queues by V. E. Beneš. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):325–326, 1964. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401878.

Linder:1966:BRP

[Lin66a] A. Linder. Book review: Principles of Statistics by M. G. Bulmer. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(1):111–112, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401539.

Linsman:1966:BRC

[Lin66b] M. Linsman. Book review: Comprendre et organiser le traitement automatique de l’information by J. Bernard. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(3):446, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401262.

Linder:1967:BRD

[Lin67] A. Linder. Book review: The Design of Experiments by R. A. Fisher. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):199, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401409.

Linder:1969:BRM

[Lin69] A. Linder. Book review: Mathematische Statistik by H. Witting. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):235, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402311.

Linder:1971:BRE

[Lin71a] A. Linder. Book review: Experimental Design; Selected Papers by F. Yates. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):248–249, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402191.

Linder:1971:WW

[Lin71b] A. Linder. Walter Wegmüller, 1910–1970. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):123, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402257.

Livi:1936:SFD

[Liv36] Livio Livi. Sulla fecondità della donna coniugata secondo l’età. (Italian) [On the fertility of married women according to age]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(4):380–387, January 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401216.

Lovasich:1970:SAR

[LNS^{+}70] Jeanne L. Lovasich, Jerzy Neyman, Elizabeth L. Scott, Jerome A. Smith, M. E. Kauffman, Lyle D. Calvin, F. Mosteller, B. C. V. Oddie, and R. Bartoszynski. Statistical aspects of rain stimulation: Problems and prospects. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):155–170, 1970. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402331.

Logan:1962:SS

[Log62] W. P. D. Logan. Satya Swaroop, 1913–1962. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(1):115, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402087.

Lopes:1956:MAT

[Lop56] W. Lopes. Mario Augusto Teixeira de Freitas, 1890–1956. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 24(1/3):153–154, 1956. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401291.

Louwes:1967:CAA

[Lou67] S. L. Louwes. Cost allocation in agricultural surveys. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(3):264–290, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401796.

Loyo:1969:JDB

[Loy69] G. Loyo. Juan de Dios Bojorquez, 1892–1967. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):238, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402313.

Loynes:1970:BRS

[Loy70] R. M. Loynes. Book review: The Single Server Queue by J. W. Cohen. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(3):416–417, 1970. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402210.

Landry:1938:ORP

[LS38] M. A. Landry and C. A. Verrijn Stuart. Observations sur le rapport présenté par M. le Prof. Winkler: “La baisse de la natalité, ses causes, et ses conséquences Économiques et sociales”. (French) [Observations on the report presented by Prof. Winkler: “The decline in the birth rate, its causes, and its economic and social consequences”]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(4):553–563, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400885.

Lellouch:1971:LTE

[LS71] J. Lellouch and D. Schwartz. L’essai therapeutique: éthique individuelle ou éthique collective?. (French) [Therapeutic trial: individual ethics or collective ethics?]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):127–136, 1971. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402169.

Lukacs:1970:CTC

[Luk70] Eugene Lukacs. Characterization theorems for certain stochastic processes. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(3):333–343, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402199.

Lukacs:1971:BRA

[Luk71] E. Lukacs. Book review: Approximations to the Poisson, Binomial and Hypergeometric Distribution Functions by W. Molenaar. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):113, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402246.

Lutjohann:1965:BRS

[Lüt65] H. Lütjohann. Book review: Schätzprobleme in der Ökonometrie by D. Lüdeke. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(2):326–327, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402041.

Macura:1970:BRD

[Mac70] M. Macura. Book review: Demographic Analysis by B. Benjamin. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):296–297, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402155.

Madle:1946:BRA

[Mad46] A. Madlé. Book review: Allgemeine Methodenlehre der Statistik by F. Klezl-Norberg. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 14(1/4):121– 122, 1946. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401150.

Madhava:1956:SAF

[Mad56] K. B. Madhava. Sequential approach in factorial designs. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 24(1/3):64–72, 1956. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401280.

Mahajan:1969:SAU

[Mah69] B. M. Mahajan. On some aspects of unadjusted and adjusted “R^{2}” estimated from Indian consumption data. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(3):294–302, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402119.

Malinvaud:1961:EPM

[Mal61] Edmond Malinvaud. Estimation et prévision dans les modèles economiques autoregressifs. (French) [Estimation and forecasting in autoregressive economic models]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(2):1–32, 1961. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401655.

Malinvaud:1968:BRS

[Mal68] E. Malinvaud. Book review: Stochastically Dependent Equations: An Introductory Text for Econometricians by P. R. Fisk. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):354, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401482.

Malleus:1969:BRL

[Mal69] Michèle Malléus. Book review: Leçons sur les problèmes de statistique analytique by Yu. V. Linnik. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(1):107–109, 1969. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402103.

Manceau:1968:SSR

[Man68a] J. N. Manceau. Sample survey of the registration areas of deaths and fetal deaths in the Brazilian northeast region. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):43–67, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1401340.

Mantel:1968:BRS

[Man68b] N. Mantel. Book review: Statistics for Biologists by R. C. Campbell. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):231–232, 1968. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401614.

March:1933:PCC

[Mar33] M. Lucien March. Parallélisme, corrélation, causalité. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(1):9–22, 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401081.

Marshall:1951:CTC

[Mar51] H. Marshall. Census taking in Canada. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(3):214–230, 1951. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401995.

Martin:1961:BRM

[Mar61] L. Martin. Book review: Méthodes statistiques en biologie et en agronomie by A. Vessereau. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(1):151, 1961. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401562.

Martel:1964:ESC

[Mar64] L. Martel. Étude statistique de la courbe de lumière de l’étoile variable SS Cygni. (French) [Statistical study of the light curve of the variable star SS Cygni]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(1/2):17–51, 1964. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401972.

Martin:1966:BRS

[Mar66] L. Martin. Book review: Statistical Method in Biological Assay by D. J. Finney. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):272–275, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401778.

Maritz:1968:BRH

[Mar68] J. S. Maritz. Book review: Handbook of the Poisson Distribution by F. A. Haight. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):355, 1968.

CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401484.

Martin:1969:BRP

[Mar69] L. Martin. Book review: Principles of Medical Statistics by Austin Bradford. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):222–225, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402299.

Maritz:1971:BRM

[Mar71] J. S. Maritz. Book review: Measuring Uncertainty: An Elementary Introduction to Bayesian Statistics by S. A. Schmitt. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):115–116, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402250.

Masuyama:1967:CCB

[Mas67] M. Masuyama. Construction cyclique de blocs incomplets partiellement èquilibres. (French) [Cyclic construction of partially balanced incomplete blocks]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):107–124, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401394.

Matern:1962:UIS

[Mat62] B. Matérn. On the use of information on supplementary variables in estimating a distribution. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30 (2):121–135, 1962. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401894.

McCarthy:1963:SOS

[McC63] P. J. McCarthy. Some observations on sampling in the construction of price indexes. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(2):182–194, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401372.

McCarthy:1966:BRS

[McC66] M. D. McCarthy. Book review: Statistics in Theory and Practice, with Special Reference to Published Statistics by L. R. Connor, A. J. H. Morrell. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):271–272, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401776.

McCarthy:1969:PRH

[McC69] P. J. McCarthy. Pseudo-replication: Half samples. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(3):239–264, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1402116.

Mason:1970:SPC

[MCC70] B. J. Mason, Jacques Cocheme, and Lyle D. Calvin. The scientific problems of cloud seeding. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38 (1):149–154, 1970. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402330.

Medin:1965:CYE

[Med65] K. Medin. Crop yield estimation and crop insurance in Sweden. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(3):414–442, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401703.

Meerwarth:1935:BRM

[Mee35] Rudolf Meerwarth. Beiträge zur repräsentativen Methode. (German) [Contributions to the representative method]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(4):363–376, January 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400818.

Meerwarth:1940:BRS

[Mee40] R. Meerwarth. Book review: Die Statistik in Deutschland nach ihrem heutigen Stand. Ehrengabe für Friedrich Zahn by Friedrich Burgdörfer. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 8(3/4):235–237, 1940. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401161.

Menges:1970:BRO

[Men70] G. Menges. Book review: Optimization Techniques in Quantitative Economic Models by J. K. Sengupta, K. A. Fox. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(3):423, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402216.

Methorst:1933:APFa

[Met33a] H. W. Methorst. Avant-propos. (French) [Foreword]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(2):5, July 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401168.

Methorst:1933:APFb

[Met33b] H. W. Methorst. Avant-propos. (French) [Foreword]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(3):5, October 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1400772.

Methorst:1935:APF

[Met35] H. W. Methorst. Avant-propos. (French) [Foreword]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(4):361, January 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400817.

Methorst:1945:APF

[Met45] H. W. Methorst. Avant-propos. (French) [Foreword]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 13(1/4):4, 1945. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1400972.

Mewes:1965:BRA

[Mew65] B. Mewes. Book review: Annuaire statistique 1963. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(3):555–556, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1401708.

Meyer:1941:KKK

[Mey41] Maximilian Meyer. Krebssterbe-, krebskranken-, krebsvorsorgestatistik. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 9(3/4):122–132, 1941. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400874.

Meyer:1942:CML

[Mey42] Maximilian Meyer. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 10 (3/4):184–200, 1942. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400897.

Mayer:1950:CDE

[MH50] Albert J. Mayer and Philip M. Hauser. Class differentials in expectation of life at birth. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18 (3/4):197–200, 1950. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401037.

Michalopoulos:1938:ORP

[Mic38] M. Jean Michalopoulos. Observations sur le rapport présenté par M. A. Molinari: “Statistiques du tourisme international”. (French) [Observations on the report presented by Mr. A. Molinari: “Statistics of international tourism”]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(4):564–565, 1938. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400886.

Mihoc:1969:BRM

[Mih69] Gh. Mihoc. Book review: Markov Processes and Potential Theory by J. Chover. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):227–228, 1969.

CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402302.

Minaru:1969:BRI

[Min69] L. Mináru. Book review: Inventory Control Abstracts, 1953–1965 by S. Eilon, W. Lampkin. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37 (2):229–230, 1969. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402304.

Mitra:1971:BRE

[Mit71] S. K. Mitra. Book review: Elements of Continuous Multivariate Analysis by A. P. Dempster. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39 (1):107–108, 1971. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402239.

Matheron:1970:SAG

[MK70] G. Mathéron and W. C. Krumbein. Structures aléatoires et gèologie mathèmatique. (French) [Random structures and mathematical geology]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):1–11, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402318.

Martin-Lof:1969:AR

[ML69] P. Martin-Löf. Algorithms and randomness. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(3):265–272, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402117.

Mokken:1962:BRS

[Mok62] R. J. Mokken. Book review: Studies in Item Analysis and Prediction by H. Solomon. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(1):86– 87, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1402076.

Molinari:1936:OPP

[Mol36] A. Molinari. Observations et propositions pour une statistique internationale du tourisme. (French) [Observations and proposals for international tourism statistics]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4 (2):243–259, July 1936. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400728.

Mortara:1934:SDS

[Mor34] Giorgio Mortara. Sulle disuguaglianze statistiche. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(1):72–77, April 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401015.

Morgenroth:1935:CML

[Mor35a] W. Morgenroth. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2 (4):389–408, January 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400820.

Mortara:1935:SSD

[Mor35b] Giorgio Mortara. Sulle statistiche degli impianti elettrici. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(4):377–388, January 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400819.

Mortara:1941:QPR

[Mor41] Giorgio Mortara. Quelques possibilités de reconstruction du mouvement d’une population à l’aide des données des recensements démographiques. (French) [Some possibilities for reconstructing the movement of a population using demographic census data]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 9(1/2):21–35, 1941. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400942.

Moran:1966:BRS

[Mor66] P. A. P. Moran. Book review: Studies in the Theory of Random Processes by A. V. Skorohod. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34 (2):281–282, 1966. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401784.

Morice:1969:BRT

[Mor69] E. Morice. Book review: Traité de fiabilité by M. Schwob, G. Perrache. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(3):328, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402131.

Morice:1971:BRTc

[Mor71a] E. Morice. Book review: Tables commentées des épreuves séquentielles exhaustive et non exhaustive by M. Dumas. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(3):405–406, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402279.

Morice:1971:BRTa

[Mor71b] E. Morice. Book review: Théorie et méthodes statistiques; Vol. 1 by P. Dagnélie. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):105–106, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1402236.

Morice:1971:BRTb

[Mor71c] E. Morice. Book review: Théorie et méthodes statistiques; Vol. 2: Les méthodes de l’inférence statistique by P. Dagnélie. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):106–107, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402237.

Morucci:1971:BRA

[Mor71d] B. Morucci. Book review: Advanced Theory of Characteristic Functions by B. Ramachandran. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):114–115, 1971. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402249.

Moser:1951:IB

[Mos51] C. A. Moser. Interview bias. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(1):28–40, 1951. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401500.

Moser:1954:BRT

[Mos54] C. A. Moser. Book review: Theorie und Technik des Stichprobenverfahrens by H. Kellerer. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22 (1/3):173, 1954. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401928.

Mosteller:1963:CCT

[Mos63] F. Mosteller. Continental Classroom’s television course in probability and statistics. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(2):153–162, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401370.

Mosteller:1971:BRT

[Mos71a] F. Mosteller. Book review: The Teaching of Probability and Statistics by L. Råde. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(3):407–408, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1402282.

Mosteller:1971:J

[Mos71b] Frederick Mosteller. The jackknife. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(3):363–368, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402273.

Mosteller:1971:JAS

[Mos71c] Frederick Mosteller. The Joint American Statistical Association– National Council of Teachers of Mathematics Committee on the Curriculum in Statistics and Probability. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(3):340–342, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402267.

Makinson:1971:STR

[MPDoA71] G. J. Makinson, G. P. Patil, Norman R. Draper, and The Catholic University of America. Statistical training and research / formation et recherche statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(3):399–400, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402277.

Mehta:1970:FSD

[MS70] J. S. Mehta and P. A. V. B. Swamy. The finite sample distribution of Theil’s mixed regression estimator and a related problem. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):202–209, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402143.

Moran:1970:MRM

[MSC^{+}70] P. A. P. Moran, M. Schüepp, Lyle D. Calvin, W. A. O’N. Waugh, A. C. Atkinson, and A. M. Walker. The methodology of rainmaking experiments. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):105– 119, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1402327.

Methorst:1934:RRC

[MT34] H. W. Methorst and J. Tinbergen. Les recherches relatives à la conjoncture au Bureau Central de Statistique des Pays-Bas. (French) [Research relating to conjuncture at the Central Statistical Office of the Netherlands]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(1):37–55, April 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401013.

Muhsam:1965:BRH

[Muh65] H. V. Muhsam. Book review: Human Capital. A Theoretical and Empirical Analysis, with Special Reference to Education by G. S. Becker. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(3):556–557, 1965. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401709.

Muhsam:1969:BRH

[Muh69] H. V. Muhsam. Book review: Health and Vital Statistics by B. Benjamin. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):221–222, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402298.

Muilwijk:1965:BRE

[Mui65] J. Muilwijk. Book review: Elementi di calcolo delle probabilità by A. Longo. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(2):326, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402040.

Myburgh:1963:JRH

[Myb63] C. A. L. Myburgh. J. R. H. Shaul, 1897–1963. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(3):455–456, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401458.

Naddeo:1968:CIF

[Nad68] A. Naddeo. Confidence intervals for the frequency function and the cumulative frequency function of a sample drawn from a discrete random variable. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):313–318, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401472.

Nair:1967:BRP

[Nai67] K. R. Nair. Book review: Proceedings of the Second Conference of the International Advisory Group of Forest Statisticians, Stockholm 1965 by B. Matérn. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(3):328– 329, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401809.

Nair:1971:BRF

[Nai71] K. R. Nair. Book review: Forest Biometrics by M. Prodan. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):247, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402189.

Nelson:1970:BRS

[Nel70] L. S. Nelson. Book review: Statistics for Experimentalists by B. E. Cooper. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(3):417–418, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402211.

Neuts:1967:BRG

[Neu67] M. F. Neuts. Book review: Green’s Function Methods in Probability Theory by J. Keilson. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(1):96, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401646.

Neyman:1957:IBB

[Ney57] J. Neyman. “Inductive behavior” as a basic concept of philosophy of science. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):7–22, 1957. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401671.

Neyman:1962:TBT

[Ney62] J. Neyman. Two breakthroughs in the theory of statistical decision making. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(1):11–27, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402069.

Neyman:1970:P

[Ney70] J. Neyman. Preface. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):??, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402317.

Nicholson:1954:SLC

[Nic54] J. L. Nicholson. A survey of the living conditions of Arab families in Tripolitania in November-December 1950. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22(1/3):68–84, 1954. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401921.

Nixon:1935:SCH

[Nix35] J. W. Nixon. The size, constitution and housing standards of the family in England and Wales, 1911 to 1931. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(2):142–162, July 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1400853.

Nixon:1936:CFB

[Nix36] J. W. Nixon. Committee on Family Budgets. Preliminary report. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(1):45–61, April 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401132.

Nixon:1938:SAC

[Nix38] J. W. Nixon. On the statistics available concerning the occupied population of the world and its distribution. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(3):377–393, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401192.

Nixon:1952:SCH

[Nix52] J. W. Nixon. On the size, constitution and housing standards of households in England and Wales, 1931 and 1951. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 20(2):135–147, 1952. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1401818.

Nixon:1953:BRW

[Nix53] J. W. Nixon. Book review: Wages and Salaries in the United Kingdom, 1920–1938 by A. L. Chapman. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(3):196, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401428.

Nixon:1954:BRB

[Nix54] J. W. Nixon. Book review: British Post-War Migration. National Institute of Economic and Social Research by J. Isaac. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22(1/3):174–175, 1954. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401930.

Nixon:1955:LH

[Nix55] J. W. Nixon. Liebmann Hersch, 1882–1955. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 23(1/3):144, 1955. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402024.

Nixon:1957:BRI

[Nix57] J. W. Nixon. Book review: Der internationale Vergleich der Lebenshaltungskosten; eine methodologische Untersuchung dargestellt an Deutschland–Frankreich by A. Jacobs. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):152, 1957. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401678.

Nixon:1958:BRT

[Nix58a] J. W. Nixon. Book review: Trends in Employment in the Service Industries, 1870–1950 by G. J. Stigler. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):140, 1958. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401585.

Nixon:1958:CFR

[Nix58b] J. W. Nixon. Charles Frederick Roos, 1901–1958. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):171, 1958. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1401595.

Nixon:1961:BRH

[Nix61a] J. W. Nixon. Book review: Historical Statistics of the United States; Colonial Times to 1957. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(1):147–148, 1961. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401557.

Nixon:1961:HS

[Nix61b] J. W. Nixon. Hans Staehle, 1902–1961. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(2):94–95, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401667.

Nixon:1961:SDC

[Nix61c] J. W. Nixon. Some demographic characteristics of Protestants and Catholics in Switzerland. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(3):13–28, 1961. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401947.

Nixon:1962:AL

[Nix62a] J. W. Nixon. Alexander Loveday, 1888–1962. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(1):114, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402086.

Nixon:1962:BRC

[Nix62b] J. W. Nixon. Book review: Canadian Political Science Association: Conference on Statistics 1960 by E. J. Beach, J. C. Welldon. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(3):401–403, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401750.

Nixon:1964:BRC

[Nix64] J. W. Nixon. Book review: Conference on Statistics 1961 by Canadian Political Science Association. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):329–330, 1964. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401882.

Nixon:1965:BRC

[Nix65] J. W. Nixon. Book review: Conferences on Statistics 1962 and 1963. Canadian Political Science Association by J. Henripin, A. Asimakopulos. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(3):559–560, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401712.

Nixon:1967:BRC

[Nix67] J. W. Nixon. Book review: Conference on Statistics 1964 by Canadian Political Science Association, S. Ostry, T. K. Rymes. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(3):321–322, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401802.

Norris:1966:BRM

[Nor66] K. P. Norris. Book review: Management Control Manual for the Textile Industry by N. L. Enrick. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):272, 1966. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401777.

Nordbotten:1967:BRO

[Nor67] S. Nordbotten. Book review: OMNITAB, a Computer Program for Statistical and Numerical Analysis by J. Hilsenrath, G. G. Ziegler, C. G. Messina, P. J. Walsh, R. J. Herbold. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):203–204, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401414.

Nordbotten:1970:IDF

[Nor70] Svein Nordbotten. Individual data files and their utilization in socio-demographic model building in the Norwegian central bureau of statistics. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):193–201, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1402142.

Neyman:1969:SM

[NSW69] J. Neyman, E. L. Scott, and M. A. Wells. Statistics in meteorology. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):119–148, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402287.

Noda:1970:BME

[NT70] Kazuo Noda and Yasushi Taga. Bayes and minimax estimation methods for the optimum decomposition of a sample space based on prior information. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(3):344– 350, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1402200.

Neudecker:1966:NAS

[NvdP66] H. Neudecker and C. van de Panne. Note on the asymptotic standard errors of latent roots of econometric equation systems. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(1):43–47, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401532.

Naylor:1968:SME

[NWW68] Th. H. Naylor, K. Wertz, and Th. Wonnacott. Some methods for evaluating the effects of economic policies using simulation experiments. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):184–200, 1968. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401606.

Nystrom:1936:OPI

[Nys36] Bertil Nyström. Observations on the possibility of improving the international comparability of building and housing statistics. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(1):71–85, April 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401134.

Nystrom:1938:CBH

[Nys38] Bertil Nyström. Commission on Building and Housing Statistics. Preliminary report. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(2):251–263, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1400743.

Oakland:1971:BRF

[Oak71] G. B. Oakland. Book review: Fault Diagnosis of Digital Systems by H. Y. Chang, E. G. Manning, G. Metze. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):238, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402179.

Office:1961:UNS

[Off61] U. N. Statistical Office. United Nations Statistical Office publications recently issued. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(2):86– 88, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401663.

Ohlsson:1961:BRS

[Ohl61] P. T. J. Ohlsson. Book review: Studies in Income and Wealth, Vol. 25. Output, Input, and Productivity Measurement. Papers Presented at the 1958 Meeting of the Conference on Research in Income and Wealth. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(3):95–96, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401960.

Omaboe:1967:BRF

[Oma67] E. N. Omaboe. Book review: Feeding India’s Growing Millions by P. V. Sukhatme. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(1):101–102, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401650.

Onicescu:1936:DSN

[Oni36] Octav Onicescu. La dépendance statistique, la notion de chaîne et ses applications. (French) [Statistical dependence, the notion of chain and its applications]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(3):349–354, October 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400803.

Osborn:1966:MIO

[Osb66] E. Osborn. A modified input–output matrix for Northern Rhodesia. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(1):48–63, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401533.

Prais:1954:GOR

[PA54] S. J. Prais and J. Aitchison. The grouping of observations in regression analysis. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22(1/3):1–22, 1954. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401916.

Paillat:1958:BRD

[Pai58] P. Paillat. Book review: The Demography of Ceylon by N. K. Sarkar. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):139–140, 1958. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401584.

Panse:1963:BRI

[Pan63] V. G. Panse. Book review: An Introduction to Statistical Science in Agriculture by D. J. Finney. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(1):141–143, 1963. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401736.

Pardubsky:1962:AIS

[Par62] B. Pardubský. Application de l’intégrale stochastique à la mesure de la rugosité de surface. (French) [Application of the stochastic integral to the measurement of surface roughness]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(3):352–360, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401744.

Parenti:1966:BRS

[Par66] G. Parenti. Book review: Statistical Methods by S. Szulc. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):282–283, 1966. CODEN ISTRDP.

ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401785.

Parzen:1967:RSA

[Par67] E. Parzen. The role of spectral analysis in time series analysis. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):125–141, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401395.

Parenti:1970:LL

[Par70] G. Parenti. Livio Livi, 1891–1969. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):190–191, 1970. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402342.

Patil:1970:STR

[PBW70] G. P. Patil, G. E. P. Box, and M-L Widén. Statistical training and research / formation et recherche statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(3):402–408, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1402205.

Pearson:1962:BRH

[Pea62] E. S. Pearson. Book review: Handliche Sammlung mathematischstatistischer Tafeln by A. Linder. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(3):403, 1962. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401751.

Pearson:1971:DH

[Pea71] E. S. Pearson. David Heron, 1881–1969. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):122–123, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402256.

Pedroni:1969:BRG

[Ped69] F. Pedroni. Book review: Graphical Rational Patterns; A New Approach to Graphical Presentation of Statistics by R. Bachi. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):221, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402297.

Permanent:1941:BRS

[Per41] Office Permanent. Book review: Statistical Activities of the American Nations, 1940. A Compendium of the Statistical Services and Activities in 22 Nations of the Western Hemisphere, Together with Information concerning Statistical Personnel in These Nations by E. Phelps. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 9(3/4):167–169, 1941. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400879.

Perez:1971:BRM

[Pér71] A. Vegas Pérez. Book review: Mathematics and Statistics for Economists by G. Tintner, C. B. Millham. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):118–119, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402253.

Peter:1966:BRA

[Pét66] G. Péter. Book review: Zur Anwendung statistischer Methoden. Festschrift zum 75 by Felix Burkhardt. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):283–284, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401787.

Peyrot:1961:BRT

[Pey61] F. Peyrot. Book review: Tables of the Hypergeometric Probability Distribution by G. J. Lieberman, D. B. Owen. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(3):88, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401952.

Peyrot:1962:BRH

[Pey62] F. Peyrot. Book review: Handbook of Statistical Tables by D. B. Owen. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(3):404–405, 1962. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401753.

Peyrot:1963:BRT

[Pey63] F. Peyrot. Book review: Tables of Random Permutations by L. E. Moses, R. V. Oakford. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(2):261, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401380.

Pfanzagl:1968:BRB

[Pfa68] J. Pfanzagl. Book review: Basic Concepts of Statistical Inference; Laws and Chance Variations by E. Sverdrup. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):239, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401624.

Pfouts:1966:AAI

[Pfo66] R. W. Pfouts. An axiomatic approach to index numbers. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):174–185, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401766.

Pietra:1937:CDC

[Pie37] Gaetano Pietra. De la concordance et du coefficient de correlation. (French) [Concordance and the correlation coefficient]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(4):500–515, January 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401058.

Pietra:1940:SMI

[Pie40] Gaetano Pietra. Sur la statistique methodologique italienne. (French) [On Italian methodological statistics]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 8(3/4):163–194, 1940. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401158.

Pietra:1944:SMI

[Pie44] Gaetano Pietra. La statistique méthodologique italienne de 1939 à 1942. (French) [Italian methodological statistics from 1939 to 1942]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 12(1/4):36–45, 1944. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401070.

Platzer:1934:SFG

[Pla34] H. Platzer. Zur Statistik der Fertigwarenpreise. (German) [To the statistics of finished goods prices]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(4):31–37, January 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401098.

Pleasants:1967:BRE

[Ple67] P. A. B. Pleasants. Book review: Elementary Methods in the Analytic Theory of Numbers by A. O. Gel’fond, Yu. V. Linnik. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):201, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401411.

Prais:1953:NHE

[Pra53] S. J. Prais. A note on heteroscedastic errors in regression analysis. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(1/2):28–29, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401514.

Pratschke:1971:AUR

[Pra71] John L. Pratschke. Adjusted and unadjusted “R^{2}”: Further evidence from Irish data. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):4–8, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1402224.

Pressat:1968:BRM

[Pre68] R. Pressat. Book review: Les mesures de la natalité; quelques applications à la Belgique by G. Wunsch. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):114, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401364.

Pribram:1936:HAD

[Pri36] Karl Pribram. How to ascertain the definition of some notions which are fundamental to business cycle analysis? Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(2):212–231, July 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400725.

Pribram:1960:EDD

[Pri60] K. Pribram. Edward Dana Durand, 1871–1960. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(1/2):118–120, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401854.

Ptoukha:1934:MPC

[Pto34] Michel Ptoukha. Méthodes pour calculer les taux de mortalité infantile selon les mois de l’année. (French) [Methods for calculating infant mortality rates by month of the year]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(2):178–189, July 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1400996.

Puri:1965:CIT

[Pur65] M. L. Puri. On the combination of independent two sample tests of a general class. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(2):229–241, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1402029.

Puri:1969:VET

[Pur69] M. L. Puri. The van Elteren W test and non-null hypothesis. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):166–175, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402289.

Puri:1970:STR

[Pur70] M. L. Puri. Statistical training and research / formation et recherche statistiques. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):177– 180, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1402333.

Putter:1951:MDE

[Put51] Joseph Putter. Sur une méthode de double échantillonnage pour estimer la moyenne d’une population Laplacienne stratifiée. (French) [On a double sampling method for estimating the mean of a stratified Laplacian population]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19 (3):231–238, 1951. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401996.

Quade:1971:DMT

[Qua71a] Dana Quade. Development of materials and techniques for the instructional use of computers in statistics courses (a project supported by the national science foundation). Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(3):361–362, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402272.

Quade:1971:UCM

[Qua71b] Dana Quade. On using a conversational mode computer in an intermediate statistical analysis course. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(3):343–345, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402268.

Quenouille:1951:VDM

[Que51] M. H. Quenouille. The variate-difference method in theory and practice. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(2):121–129, 1951. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401239.

Qvist:1966:BRM

[Qvi66] B. Qvist. Book review: Monte Carlo Methods by J. M. Hammersley, D. C. Handscomb. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(1):113– 114, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401542.

Rahts:1933:SSG

[Rah33] J. Rahts. Ueber den Schluss der Sterbetafeln. (German) [On the end of the death tables]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1 (2):19–29, July 1933. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401170.

Ramsey:1970:NSE

[Ram70] J. B. Ramsey. A note on single equation estimators in simultaneous equation systems. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):236–243, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1402147. See cor-

rection [Ano70c].

Rao:1958:BRM

[Rao58] C. R. Rao. Book review: Multivariate Correlational Analysis by Ph. H. Dubois. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):141, 1958. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401587.

Rao:1961:BRC

[Rao61] C. R. Rao. Book review: Contributions to Probability and Statistics. Essays in Honor of Harold Hotelling by I. Olkin, S. G. Ghurye, W. Hoeffding, W. G. Madow, H. B. Mann. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(2):83–85, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401661.

Rao:1963:FPP

[Rao63a] C. R. Rao. Felicitations to Professor P. C. Mahalanobis on his seventieth birthday. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(1):1, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401724.

Rao:1963:RTF

[Rao63b] C. R. Rao. Report of the Treasurer for the fiscal years 1961 and 1962. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(3):478–486, 1963. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401462.

Rao:1966:CSR

[Rao66] J. N. K. Rao. On the comparison of sampling with and without replacement. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):125–138, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401762.

Rao:1968:BRA

[Rao68] C. R. Rao. Book review: The Advanced Theory of Statistics; Vol. 3: Design and Analysis, and Time Series by M. G. Kendall, A. Stuart. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):109, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401356.

Rao:1969:BRE

[Rao69] J. N. K. Rao. Book review: Exercises in Probability and Statistics by N. A. Rahman. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(3):326–327, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1402129.

Renyi:1960:CJ

[Rén60] A. Rényi. Charles Jordan, 1871–1959. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(1/2):117–118, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401853.

Renyi:1961:BRI

[Rén61] A. Rényi. Book review: Introduction to Probability and Statistics by L. Alder, E. B. Roessler. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(1):150, 1961. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401561.

Renyi:1963:BRG

[Rén63] A. Rényi. Book review: Geometrical Probability by M. G. Kendall, P. A. P. Moran. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(3):449–450, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401453.

Renyi:1965:FIT

[Rén65] A. Rényi. On the foundations of information theory. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(1):1–14, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401301.

Renyi:1967:BRI

[Rén67] A. Rényi. Book review: Introduction to Probability and Statistics from a Bayesian Viewpoint, 1. Probability 2. Inference by D. V. Lindley. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(1):96–99, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401647.

Revesz:1966:BRE

[Rév66] P. Révész. Book review: Ergodic Theory and Information by P. Billingsley. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(3):446–447, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401263.

Revesz:1970:BRS

[Rév70] P. Révész. Book review: Stochastic Approximation by M. T. Wasan. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):188–189, 1970. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402340.

Rice:1936:CML

[Ric36] Stuart A. Rice. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(4):427–447, January 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401221.

Rice:1945:CML

[Ric45a] Stuart A. Rice. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 13(1/4):62–93, 1945. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1400977.

Rice:1945:II

[Ric45b] Stuart A. Rice. ISI, IASI and UN. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 13(1/4):5–8, 1945. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400973.

Rice:1947:AMA

[Ric47a] Stuart A. Rice. Avis aux membres / announcements to members. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 15(1/4):137–169, 1947. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401212.

Rice:1947:TFS

[Ric47b] Stuart A. Rice. The Twenty-Fifth Session. Washington, D. C., September 6–18, 1947. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 15(1/4): 1–7, 1947. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401204.

Rice:1949:TSS

[Ric49] Stuart A. Rice. The Twenty-Sixth Session. Berne, September 5–10, 1949. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 17(3/4):113–118, 1949. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400862.

Rice:1951:TSS

[Ric51] Stuart A. Rice. The Twenty-Seventh Session. New Delhi, December 5–11; Calcutta, December 16–18, 1951. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(3):209–213, 1951. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401994.

Rice:1953:TES

[Ric53] Stuart A. Rice. The Twenty-Eighth Session. Rome, September 6–12, 1953. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(3):119–123, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401423.

Richter:1956:AVV

[Ric56] H. Richter. Average value and volume indices of imports by areas of origin and by commodity categories of metropolitan OEEC member countries combined. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 24(1/3): 27–51, 1956. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401278.

Rice:1959:IAS

[Ric59a] Stuart A. Rice. The inter American statistical institute at age nineteen. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 27(1/3):1–7, 1959. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402054.

Rice:1959:WFO

[Ric59b] Stuart A. Rice. William Fielding Ogburn, 1886–1959. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 27(1/3):96–97, 1959. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1402064.

Rice:1961:WFW

[Ric61] Stuart A. Rice. Walter F. Willcox. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(1):1–3, 1961. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401551.

Rice:1964:WFW

[Ric64] Stuart A. Rice. Walter Francis Willcox March 28, 1861–October 30, 1964. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):340–346, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401885.

Rice:1966:SCI

[Ric66] S. A. Rice. The second crisis of the international statistical institute. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(3):297–308, 1966. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401251.

Rice:1968:TOW

[Ric68] S. A. Rice. Toward “One world in statistics”: Dips and gaps in an upward trend. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):245–251, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401466.

Rider:1964:BRS

[Rid64] P. R. Rider. Book review: The Statistical Basis of Acceptance Sampling, An Introduction for Industrial Executives by S. K. Ekambaram. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):324, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401875.

Rios:1965:BRM

[Rio65] Sixto Rios. Book review: Multistage Inventory Models and Techniques by Herbert E. Scarf, Dorothy M. Gilford, Maynard W. Shelley. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(2):329–331, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402044.

Rivet:1953:ML

[Riv53] R. Rivet. Max Lazard, 1875–1953. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(3):210, 1953. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401438.

Rjabushkin:1965:VSN

[Rja65] T. V. Rjabushkin. Vasily Sergejevitch Nemtchinov, 1894–1964. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(2):344–345, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402050.

Read:1970:GAI

[RMK^{+}70a] W. A. Read, Matthias Matschinski, D. G. Kendall, F. Chayes, D. M. Hawkins, E. L. Scott, E. H. Timothy Whitten, and Graham Lea. Geological applications of iterative-fit trend-surface analysis (abstract). Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):19–20 + 32–34, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1402320.

Robinson:1970:SFG

[RMK^{+}70b] Joseph E. Robinson, Matthias Matschinski, D. G. Kendall, F. Chayes, D. M. Hawkins, W. A. Read, E. L. Scott, E. H. Timothy Whitten, and Graham Lea. Spatial filtering of geological data. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):21–34, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402321.

Robbins:1963:EBA

[Rob63] H. Robbins. The empirical Bayes approach to testing statistical hypotheses. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(2):195–208, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401373.

Romaniuk:1963:ESS

[Rom63] K. Romaniuk. Edward Szturm de Sztrem, 1885–1962. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(3):454–455, 1963. CODEN ISTRDP.

ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401457.

Romaniuk:1966:OL

[Rom66] K. Romaniuk. Oskar Lange, 1904–1965. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):294, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401790.

Roppert:1970:BRI

[Rop70] J. Roppert. Book review: Introduction to Probability Theory and Statistical Inference by H. J. Larson. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):304–305, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402162.

Roth:1963:ETC

[Rot63] G. J. Roth. The equilibrium of traffic on congested streets. An economic approach. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(3):386–396, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401446.

Roy:1964:FD

[Roy64] René Roy. François Divisia, 1889–19964. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):348–351, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401888.

Roy:1969:BRC

[Roy69] J. Roy. Book review: Computer Programming and Related Mathematics by R. V. Andree. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):220, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1402296.

Rosenblatt-Roth:1968:BRM

[RR68] M. Rosenblatt-Roth. Book review: Martingales et processus de Markov by G. A. Hunt. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):234– 235, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401618.

Rao:1965:FIT

[RRK65] C. R. Rao, A. Rényi, and D. G. Kendall. [On the foundations of information theory]: Discussion. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33 (1):14, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401302.

Roy:1968:ITE

[RS68] S. N. Roy and J. N. Srivastava. Inference on treatment effects in incomplete block designs. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):1–6, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401335.

Rueff:1954:CR

[Rue54] J. L. Rueff. Charles Rist, 1874–1955. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22(1/3):193–194, 1954. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401941.

Rutledge:1965:BRT

[Rut65] R. W. Rutledge. Book review: Tables for Normal Sampling with Unknown Variances: The Student Distribution and Economically Optimum Sampling Plans by J. Bracken, A. Schleifer, Jr. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(3):557–558, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401710.

S:1934:BRA

[S.34] F. S. Book review: 50 Années de l’Institut International de Statistique by Friedrich Zahn. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(2):232, July 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401001.

Saenger:1935:MIG

[Sae35] Konrad Saenger. Über die möglichkeit einer internationalen gemüsestatistik. (German) [On the possibility of international vegetable statistics]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(1):40–59, April 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401048.

Saenger:1936:BRB

[Sae36] K. Saenger. Book review: Beiträge zur deutschen Statistik. Festgabe für Franz Zizek zur 60. Wiederkehr seines Geburtstages by Paul Flaskämper, Adolf Blind. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(3):432–435, October 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400810.

Sagoroff:1958:BRA

[Sag58] S. Sagoroff. Book review: Allgemeine Statistik by O. Most. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):136–137, 1958. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401580.

Sarndal:1970:CEI

[Sär70] Carl-Erik Särndal. A class of explicata for “Information” and “Weight of evidence”. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):223– 235, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1402146.

Sastry:1971:BRM

[Sas71] N. S. R. Sastry. Book review: Monetary Statistics of the United States by M. Friedman, A. J. Schwartz. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):242–243, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402185.

Sauvy:1948:MDC

[Sau48] Alfred Sauvy. Sur la mesure du coût de la vie et du pouvoir d’achat. (French) [On measuring the cost of living and purchasing power]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 16(1/4):36–44, 1948. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401231.

Sauvy:1950:OAR

[Sau50] A. Sauvy. [Observations au rapport de M. Sauvy sur les développements récents des études de statistique démographique]: Réponse. (French) [Observations to the report of Mr. Sauvy on recent developments in demographic statistics studies]: Response. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18(1/2):77, 1950. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400764.

Saunders:1963:BRR

[Sau63a] C. T. Saunders. Book review: The Rate and Direction of Inventive Activity. A Conference Report. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(2):260–261, 1963. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401379.

Sauvy:1963:LSR

[Sau63b] A. Sauvy. L’estimation sans rigueur. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(1):2–10, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401725.

Sauvy:1968:MDN

[Sau68] A. Sauvy. Masculinité des derniers-nés d’une famille et distribution de familles selon les sexes. (French) [Masculinity of the last-born children of a family and distribution of families according to sexes]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):252–259, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401467.

Sauvy:1971:GC

[Sau71] A. Sauvy. Gabriel Chevry, 1907–1970. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):121–122, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402255.

Savorgnan:1933:SDN

[Sav33] Franco Savorgnan. La statistica delle nascite secondo l’ordine di generazione. (Italian) [Statistics of births according to the order of generation]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(1):23–38, 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401082.

Savorgnan:1936:BRI

[Sav36] F. Savorgnan. Book review: Istituto Centrale di Statistica. Decennale 1926–1936. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(3):431–432, October 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400809.

Savorgnan:1941:CSG

[Sav41] F. Savorgnan. Comment s’éteint une génération. (French) [How a generation dies]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 9(3/4):93–103, 1941. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1400871.

Savage:1964:CTR

[Sav64] I. Richard Savage. Contributions to the theory of rank order statistics: Applications of lattice theory. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(1/2):52–64, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401973.

Saxer:1966:BRT

[Sax66] W. Saxer. Book review: Théorie des graphes et structures sociales by C. Flament. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):275– 276, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401779.

Schafer:1936:DPE

[Sch36] E. Schäfer. Directives pour l’élaboration des statistiques criminelles dans les divers pays. Rapport au nom de la commission mixte constituée par l’Institut International de Statistique et la Commission Internationale Pénale et Pénitentiaire. (French) [Guidelines for the compilation of criminal statistics in the various countries. Report on behalf of the Joint Commission constituted by the International Statistical Institute and the International Penal and Prison Commission]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(2):195–205, July 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400723.

Schmetterer:1959:BRI

[Sch59] L. Schmetterer. Book review: An Introduction to Combinatorial Analysis by J. Riordan. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 27(1/3): 71, 1959. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1402058.

Schneider:1965:BRM

[Sch65] E. Schneider. Book review: Méthodes statistiques de l’économétrie by E. Malinvaud. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(3):562–563, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401717.

Schubnell:1966:BRN

[Sch66] H. Schubnell. Book review: New Estimates of Fertility and Population in the United States by A. J. Coale, M. Zelnik. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(3):447, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401264.

Schmetterer:1967:BRT

[Sch67] L. Schmetterer. Book review: The Theory of Stochastic Processes by D. R. Cox, H. D. Miller. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):195, 1967. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401404.

Schmutz:1968:BRG

[Sch68a] B. Schmutz. Book reviews: Grundriss der Industriestatistik, I by G. Forbrig; Die Industriestatistik der Bundesrepublik Deutschland by K. Werner. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):105–106, 1968.

CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401352.

Schweitzer:1968:BRM

[Sch68b] T. T. Schweitzer. Book review: The Measurement of Consumers’ Expenditure and Behaviour in the United Kingdom, 1920–1938; Vol. II by R. Stone, D. A. Rowe. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):238–239, 1968. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401623.

Schattan:1971:BRE

[Sch71a] S. Schattan. Book review: Economic Forecasts and Expectations— Analysis of Forecasting Behavior and Performance by J. Mincer. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):245–247, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402188.

Schwartz:1971:OCA

[Sch71b] D. Schwartz. Operation of the centre for applied statistics in medicine and medical biology. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(3):346–347, 1971. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402269.

Scott:1967:SDM

[Sco67a] C. Scott. Sampling for demographic and morbidity surveys in Africa. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):154–171, 1967. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401397.

Scott:1967:BRE

[Sco67b] E. L. Scott. Book review: The Elements of Stochastic Processes with Applications to the Natural Sciences by N. T. J. Bailey. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(1):92, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401640.

Sukhatme:1970:NAC

[SEM^{+}70] P. V. Sukhatme, N. Erus, J. M. P. Memoria, Colin Clark, and J. M. Hammersley. Need of an assured and controlled supply of water for improving agricultural production. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):120–139, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402328.

Sen:1971:RSP

[Sen71] Pranab Kumar Sen. Robust statistical procedures in problems of linear regression with special reference to quantitative bio-assays, I. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):21–38, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402226.

Shaul:1956:INF

[Sha56] J. R. H. Shaul. An index number of factor loadings derived from factor analysis of attributes. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 24 (1/3):11–16, 1956. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401276.

Shaul:1963:BRA

[Sha63] J. R. H. Shaul. Book review: African Studies in Income and Wealth. International Association for Research in Income and Wealth by L. H. Samuels. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31 (2):263–264, 1963. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401384.

Sheps:1971:RMP

[She71] Mindel C. Sheps. A review of models for population change. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):185–196, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402173.

Shirras:1937:INI

[Shi37] G. Findlay Shirras. India’s national income. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(4):467–483, January 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401056.

Shrikhande:1965:BRM

[Shr65] S. S. Shrikhande. Book review: The Method of Paired Comparisons by H. A. David. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(1):173, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401312.

Siddiqui:1964:DSS

[Sid64] M. M. Siddiqui. Distribution of Student’s t in samples from a rectangular universe. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):242–250, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401869.

Silberman:1954:SSP

[Sil54] L. Silberman. The social survey of Port Louis (Mauritius): a survey in an underdeveloped territory. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22(1/3):85–94, 1954. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401922.

Sillitto:1964:NTT

[Sil64] G. P. Sillitto. Note on Takeuchi’s table of difference sets generating balanced incomplete block designs. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):251, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401870.

Schwartz:1961:TMP

[SL61] D. Schwartz and Ph. Lazar. Taux de mortalité par une cause donnée de décès en tenant compte des autres causes de décès ou de disparition. (French) [Mortality rate from a given cause of death taking into account other causes of death or disappearance]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(3):44–56, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401949.

Snyder:1934:GWT

[Sny34a] Carl Snyder. Growth of world trade vs. basic production. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(1):26–36, April 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401012.

Snyder:1934:NMT

[Sny34b] Carl Snyder. New measurements of trade, and of economic growth. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(4):5–15, January 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401095.

Snyder:1934:SRT

[Sny34c] Carl Snyder. On the statistical relation of trade, credit and prices. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(3):278–291, October 1934. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400986.

Somermeijer:1962:BRH

[Som62] W. H. Somermeijer. Book review: Household Survey Carried out in the Urban Areas of Greece during 1957–58. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(1):84, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402074.

Somermeyer:1967:BRE

[Som67a] W. H. Somermeyer. Book review: Essays on Econometrics and Planning by C. R. Rao. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(3):323– 325, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401804.

Somers:1967:BRE

[Som67b] R. H. Somers. Book review: Elementary Applied Statistics: For Students in Behavioral Science by L. C. Freeman. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):199–200, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1401410.

Sorin:1971:DCD

[Sor71] B. Sorin. La distribution combinatoire du nombre de différences premières positives. (French) [The combinatorial distribution of the number of positive first differences]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):9–20, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402225.

Spjotvoll:1968:CIT

[Spj68] E. Spjøtvoll. Confidence intervals and tests for variance ratios in unbalanced variance components models. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):37–42, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401339.

Sporea:1965:SM

[Spo65] C. Sporea. Sabin Manuila, 1894–1964. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(3):568–569, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401721.

Sprott:1971:BRC

[Spr71] D. A. Sprott. Book review: Combination of One-Sided Statistical Tests by J. Oosterhoff. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):114, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1402248.

Srivastava:1970:BRT

[Sri70] J. N. Srivastava. Book review: The Theory of Random Clumping by S. A. Roach. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):187–188, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1402339.

Shiskin:1963:BCI

[SS63] J. Shiskin and C. T. Saunders. Business cycle indicators: The known and the unknown. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31 (3):361–385, 1963. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401445.

Sengupta:1971:RSL

[ST71] Jati K. Sengupta and Gerhard Tintner. A review of stochastic linear programming. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):197–223, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1402174.

Stephan:1957:BRI

[Ste57] F. F. Stephan. Book review: The Industrial Mobility of Labor as a Probability Process by I. Blumen, M. Kogan, Ph. J. McCarthy. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):157–158, 1957. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401685.

Stephan:1960:BRS

[Ste60] F. F. Stephan. Book review: Survey Methods in Social Investigation by C. A. Moser. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(3):296, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401325.

Stefanov:1967:BRF

[Ste67] I. Stefanov. Book review: Les formules d’indices de prix. Calculs numériques et commentaires théoriques by J. Fourastié. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(1):95, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401644.

Stone:1948:AMD

[Sto48] Richard Stone. The analysis of market demand. An outline of methods and results. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 16(1/4): 23–35, 1948. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401230.

Stone:1950:BRN

[Sto50] Richard Stone. Book review: National Income Statistics of Various Countries, 1938–1947. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18 (1/2):130–131, 1950. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400767.

Stone:1956:BRC

[Sto56] R. J. N. Stone. Book review: Consumption and Business Fluctuations. National Bureau of Economic Research by R. P. Mack. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 24(1/3):141–142, 1956. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401285.

Stone:1958:BRI

[Sto58] R. Stone. Book review: The Income of Nations; Theory, Measurement and Analysis: Past and Present by P. Studenski. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):142, 1958. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401589.

Stone:1968:GDU

[Sto68] R. Stone. The generation, distribution and use of income. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):148–157, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401603.

Stuart:1933:P

[Stu33] C. A. Verrijn Stuart. Préface. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(1): 7–8, 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401080.

Stuart:1934:XSIb

[Stu34a] C. A. Verrijn Stuart. La XXIIe session de l’Institut International de Statistique, Londres 1934. (French) [The XXIInd session of the International Statistical Institute, London 1934]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(2):190–198, July 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1400997.

Stuart:1934:XSIa

[Stu34b] C. A. Verrijn Stuart. XXIe session de l’Institut International de Statistique. (French) [XXIst session of the International Statistical Institute]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(4):24–30, January 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401097.

Stuart:1935:CML

[Stu35] C. A. Verrijn Stuart. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3 (3):301–310, October 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400796.

Stuart:1936:XSI

[Stu36] C. A. Verrijn Stuart. La XXIIIème session de l’Institut International de Statistique, athènes 1936. (French) [The XXIII session of the International Institute of Statistics, Athens 1936]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(3):367–377, October 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400805.

Stuart:1938:MPC

[Stu38] C. A. Verrijn Stuart. La mortalité par cancer aux Pays-Bas. (French) [Cancer mortality in The Netherlands]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(2):219–228, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1400739.

Stuart:1941:BRS

[Stu41] C. A. Verrijn Stuart. Book review: Studies in American Demography by Walter F. Willcox. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 9(3/4):165, 1941. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400877.

Stuart:1964:MSP

[Stu64] A. Stuart. Multistage sampling with preliminary random stratification of first-stage units. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32 (3):193–201, 1964. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401864.

Stone:1950:USM

[SUD50] Richard Stone, J. E. G. Utting, and J. Durbin. The use of sampling methods in national income statistics and social accounting. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18(1/2):21–44, 1950. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400757.

Sukhatme:1952:MOE

[Suk52] P. V. Sukhatme. Measurement of observational errors in surveys. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 20(2):121–134, 1952. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401817.

Sukhatme:1969:VGP

[Suk69] P. V. Sukhatme. Vinayak Govind Panse, 1906–1969. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(3):347–348, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1402139.

Sunter:1971:BRP

[Sun71] A. B. Sunter. Book review: Practical Business Statistics by F. E. Croxton, D. J. Cowden, B. W. Bolch. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):240–241, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402182.

Sverdrup:1963:BRC

[Sve63] E. Sverdrup. Book review: Combinatorial Chance by F. N. David, D. E. Barton. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(1):143–144, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401738.

Sverdrup:1966:PSD

[Sve66] E. Sverdrup. The present state of the decision theory and the Neyman–Pearson theory. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34 (3):309–333, 1966. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401252.

Sykes:1968:BRS

[Syk68] Z. M. Sykes. Book review: Some Theory of Sampling by W. E. Deming. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):103–104, 1968. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401350.

Szameitat:1964:BRG

[Sza64a] K. E. J. W. Szameitat. Book review: A Guide to Algol Programming by D. D. McCracken. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32 (3):327–328, 1964. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401880.

Szameitat:1964:BRI

[Sza64b] K. E. J. W. Szameitat. Book review: An Introduction to Symbolic Programming by P. Wegner. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32 (3):326–327, 1964. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401879.

T:1936:BRN

[T.36] J. T. Book review: Narrenspiegel der Statistik by Ernst Wagemann. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(4):473, January 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401223.

Tagliacarne:1957:BRR

[Tag57] G. Tagliacarne. Book review: Rilevazioni statistiche by B. Barberi. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):155–156, 1957. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401683.

Takeuchi:1962:TDS

[Tak62] K. Takeuchi. A table of difference sets generating balanced incomplete block designs. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(3):361–366, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401745.

Takacs:1964:CMT

[Tak64] L. Takács. Combinatorial methods in the theory of queues. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):207–219, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401866.

Taylor:1968:PCE

[Tay68] L. D. Taylor. Personal consumption expenditure in Sweden 1931– 1958. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):19–36, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401338.

Theil:1962:FFE

[TB62] H. Theil and J. C. G. Boot. The final form of econometric equation systems. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(2):136–152, 1962. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401895.

Teghem:1962:BRE

[Teg62] J. Teghem. Book review: Exploitation des relevés expérimentaux by G. Coulmy. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(2):271, 1962.

CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401901.

Tekse:1965:DGD

[Tek65] K. Tekse. Describing the geographical distribution of the population. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(2):259–269, 1965. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402031.

Theil:1951:ETS

[The51] H. Theil. Estimates and their sampling variance of parameters of certain heteroscedastic distributions. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(2):141–147, 1951. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401241.

Theil:1952:TSE

[The52] H. Theil. On the time shape of economic microvariables and the Munich business test. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 20(2):105– 120, 1952. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401816.

Theil:1953:NHE

[The53] H. Theil. [A note on heteroscedastic errors in regression analysis]: Comment. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(1/2):29, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401515.

Theil:1955:MCB

[The55] H. Theil. The measurement of consumers’ behaviour. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 23(1/3):36–46, 1955. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1402009.

Theil:1957:SEE

[The57] H. Theil. Specification errors and the estimation of economic relationships. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):41–51, 1957. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401673.

Theodorescu:1969:BRM

[The69] R. Theodorescu. Book review: Les modèles stochastiques d’apprentissage by H. Rouanet. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):235, 1969. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402310.

Thirring:1933:BFF

[Thi33] Gustav Thirring. Der Begriff des “Fremden” in der Fremdenverkehrsstatistik. (German) [The concept of the “foreigner” in tourism statistics]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(3):40–46, October 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400775.

Thionet:1954:RFC

[Thi54] P. Thionet. Recherche d’une fonction de coût applicable aux enquêtes par sondage de l’I.N.S.E.E.. (French) [Search for a cost function applicable to I.N.S.E.E. sample surveys]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 22(1/3):33–47, 1954. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1401918.

Thionet:1959:LRS

[Thi59] P. Thionet. L’ajustement des résultats des sondages sur ceux des dénombrements. (French) [Adjusting survey results to those of enumerations]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 27(1/3):8–25, 1959. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402055.

Thirring:1963:ADb

[Thi63a] L. Thirring. Alexandre Dobrovits, 1889–1963. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(2):272–273, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401390.

Thirring:1963:ADa

[Thi63b] L. Thirring. Aloys Dolányi, 1877–1963. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(2):270–271, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401389.

Thirring:1963:BRA

[Thi63c] L. Thirring. Book review: Abstract of British Historical Statistics by B. R. Mitchell, Phyllis Deane. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(3):451–452, 1963. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401455.

Thirring:1963:DL

[Thi63d] L. Thirring. Dezsö Laky, 1887–1962. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(2):267–269, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401388.

Thirring:1964:BRH

[Thi64] L. Thirring. Book review: Hatvani István professzor (1718–1786) és a magyar statisztikai tudomány kezdetei (Le professeur Étienne Hatvani (1718–1786) et les origines de la statistique hongroise) by R. Horváth. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(1/2):183, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401986.

Thionet:1966:CTN

[Thi66a] P. Thionet. Sur certains tests non paramètriques bien connus. (French) [On some well-known non-parametric tests]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(1):13–26, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1401530.

Thirring:1966:GB

[Thi66b] L. Thirring. Gyula Barsy, 1909–1965. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):294–296, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401791.

Thionet:1968:CEA

[Thi68] P. Thionet. Contribution a l’étude des applications statistiques du problème du scrutin. (French) [Contribution to the study of statistical applications of the voting problem]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):296–308, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401470.

Thorp:1969:OGS

[Tho69] E. O. Thorp. Optimal gambling systems for favorable games. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(3):273–293, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402118.

Torrens-Ibern:1966:BRM

[TI66] J. Torrens-Ibern. Book review: Management Goals and Accounting for Control by Y. Ijiri. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(3):451– 452, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401269.

Tiao:1968:BRA

[Tia68] G. C. Tiao. Book review: Applied Regression Analysis by N. R. Draper, H. Smith. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):104–105, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401351.

Tiemann:1969:BRB

[Tie69] A. E. Tiemann. Book review: Balance of Payments Adjustment Policies; Japan, Germany, and the Netherlands by M. Michaely. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(3):325–326, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402128.

Tikkiwal:1967:TMS

[Tik67] B. D. Tikkiwal. Theory of multiphase sampling from a finite or an infinite population on successive occasions 1, 2. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(3):247–263, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401795.

Tinbergen:1936:DSP

[Tin36] J. Tinbergen. Sur la détermination statistique de la position d’équilibre cyclique. (French) [On the statistical determination of the cyclic equilibrium position]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(2): 173–188, July 1936. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400721.

Tinbergen:1941:UIE

[Tin41] J. Tinbergen. Unstable and indifferent equilibria in economic systems. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 9(1/2):36–50, 1941. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400943.

Tinbergen:1942:PDP

[Tin42] J. Tinbergen. Professor Douglas’ production function. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 10(1/2):37–48, 1942. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1401184.

Tinbergen:1944:CCE

[Tin44] J. Tinbergen. Colin Clark’s “Economics of 1960”. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 12(1/4):1–4, 1944. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1401067.

Tinbergen:1947:BRM

[Tin47a] J. Tinbergen. Book review: Des mouvements économiques généraux by Léon H. Dupriez. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 15 (1/4):127–129, 1947. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401209.

Tinbergen:1947:QEI

[Tin47b] J. Tinbergen. Quelques estimations de l’influence des contingentements 1933–1938 sur l’emploi aux Pays-Bas. (French) [Some estimates of the influence of the 1933–1938 quotas on employment in the Netherlands]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 15(1/4): 8–23, 1947. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401205.

Tinbergen:1950:BRE

[Tin50] J. Tinbergen. Book review: Economic Fluctuations in the United States, 1921–1941 by Lawrence R. Klein. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 18(3/4):260–262, 1950. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401041.

Tornqvist:1963:BRO

[Tör63a] L. Törnqvist. Book review: Ökonometrie by Professor Menges. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(1):143, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401737.

Tornqvist:1963:TRS

[Tör63b] L. Törnqvist. The theory of replicated systematic cluster sampling with random start. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(1):11–23, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401726.

Tornqvist:1968:BRC

[Tör68] L. Törnqvist. Book review: Consumer Demand in the United States, 1929–1970; Analyses and Projections by H. S. Houthakker, L. D. Taylor. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):234, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401617.

Tracy:1968:BRS

[Tra68] D. S. Tracy. Book review: Symmetric Function and Allied Tables by F. N. David, M. G. Kendall, D. E. Barton. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):351–352, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401479.

Theil:1956:EWM

[TvY56] H. Theil and J. van Yzeren. On the efficiency of Wald’s method of fitting straight lines. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 24(1/3): 17–26, 1956. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401277.

Ugge:1969:MB

[Ugg69] A. Uggè. Marcello Boldrini, 1890–1969. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(3):335–343, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402136.

Uppuluri:1968:BRI

[Upp68] V. R. Rao Uppuluri. Book review: Information Theory by R. B. Ash. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):95, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401344.

Vadja:1959:BRS

[Vad59] S. Vadja. Book reviews: Studies in Mathematical Theory of Inventory and Production by K. J. Arrow, S. Karlin, H. Scarf; Studies in Linear and Non Linear Programming by K. J. Arrow, L. Hurwicz, H. Uzawa. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 27(1/3):72, 1959. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402060.

Vajda:1965:BRG

[Vaj65] S. Vajda. Book review: Game Theory and Related Approaches to Social Behavior: Selections by Martin Shubik. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(2):323, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402036.

W:1934:BRE

[vBW34] N. L. J. v. B. W. Book review: Economic and Social Investigations in Manchester 1833–1933. A Centenary History of the Manchester Statistical Society by T. S. Ashton. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2 (2):234–235, July 1934. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401003.

vandenBrink:1961:BRD

[vdB61] T. van den Brink. Book review: Demographic and Economic Change in Developed Countries by Universities National Bureau Committee for Economic Research. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(3):94–95, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401959.

vandenBrink:1962:BRA

[vdB62] T. van den Brink. Book review: Armengaud, A. Histoire générale de la population mondiale by M. R. Reinhard. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(2):276–277, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401908.

vandenBrink:1963:BRG

[vdB63] T. van den Brink. Book review: Grundriss der Sozialwissenschaftlichen Statistik. Teil II: Besondere Statistik, Band 1: Bevölkerungsstatistik by P. Flaskämper. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(2):263, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401383.

vandenBrekel:1968:BRP

[vdB68] J. C. van den Brekel. Book review: Pratique de la démographie; Trente sujets d’analyse by R. Pressat. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(2):237–238, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401621.

vandenBeld:1969:BRN

[vdB69] C. A. van den Beld. Book review: National Economic Planning by M. F. Millikan. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):232–234, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1402307.

vandePanne:1965:BRE

[vdP65] C. van de Panne. Book review: Elements of Linear Programming with Economic Applications by R. C. Geary, M. D. McCarthy. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(3):561, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401714.

vandePanne:1970:BRO

[vdP70] C. van de Panne. Book review: Optimizing Nonlinear Models of Certain Transportation and Inventory Systems by S. Erlander. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(3):418–419, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1402213.

vanEeden:1955:STT

[vE55] Constance van Eeden. A sequential test with three possible decisions for comparing two unknown probabilities, based on groups of observations. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 23(1/3):20–28, 1955. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402007.

vanEeden:1961:SAP

[vE61] Constance van Eeden. Some approximations to the percentage points of the non-central T -distribution. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(1):4–31, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401552.

Venekamp:1968:BRM

[Ven68] P. E. Venekamp. Book review: Möglichkeiten und Grenzen der Erstellung und Auswertung regionaler Input-Output Tabellen by G. Strassert. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):362–363, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401493.

Verstege:1958:BRR

[Ver58] J. Ch. W. Verstege. Book review: Regionale Struktur- und Wirtschaftsforschung, Aufgaben und Methoden by O. Boustedt, H. Ranz. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):139, 1958. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401583.

Verstege:1961:BRS

[Ver61] J. Ch. W. Verstege. Book review: Stadtregionen in der Bundesrepublik Deutschland. Band XIV. Raum und Bevölkerung 1. Forschungsund Sitzungsbericht der Akademie für Raumforschung und Landesplanung. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 29(2):83, 1961. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401660.

Verstege:1963:BRP

[Ver63] J. C. W. Verstege. Book review: Pratique des enquêtes statistiques by G. R. Chevry. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 31(1):141, 1963. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401735.

Vessereau:1967:BRE

[Ves67] A. Vessereau. Book review: Exercices programmés de statistique à l’usage des médecins et des biologistes by A. J. Valleron, Ph. Lazar. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):208–209, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401419.

Vessereau:1968:BRD

[Ves68a] A. Vessereau. Book review: Dictionnaire de Statistique by E. Morice. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):359, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401488.

Vessereau:1968:BREc

[Ves68b] A. Vessereau. Book review: Eléments de probabilité et statistique à l’usage des étudiants en biologie humaine et générale by P. Lazar, D. Schwartz. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):111–112, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401360.

Vessereau:1968:BREb

[Ves68c] A. Vessereau. Book review: Eléments de théorie moderne des probabilités by Hérault. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):107–108, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401355.

Vessereau:1968:BREa

[Ves68d] A. Vessereau. Book review: Exercices de calcul des probabilités by G. Calot. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):99–100, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401348.

Vessereau:1968:BRI

[Ves68e] A. Vessereau. Book review: Incertitudes et décisions industrielles by J. Mothes. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):112–113, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401361.

Vessereau:1968:BRP

[Ves68f] A. Vessereau. Book review: Prévisions et décisions statistiques dans l’entreprise by J. Mothes. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):359–360, 1968. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401489.

Vessereau:1968:BRS

[Ves68g] A. Vessereau. Book review: Statistiques commentées by G. Reeb, A. Fuchs. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):360–361, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401491.

Vessereau:1970:BRL

[Ves70a] A. Vessereau. Book review: L’essai thérapeutique chez l’homme by D. Schwartz, R. Flamant, J. Lellouch. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(3):421–423, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402215.

Vessereau:1970:BRS

[Ves70b] A. Vessereau. Book review: Structures génétiques des populations by A. Jacquard. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(3):419–421, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1402214.

vHofsten:1966:PRC

[vH66] E. v. Hofsten. Population registers and computers: New possibilities for the production of demographic data. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):186–194, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401767.

Vinci:1936:DRD

[Vin36] Felice Vinci. La dimostrazione razionale della curva normale di probabilità e il libero arbitrio. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(2): 189–194, July 1936. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400722.

Vinci:1939:NRG

[Vin39] Felice Vinci. Una nuova rappresentazione grafica. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(2/3):138–144, 1939. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1401027.

Vincent:1956:BRL

[Vin56] P. Vincent. Book review: Life Tables from Limited Data; A Demographic Approach by G. J. Stolnitz. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 24(1/3):142–143, 1956. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401286.

Voight:1970:NHC

[Voi70] Robert B. Voight. The new haven census use study: An experience in the practical applications of statistical data. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(3):369–390, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1402202.

Vos:1964:SST

[Vos64] J. W. E. Vos. Sampling in space and time. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):226–241, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401868.

Vos:1965:BRF

[Vos65] J. W. E. Vos. Book review: Factor Analysis as a Statistical Method by D. N. Lawley, A. E. Maxwell. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(3):562, 1965. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401716.

Vos:1969:BRF

[Vos69] J. W. E. Vos. Book review: Formulae and Tables for Statistical Work by C. R. Rao, S. K. Mitra, A. Matthai. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):234–235, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402309.

Vos:1971:BRN

[Vos71] J. W. E. Vos. Book review: New Developments in Survey Sampling by N. L. Johnson, H. Smith, Jr. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):110, 1971. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402242.

vanRyzin:1968:BRP

[vR68] J. van Ryzin. Book review: Probability Theory by K. Krickeberg. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):111, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401359.

vanReeken:1970:BRS

[vR70] A. J. van Reeken. Book review: Systems of Frequency Curves by W. P. Elderton, N. L. Johnson. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):299–300, 1970. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1402158.

Vranic:1964:ADR

[Vra64] V. Vranić. On an application of duality in representing correlation. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(1/2):65–71, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401974.

vanZanten:1933:QCC

[vZ33] J. H. van Zanten. Quelques considérations sur la comparaison internationale des indices du coût de la vie et des salaires réels. (French) [Some considerations on the international comparison of cost of living indices and real wages]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(3): 20–39, October 1933. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400774.

vanZanten:1936:LSA

[vZ36] J. H. van Zanten. L’uniformité dans la statistique des accidents de la circulation. (French) [Uniformity in traffic accident statistics].

Revue de l’Institut international de statistique = Review of the International Statistical Institute, 4(2):260–265, July 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400729.

vanZanten:1938:BCD

[vZ38] J. H. van Zanten. Les bases de comparaison du nombre des accidents de la circulation routière. (French) [Bases for comparing the number of road traffic accidents]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(2):264–287, 1938. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1400744.

vanZwet:1966:BRE

[vZ66] W. R. van Zwet. Book review: Expected Values of Discrete Random Variables and Elementary Statistics by A. L. Edwards. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(3):448–449, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401266.

Wagemann:1938:SHM

[Wag38] Ernst Wagemann. Statistik als Hilfsmittel der medizinischen Forschung. Mit einem Beitrag über “Statistik der Kreislaufkrankheiten” von Dr. Boehm. (German) [Statistics as an aid to medical research. With an article on “Statistics of circulatory diseases” by Dr. Boehm]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(2):294– 304, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400746.

Wahlund:1956:GD

[Wah56] S. Wahlund. Gunnar Dahlberg, 1893–1956. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 24(1/3):160–161, 1956. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401295.

Walldin:1962:SPS

[Wal62] K. E. Walldin. Stochastic processes with stationary and independent increments, I. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(3):293–335, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401742.

Walsh:1966:BRC

[Wal66] J. E. Walsh. Book review: Convex Transformations of Random Variables by W. R. van Zwet. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(1):116, 1966. CODEN ISTRDP. ISSN 0373-1138 (print),

2212-1846 (electronic). URL https://www.jstor.org/stable/
1401546.

Wallace:1967:BRA

[Wal67] D. L. Wallace. Book review: The Analysis of Variance by A. Huitson. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(2):204–205, 1967. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401415.

Wegmuller:1957:AK

[Weg57] W. Wegmüller. Albert Koller, 1896–1957. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):190–191, 1957. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1401694.

Wegmuller:1964:BRM

[Weg64] W. Wegmüller. Book review: Méthodes statistiques dans les sciences humaines by P. Pèpe, M. Tisserand-Perrier. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(1/2):185, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1401988.

Werthamer:1967:BRQ

[Wer67] N. R. Werthamer. Book review: Quantum Field Theoretical Methods in Statistical Physics by A. A. Abrikosov, L. P. Gor’kov, I. Ye. Dzyaloshinskii. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(3):321, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401801.

Westergaard:1934:SM

[Wes34] Herald Westergaard. On statistical methods. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(2):161–177, July 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1400995.

Wetherill:1970:BRM

[Wet70] G. B. Wetherill. Book review: Mathematics for Introductory Statistics: A Programmed Review by A. R. Baggaley. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(2):294, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402153.

Weyr:1938:PWS

[Wey38] František Weyr. Das Problem der Willensfreiheit in der Statistik. Bemerkungen zu dem gleichnamigen Aufsatze von Univ. Prof. Dr.Wilhelm Winkler, Wien. (German) [The problem of free will in statistics. Remarks on the essay of the same name by Univ. Prof. Dr. Wilhelm Winkler, Vienna]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(4):329–334, January 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400907.

Whittle:1966:BRI

[Whi66] P. Whittle. Book review: Introductory Engineering Statistics by I. Guttman, S. S. Wilks. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(3):450, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401268.

Whittle:1969:BRW

[Whi69] P. Whittle. Book review: S. S. Wilks’ Collected Papers by T. W. Anderson, S. S. Wilks. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):220, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1402295.

Wurzburger:1937:CML

[WI37] Eugen Würzburger and Philip J. Idenburg. Communications sur la méthodologie, la législation, l’organisation et l’administration statistiques. (French) [Communications on statistical methodology, legislation, organization and administration]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(3):227–239, October 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400954.

Widen:1967:BRR

[Wid67] L. Widén. Book review: Regional Model Life Tables and Stable Populations by A. J. Coale, P. Demeny. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(1):93–94, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401642.

Wijngaard:1969:BRP

[Wij69] J. Wijngaard. Book review: Problems of Stocks and Storage by A. J. H. Morrell. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(2):234, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402308.

Willcox:1935:IMM

[Wil35] Walter F. Willcox. An improved method of measuring public health in the United States. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(1):5–13, April 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1401046.

Willcox:1936:DS

[Wil36] Walter F. Willcox. Definitions of statistics. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(4):388–399, January 1936. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401217.

Willcox:1937:PCM

[Wil37] Walter F. Willcox. The population of China and its modern increase. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(1):3–14, April 1937. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400921.

Willcox:1938:FS

[Wil38] Walter F. Willcox. The founder of statistics. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(4):321–328, January 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400906.

Willcox:1955:GFS

[Wil55] Walter F. Willcox. George Findlay Shirras, 1885–1955. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 23(1/3):144–145, 1955. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402025.

Wilks:1959:BRA

[Wil59] S. S. Wilks. Book review: The Advanced Theory of Statistics. I. Distribution Theory by M. G. Kendall, A. Stuart. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 27(1/3):71–72, 1959. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1402059.

Williams:1968:BRS

[Wil68] E. J. Williams. Book review: Statistics in Biology; Vol. 1 by C. I. Bliss. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):96–99, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401346.

Winkler:1937:PWS

[Win37] Wilhelm Winkler. Das Problem der Willensfreiheit in der Statistik. (German) [The problem of free will in statistics]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(2):115–131, July 1937. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1400830.

Winkler:1938:GSU

[Win38a] Wilhelm Winkler. Der Geburtenrückgang, seine Ursachen und seine wirtschaftlichen und sozialen Wirkungen. (German) [The decline in birth rates, its causes and its economic and social effects]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(1):56–102, 1938. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400842.

Winkler:1938:EAN

[Win38b] Wilhelm Winkler. Einige alte und neue Masze des natürlichen Bevölkerungswachstums. (German) [Some old and new measures of natural population growth]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(1): 25–49, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846

(electronic). URL https://www.jstor.org/stable/1400840.

Winkler:1939:ORP

[Win39] Wilhelm Winkler. Observations sur le rapport présenté par M. le Prof. Winkler: “La baisse de la natalité, ses causes, et ses conséquences économiques et sociales”: Schlusswort des Referenten. (French) [Observations on the report presented by Prof. Winkler: “The decline in the birth rate, its causes, and its economic and social consequences”: Response]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 7(4):259–268, 1939. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1400964.

Winkler:1942:SBZ

[Win42] Wilhelm Winkler. Die stationäre Bevölkerung. Zugleich ein Beitrag zur Sterblichkeitsmessung. (German) [Stationary population. At the same time a contribution to mortality measurement]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 10(1/2):49–74, 1942. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401185.

Winkler:1944:LBB

[Win44] Wilhelm Winkler. Die Lebensjahre einer Bevölkerung. Bemerkungen zu Liebmann Hersch’s “Potentiel-vie”. (German) [The years of life of a population. Comments on Liebmann Hersch’s “Potentielvie”]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 12(1/4):5–22, 1944. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401068.

Winkler:1957:BRW

[Win57] W. Winkler. Book review: Wirtschaftskunde der Bundesrepublik Deutschland. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 25(1/3):152–153, 1957. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401679.

Winkler:1971:SS

[Win71] W. Winkler. Slawtscho Sagoroff, 1898–1970. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):123–125, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402258.

Wise:1971:BRM

[Wis71] M. E. Wise. Book review: Mathematik für Mediziner und Biologen by G. Fuchs. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(1):109–110, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402241. See correction

[Ano71a].

Wold:1951:BRR

[Wol51] H. Wold. Book review: Recherches théoriques modernes sur le calcul des probabilités. I: Généralités sur les Probabilités. Eléments aléatoires by M. Fréchet. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 19(1):109, 1951. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401503.

Wold:1955:BRI

[Wol55] H. Wold. Book review: An Introduction to Stochastic Processes, with Special Reference to Methods and Applications by M. S. Bartlett. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 23(1/3):125, 1955. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402013.

Wold:1958:CSI

[Wol58] H. O. A. Wold. A case study of interdependent versus causal chain systems. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 26(1/3):5–25, 1958. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401568.

Wold:1967:BRR

[Wol67] H. Wold. Book review: Random Processes and the Growth of Firms. A Study of the Pareto Law by J. Steindl. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 35(3):329–330, 1967. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor. org/stable/1401811. See correction [Ano68d].

Wold:1968:BRE

[Wol68] H. O. Wold. Book review: Econometric Models and Methods by C. F. Christ. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(1):100–103, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401349.

Wold:1970:RFJ

[Wol70] Herman Wold. Ragnar Frisch and Jan Tinbergen, winners of the first Alfred Nobel Memorial Prize in Economic Science. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 38(1):??, 1970. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1402316.

Woolsey:1968:BRA

[Woo68] T. D. Woolsey. Book review: After Eighty Years by L. I. Dublin. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 36(3):353, 1968. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401481.

Wurzburger:1935:WLS

[Wür35] Eugen Würzburger. Was lehrt die Sterblichkeitsstatistik?. (German) [What do mortality statistics teach?]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 3(2):129–141, July 1935. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.

org/stable/1400852.

Wagemann:1938:KAW

[WW38] Ernst Wagemann and R. Wagenführ. Kurzfristig anfallende Wirtschaftsstatistiken und ihre Bedeutung für die Wirtschaftsbeobachtung in Deutschland. (German) [Short-term economic statistics and their significance for economic observation in Germany]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(2):305–315, 1938. CO-DEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400747.

Yates:1962:SRA

[Yat62] F. Yates. Sir Ronald Aylmer Fisher (1890–1962). Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(2):280–282, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1401912.

Yoshizawa:1971:TS

[Yos71] T. Yoshizawa. The teaching of statistics. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(3):348–350, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1402270.

Zahn:1933:MIF

[Zah33] F. Zahn. Un mot d’introduction. (French) [A word of introduction]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(1):5–6, 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401079.

Zahn:1934:WMS

[Zah34] Friedrich Zahn. Der Wert des Menschen in der Statistik. (German) [The value of people in statistics]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 2(3):265–277, October 1934. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400985.

Zarkovic:1956:ISC

[Žar56] S. S. Žarković. An illustration of some characteristic situations in the application of the difference estimate. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 24(1/3):52–63, 1956. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1401279.

Zarkovic:1962:BRF

[Zar62a] S. S. Zarkovic. Book review: Family Living Studies. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(1):84–86, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1402075.

Zarkovich:1962:BRC

[Zar62b] S. S. Zarkovich. Book review: Curso de muestreo y aplicaciones by F. Azorin Poch. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 30(3):405–406, 1962. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (elec-

tronic). URL https://www.jstor.org/stable/1401754.

Zaremba:1965:KDC

[Zar65] S. K. Zaremba. Kurtosis and determinate components in linear processes. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 33(3):378–413, 1965. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1401702.

Zarkovich:1966:BRS

[Zar66] S. S. Zarkovich. Book review: Survey Sampling by L. Kish. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 34(2):276–277, 1966. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401780.

Zarkovich:1969:BRS

[Zar69] S. S. Zarkovich. Book review: Sampling Theory and Methods by M. N. Murthy. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(1):110, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1402105.

Zahn:1942:LIS

[ZH42] Friedrich Zahn and Wilhelm Henninger. Die Langlebigkeit in der internationalen Statistik. (German) [Longevity in international statistics]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 10(3/4):119–138, 1942. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic).

URL https://www.jstor.org/stable/1400894.

Zimmerman:1969:BRT

[Zim69] L. J. Zimmerman. Book review: The Theory and Design of Economic Development by I. Adelman, E. Thorbecke. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 37(3):321–322, 1969. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.

jstor.org/stable/1402123.

Zizek:1933:LGS

[Ziz33] Franz Zizek. Der logische Grundcharakter der statistischen Zahlen. (German) [The basic logical character of statistical numbers]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 1(3):7–19, October 1933. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1400773.

Zizek:1938:VBS

[Ziz38] Franz Zizek. Die verschiedenen Begriffe von “Statistik”. (German) [The different terms of “statistics”]. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 6(4):519–552, 1938. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://www.jstor.org/

stable/1400884.

Zwingli:1953:BRB

[Zwi53] U. Zwingli. Book review: Bibliography on Income and Wealth. Cambridge, Bowes and Bowes, 1952-’53 by D. Creamer, Ph. Deane. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 21(1/2):82, 1953. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401519.

Zwingli:1960:BRW

[Zwi60] U. Zwingli. Book review: The World’s Metropolitan Areas. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28(3):297–298, 1960. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1401327.

Zwingli:1971:BRW

[Zwi71] U. Zwingli. Book review: World Urbanization 1950–1970; Vol. 1. Basic Data for Cities, Countries and Regions by K. Davis. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 39(2):241–242, 1971. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL https://

www.jstor.org/stable/1402183.